### v2

# Dual-Directed Algorithm Design for Efficient Pure Exploration

Abstract
We consider pure-exploration problems in the context of stochastic sequential adaptive experiments with a finite set of alternatives. The central objective is to answer a query regarding the alternatives with high confidence while minimizing measurement efforts. One canonical example is identifying the best-performing alternative, a problem known as ranking and selection in simulation or best-arm identification in machine learning. We formulate the problem complexity measure as a maximin optimization problem for the static fixed-budget, fixed-confidence, and posterior convergence rate settings. By incorporating dual variables directly into the analysis, we derive necessary and sufficient conditions for an allocation’s optimality. The introduction of dual variables allows us to sidestep the combinatorial complexity that arises when considering only primal variables. These optimality conditions enable the extension of the top-two algorithm design principle to more general pure-exploration problems. Moreover, our analysis yields a straightforward and effective information-directed selection rule that adaptively chooses from a candidate set based on the informational value of the candidates. We demonstrate the broad range of contexts in which our design principle can be implemented. In particular, we establish that, when combined with information-directed selection, top-two Thompson sampling achieves asymptotic optimality in Gaussian best-arm identification, resolving a notable open question in the pure-exploration literature. Additionally, our algorithm attains optimality in $\varepsilon$-best-arm identification (or ranking and selection with a probability of good selection guarantee) and thresholding bandits. Furthermore, our results provide a general principle for adapting Thompson sampling to general pure-exploration problems. Extensive numerical experiments highlight the efficiency of our proposed algorithms compared to existing methods.
1 Introduction
In stochastic sequential adaptive experiments, a decision-maker (DM) often gathers noisy data from a finite set of designs or alternatives. A prevalent objective is to allocate measurement efforts efficiently and identify the top-performing option with high confidence while minimizing the required measurements. This problem is widely recognized as the ranking and selection (R&S) problem, which is frequently referred to as the best-arm-identification (BAI) problem in the machine learning literature. R&S problem serves as a critical abstraction for numerous applications. For instance, it is used to determine optimal hyperparameters for algorithms (Shang et al., 2019). In large language models, it helps identify the best beam search direction for automatic prompt optimization (Pryzant et al., 2023). Recently, it has been applied in brain-computer interface systems to interpret brain signals under various stimulus paradigms (Zhou et al., 2023). For a detailed discussion on the connection between BAI and R&S, see Glynn and Juneja (2015).
A defining feature of R&S and BAI problems is their focus on end outcomes. In these settings, the DM is not penalized for losses incurred during the adaptive trials. Instead, the primary concern lies in the accuracy of the final decision, which often carries significant consequences. For instance, the final decision might result in the deployment of a “winning” feature, the construction of infrastructure, or the approval of a new drug. The objective is to accurately identify the best alternative by the end of the experiment, as resources are typically committed to the selected option. Unlike scenarios that prioritize immediate gains from exploiting high-performing alternatives, these problems emphasize exploration to ensure an informed final decision. Consequently, such sequential decision problems are often referred to as pure-exploration problems.
The study of pure-exploration problems dates back to the 1950s in the context of R&S (Bechhofer, 1954). Since then, a wide range of algorithms have been developed to address the canonical R&S problem in both the simulation and statistical learning communities. Examples include elimination algorithms (Kim and Nelson, 2001; Fan et al., 2016; Bubeck et al., 2013), Optimal Computing Budget Allocation (OCBA) algorithms (Chen et al., 2000; Gao and Chen, 2015; Wu and Zhou, 2018b), knowledge-gradient algorithms (Frazier et al., 2008), and UCB-type algorithms (Kalyanakrishnan et al., 2012; Kaufmann and Kalyanakrishnan, 2013). For a comprehensive overview of R&S, we recommend the survey articles by Fu and Henderson (2017); Hong et al. (2021); Fan et al. (2024).
1.1 Motivations
A unified approach to different problem formulations.
Various formulations have been developed to evaluate algorithm performance in pure exploration.
The fixed-budget problem is often framed as a stochastic dynamic programming problem. However, the complexity of solving such problems exactly has led to a focus on approximate solutions, such as restricting analysis to procedures based on fixed static allocations. Building on OCBA (Chen et al., 2000), Glynn and Juneja (2004) provide a large-deviation characterization of the probability of incorrect selection (PICS) for static-allocation procedures, identifying an optimal static allocation that maximizes the exponential decay rate of PICS. This approach introduces a notion of problem complexity as the optimal decay exponent. Subsequent works, such as Chen and Ryzhov (2023), build upon this framework by designing dynamic procedures that converge to the optimal static allocation.
The fixed-precision formulation, also referred to as the fixed-confidence setting in the bandit literature, traces its origins to Chernoff (1952)’s seminal work on sequential binary hypothesis testing. R&S and BAI extend this to multiple hypotheses with sequentially chosen experiments. In this setting, the DM determines when to terminate the experiment and propose an estimated top performer. The goal is to minimize the expected sample size while guaranteeing a pre-specified probability of correct selection (PCS). Garivier and Kaufmann (2016) provided matching lower and upper bounds on the sample complexity, framing it as an optimization problem that prescribes the optimal sample allocation ratio.
The posterior convergence rate setting (Russo, 2020) maintains a posterior distribution over the unknown parameters and evaluates the algorithms by the posterior PICS. Interestingly, this approach operates within a frequentist framework, where the problem instance is fixed, yet it adopts a Bayesian objective to evaluate performance. Russo (2020) identifies the optimal decay exponent for posterior PICS as the complexity measure in this setting.
Remarkably, for Gaussian noise, the problem complexity optimization problems across the three formulations are identical, as discussed in Section 2.5.2. Despite extensive research on individual formulations of R&S and BAI, a unified algorithmic framework capable of addressing all three major formulations simultaneously remains an open challenge.
A unified approach to general pure exploration problems.
While selecting the best alternative is a common example, many other problems fall under the umbrella of pure-exploration in sequential adaptive experiments. These problems broadly aim to resolve an exploration query with high confidence while minimizing sampling resources. For instance, the DM may want to identify alternatives sufficiently close to the best-performing one, often referred to as R&S with a probability of good selection (PGS) guarantee or $\varepsilon$-BAI in the machine learning literature. Other objectives include identifying all alternatives exceeding a specified performance threshold (Locatelli et al., 2016), selecting the best-$k$ alternatives (Chen et al., 2008), or determining whether the mean performance lies within a convex feasible set (Qiao and Tewari, 2023).
Despite scattered studies across the literature addressing specific exploration tasks, developing a unified algorithmic framework capable of handling such diverse queries remains an open and critical challenge.
Generalizing top-two and Thompson sampling algorithms.
A recent class of algorithms, known as top-two algorithms, initially introduced by Russo (2020), has gathered increasing attention. These algorithms employ a two-phase approach for selecting the next alternatives to sample. In the first phase, a pair of top-two candidates – referred to as the leader and the challenger – are identified. The second phase, called the selection step, involves sampling one of these candidates, often based on a potentially biased coin toss. [^1] We use the term “sampling” specifically to refer to the act of picking an alternative to draw a sample from, and reserve the term “selection” to describe the step of determining which candidate should be sampled. The simplicity and effectiveness of top-two algorithms have led to their widespread adoption, with notable contributions from Qin et al. (2017), Shang et al. (2020), and Jourdan et al. (2022).
However, a notable limitation of these algorithms is their reliance on a tuning parameter to achieve optimality, which necessitates additional adaptive tuning steps. Whether simple, parameter-free algorithms that are provably optimal can be developed remains an open question, see Russo (2020, Section 8).
Thompson sampling (TS) has demonstrated remarkable empirical success in minimizing regret for stochastic bandits, gaining significant attention from academia and industry (Russo et al., 2018). However, its inherent focus on maximizing cumulative rewards leads to excessive exploitation of the estimated best alternative, resulting in insufficient exploration of perceived suboptimal options to confidently verify whether the perceived best alternative is indeed the best. Consequently, modifications to TS are necessary for its effective application to pure exploration. Adapting TS for pure exploration could have an immediate impact in areas where TS is already widely used, such as A/B testing (Graepel et al., 2010), news article recommendation (Chapelle and Li, 2011), and revenue management (Ferreira et al., 2018). Several studies have proposed TS variants for pure exploration, beginning with the Top-Two Thompson Sampling (TTTS) algorithm (Russo, 2020). Other adaptations include the Murphy sampling algorithm introduced by Kaufmann et al. (2018), which modifies classical TS to test whether the lowest mean falls below a given threshold.
However, the existing literature on top-two algorithms has primarily concentrated on BAI, leaving general pure-exploration queries largely unaddressed. Similarly, for TS, there is a notable absence of unified guidance on systematically tailoring TS to effectively handle general pure-exploration queries.
1.2 Contributions and outline
This paper proposes a unified algorithm design framework that accommodates diverse problem formulations and a wide range of pure exploration queries. Our main contributions are summarized as follows.
  1. 1.
    We identify common structural elements across various problem formulations and exploration queries, formulating the problem complexity into a distinctive non-smooth maximin convex optimization problem. This formulation characterizes the inherent difficulty of each query and serves as the foundation for our unified algorithmic framework. Guided by the shared structure of these problem complexity optimization problems, our framework seamlessly adapts to diverse pure exploration tasks.
  2. 2.
    We formulate the necessary and sufficient conditions for the optimality of problem complexity optimization problems using dual variables, offering a novel perspective on existing algorithms and laying the groundwork for generalizations. Specifically, our use of dual variables sheds new light on the foundations of top-two algorithms and variants of Thompson Sampling, revealing their basis in the complementary slackness condition of the Karush-Kuhn-Tucker (KKT) conditions. This enables a natural generalization of the top-two principle to diverse pure-exploration problems. Furthermore, the dual perspective supplies top-two algorithms with an adaptive and provably optimal selection step, solving an open problem in Russo (2020).
  3. 3.
    Our design principle is highly flexible and easily adaptable to practical settings with minimal modifications. We demonstrate the versatility of our algorithm framework through its application to Gaussian noises with unequal and unknown variances, correlated rewards using common random numbers, and general pure-exploration problems. Additionally, we offer a systematic and accessible approach to tailoring Thompson sampling for general pure-exploration problems.
  4. 4.
    We provide theoretical guarantee for the opimality of our algorithms in various formulations and exploration queries. In particular, we show that our algorithm is optimal for R&S problems with Gaussian noise under all three major formulations – fixed-budget, fixed-precision, and posterior convergence rate. We further establish optimality of our framework in $\varepsilon$-BAI (PGS) and thresholding bandits.
  5. 5.
    We conduct extensive numerical studies to showcase the compelling performance of our unified algorithm across various settings. These experiments encompass a wide range of scenarios, including the fixed-confidence and fixed-budget settings, large-scale problems, applications with common random numbers, and general exploration tasks beyond conventional R&S and BAI. This comprehensive evaluation highlights the performance and robustness of our unified algorithm.
The rest of the paper is organized as follows. In Section 2, we present a teaser study on indifference-zone-free R&S (BAI) with Gaussian noise, demonstrating its optimality in posterior convergence rate and static fixed-budget settings. Section 3 introduces the problem formulation for general pure exploration and provides detailed explanations of our unified algorithm design principle, where we also present optimality for Gaussian BAI in fixed-confidence setting. In Section 4, we discuss extensions beyond the base cases, including exploration queries with multiple correct answers (probability of good selection guarantees) and Gaussian noise with unknown variances. Finally, in Section 5, we conduct an extensive numerical study across various scenarios and discuss practical considerations for algorithm implementation.
2 A teaser: ranking and selection in Gaussian models
To motivate our algorithm design principle, we focus on ranking and selection in Gaussian models with known but potentially unequal variances. [^2] The case of unknown variances is discussed in Section 4, where we also explore an alternative formulation involving the probability of good selection. We provide a detailed analysis of our algorithm within the canonical indifference-zone-free ranking and selection (R&S) problem, also known as the best arm identification (BAI) problem for unstructured bandits in the bandit literature. Our analysis yields optimal algorithms for R&S in the posterior convergence rate setting (Russo, 2020).
2.1 Problem formulation
There are $K$ systems with unknown performance, and a DM aims to identify the best system after sequentially allocating a finite sampling budget $T$ among these $K$ systems. For any system $i\in[K]\triangleq\{1,2,\dots,K\}$, its performance is characterized by an unknown parameter $\theta_{i}\in\mathbb{R}$. At each time step $t\in\mathbb{N}_{0}\triangleq\{0,1,\ldots\}$, the DM selects a system $I_{t}\in[K]$, and then observes a random performance $Y_{t+1,I_{t}}=\theta_{I_{t}}+W_{t,I_{t}}$, where $W_{t,I_{t}}\sim N(0,\sigma_{I_{t}}^{2})$ is Gaussian with a system-dependent variance $\sigma_{I_{t}}^{2}$, drawn independently across time. We assume that the variances $(\sigma_{1}^{2},\ldots,\sigma_{K}^{2})$ are known, and there is a unique best system, denoted by $I^{*}\triangleq\operatorname*{argmax}_{i\in[K]}\theta_{i}$. The goal is to identify this unknown best system $I^{*}$ efficiently.
While we study this problem in a frequentist setting, where the unknown mean parameters $(\theta_{1},\ldots,\theta_{k})$ are fixed but unknown, in this section, we focus on Bayesian sampling rules, which involve maintaining posterior beliefs over the unknown mean parameters. Our objective is to minimize the posterior probability that the best system $I^{*}$ is not the best, an objective popularized by Russo (2020). The performance of an algorithm is evaluated based on the rate at which this posterior probability decreases to zero as the number of samples grows. This can be understood as a “posterior fixed-budget setting.” Although this setting is less common in the simulation literature, it isolates the core principle of our algorithm design.
Prior and posterior distributions.
For each system $i\in[K]$, we assume that the prior over its mean performance $\theta_{i}$ is an uninformative Gaussian distribution (with infinite prior variance), independent of the priors over other systems’ mean performances. At the beginning of round $t\in\mathbb{N}_{0}$, the information available to the DM is $\mathcal{H}_{t}\triangleq(I_{0},Y_{1,I_{0}},\dots,I_{t-1},Y_{t,I_{t-1}})$. Due to the uninformativeness and independence of the priors, the posterior distribution over $\theta_{i}$ is $N(\theta_{t,i},\sigma^{2}_{t,i})$, where the posterior mean and variance are given by:
$\theta_{t,i}\triangleq\frac{\sum_{\ell=0}^{t-1}{\mathds{1}(I_{\ell}=i)Y_{\ell+1,I_{\ell}}}}{N_{t,i}}\quad\text{and}\quad\sigma^{2}_{t,i}\triangleq\frac{\sigma_{i}^{2}}{N_{t,i}}$
with $N_{t,i}\triangleq\sum_{\ell=0}^{t-1}\mathds{1}(I_{\ell}=i)$ being the number of samples allocated to system $i$ before time $t$. We denote the posterior distribution over $(\theta_{1},\ldots,\theta_{K})$ by $\Pi_{t}$.
Let $\Theta_{I^{*}}\triangleq\{(\vartheta_{1},\ldots,\vartheta_{K})\in\mathbb{R}^{K}:\vartheta_{I^{*}}>\vartheta_{j},\forall j\neq I^{*}\}$ be the set of parameter vectors such that the true best system $I^{*}$ is indeed the unique best. Our objective is then to minimize $\Pi_{t}(\Theta^{c}_{I^{*}})$ where the set $\Theta^{c}_{I^{*}}$ is the complement of $\Theta_{I^{*}}$. This quantity is the posterior probability that the true best system $I^{*}$ is not the best. For a well-behaved algorithm (e.g., uniform allocation), this posterior probability typically converges to zero exponentially fast as $t\to\infty$. We are seeking sampling rules that maximize the asymptotic exponent of the exponential decay, characterized by $\liminf_{t\to\infty}-\frac{1}{t}\log\Pi_{t}(\Theta^{c}_{I^{*}})$.
Instance-optimality.
This section focuses on a fixed mean parameter vector $(\theta_{1},\ldots,\theta_{K})$, so the notion of optimality is tailored to this specific instance. An allocation rule is deemed optimal if it achieves the maximal asymptotic rate of exponential decay for the posterior probability of the set $\Theta^{c}_{I^{*}}$, as defined below.
Definition 1 (Instance-optimal sampling rules).
A sampling rule $\textup{{A}}^{*}$ is optimal if, for any other non-anticipating sampling rule A,
$\liminf_{t\to\infty}\frac{-\log\Pi_{t}^{\textup{{A}}^{*}}(\Theta^{c}_{I^{*}})}{-\log\Pi_{t}^{\textup{{A}}}(\Theta^{c}_{I^{*}})}\geq 1,$
where the superscript over $\Pi_{t}$ denotes the dependence on the sampling rule under which the information is collected.
A sufficient condition for optimality.
We provide a sufficient condition on a sampling rule to be optimal in terms of its asymptotic allocation among all systems. Denote $p_{t,i}\triangleq\frac{N_{t,i}}{t}$ be the fraction of samples allocated to arm $i\in[K]$, and let $\bm{p}_{t}\triangleq(p_{t,1},\ldots,p_{t,K})$ be the empirical allocation vector.
Theorem 1 (A sufficient condition for optimality).
A sampling rule is (instance-)optimal if $\bm{p}_{t}$ converges almost surely to a unique probability vector $\bm{p}^{*}$ satisfying
  1. 1.
    information balance:
    $\frac{(\theta_{I^{*}}-\theta_{j})^{2}}{2\left(\sigma_{I^{*}}^{2}/p_{I^{*}}+\sigma_{j}^{2}/p_{j}\right)}=\frac{(\theta_{I^{*}}-\theta_{j^{\prime}})^{2}}{2\left(\sigma_{I^{*}}^{2}/p_{I^{*}}+\sigma_{j^{\prime}}^{2}/p_{j^{\prime}}\right)},\quad\forall j,j^{\prime}\neq I^{*},$ (1)
  2. 2.
    stationarity:
    $p^{*}_{I^{*}}=\sum_{j\neq I^{*}}h^{j}_{I^{*}}(\bm{p}^{*})\cdot\frac{p_{j}^{*}}{1-h^{j}_{I^{*}}(\bm{p}^{*})},\quad\text{where}\quad h^{j}_{I^{*}}(\bm{p}^{*})\triangleq\frac{\sigma_{I^{*}}^{2}/p^{*}_{I^{*}}}{\sigma_{I^{*}}^{2}/p^{*}_{I^{*}}+\sigma_{j}^{2}/p^{*}_{j}},\quad\forall j\neq I^{*}.$ (2)
One may observe that the stationarity condition (2) is a rewrite of the conventional “balance of sum-of-squares”
$\frac{(p^{*}_{I^{*}})^{2}}{\sigma^{2}_{I^{*}}}=\sum_{j\neq I^{*}}\frac{(p^{*}_{j})^{2}}{\sigma^{2}_{j}},$ (3)
which is observed in various settings, e.g., the static fixed-budget setting by Glynn and Juneja (2004, Theorem 1) and Chen and Ryzhov (2023, Lemma 1), the fixed-confidence setting by Bandyopadhyay et al. (2024, Theorem 2.1) and the posterior convergence setting by Garivier and Kaufmann (2016, Theorem 5).
We now demonstrate how the information balance condition (1) and the stationarity condition (2) in Theorem 1 guide the design of sampling rules.
2.2 Top-two design principle as information balance condition
We show that the information balance condition in (1) naturally gives rise to the well-known top-two design principle introduced by Russo (2020). To illustrate this connection, we examine the Top-Two Thompson Sampling (TTTS) algorithm proposed in their work. This condition also underpins a series of related “top-two” algorithms, including those introduced in Qin et al. (2017); Shang et al. (2020); Jourdan et al. (2022); Bandyopadhyay et al. (2024).
2.2.1 Top-two Thompson sampling
The TTTS algorithm maintains a posterior distribution $\Pi_{t}$ over the true mean vector $\bm{\theta}=(\theta_{1},\ldots,\theta_{K})$. In each iteration $t$, TTTS first generate a sample $\tilde{\bm{\theta}}=(\tilde{\theta}_{1},\ldots,\tilde{\theta}_{K})$ from the posterior distribution $\Pi_{t}$ and identifies one best system under this sample, denoted by $I_{t}^{(1)}\in\operatorname*{argmax}_{i\in[K]}\tilde{\theta}_{i}$. The system $I_{t}^{(1)}$ is referred to as the leader. TTTS then repeatedly samples $\tilde{\bm{\theta}}^{\prime}=(\tilde{\theta}^{\prime}_{1},\ldots,\tilde{\theta}^{\prime}_{K})$ from $\Pi_{t}$ until another system $I_{t}^{(2)}\in\operatorname*{argmax}_{i\in[K]}\tilde{\theta}^{\prime}_{i}$ differs from $I_{t}^{(1)}$. We refer to this system $I_{t}^{(2)}$ as the challenger, and the pair $(I_{t}^{(1)},I_{t}^{(2)})$ is referred to as the top-two candidates.
Algorithm 1 TTTS sub-routine for best-arm identification (Russo, 2020)
1:Posterior distribution $\Pi_{t}$.
2:Sample $\tilde{\bm{\theta}}=\left(\tilde{\theta}_{1},\ldots,\tilde{\theta}_{K}\right)\sim\Pi_{t}$ and define $I_{t}^{(1)}\in\operatorname*{argmax}_{i\in[K]}\tilde{\theta}_{i}$
3:Repeatedly sample $\tilde{\bm{\theta}}^{\prime}=\left(\tilde{\theta}^{\prime}_{1},\ldots,\tilde{\theta}^{\prime}_{K}\right)\sim\Pi_{t}$ until $I_{t}^{(2)}\in\operatorname*{argmax}_{i\in[K]}\tilde{\theta}^{\prime}_{i}$ differs from $I_{t}^{(1)}$
4:Return $\left(I_{t}^{(1)},I_{t}^{(2)}\right)$
Given the top-two candidates $(I_{t}^{(1)},I_{t}^{(2)})$, TTTS measures only one of them using a potentially biased coin. For example, TTTS measures the system $I_{t}^{(1)}$ with probability $\beta$ and $I_{t}^{(2)}$ with probability $1-\beta$, where $\beta$ is the coin bias, typically set to $\beta=0.5$. A prominent open question in the top-two algorithm literature is how to tune (or select) the parameter $\beta$ to achieve the exact optimality. In Section 2.3, we introduce an optimal, adaptive selection rule called information-directed selection (IDS), which dynamically determines the coin bias at each iteration. Before doing so, we establish the connection between the TTTS sub-routine and the information balance condition (1).
2.2.2 Top-two algorithms and information balance condition
The idea of TTTS and other top-two algorithms arises naturally from the information balance condition (1). Recall that the key step in TTTS is sampling repeatedly from the posterior to obtain a challenger. We motivate TTTS by heuristically arguing that a system $j$ appears as a challenger, given that a system $i$ is the leader, asymptotically only when the following function is minimized
$C_{j}(\bm{p})=C_{j}(\bm{p};\bm{\theta})\triangleq\frac{(\theta_{I^{*}}-\theta_{j})^{2}}{2\left(\sigma_{I^{*}}^{2}/p_{I^{*}}+\sigma_{j}^{2}/p_{j}\right)},$ (4)
where $\bm{p}=(p_{1},\ldots,p_{K})$ is the empirical allocation of samples among the systems. Suppose $i$ is the leader, then $j$ appears as the challenger only when the posterior sample satisfies $\tilde{\theta}^{\prime}_{j}>\tilde{\theta}^{\prime}_{i}$. When the context is clear, we will omit the dependence of $C_{j}(\bm{p};\bm{\theta})$ on the mean parameter vector $\bm{\theta}$. By Russo (2020, Proposition 5), this occurs under the posterior distribution with an asymptotic probability proportional to $\exp\left(-tC_{j}(\bm{p})\right)$. The challenger is the system with the highest posterior sample, which, when $t$ is large, corresponds to the system with the smallest $\exp\left(-tC_{j}(\bm{p})\right)$. Thus, TTTS asymptotically chooses the $j$ with the smallest $C_{j}(\bm{p})$ as the challenger. Since $C_{j}(\bm{p})$ is a monotonically increasing function in $\bm{p}$, TTTS satisfies the information balance condition by maintaining equality among the $C_{j}(\bm{p})$.
2.3 Information-direction selection among top-two via stationarity condition
More importantly, we demonstrate how the stationarity condition in (2) allows us to address a significant open problem identified in Russo (2020): achieving exact optimality rather than merely $\beta$-optimality, a form of restricted optimality defined in Russo (2020).
The original TTTS algorithm adopts a non-adaptive tuning parameter $\beta$ to decide which of the top-two candidates to select in each round. An additional tuning step is required to guarantee that $\beta$ converges to the optimal $\beta^{*}$. However, the main challenge is that $\beta^{*}$ depends on the unknown mean vector $(\theta_{1},\ldots,\theta_{K})$.
We propose the information-directed selection (IDS) for top-two algorithms under Gaussian distribution. IDS takes the current sample allocation $\bm{p}_{t}$ and the current top-two candidates $\left(I_{t}^{(1)},I_{t}^{(2)}\right)$ as input and returns a selected candidate. IDS is hyperparameter-free and guarantees asymptotic optimality.
Algorithm 2 Information-directed selection (IDS) for Gaussian R&S with heterogeneous variances
1:Sample allocation $\bm{p}$ and top-two candidates $\left(i,j\right)$.
2:Return $i$ with probability $h_{i}^{j}(\bm{p})=\frac{\sigma_{i}^{2}/p_{i}}{\sigma_{i}^{2}/p_{i}+\sigma_{j}^{2}/p_{j}}$ and return $j$ with probability $1-h_{i}^{j}(\bm{p})=\frac{\sigma_{j}^{2}/p_{j}}{\sigma_{i}^{2}/p_{i}+\sigma_{j}^{2}/p_{j}}$.
We now argue that IDS naturally arises from the stationarity condition in (2). To see this, note that TTTS selects a system using a two-step procedure: it first identifies a pair of candidates and then measures one from the pair. This procedure naturally provides a law of total probability formula by conditioning on the choice of the challenger. In particular, the left-hand side of (2), representing the proportion of samples allocated to the best system $I^{*}$, can be decomposed into a summation of terms $h^{j}_{I^{*}}(\bm{p}^{*})\cdot\frac{p_{j}^{*}}{1-h^{j}_{I^{*}}(\bm{p}^{*})}$. By IDS, $h^{j}_{I^{*}}(\bm{p}^{*})$ is interpreted as the probability that system $I^{*}$ is sampled, given that $j$ is proposed as the challenger. Let $\mu^{*}_{j}$ denote the probability that $j$ is chosen as the challenger. We argue that $\mu^{*}_{j}=\frac{p_{j}^{*}}{1-h^{j}_{I^{*}}(\bm{p}^{*})}$. Specifically, system $j$ is sampled only if it appears as the challenger (with probability $\mu^{*}_{j}$) and is selected by IDS (with probability $1-h^{j}_{I^{*}}(\bm{p}^{*})$). Therefore, the relationship $p_{j}^{*}=\mu^{*}_{j}(1-h^{j}_{I^{*}}(\bm{p}^{*}))$ must hold. As a sanity check, note that the balance of sum-of-squares in (3) ensures $\sum_{j\neq I^{*}}\mu_{j}^{*}=1$, making $\bm{\mu}$ is a valid probability vector:
$\sum_{j\neq I^{*}}\mu^{*}_{j}=\sum_{j\neq I^{*}}\frac{p^{*}_{j}}{1-h^{j}_{I^{*}}(\bm{p}^{*})}=\sum_{j\neq I^{*}}\frac{p^{*}_{j}}{\frac{\sigma_{j}^{2}/p^{*}_{j}}{\sigma_{i}^{2}/p^{*}_{i}+\sigma_{j}^{2}/p^{*}_{j}}}=\sum_{j\neq I^{*}}p^{*}_{j}+p^{*}_{I^{*}}\cdot\sum_{j\neq I^{*}}\frac{(p^{*}_{j})^{2}/\sigma^{2}_{j}}{(p^{*}_{I^{*}})^{2}/\sigma^{2}_{I^{*}}}=\sum_{i\in[K]}p_{i}=1,$
where the second-to-last inequality follows from (3).
Remark 1 (Selection versus tuning).
Our IDS inherently differs from the standard approach of adaptive $\beta$-tuning (e.g., in Russo 2020). In that approach, the same parameter $\beta$ is used regardless of the proposed top-two candidates, even though the value of $\beta$ may be updated over time. Hence, we regard this method as a hyperparameter tuning technique. In contrast, our information-directed selection selects from the proposed candidates using the respective selection function $h^{I_{t}^{(2)}}_{I_{t}^{(1)}}(\bm{p}_{t})$. The selection functions are calculated using different pairs of candidates at different time steps. Hence, the selection functions at different time steps cannot be regarded as the same hyperparameter.
2.4 Optimality of TTTS-IDS
For theoretical analysis, we focus on the TTTS-IDS algorithm, i.e., the TTTS algorithm but replacing the selection step based on a tuning parameter $\beta$ by IDS (Algorithm 2). We prove that under Gaussian distribution, TTTS-IDS satisfies the sufficient condition for optimality in Theorem 1. This resolves an open problem in the pure exploration literature; see Russo (2020, Section 8). The proof is provided in Appendix D.
Theorem 2.
TTTS-IDS is optimal in the sense of Definition 1. In particular,
$\lim_{t\to\infty}-\frac{1}{t}\log\Pi_{t}^{\textup{{TTTS-IDS}}}(\Theta^{c}_{I^{*}})=\max_{\bm{p}\in\mathcal{S}_{K}}\min_{j\neq I^{*}}C_{j}(\bm{p}),$
for $C_{j}(\bm{p})$ defined in (4), where $\mathcal{S}_{K}$ is the $(K-1)$-dimensional probability simplex.
We remark that TTTS is arguably the most challenging version to establish optimality, since its leader and challenger are both random (even given the history). Our optimality result applies to other algorithms following the design principles in Section 3.4, e.g., those in Russo (2020); Qin et al. (2017); Shang et al. (2020); Jourdan et al. (2022).
Proof sketch.
Theorem 1 relies on a generalization of Qin et al. (2017, Theorem 2) to heterogeneous variances. Following the same analysis, we can establish the posterior convergence rate of any sampling rule that has empirical allocation converging almost surely to a deterministic probability vector.
Lemma 1.
If a sampling rule A satisfies that its empirical allocation $\bm{p}_{t}$ converges almost surely to some allocation $\bm{p}$, then
$\lim_{t\to\infty}-\frac{1}{t}\log\Pi^{\textup{{A}}}_{t}(\Theta^{c}_{I^{*}})=\min_{j\neq I^{*}}C_{j}(\bm{p}),$
for $C_{j}(\bm{p})$ defined in (4).
It turns out that the optimal allocation $\bm{p}^{*}$ in Theorem 1 is the unique solution to
$\Gamma^{*}_{\bm{\theta}}\triangleq\max_{\bm{p}\in\mathcal{S}_{K}}\min_{j\neq I^{*}}C_{j}(\bm{p}).$ (5)
Combining this with Lemma 1 completes the proof of Theorem 1.
Furthermore, with the uniqueness of $\bm{p}^{*}$ established in Theorem 1, Lemma 1 also implies that if a sampling rule results in $\bm{p}_{t}$ converging to any $\bm{p}\neq\bm{p}^{*}$, then it cannot be optimal.
To establish Theorem 2, it suffices to show that its almost sure convergence to the optimal allocation.
Lemma 2.
Under TTTS-IDS, $\bm{p}_{t}$ converges almost surely to the allocation $\bm{p}^{*}$ in Theorem 1.
We present complete proofs in Appendix A.
2.5 Discussions
2.5.1 Novel optimality conditions by incorporating dual variables.
The literature has explored the optimality conditions in Theorem 1, such as in Glynn and Juneja (2004); Garivier and Kaufmann (2016), where these conditions are directly expressed in terms of the decision variable $\bm{p}$.
Our approach takes this a step further by directly incorporating the dual variable into the optimality condition. We begin by defining the optimal dual variable as
$\mu^{*}_{j}=\frac{p_{j}^{*}}{1-h^{j}_{I^{*}}(\bm{p}^{*})},\quad\forall j\neq I^{*},$
which plays a central role in deriving IDS in Section 2.3. These dual variables correspond to the inequality constraints (6) in the following reformulation of the optimal allocation problem (5):
$\Gamma^{*}_{\bm{\theta}}=\max_{\phi,\bm{p}}$ $\hskip 10.0pt\phi,$
$\mathrm{s.t.}$ $\hskip 10.0pt\sum_{i\in[K]}p_{i}-1=0,\quad p_{i}\geq 0,\quad\forall i\in[K],$
$\hskip 10.0pt\phi-C_{j}(\bm{p})\leq 0,\quad\forall j\neq I^{*}.$ (6)
The conditions in Theorem 1 are derived directly from the Karush-Kuhn-Tucker (KKT) conditions, where the information balance condition (1) corresponds to the complementary slackness condition, and (2) represents the stationarity condition; see Theorem 4 for details.
The novelty of incorporating dual variables lies in their significant potential for generalization, as demonstrated in Section 3. For instance, consider the information balance condition. As noted in Zhang et al. (2023, Table 1), the information balance for identifying the best-$k$ alternatives involves a complex combinatorial structure, making it nearly impossible to explicitly generalize this condition for such tasks. Appendix F.3 provides a case study highlighting these complexities.
However, by leveraging dual variables, we can bypass these combinatorial challenges and instead enforce the information balance implicitly through the complementary slackness condition:
$\mu^{*}_{j}\cdot\left(\min_{j\neq I^{*}}C_{j}(\bm{p}^{*})-C_{j}(\bm{p}^{*})\right)=0,\quad\forall j\neq I^{*}.$
This condition holds for a broad range of problems beyond R&S; as demonstrated in Theorem 4.
Furthermore, the dual variables provide an alternative interpretation of the optimal allocation problem as a zero-sum game between the DM and an adversarial skeptic, see Remark 9.
2.5.2 Optimality in alternative performance criteria
The same optimal allocation $\bm{p}^{*}$ arises in other performance criteria studied in the literature, and thus our algorithm is also optimal in such settings.
Static-allocation fixed-budget setting.
Consider the frequentist probability of correct selection (PCS), defined as $\mathbb{P}_{\bm{\theta}}\left(\widehat{I}_{T}^{*}=I^{*}\right)$ for a fixed budget $T$, where the probability is taken over the randomness in both observations and the algorithm and $\widehat{I}_{T}^{*}$ is the output of the algorithm at the end of horizon. Due to the dynamic nature of typical algorithms – the decision at later stages depends on earlier observations – analyzing PCS in the fixed-budget setting is challenging (Hong et al., 2021; Komiyama, 2022). To address this, a common framework in the simulation literature, initiated by Glynn and Juneja (2004) and followed by Chen and Ryzhov (2023), focuses on PCS under a static algorithm. These static algorithms allocate samples proportionally to a fixed probability vector $\bm{p}\in\mathcal{S}_{K}$. The framework characterizes the large deviation rate of the probability of incorrect selection under a static algorithm as:
$-\lim_{T\to\infty}\frac{1}{T}\log\left(1-\mathbb{P}_{\bm{p},\bm{\theta}}\left(\widehat{I}_{T}^{*}=I^{*}\right)\right)=\tilde{\Gamma}_{\bm{\theta}}(\bm{p}),$
for some large deviation rate function $\tilde{\Gamma}_{\bm{\theta}}(\bm{p})$, where $\mathbb{P}_{\bm{p},\bm{\theta}}$ denotes the probability under system parameter $\bm{\theta}$ and static allocation $\bm{p}$. For Gaussian noise, the large deviation rate function $\tilde{\Gamma}_{\bm{\theta}}(\bm{p})=\Gamma_{\bm{\theta}}(\bm{p})$. Glynn and Juneja (2004) proceeds to maximize the large deviation rate, leading to the same optimal allocation problem as in (5). [^3] For general distributions beyond Gaussian, the expression of $C_{j}$ in the optimal allocation for static fixed-budget and posterior convergence rate settings differs. However, the maximin structure remains unchanged, and our analysis extends to these cases. A common goal of fixed-budget algorithms is to ensure the dynamic allocation converges almost surely to the optimal static allocation $\bm{p}^{*}$. Our algorithm achieves this goal, as guaranteed by Lemma 2.
Fixed-confidence setting.
The fixed-confidence formulation for best-arm identification has also been studied in the machine learning literature, yielding the same $\bm{p}^{*}$; see Garivier and Kaufmann (2016, Lemma 4 and Theorem 5). In this setting, the objective is to design an algorithm that stops in finite time and recommends the best system with a confidence guarantee of $1-\delta$ for some confidence level $\delta$, while minimizing the total number of samples used. We address this objective for the general pure exploration problem in Section 3, where we also prove that our algorithm achieves optimality in R&S with Gaussian noise under this fixed-confidence formulation.
3 General pure-exploration problems
We now broaden our focus from traditional ranking and selection problems to address a wider range of pure exploration queries relevant to decision-making. To align with the common terminology in machine learning literature on pure exploration, we adopt terms such as “arm” instead of “system” and “reward” instead of “performance.” Notably, we will show that many formulations commonly encountered in simulation literature, such as the PGS and the indifference-zone (IZ) approach, have clear counterparts within the pure exploration framework.
We consider a stochastic multi-armed bandit problem with $K$ arms, denoted by $i\in[K]$. The rewards from arm $i$ are independent and identically distributed random variables following a distribution $F_{\theta_{i}}$, parameterized by the mean $\theta_{i}$ and belonging to the single-parameter exponential family. We assume that the rewards from different arms are mutually independent. A problem instance is defined by a fixed but unknown vector $\bm{\theta}\triangleq(\theta_{1},\ldots,\theta_{K})\in\mathbb{R}^{K}$. The DM’s goal is to answer a exploration query regarding $\bm{\theta}$ by adaptively allocating the sampling budget across the arms and observing the noisy rewards from the selected arms.
Exploration query.
An exploration query specifies a question regarding the unknown mean vector $\bm{\theta}$. We assume that the query has a unique answer, denoted by $\mathcal{I}(\bm{\theta})$. [^4] Exploration queries with multiple correct answers require more care, see Section 4.1 for an example. Several notable exploration queries have been explored in the literature. In the indifference-zone-free Ranking and Selection (R&S) problem (Fan et al., 2016) discussed in Section 2, the query identifies the unique best arm as $\mathcal{I}({\bm{\theta}})=\operatorname*{argmax}_{i\in[K]}\theta_{i}$. Alternatively, one might seek to identify the exact set of best-$k$ arms, a problem referred to as best-$k$ identification, where $\mathcal{I}({\bm{\theta}})=\operatorname*{argmax}_{\mathcal{S}\subseteq[K],|\mathcal{S}|=k}\left\{\sum_{i\in\mathcal{S}}\theta_{i}\right\}$.
Other exploration queries include R&S under the indifference-zone (IZ) assumption (Bechhofer, 1954), or identifying all $\varepsilon$-good arms (Mason et al., 2020), where $\mathcal{I}(\bm{\theta})=\{i\in[K]:\theta_{i}\geq\max_{i\in[K]}\theta_{i}-\varepsilon\}$. Another example is the thresholding bandit problem (Lai and Liao, 2012; Locatelli et al., 2016), where the goal is to identify all arms with mean rewards exceeding a threshold $\mathfrak{T}$, represented by $\mathcal{I}({\bm{\theta}})=\{i\in[K]:\theta_{i}>\mathfrak{T}\}$.
Remark 2 (Extensions).
Our analysis extends beyond the base case to accommodate: (i) reward distributions outside the single-parameter exponential family, and (ii) exploration queries with multiple correct answers. We present these extensions in Section 4. Specifically, in Section 4.1, we consider the problem of identifying one of the $\varepsilon$-good alternatives, known as $\varepsilon$-best-arm identification ($\varepsilon$-BAI, Eckman and Henderson 2018; Jourdan et al. 2023a). This problem aligns with R&S under a PGS guarantee. The reason $\varepsilon$-BAI is compatible with our framework is that it can be reformulated as a modified pure-exploration problem with a unique answer. In Section 4.2, we examine the Gaussian case with unknown variances – a scenario in which the reward distribution no longer belongs to a single-parameter exponential family. Similar analyses are possible for other distributions; for instance, Agrawal et al. (2020) provides an analysis for heavy-tailed distributions. Importantly, our algorithm can be readily extended to accommodate these cases.
Fixed-confidence performance criteria.
We focus on the fixed-confidence setting, which aims to identify the correct answer with a PCS guarantee of at least $1-\delta$ for a fixed $\delta\in(0,1)$, while minimizing the sampling budget. In addition to the sampling rule discussed in Section 2, an algorithm requires both stopping and decision rules. At each time step $t\in\mathbb{N}_{0}$, the DM first determines whether to stop according to a stopping rule $\tau_{\delta}$. If stopping occurs at $t=\tau_{\delta}$, the DM provides an estimate ^$\widehat{\mathcal{I}}_{\tau_{\delta}}$ of the answer $\mathcal{I}(\bm{\theta})$ based on a decision rule. If not, she select an arm $I_{t}$ according to a sampling rule and observe the reward from arm $I_{t}$, denoted by $Y_{t+1,I_{t}}$. This procedure is summarized in Algorithm 3 where the information that DM has at the beginning of time $t$ is denoted by $\mathcal{H}_{t}=(I_{0},Y_{1,I_{0}},\dots,I_{t-1},Y_{t,I_{t-1}})$.
Algorithm 3 A template for anytime pure exploration algorithms
1:A stopping rule (stop), a decision rule (decision), a sampling rule (sample), and $\mathcal{H}_{0}=\varnothing$.
2:for $t=0,1,\dots$ do
3:     If $\textup{{stop}}(\mathcal{H}_{t})$ return ^$\widehat{\mathcal{I}}=\textup{{decision}}(\mathcal{H}_{t})$.
4:     Run sampling rule $I_{t}=\textup{{sample}}(\mathcal{H}_{t})$.
5:     Select arm $I_{t}$, observe reward $Y_{t+1,I_{t}}$, and update history $\mathcal{H}_{t+1}=\mathcal{H}_{t}\cup\{I_{t},Y_{t+1,I_{t}}\}$.
6:end for
We focus on the set of problem instances under a mild assumption. Specifically, let $\overline{\Theta}$ denote the set of parameters that yield a unique answer to the exploration query. Let $\mathfrak{I}$ be the collection of all possible correct answers, and for each $\mathcal{I}\in\mathfrak{I}$, let $\Theta_{\mathcal{I}}$ denote the set of parameters with $\mathcal{I}$ as the correct answer, i.e., $\Theta_{\mathcal{I}}\triangleq\{\bm{\vartheta}\in\overline{\Theta}:\mathcal{I}(\bm{\vartheta})=\mathcal{I}\}.$ We assume that each $\Theta_{\mathcal{I}}$ is an open set. This assumption rules out scenarios that are statistically unidentifiable. Our focus is on the parameter space $\Theta\triangleq\cup_{\mathcal{I}\in\mathfrak{I}}\Theta_{\mathcal{I}}$. We note that this assumption is generally unrestrictive; for instance, in R&S, we have $\overline{\Theta}=\Theta$, i.e., meaning we consider all problem instances with a unique solution.
Definition 2 ($\delta$-correct).
A policy is said to be $\delta$-correct if the probability of correct selection upon stopping is at least $1-\delta$, i.e.,
$\mathbb{P}_{\bm{\theta}}\left(\tau_{\delta}<\infty,\,\,\widehat{\mathcal{I}}_{\tau_{\delta}}=\mathcal{I}(\bm{\theta})\right)\geq 1-\delta,\quad\hbox{for anyproblem instance }\bm{\theta}\in\Theta.$
Let $\mathcal{A}$ denote the class of policies that are $\delta$-correct for any $\delta>0$.
The objective is to design a policy $\pi$ that minimizes the expected sample complexity $\mathbb{E}^{\pi}_{\bm{\theta}}[\tau_{\delta}]$ over $\mathcal{A}$, where $\tau_{\delta}\triangleq\min\{t\geq 0:\textup{{stop}}(\mathcal{H}_{t})=\textup{{True}}\}$ represents the number of samples collected when the policy stops. Specifically, we say that a policy is optimal in fixed confidence if it achieves universal efficiency, as defined below.
Definition 3 (Universal efficiency in fixed confidence).
A policy $\pi^{*}$ is universally efficient if $\pi^{*}$ is $\delta$-correct (i.e., $\pi^{*}\in\mathcal{A}$) and for any other $\pi\in\mathcal{A}$, we have
$\limsup_{\delta\to 0}\frac{\mathbb{E}^{\pi^{*}}_{\bm{\theta}}[\tau_{\delta}]}{\mathbb{E}^{\pi}_{\bm{\theta}}[\tau_{\delta}]}\leq 1,\quad\hbox{for any problem instance }\bm{\theta}\in\Theta.$
Compared with the “pointwise” optimality in Definition 1, Definition 3 is stronger, as it requires optimality to hold universally for all $\bm{\theta}$ under a $\delta$-correctness constraint that must be satisfied for all $\delta$ and all $\bm{\theta}\in\Theta$.
3.1 Preliminary
We now introduce key concepts and notation that form the foundation of our algorithm design principles.
The culprits.
For an algorithm to correctly answer the pure exploration query, it must distinguish between problem instances that yield different answers. For a given problem instance $\bm{\theta}\in\Theta$, the intrinsic complexity of the pure exploration query in the fixed-confidence setting is closely related to the alternative set, defined as the set of parameters that yield a correct answer different from $\mathcal{I}(\bm{\theta})$:
$\mathrm{Alt}(\bm{\theta})\triangleq\{\bm{\vartheta}\in\Theta:\mathcal{I}(\bm{\vartheta})\neq\mathcal{I}(\bm{\theta})\}.$
To formulate a maximin problem as in (5), we rely on the following assumption.
Assumption 1 (Wang et al. 2021).
$\mathrm{Alt}(\bm{\theta})$ is a finite union of convex sets. Namely, there exists a finite set $\mathcal{X}(\bm{\theta})$ and a collection of convex sets $\{\mathrm{Alt}_{x}(\bm{\theta}):x\in\mathcal{X}(\bm{\theta})\}$ such that $\mathrm{Alt}(\bm{\theta})=\cup_{x\in\mathcal{X}(\bm{\theta})}\mathrm{Alt}_{x}(\bm{\theta})$.
This condition is indeed satisfied by commonly encountered exploration queries; see Appendix G.
We refer to the collection $\mathcal{X}(\bm{\theta})$ as the set of culprits. These culprits represent the fundamental causes that lead the correct answer under an alternative problem instance $\bm{\vartheta}$ to deviate from $\mathcal{I}(\bm{\theta})$. Thus, the culprit set is determined by both the exploration query and the specific problem instance $\bm{\theta}$.
Example 1 (Ranking and selection).
For R&S, the culprit set is defined as $\mathcal{X}(\bm{\theta})=[K]\backslash\{I^{*}(\bm{\theta})\}$, where each $j\in\mathcal{X}(\bm{\theta})$ represents a scenario in which, under an alternative problem instance $\bm{\vartheta}\in\Theta$, system $j$ outperforms the best system $I^{*}(\bm{\theta})$ identified under $\bm{\theta}$. In such scenarios, $I^{*}(\bm{\theta})$ is an incorrect selection due to the existence of the culprit $j$. The corresponding alternative set for each culprit $j\in\mathcal{X}(\bm{\theta})$ is $\mathrm{Alt}_{j}(\bm{\theta})=\{\bm{\vartheta}\in\Theta:\vartheta_{j}\geq\vartheta_{I^{*}(\bm{\theta})}\}$.
Example 2 (Best-$k$ identification).
In the best-$k$ identification problem, the goal is to identify the set of $k$ arms with the largest means. The exploration query can be broken down into verifying whether any top arm $i\in\mathcal{I}(\bm{\theta})$ has a higher mean than any bottom arm $j\in\mathcal{I}^{c}(\bm{\theta})$. If there exist pairs $(i,j)\in\mathcal{X}(\bm{\theta})\triangleq\mathcal{I}(\bm{\theta})\times\mathcal{I}^{c}(\bm{\theta})$ such that $\vartheta_{j}>\vartheta_{i}$, then either $i\notin\mathcal{I}(\bm{\vartheta})$ or $j\in\mathcal{I}(\bm{\vartheta})$, leading to $\mathcal{I}(\bm{\vartheta})\neq\mathcal{I}(\bm{\theta})$. A culprit in this context is specified by a pair of arms $(i,j)\in\mathcal{X}(\bm{\theta})$, with the alternative set $\mathrm{Alt}_{i,j}(\bm{\theta})=\{\bm{\vartheta}\in\Theta:\vartheta_{j}>\vartheta_{i}\}$.
Example 3 (Thresholding bandits).
In thresholding bandits, the goal is to identify the set of arms with means above a threshold $\mathfrak{T}$, i.e., $\mathcal{I}(\bm{\theta})\triangleq\{i\in[K]:\theta_{i}>\mathfrak{T}\}$. The exploration query can be broken down into comparing each arm $i\in[K]$ with the threshold $\mathfrak{T}$. A culprit is then specified by an arm $i\in\mathcal{X}(\bm{\theta})\triangleq[K]$ with the corresponding alternative set $\mathrm{Alt}_{i}(\bm{\theta})=\{\bm{\vartheta}\in\Theta:(\mathfrak{T}-\vartheta_{i})(\mathfrak{T}-\theta_{i})<0\}$ This set indicates that arm $i$ is on opposite sides of $\mathfrak{T}$ under $\bm{\vartheta}$ and $\bm{\theta}$, resulting in an incorrect answer.
The generalized Chernoff information $C_{x}$.
To correctly answer the pure-exploration query, every potential culprit must be ruled out. From a hypothesis testing perspective, this involves testing the hypotheses $H_{0,x}:\bm{\theta}\in\mathrm{Alt}_{x}(\bm{\theta}_{t})$ versus $H_{1,x}:\bm{\theta}\notin\mathrm{Alt}_{x}(\bm{\theta}_{t})$, for all $x\in\mathcal{X}(\bm{\theta}_{t})$, where $\bm{\theta}_{t}$ represents the sample mean at time $t$ and $H_{1,x}$ corresponds to the scenario in which $x$ is not a culprit. For instance, in R&S, rejecting $H_{0,x}$ indicates sufficient evidence to support the desired ordering $\theta_{I^{*}(\bm{\theta}_{t})}>\theta_{x}$.
To this end, we utilize the generalized log-likelihood ratio test (GLRT) statistic, as proposed in Kaufmann and Koolen (2021). Let $\mathcal{S}_{K}$ denote the probability simplex of dimension $K-1$ and let $\bm{p}\in\mathcal{S}_{K}$ represent a generic allocation vector, indicating the proportion of samples allocated to the $K$ arms. If an algorithm allocates according to $\bm{p}_{t}$ at time $t$, then the GLRT statistic given by
$\ln\frac{\sup_{\bm{\vartheta}\in\Theta}L(\bm{\vartheta})}{\sup_{\bm{\vartheta}\in\mathrm{Alt}_{x}(\bm{\theta}_{t})}L(\bm{\vartheta})}=\inf_{\bm{\vartheta}\in\mathrm{Alt}_{x}(\bm{\theta}_{t})}\ln\frac{L(\bm{\theta}_{t})}{L(\bm{\vartheta})}=t\cdot\inf_{\bm{\vartheta}\in\mathrm{Alt}_{x}(\bm{\theta}_{t})}\sum_{i\in[K]}p_{i}d(\theta_{t,i},\vartheta_{i}),$
where $d(\cdot,\cdot)$ denotes the Kullback-Leibler (KL) divergence. The first equality above leverages the fact that the unconstrained maximum likelihood estimator for the mean of a canonical single-parameter exponential family is the sample mean.
We extend $C_{j}$ in (4) as the population version of the GLRT statistic, normalized by the sample size $t$,
$C_{x}(\bm{p})=C_{x}(\bm{p};\bm{\theta})\triangleq\inf_{\bm{\vartheta}\in\mathrm{Alt}_{x}(\bm{\theta})}\sum_{i\in[K]}p_{i}d(\theta_{i},\vartheta_{i}),\quad\text{for all }x\in\mathcal{X}(\bm{\theta}).$ (7)
We refer to $C_{x}$ as the generalized Chernoff information, extending a classical concept in hypothesis testing (Cover and Thomas, 2006). When the context is clear, we will omit the dependence of $C_{x}(\bm{p};\bm{\theta})$ on the problem instance $\bm{\theta}$. The following result establishes the basic properties of $C_{x}$.
Lemma 3.
Assume that the reward distribution follows a canonical single-parameter exponential distribution parameterized by its mean. Under Assumption 1, for each culprit $x\in\mathcal{X}$:
  1. 1.
    There exists a unique $\bm{\vartheta}^{x}\in\mathrm{Alt}_{x}(\bm{\theta})$ that achieves the infimum in (7).
  2. 2.
    The functions $C_{x}(\bm{p})$ are differentiable concave functions, with $\frac{\partial C_{x}(\bm{p})}{\partial p_{i}}=d(\theta_{i},\vartheta^{x}_{i})$ for all $i\in[K].$
  3. 3.
    The functions $C_{x}(\bm{p})$ satisfy the following partial differential equation
    $C_{x}(\bm{p})=\sum_{i\in[K]}p_{i}\frac{\partial C_{x}(\bm{p})}{\partial p_{i}}.$ (8)
When the sampling effort is distributed according to $\bm{p}$, the function $C_{x}$ quantifies the average information gain per sample needed to assert $H_{1,x}$, i.e., to reject culprit $x$. The partial differential equation (PDE) in (8) shows that this information gain can be effectively decomposed into contributions from the sample allocations to individual arms; see the discussion in Remark 5.
The optimal allocation problem.
Recall that $H_{0,x}$ must be rejected for all culprits $x\in\mathcal{X}(\bm{\theta})$ to correctly answer the pure-exploration query. Furthermore, $C_{x}(\bm{p})$ represents the information gathered to reject $H_{0,x}$ under sample allocation $\bm{p}$ – the larger $C_{x}(\bm{p})$ is, the easier it is to reject $H_{0,x}$. By taking the minimum over all culprits $x\in\mathcal{X}(\bm{\theta})$, we define
$\Gamma_{\bm{\theta}}(\bm{p})\triangleq\min_{x\in\mathcal{X}(\bm{\theta})}C_{x}(\bm{p}),$
which corresponds to the GLRT statistic for the leading culprit, i.e., the culprit that is hardest to reject. To maximize information gathering, we should follow an allocation $\bm{p}$ that maximizes $\Gamma_{\bm{\theta}}(\bm{p})$. This leads to the following optimal allocation problem:
$\Gamma^{*}_{\bm{\theta}}\triangleq\max_{\bm{p}\in\mathcal{S}_{K}}\Gamma_{\bm{\theta}}(\bm{p})=\max_{\bm{p}\in\mathcal{S}_{K}}\min_{x\in\mathcal{X}(\bm{\theta})}C_{x}(\bm{p}),$ (9)
which generalizes (5) in the context of R&S with Gaussian noise.
Stopping and decision rules.
A fixed-confidence algorithm requires three components: stopping, decision, and sampling rules. We follow standard practices in the literature for the stopping and decision rules, allowing us to focus on designing optimal sampling rules.
For the stopping rule, Kaufmann and Koolen (2021) provide the Chernoff stopping rule based on the GLRT for exponential family bandits:
$\tau_{\delta}=\inf\left\{t\in\mathbb{N}_{0}\,:\,t\cdot\Gamma_{\bm{\theta}_{t}}(\bm{p}_{t})>\gamma(t,\delta)\right\},$ (10)
where $\gamma(t,\delta)$ is a threshold function. Kaufmann and Koolen (2021, Proposition 15) provide a $\gamma(t,\delta)$ ensuring that, combined with any sampling rule, the stopping rule in (10) is $\delta$-correct for any pure-exploration query with a unique correct answer. For completeness, we rephrase their results in Appendix B.1.
We adopt a specific decision rule
$\textup{{decision}}(\mathcal{H_{\tau_{\delta}}})=\mathcal{I}(\bm{\theta}_{\tau_{\delta}}),$ (11)
where $\mathcal{I}(\bm{\theta}_{\tau_{\delta}})$ is the correct answer based on the sample mean $\bm{\theta}_{\tau_{\delta}}$ upon stopping. [^5] In cases where multiple correct answers exist under $\bm{\theta}_{\tau_{\delta}}$, ties are broken arbitrarily. For simplicity, we abuse the definition $\mathcal{I}(\bm{\theta}_{\tau_{\delta}})$ to represent one of the correct answers.
3.2 A sufficient condition of sampling rules for fixed-confidence optimality
We now present a sufficient condition for a sampling rule to achieve universal efficiency in the fixed-confidence setting. This condition requires the allocation of sampling effort to converge “rapidly” to the optimal solution of (9). We refer to this mode of convergence as $\mathcal{L}$-convergence, which is stronger than almost sure convergence. Specifically, $\mathcal{L}$-convergence ensures that for any fixed neighborhood of the target, the sequence remains within that neighborhood indefinitely after a random time with finite expectation.
Let the space $\mathcal{L}$ consist of all measurable $X$ with $\|X\|_{1}<\infty$ where $\|X\|_{1}\triangleq\mathbb{E}\left[|X|\right]$ is the $1$-norm.
Definition 4 ($\mathcal{L}$-convergence).
For a sequence of real valued random variables $\{X_{t}\}_{t\in\mathbb{N}_{0}}$ and a scalar $x\in\mathbb{R}$, we say $X_{t}\xrightarrow{\mathcal{L}}x$ if
for $\text{for all }\epsilon>0\,\text{ there exists }\,T\in\mathcal{L}\text{ such that for all }t\geq T,\,\,|X_{t}-x|\leq\epsilon.$
For a sequence of random vectors $\{\bm{X}_{t}\}_{t\in\mathbb{N}_{0}}$ taking values in $\mathbb{R}^{d}$ and a vector $\bm{x}\in\mathbb{R}^{d}$, where $\bm{X}_{t}=(X_{t,1},\ldots,X_{t,d})$ and $\bm{x}=(x_{1},\ldots,x_{d})$, we say $\bm{X}_{t}\xrightarrow{\mathcal{L}}\bm{x}$ if $X_{t,i}\xrightarrow{\mathcal{L}}x_{i}$ for all $i\in[d]$.
It is straightforward to observe that $\mathcal{L}$-convergence implies almost sure convergence.
Corollary 1.
If $\bm{X}_{t}\xrightarrow{\mathcal{L}}\bm{x}$, then $\bm{X}_{t}\xrightarrow{\mathrm{a.s.}}\bm{x}$.
Universal efficiency requires controlling the expectation of the stopping time $\tau_{\delta}$ in (10). This is achieved by establishing a random time $T_{\varepsilon}\in\mathcal{L}$ such that $|\Gamma_{\bm{\theta}_{t}}(\bm{p}_{t})-\Gamma^{*}|<\epsilon$ for all $t>T_{\epsilon}$. Since $T_{\varepsilon}\in\mathcal{L}$, the expectation $\mathbb{E}^{\pi}_{\bm{\theta}}[\tau_{\delta}]$ is well-controlled. Notably, $\Gamma_{\bm{\theta}_{t}}(\bm{p}_{t})\xrightarrow{\mathcal{L}}\Gamma^{*}$ is implied by $\bm{p}_{t}\xrightarrow{\mathcal{L}}\bm{p}^{*}$. We formalize this sufficient condition below, with the proof deferred to Appendix B. This condition generalizes the result in Qin et al. (2017), which was specifically tailored to BAI.
Theorem 3 (A sufficient condition for optimality).
Suppose that for any $\bm{\theta}\in\Theta$, the optimal allocation problem (9) admits a unique, component-wise strictly positive optimal allocation $\bm{p}^{*}(\bm{\theta})$. A policy $\pi$ is universally efficient (Definition 3) if the following conditions are satisfied:
  1. 1.
    this policy $\pi$ employs the stopping rule (10) and the decision rule (11), and
  2. 2.
    this policy $\pi$ uses a sampling rule satisfying that for any $\bm{\theta}\in\Theta$, $\bm{p}_{t}\xrightarrow{\mathcal{L}}\bm{p}^{*}(\bm{\theta})$.
In particular,
$\lim_{\delta\to 0}\frac{\mathbb{E}^{\pi}_{\bm{\theta}}[\tau_{\delta}]}{\log(1/\delta)}=(\Gamma^{*}_{\bm{\theta}})^{-1},\quad\forall\bm{\theta}\in\Theta.$
Remark 3 (Uniqueness and positivity of optimal solution).
For BAI, Gaussian $\epsilon$-BAI, and thresholding bandits, the existence of a unique optimal solution $\bm{p}^{*}(\bm{\theta})>0$ for any $\bm{\theta}\in\Theta$ has been established by Garivier and Kaufmann (2016), Jourdan et al. (2023a), and Ménard (2019), respectively. To the best of our knowledge, we are the first to prove the uniqueness and positivity of the optimal solution for the best-$k$ identification problem with single-parameter exponential family distributions; see Proposition 9.
Remark 4 (Other performance criteria).
We discussed two alternative performance criteria in Theorem 1 and Section 2.5.2, both focusing on the almost sure convergence of the empirical allocation $\bm{p}_{t}\stackrel{{\scriptstyle a.s.}}{{\to}}\bm{p}^{*}$. For the posterior convergence rate setting (Russo, 2020), the optimal allocation aligns with (9), even for other single-parameter exponential family distributions. For the static-allocation fixed-budget setting (Glynn and Juneja, 2004) under Gaussian distributions, the optimal allocation also coincides with (9) due to the symmetry of the KL divergence. For other distributions, swapping the parameters in the KL divergence in (7) is necessary, but the properties of $C_{x}$ in (8) and the maximin structure of (9) remain valid, leading to a similar sufficient condition. Our sufficient condition in Theorem 3, which requires the $\mathcal{L}$-convergence of the sample allocation, is stronger than almost sure convergence. Consequently, this condition also implies optimality in the posterior convergence rate and the large deviations rate of static allocations.
3.3 Optimality conditions for the optimal allocation problem
We define selection functions that generalize $h^{j}_{i}(\bm{p})$ in Algorithm 2 to general pure-exploration problems:
$h_{i}^{x}(\bm{p})=h_{i}^{x}(\bm{p};\bm{\theta})\triangleq\frac{p_{i}\frac{\partial C_{x}(\bm{p})}{\partial p_{i}}}{C_{x}(\bm{p})}=\frac{p_{i}d(\theta_{i},\vartheta^{x}_{i})}{\sum_{k\in[K]}p_{k}d(\theta_{k},\vartheta^{x}_{k})},\quad\hbox{for}\ x\in\mathcal{X}(\bm{\theta})\ \hbox{and}\ i\in[K].$ (12)
Following from the PDE characterization of $C_{x}$ in Lemma 3, $\bm{h}^{x}(\bm{p})=\{h_{i}^{x}(\bm{p}):i\in[K]\}\in\mathcal{S}_{K}$ forms a valid probability vector for any $x\in\mathcal{X}(\bm{\theta})$. In fact, $\bm{h}^{x}(\bm{p})$ can be interpreted as the probability distribution over the arms to be sampled when $x$ is identified as the leading culprit.
We derive the following optimality conditions for (9) in the cases where the optimal solution $\bm{p}^{*}$ is component-wise strictly positive. These conditions extend those presented in Theorem 1. The detailed proof can be found in Appendix C.
Theorem 4.
Assume that $\bm{p}^{*}$ is component-wise strictly positive. A feasible solution $(\phi,\bm{p})$ to (9) is optimal if and only if $\phi=\min_{x\in\mathcal{X}(\bm{\theta})}C_{x}(\bm{p})$ and there exist dual variables $\bm{\mu}\in\mathcal{S}_{|\mathcal{X}(\bm{\theta})|}$ such that
$p_{i}=\sum_{x\in\mathcal{X}(\bm{\theta})}\mu_{x}h_{i}^{x},\quad\forall i\in[K],$ (13a)
$\phi=\min_{x\in\mathcal{X}(\bm{\theta})}C_{x}(\bm{p}),\quad\text{and}\quad\sum_{x\in\mathcal{X}(\bm{\theta})}\mu_{x}(\phi-C_{x}(\bm{p}))=0.$ (13b)
The condition (13) is a equivalent form of the KKT condition for optimality, where (13a) is the stationarity condition and (13b) is the complementary slackness condition.
The following result extends the information balance for BAI in (1). It is a direct corollary of (13) and the assumption that $\bm{p}^{*}$ is strictly positive. We defer the proof to Appendix J.4.
Corollary 2 (Information balance).
Assume that $\bm{p}^{*}$ is component-wise strictly positive, then
$\Gamma_{\bm{\theta}}^{*}=\min_{x\in\mathcal{X}_{i}}C_{x}(\bm{p}^{*}),\quad\hbox{for all }i\in[K],\quad\textrm{where}\quad\mathcal{X}_{i}\triangleq\left\{x\in\mathcal{X}(\bm{\theta}):\frac{\partial C_{x}(\bm{p})}{\partial p_{i}}>0\right\}.$
Surprisingly, this information balance condition proves to be too crude to guarantee sufficiency for optimality in pure-exploration problems beyond BAI, even when combined with the correct overall balance condition; see Appendix F.1 for a counterexample. To address this challenge, we introduce dual variables that implicitly encode a more refined information balance.
3.4 Dual-directed algorithm design
Based on Theorem 4, a sampling rule can be immediately derived by tracking the two sets of conditions in (13). Our general algorithmic design framework follows the same idea as that discussed in Section 2.2 and Section 2.3. In particular, our algorithm takes two steps: it first identifies a leading culprit; and then selects an arm that provides the most information in curbing the identified leading culprit.
We first discuss how the leading culprit is detected. Recall that smaller the value of $C_{x}$, the less information we have to rule out $x$ as the root cause of an incorrect answer. Additionally, the dual variable $\mu_{x}$ can be interpreted as the proportion of time $x$ is detected as the leading culprit. In our algorithm, the culprit $x$ with the smallest (empirical estimation of) $C_{x}$ is detected as the leading culprit $x_{t}$. This step is guided by the complementary slackness conditions in (13b), which indicate that only the culprit $x$ with the smallest $C_{x}$ has $\mu_{x}>0$ and therefore requires attention.
Once $x_{t}$ has been detected as the leading culprit, the algorithm employs an information-directed selection (IDS) rule to select an arm for sampling, specified by $\bm{h}^{x_{t}}$ defined in (12). This procedure is formalized in Algorithm 4, which extends the IDS rule from Algorithm 2 for BAI to general pure-exploration problems The IDS rule is designed to depend on the detected culprit $x_{t}$, aiming to maximize the average information gain required to rule out the leading culprit. In particular, this process satisfies the stationarity conditions in (13a), where the sampling probability of arm $i$ is decomposed into a law-of-total-probability formula. Here, $h^{x}_{i}$ specifies the probability of sampling $i$ given that $x$ is detected as the leading culprit.
Algorithm 4 Information-directed selection (IDS) for pure exploration
1:Sample allocation $\bm{p}$, an estimate $\widehat{\bm{\theta}}$ and the detected leading culprit $x$.
2:Return $I\in[K]$ sampled from the distribution $\bm{h}^{x}(\bm{p};\widehat{\bm{\theta}})$.
Remark 5 (Information-directed selection).
$C_{x}(\bm{p})$ is defined as the population version of the GLRT statistic, normalized by the sample size $t$, for testing $H_{0,x}:\bm{\theta}\in\mathrm{Alt}_{x}(\bm{\theta}_{t})$ versus $H_{1,x}:\bm{\theta}\notin\mathrm{Alt}_{x}(\bm{\theta}_{t})$. Thus, it quantifies the information gathered per unit sample under the allocation vector $\bm{p}$ to determine whether the culprit $x$ is responsible for an incorrect answer. The PDE characterization of $C_{x}(\bm{p})$ in (8) reveals a decomposition of this information. To interpret, $\frac{\partial C_{x}(\bm{p})}{\partial p_{i}}$ represents the information gain per unit allocation to arm $i$ while $p_{i}\frac{\partial C_{x}(\bm{p})}{\partial p_{i}}$ quantifies the information gain per unit sample attributed to allocating a $p_{i}$ proportion of the samples to arm $i$. The selection function $h^{x}_{i}(\bm{p})=\frac{p_{i}\frac{\partial C_{x}(\bm{p})}{\partial p_{i}}}{C_{x}(\bm{p})}$ suggests that sampling should be done proportionally to the information gain contributed by each arm when determining the allocation among the candidates. Moreover, the selection functions $h^{x}_{i}(\bm{p})$ identify a set of candidates whose allocation probabilities are positive under a given culprit $x$, defined as $\mathfrak{C}_{x}\triangleq\{i\in[K]:\frac{\partial C_{x}(\bm{p})}{\partial p_{i}}>0\}$. This result is intuitive because only the arms in $\mathfrak{C}_{x}$ contribute to the information gain relevant to the culprit $x$. For the specific case of BAI, one can verify that the cardinality of $\mathfrak{C}_{x}$ is always two, leading to the well-known class of top-“two” algorithms.
Remark 6 (The role of the dual variables).
The novelty of our algorithm design framework lies in the incorporation of dual variables. Existing constant-optimal algorithms typically fall into two categories. The first solves the optimal allocation problem using plug-in estimators in each iteration and tracks the solution, as pioneered by Garivier and Kaufmann (2016) and extended in works like Kato and Ariu (2021); Al Marjani et al. (2022). The second constructs optimality conditions based solely on the allocation vector $\bm{p}$ (e.g., (1) and (2)), then tracks these conditions, as demonstrated in Chen and Ryzhov (2023); Bandyopadhyay et al. (2024). Our approach is more fundamental, as it introduces dual variables into the optimality conditions. The dual variables encode the information balance structure described in Corollary 2, which is critical for addressing general pure exploration queries. Unlike approaches that require explicit characterization of the optimal information balance structure – often infeasible without knowing the true parameter $\bm{\theta}$ (Zhang et al., 2023) – dual variables allow us to bypass this complexity entirely. Moreover, dual variables enable a law-of-total-probability-type stationarity condition (13a), naturally leading to the IDS rule, which is optimal across various pure exploration tasks.
3.5 A modular template for proposed algorithms
Our algorithm design principle focuses on adaptively acquiring information that most effectively curbs the leading culprit. Accordingly, we name it Adaptive Culprit Curbing (ACC). We now introduce a modular template for the dual-directed ACC algorithms.
Our primary focus is on the ACC sampling rules, which determine the arm to be sampled at each time step based on the observed history. These rules are independent of the total budget of samples $T$ or the confidence parameter $\delta$, making them any-time. In the fixed-confidence setting, the ACC sampling rule is coupled with a stopping rule to ensure $\delta$-correctness. As detailed in Section 3.1, we adopt standard stopping and decision rules from the literature for this setting. In contrast, the fixed-budget and posterior convergence rate settings terminate at a pre-specified budget $T$ and are evaluated based on their respective PCS measures. This versatility highlights the adaptability of the ACC framework across various performance criteria.
We propose the ACC sampling rules that follow three steps: first, estimate (est) the unknown parameters using $\widehat{\bm{\theta}}$, which also provides the most plausible answer $\mathcal{I}(\widehat{\bm{\theta}})$ to the exploration query; next, detect (det) the leading culprit $x$ that could make $\mathcal{I}(\widehat{\bm{\theta}})$ incorrect; and finally, select (sel) one arm to sample based on the detected leading culprit $x$. A template for our sampling rule is summarized in Algorithm 5.
Algorithm 5 A modular template for Adaptive Culprit Curbing (ACC) sampling rule
1:Current history $\mathcal{H}$, an estimation rule (est), a detection rule (det), and a selection rule (sel).
2:Run estimation rule to obtain $\widehat{\bm{\theta}}=\textup{{est}}(\mathcal{H})$.
3:Run detection rule to obtain the leading culprit $x=\textup{{det}}(\widehat{\bm{\theta}},\mathcal{H})$.
4:Return $I=\textup{{sel}}\left(x,\widehat{\bm{\theta}},\mathcal{H}\right)$.
We will follow the naming convention of est-det-sel for our proposed ACC algorithms. Below, we elaborate on the options for the three sub-routines required by ACC. We provide explicit formulas of $\mathcal{X},C_{x}$ and $\bm{h}^{x}$ for a wide range of examples in Appendix G.
3.5.1 Estimation (est) rules.
We consider two estimation rules. The empirical-best (EB) rule estimates the means using the empirical mean $\bm{\theta}_{t}$ based on the history $\mathcal{H}_{t}$ at time $t$. The Thompson-sampling (TS) rule estimates the means by drawing a sample $\tilde{\bm{\theta}}_{t}$ from the posterior distribution $\Pi_{t}$ at time $t$.
The estimate $\widehat{\bm{\theta}}$ returned by the estimation rule determines the perceived best answer $\mathcal{I}(\widehat{\bm{\theta}})$, culprit set $\mathcal{X}(\widehat{\bm{\theta}})$ and the decomposed alternative set $\mathrm{Alt}_{x}(\widehat{\bm{\theta}})$. Along with the sample allocation $\bm{p}$, this determines the estimated generalized Chernoff information $C_{x}(\bm{p},\widehat{\bm{\theta}})$.
3.5.2 Detection (det) rules.
We propose three detection rules.
  1. 1.
    The KKT detection rule (Algorithm 6) explicitly follows the complementary slackness conditions (13b) from our KKT analysis. It determines the leading culprit as the $x$ that minimizes $C_{x}(\bm{p},\widehat{\bm{\theta}})$.
  2. 2.
    The TS detection rule (Algorithm 7) generalizes the TTTS algorithm (Algorithm 1) for BAI to more general pure-exploration problems. It identifies an alternative problem instance with a different answer (and hence, a culprit) that is most likely to emerge from the posterior distribution. This framework offers a systematic approach to adapting the classical TS algorithm for pure exploration problems.
  3. 3.
    The Posterior-Probability-Sampling (PPS) detection rule (Algorithm 8) samples the leading culprit $x$ with probability proportional to $\Pi(\mathrm{Alt}_{x}(\widehat{\bm{\theta}}))$, where $\Pi$ denotes the current posterior distribution.
Following the same rationale as in Section 2.2.2, Russo (2020, Proposition 5) establish that the TS and PPS detection rules are asymptotically equivalent to the KKT rule, as $t\to\infty$.
Algorithm 6 KKT detection rule for pure exploration
1:Current sample allocation $\bm{p}$ and an estimate $\widehat{\bm{\theta}}$.
2:Return the leading culprit $x=\operatorname*{argmin}_{x\in\mathcal{X}(\widehat{\bm{\theta}})}C_{x}(\bm{p},\widehat{\bm{\theta}})$, breaking ties arbitrarily.
Algorithm 7 TS detection rule for pure exploration
1:Estimation $\widehat{\bm{\theta}}$ and posterior distribution $\Pi$.
2:Repeatedly sample $\tilde{\bm{\theta}}\sim\Pi$ until $\tilde{\bm{\theta}}\in\mathrm{Alt}(\widehat{\bm{\theta}})$.
3:Return the leading culprit $x\in\{x^{\prime}\in\mathcal{X}(\widehat{\bm{\theta}}):\tilde{\bm{\theta}}\in\mathrm{Alt}_{x^{\prime}}(\widehat{\bm{\theta}})\}$, breaking ties arbitrarily.
Algorithm 8 PPS detection rule for pure exploration
1:Estimation $\widehat{\bm{\theta}}$ and posterior distribution $\Pi$.
2:Return the leading culprit $x\in\mathcal{X}(\widehat{\bm{\theta}})$ sampled with probability proportional to $\Pi\left(\mathrm{Alt}_{x}(\widehat{\bm{\theta}})\right)$.
3.5.3 Selection (sel) rule.
For the selection rule, we follow the IDS rule as described in Algorithm 4.
Remark 7 (When does the design framework work?).
The derivation of the ACC algorithm relies only on two ingredients that lead directly to the KKT conditions in Theorem 4: (i) Assumption 1 induces the maximin convex optimization problem central to our framework, which is satisfied by many common exploration queries, as demonstrated in Appendix G; and (ii) by definition (7) using the GLRT, the generalized Chernoff information is homogeneous of degree one, satisfying $C_{x}(t\cdot\bm{p})=t\cdot C_{x}(\bm{p})$ for any $t>0$, which is equivalent to the PDE characterization in (8) by Euler’s theorem for homogeneous functions; see Apostol (1969, p. 287) for example.
3.6 Universal efficiency
In this section, we present two optimality results in the sense of Definition 3: one for Gaussian BAI and another for thresholding bandits.
3.6.1 Gaussian BAI
We focus on the TTTS sub-routine (Algorithm 1) combined with the information-directed selection sub-routine (Algorithm 2). Under the naming convention of Section 3.5, this is the TS-TS-IDS algorithm. To honor the original name proposed in Russo (2020), we refer to this algorithm as TTTS-IDS. As a theoretical guarantee, we prove that for Gaussian BAI, TTTS-IDS satisfies the sufficient condition for fixed-confidence optimality in Theorem 3. This addresses an open problem in the pure exploration literature, as noted by Russo (2020, Section 8). The proof is deferred to Appendix D.
Theorem 5.
Let $\bm{\theta}\in\Theta$ be a Gaussian instance with unique arm means. Under TTTS-IDS (with uninformative and independent prior distribution across arms), $\bm{p}_{t}\xrightarrow{\mathcal{L}}\bm{p}^{*}.$
In conjunction with Theorem 2 on the fixed-budget posterior convergence rate and the discussion of the static fixed-budget setting in Remark 4, the TTTS-IDS is optimal across all three settings.
Remark 8 (Fixed-confidence optimality of other top-two algorithms).
The analysis of TTTS-IDS is arguably the most complex among the top-two algorithms, as both the leader and challenger are random variables conditioned on the history. Our proof strategy can be extended to other top-two algorithms combined with IDS, such as TS-KKT-IDS, TS-PPS-IDS, and those presented in Russo (2020); Qin et al. (2017); Shang et al. (2020); Jourdan et al. (2022).
3.6.2 Thresholding bandits
In this section, we prove the universal efficiency of the EB-KKT-IDS algorithm for the thresholding bandit problem (TBP) discussed in Example 3. For TBP, $\mathcal{X}=[K]$ and the optimal allocation problem simplifies as
$\Gamma^{*}_{\bm{\theta}}$ $=\max_{\bm{p}\in\mathcal{S}_{K}}\min_{i\in\mathcal{X}}\inf_{\bm{\vartheta}\in\mathrm{Alt}_{i}}\sum_{i}p_{i}d(\theta_{i},\vartheta_{i})=\max_{\bm{p}\in\mathcal{S}_{K}}\min_{i\in\mathcal{X}}\inf_{(\mathfrak{T}-\vartheta_{i})(\mathfrak{T}-\theta_{i})<0}p_{i}d(\theta_{i},\vartheta_{i})=\max_{\bm{p}\in\mathcal{S}_{K}}\min_{i\in[K]}p_{i}d(\theta_{i},\mathfrak{T}),$
where the second equality follows from the fact that we can always set $\vartheta_{j}=\theta_{j}$ for $j\neq i$, and the last equality follows from the monotonicity of the KL divergence. We apply Lemma 3 to obtain that $\frac{\partial C_{i}}{\partial p_{i}}=d(\theta_{i},\mathfrak{T})$ and $\frac{\partial C_{i}}{\partial p_{j}}=0$ for any $j\neq i$. Hence, we have $h^{i}_{i}(\bm{p})=\frac{p_{i}d(\theta_{i},\mathfrak{T})}{C_{i}(\bm{p})}=1$, and $h^{i}_{j}(\bm{p})=0$, for all $j\neq i$ and $i\in[K]$.
With the explicit forms of $\mathcal{X}$, $C_{x}$, and $\bm{h}^{x}$ derived, we directly arrive at the EB-KKT-IDS algorithm for TBP, which is proven to be universally efficient. The detailed proof is provided in Appendix E.
Theorem 6.
Let $\bm{\theta}\in\Theta$ be a Gaussian instance. Under EB-KKT-IDS, $\bm{p}_{t}\xrightarrow{\mathcal{L}}\bm{p}^{*}$.
4 Extensions
In this section, we briefly discuss several important problems not covered by the main framework introduced in Section 3. The first is related to alternative performance criteria, specifically the probability of good selection (PGS), also known as $(\epsilon,\delta)$-PAC (probably approximately correct) in the learning literature. The second concerns a practical scenario in which Gaussian distributions have unknown variances.
4.1 Optimal probability of good selection guarantee: Gaussian $\varepsilon$-best-arm identification
This section demonstrates that our framework can be extended to develop optimal algorithms for a variant of the R&S problem, specifically R&S with a probability of good selection (PGS) guarantee. This formulation is related to, but distinct from, the probability of correct selection (PCS) in the indifference-zone (IZ) formulation. For a detailed discussion, see Eckman and Henderson (2018) and Hong et al. (2021). In the bandit literature, this problem is referred to as the $\varepsilon$-best-arm identification problem (Garivier and Kaufmann, 2021; Jourdan et al., 2023a), or $(\epsilon,\delta)$-PAC learning.
Consider the same $K$-armed bandits. Let $\mathcal{I}_{\varepsilon}(\bm{\theta})=\{i\in[K]:\theta_{i}\geq\theta_{I^{*}}-\varepsilon\}$ denote the set of “good enough arms”, whose gap from the best arm $I^{*}=\operatorname*{argmax}_{i}\theta_{i}$ (assuming uniqueness) is within a specified tolerance $\varepsilon$. The $\varepsilon$-BAI problem aims to find any arm $i\in\mathcal{I}_{\varepsilon}(\bm{\theta})$.
To apply our framework, the first step is to establish a lower bound on the problem complexity. Notice that the $\varepsilon$-BAI problem differs significantly from the BAI problem in that the correct answer is non-unique; any arm $i\in\mathcal{I}_{\varepsilon}$ can serve as a correct answer. When $\varepsilon$ exceeds the gap between the best and second-best arm, there may be multiple correct answers. For a discussion on pure-exploration problems with multiple correct answers, see Degenne and Koolen (2019).
In this section, we restrict our focus to Gaussian bandits with a homogeneous variance of $\sigma^{2}=1$. or Gaussian bandits, Jourdan et al. (2023a, Lemma 9) establishes the lower bound on problem complexity
$\lim_{\delta\to 0}\frac{\mathbb{E}_{\bm{\theta}}[\tau_{\delta}]}{\log(1/\delta)}\geq(\Gamma^{*}_{\bm{\theta},\varepsilon})^{-1},\hbox{ where }\Gamma^{*}_{\bm{\theta},\varepsilon}=\max_{i\in\mathcal{I}_{\varepsilon}(\bm{\theta})}\max_{\bm{p}\in\mathcal{S}_{K}}\min_{j\neq i}\frac{(\theta_{i}-\theta_{j}+\varepsilon)^{2}}{2\left(\frac{1}{p_{i}}+\frac{1}{p_{j}}\right)}=\max_{\bm{p}\in\mathcal{S}_{K}}\min_{j\neq I^{*}}\frac{(\theta_{I^{*}}-\theta_{j}+\varepsilon)^{2}}{2\left(\frac{1}{p_{I^{*}}}+\frac{1}{p_{j}}\right)}.$ (14)
To interpret the complexity lower bound, note that finding any $\varepsilon$-good arm can be achieved by asking, for each arm $i$ the question: Is arm $i$ an $\varepsilon$-good arm? This question can be further decomposed into smaller hypothesis tests $H_{0,i,j}:\theta_{j}-\varepsilon\geq\theta_{i}$ versus $H_{1,i,j}:\theta_{j}-\varepsilon<\theta_{i}$ for any $j\neq i$. To see this, if there exists an arm $i\neq j$ for which $H_{0,i,j}$ holds, then arm $i$ cannot be $\varepsilon$-good, as $\theta_{j}-\theta_{i}>\varepsilon$. Therefore, to assert that arm $i$ is $\varepsilon$-good, we need to reject $H_{0,i,j}$ for all $j\neq i$. The term $\frac{(\theta_{i}-\theta_{j}+\varepsilon)^{2}}{2\left(\frac{1}{p_{i}}+\frac{1}{p_{j}}\right)}$ represents the test statistic, where a higher value increases the likelihood of rejecting $H_{0,i,j}$. Thus, $\min_{j\neq i}\frac{(\theta_{i}-\theta_{j}+\varepsilon)^{2}}{2\left(\frac{1}{p_{i}}+\frac{1}{p_{j}}\right)}$ corresponds to the information needed to assert that $i$ is $\varepsilon$-good. The problem complexity depends on maximizing this information across all arms under their optimal allocations. For the last equality, we observe that the best arm $I^{*}$ is always the easiest to confirm as $\varepsilon$-good.
With the maximin complexity lower bound in (14), the $\varepsilon$-BAI problem falls within the scope of our framework. Specifically, we can adapt the KKT detection rule (Algorithm 6) by substituting in
$C_{j}(\bm{p};\bm{\theta})=\frac{(\theta_{I^{*}(\bm{\theta})}-\theta_{j}+\varepsilon)^{2}}{2\left(\frac{1}{p_{I^{*}(\bm{\theta})}}+\frac{1}{p_{j}}\right)},\quad\hbox{and}\quad h^{j}_{i}(\bm{p})=\frac{p_{j}}{p_{i}+p_{j}},\quad h^{j}_{j}(\bm{p})=\frac{p_{i}}{p_{i}+p_{j}},\quad\hbox{for }j\neq I^{*}(\bm{\theta}).$
This approach yields the EB-TCε algorithm studied in Jourdan et al. (2023a). Following our proof techniques for Theorem 5, Jourdan et al. (2023a) establishes the fixed-confidence optimality of EB-TCε, when combined with IDS and recommending the empirical best arm upon stopping.
Proposition 1 (Lemma 24 in Jourdan et al. 2023a).
Under EB-TCε-IDS, we have $\bm{p}_{t}\xrightarrow{\mathcal{L}}\bm{p}^{*}.$
4.2 Dealing with unknown variances
The Gaussian distributions with unknown variances no longer belong to the single-parameter exponential family. Fortunately, our framework can still be applied with appropriate modifications to the GLRT statistic.
Let $(\bm{\theta},\bm{\sigma}^{2})$ denote the unknown mean and variance vectors. The alternative set is given by $\mathrm{Alt}(\bm{\theta},\bm{\sigma}^{2})=\{(\bm{\vartheta},\bm{\varsigma}^{2})\in\Theta\times\mathbb{R}_{+}^{K}:I^{*}(\bm{\vartheta})\neq I^{*}(\bm{\theta})\},$ where $I^{*}(\bm{\theta})$ is the index of the best arm under $\bm{\theta}$. This alternative set can be decomposed into a finite union of convex sets, thus satisfying Assumption 1,
$\mathrm{Alt}(\bm{\theta},\bm{\sigma}^{2})=\cup_{x\in\mathcal{X}(\bm{\theta},\bm{\sigma}^{2})}\mathrm{Alt}_{x}(\bm{\theta},\bm{\sigma}^{2}),$
where $\mathcal{X}(\bm{\theta},\bm{\sigma}^{2})\triangleq[K]\backslash\{I^{*}(\bm{\theta})\}$ and $\mathrm{Alt}_{x}(\bm{\theta},\bm{\sigma}^{2})=\{(\bm{\vartheta},\bm{\varsigma}^{2})\in\Theta\times\mathbb{R}_{+}^{K}:\vartheta_{x}>\theta_{x}\}.$ The generalized Chernoff information is similarly defined as the population version of the GLRT statistic, with the key distinction that the likelihood ratio now involves the KL divergence for a Gaussian distribution with two parameters:
$C_{x}(\bm{p};\bm{\theta},\bm{\sigma}^{2})\triangleq\inf_{(\bm{\vartheta},\bm{\varsigma}^{2})\in\mathrm{Alt}_{x}(\bm{\theta},\bm{\sigma}^{2})}\sum_{i\in[K]}p_{i}d\left((\theta_{i},\sigma^{2}_{i}),(\vartheta_{i},\varsigma^{2}_{i})\right),\quad\text{for all }x\in\mathcal{X}(\bm{\theta},\bm{\sigma}^{2}),$ (15)
where $d\left((\theta,\sigma^{2}),(\vartheta,\varsigma^{2})\right)=\frac{1}{2}\left(\frac{(\theta-\vartheta)^{2}}{\varsigma^{2}}+\frac{\sigma^{2}}{\varsigma^{2}}-1-\log\left(\frac{\sigma^{2}}{\varsigma^{2}}\right)\right).$ Following Jourdan et al. (2023b, Lemma 11),
$C_{x}(\bm{p};\bm{\theta},\bm{\sigma}^{2})=\inf_{\vartheta\in\left[\theta_{x},\theta_{I^{*}(\bm{\theta})}\right]}\frac{1}{2}\left[{p_{I^{*}(\bm{\theta})}}\log\left(1+\frac{(\theta_{I^{*}(\bm{\theta})}-\vartheta)^{2}}{\sigma^{2}_{I^{*}(\bm{\theta})}}\right)+{p_{x}}\log\left(1+\frac{(\theta_{x}-\vartheta)^{2}}{\sigma^{2}_{x}}\right)\right].$
It is important to note that the minimizer $\vartheta$ in the definition cannot be resolved in closed form.
Compared to the known-variance case, the generalized Chernoff information takes a distinct form, revealing a fundamental difference between the two problems. Notably, the generalized Chernoff information under unknown variances is strictly smaller than that for known variances, and there are problem instances where the ratio between the two can be arbitrarily small (Jourdan et al. 2023b, Lemma 4 and 12). Crucially, substituting empirical variances into the complexity measures derived for known-variance cases, such as those in Glynn and Juneja (2004); Garivier and Kaufmann (2016), does not produce optimal results.
Fortunately, the key components underlying our algorithm framework remain intact. The first component is the maximin formulation of the optimal allocation problem, which now reads:
$\Gamma_{\bm{\theta},\bm{\sigma}^{2}}^{*}=\max_{\bm{p}\in\mathcal{S}_{k}}\min_{x\in\mathcal{X}(\bm{\theta},\bm{\sigma}^{2})}C_{x}(\bm{p};\bm{\theta},\bm{\sigma}^{2}).$
[…]
$h_{i}^{x}(\bm{p};\bm{\theta},\bm{\sigma}^{2})\triangleq\frac{p_{i}\frac{\partial}{\partial p_{i}}C_{x}(\bm{p};\bm{\theta},\bm{\sigma}^{2})}{C_{x}(\bm{p};\bm{\theta},\bm{\sigma}^{2})},\quad\hbox{for}\ x\in\mathcal{X}(\bm{\theta})\ \hbox{and}\ i\in[K],$ (16)
is still a valid probability vector since $C_{x}(\bm{p};\bm{\theta},\bm{\sigma}^{2})$ is homogeneous of degree one by definition. The KKT condition retains the same form as in Theorem 4, with appropriate substitutions to account for the unknown variances. Specifically, $C_{x}(\bm{p};\bm{\theta},\bm{\sigma}^{2})$ defined in (15) replaces $C_{x}(\bm{p};\bm{\theta})$, and $\mathcal{X}(\bm{\theta},\bm{\sigma}^{2})$ replaces $\mathcal{X}(\bm{\theta})$. Additionally, the new IDS rule is applied, using $h_{i}^{x}(\bm{p};\bm{\theta},\bm{\sigma}^{2})$ as defined in (16). The added complexity stems from the absence of closed-form expressions for the partial derivatives, though these can be efficiently computed using a bisection search. With these specifications, a KKT detection rule – and thus a complete algorithm – can be implemented directly. Similarly, Bayesian detection rules, such as TS and PPS, are straightforward to implement when paired with an appropriate conjugate prior.
5 Numerical Experiments
While best-$k$-arm identification may seem like a straightforward extension of BAI, it significantly alters the problem structure. This added complexity complicates the generalization of traditional approaches initially developed for BAI. Due to space constraints, we provide a detailed discussion of these challenges and the performance of our ACC algorithm in Appendix F.
For the numerical study, we focus on the following problem instances. For each case, we examine both Bernoulli and Gaussian bandits (with variance $\sigma^{2}=1$) and adopt an uninformative prior for our algorithms. Additional numerical experiments are collected in Appendix I.
Case ID $K,k$ $\bm{\theta}$
1 $5,2$ $(0.1,0.2,0.3,0.4,0.5)$
2 $20,5$ $\theta_{i}=0.05*i,\forall i\in[20]$
3 $15,1$ $\theta_{i}=0.3,\forall i\in[14];\theta_{15}=0.7$
4 $100,10$ $\theta_{i}=0.3,\forall i\in[90];\theta_{i}=0.7,\forall i\in[100]\backslash[90]$
5 $50,25$ $\theta_{i}=0.2,\forall i\in[10];\theta_{i}=0.5,\forall i\in[25]\backslash[10];\theta_{i}=0.8,\forall i\in[50]\backslash[25]$
5.1 Fixed-confidence setting
For benchmarks, we compare our proposed algorithms with existing ones, which are briefly summarized in Appendix H. The $m$-LinGapE and MisLid algorithms, implemented by Tirinzoni and Degenne (2022), are designed for linear bandits and therefore are not implemented for Bernoulli bandits. To ensure fair comparisons, we set the exploration rate as $\beta({t,\delta})=\log\left((\log(t)+1)/\delta\right)$ in UGapE and KL-LUCB, along with their respective stopping rules. Table 1 presents the fixed-confidence sample complexity, highlighting the superior performance of our proposed algorithms. The FWS algorithm also performs competitively, though with significantly higher computational complexity. Table 1 also reports computation times for selected cases, demonstrating our algorithm’s efficiency. Notably, computation time is primarily constrained by the stopping rule calculation, which scales as $O(k(K-k))$; see Section 5.4.3 for further discussion.
Case $\delta$ LB TTTS-IDS TS-KKT-IDS TS-PPS-IDS FWS $m$-LinGapE MisLid LMA KL-LUCB UGapE Uniform
Bernoulli 1 0.1 487 838$\pm$12 763$\pm$14 857$\pm$13 852$\pm$14 1131$\pm$13 1643$\pm$26 1639$\pm$25 1322$\pm$13
0.01 974 1349$\pm$19 1287$\pm$19 1414$\pm$20 1323$\pm$20 1683$\pm$18 2687$\pm$38 2666$\pm$38 2330$\pm$17
2 0.1 1637 3417$\pm$27 2779$\pm$25 3462$\pm$26 2964$\pm$25 6329$\pm$25 5453$\pm$48 5726$\pm$50 18465$\pm$42
0.01 3275 5007$\pm$38 4406$\pm$37 5067$\pm$38 4565$\pm$36 9314$\pm$39 8856$\pm$74 8851$\pm$73 31236$\pm$60
3 0.1 155 350$\pm$1 282$\pm$1 354$\pm$1 284$\pm$1 517$\pm$1 533$\pm$4 550$\pm$2 610$\pm$1
0.01 311 504$\pm$2 435$\pm$2 506$\pm$2 443$\pm$1 725$\pm$1 843$\pm$6 901$\pm$2 899$\pm$1
4 0.1 1114 3536$\pm$2 2755$\pm$1 3668$\pm$1 2834$\pm$3 7350$\pm$1 3859$\pm$3 4562$\pm$2 8992$\pm$1
0.01 2228 4644$\pm$2 3522$\pm$2 4716$\pm$2 3868$\pm$3 9772$\pm$1 6048$\pm$5 7161$\pm$2 11607$\pm$1
5 0.1 934 2868$\pm$2 2121$\pm$2 2951$\pm$2 2536$\pm$5 4513$\pm$2 3870$\pm$48 3419$\pm$4 6703$\pm$2
0.01 1869 3759$\pm$3 2869$\pm$3 3820$\pm$3 3341$\pm$5 6053$\pm$3 5203$\pm$5 5566$\pm$22 8745$\pm$2
Gaussian 1 0.1 2159 3919$\pm$58 3667$\pm$61 3991$\pm$61 3684$\pm$60 4327$\pm$58 4411$\pm$65 5809$\pm$63 7775$\pm$121 7618$\pm$117 5975$\pm$58
0.01 4318 6230$\pm$85 5949$\pm$89 6169$\pm$85 5974$\pm$89 6717$\pm$84 6786$\pm$91 8431$\pm$78 11932$\pm$168 11956$\pm$167 10162$\pm$80
2 0.1 9459 19424$\pm$151 16371$\pm$146 19282$\pm$148 17385$\pm$153 24526$\pm$162 24534$\pm$182 34280$\pm$152 32650$\pm$284 32828$\pm$284 101990$\pm$225
0.01 18918 29214$\pm$218 26274$\pm$215 29685$\pm$225 27227$\pm$227 33563$\pm$215 36828$\pm$270 46383$\pm$220 52323$\pm$436 50977$\pm$417 178697$\pm$322
3 0.1 647 1484$\pm$5 1181$\pm$5 1455$\pm$5 1240$\pm$7 1720$\pm$4 1964$\pm$7 1959$\pm$5 2497$\pm$44 2331$\pm$7 2617$\pm$5
0.01 1294 2152$\pm$7 1827$\pm$7 2110$\pm$7 1904$\pm$7 2477$\pm$5 2883$\pm$9 2768$\pm$6 3627$\pm$25 3757$\pm$10 3933$\pm$6
4 0.1 4605 14756$\pm$8 19101$\pm$161* 15035$\pm$6 17251$\pm$52 19585$\pm$5 25815$\pm$15 24967$\pm$5 53498$\pm$441* 66248$\pm$463* 38579$\pm$4
0.01 9210 19543$\pm$12 20279$\pm$87* 19742$\pm$8 21669$\pm$53 25595$\pm$6 32946$\pm$16 32672$\pm$6 29321$\pm$128* 31197$\pm$76* 49312$\pm$5
5 0.1 4177 12994$\pm$14 16492$\pm$191* 13113$\pm$11 19577$\pm$102 15660$\pm$11 20502$\pm$20 18676$\pm$10 55535$\pm$552* 56074$\pm$480* 31350$\pm$8
0.01 8355 17226$\pm$15 17340$\pm$111* 17291$\pm$14 21875$\pm$96 20333$\pm$14 27519$\pm$28 25050$\pm$13 25373$\pm$138* 25896$\pm$88* 40921$\pm$9
Case $\delta$ ($K,k$) Execution time per iteration in microseconds ($\upmu$s)
Gaussian 1 0.1 (5,2) 90.3 8.5 7.5 405.8 12.3 17.5 12.9 1.2 3.6 3.2
0.01 (5,2) 738.7 8.8 8.0 385.6 12.2 16.1 12.2 1.1 3.9 3.1
4 0.1 (100,10) 279.4 657.5 530.3 31295.7 12998.1 9844.7 4440.2 10.4 153.9 217.1
0.01 (100,10) 591.4 680.4 531.2 32563.8 13660.1 10323.8 4555.7 10.3 167.6 218.9
Table 1: Fixed-confidence sample complexity, averaged over $1000$ replications. Cases where the algorithm did not stop within $10^{6}$ steps for some replications are marked with *, with problem complexity recorded as $10^{6}$ in these cases.
5.2 Convergence
We rigorously establish that the empirical allocation under our proposed algorithms converges strongly to the optimal allocation for Gaussian BAI. In Figure 1, we present the empirical convergence rate of $O(1/\sqrt{t})$ to the optimal value for our algorithms in Case 1, where $\bm{\theta}=(0.5,0.4,0.3,0.2,0.1)$, for identifying the best-$2$ under both Gaussian and Bernoulli reward distributions. These results suggest that this convergence behavior extends to more general settings. We record the trajectories of $\Gamma_{\bm{\theta}}(\bm{p}_{t})$ over $1000$ independent replications and plot the mean trajectories, with shaded areas representing the first and third quartiles.
Figure 1: Convergence to the optimal value at an empirical rate of $O(t^{-1/2})$.
5.3 Advantage of IDS over $\beta$-tuning.
To isolate the effect of the selection rule, we implement ACC in Algorithm 5 with the Thompson sampling estimation rule and the PPS detection rule, comparing the IDS selection rule in Algorithm 2 with $0.5$-tuning, i.e., set $h^{j}_{i}=h^{j}_{j}=0.5$. We consider the slippage configuration where $\bm{\theta}$ is given by $\theta_{i}=0.75,i\in[k]$ and $\theta_{j}=0.5,j\in[K]\backslash[k]$ for both Gaussian and Bernoulli bandits. Instances are labeled by $(K,k)$-distribution. Table 2 reports the significant reduction in sample complexity achieved by using IDS.
$(K,k)$-dist $(500,1)$-G $(500,2)$-G $(500,3)$-G $(500,5)$-G $(500,1)$-B $(500,2)$-B $(500,3)$-B $(500,5)$-B
TS-PPS-0.5 285823 285591 283401 283180 62596 61914 62660 62734
TS-PPS-IDS 199374 216668 223860 241869 44640 47044 50436 52767
Reduction 30.2% 24.1% 21.0% 14.6% 28.7% 24.0% 19.5% 15.9%
Table 2: Fixed-confidence sample complexity with $\delta=0.001$, averaged over 100 replications.
5.4 Large-scale experiments
The simulation literature has recently seen a surge of interest in large-scale R&S problems, which involve a substantial number of alternative systems. For those interested, we recommend the recent survey by Fan et al. (2024), which provides a comprehensive overview of large-scale simulation optimization. In this section, we evaluate the performance of our proposed algorithms on several large-scale R&S problem instances, considering both fixed-confidence and fixed-budget formulations.
5.4.1 Fixed-confidence.
In this experiment, we consider the large-scale slippage configuration studied in Zhong and Hong (2022), where the means of the systems are set as $\mu_{0}=0.1$ and $\mu_{i}=0.0$ for $i=2,\dots,K$. The noisy performance of the systems is independent and follows Gaussian distributions with a known variance of $\sigma^{2}=1$. The desired confidence level is set to $\delta=0.05$.
In Table 3, we compare the fixed-confidence performance of Procedure $\mathcal{KT}_{0}$ (Zhong and Hong, 2022), Procedure $\mathcal{KN}$ (Kim and Nelson, 2001), and two proposed algorithms, TS-KKT-IDS and TTTS-IDS. The simulation results for Procedures $\mathcal{KT}_{0}$ and $\mathcal{KN}$ are taken from Zhong and Hong (2022, Table 2). Both $\mathcal{KT}_{0}$ and $\mathcal{KN}$ require the specification of an indifference zone parameter, which is set to the (unknown) true gap of $0.1$ in this experiment. In contrast, our algorithms are entirely parameter-free. For our fixed-confidence algorithms, the stopping rule (10) is applied using a heuristic threshold $\gamma(t,\delta)=\Phi^{-1}(1-\delta)$, i.e., the $1-\delta$ quantile of the standard normal distribution. All results are averaged over $1000$ independent replications.
TS-KKT-IDS TTTS-IDS EB-TS-IDS Procedure $\mathcal{KT}_{0}$ Procedure $\mathcal{KN}$
$K$ PCS $\mathbb{E}[\tau_{\delta}]/K$ PCS $\mathbb{E}[\tau_{\delta}]/K$ PCS $\mathbb{E}[\tau_{\delta}]/K$ PCS $\mathbb{E}[\tau_{\delta}]/K$ PCS $\mathbb{E}[\tau_{\delta}]/K$
$10^{2}$ $0.999$ $616\pm 16$ $1.0$ $758\pm 14$ $0.999$ $727\pm 12$ $0.963$ $1462\pm 3$ $0.978$ $884\pm 8$
$10^{3}$ $1.0$ $764\pm 14$ $1.0$ $864\pm 12$ $1.0$ $788\pm 7$ $0.965$ $1512\pm 1$ $0.989$ $1165\pm 8$
$10^{4}$ $1.0$ $937\pm 13$ $1.0$ $1015\pm 12$ $0.999$ $898\pm 5$ $0.956$ $1527\pm 1$ $0.994$ $1419\pm 9$
$10^{5}$ $1.0$ $1069\pm 14$ $1.0$ $1154\pm 13$ $1.0$ $1024\pm 3$ $0.955$ $1528\pm 1$ $0.994$ $1686\pm 8$
Table 3: Sample complexity and estimated PCS are compared for the proposed algorithms, Procedure $\mathcal{KT}_{0}$ and Procedure $\mathcal{KN}$ across different numbers of systems in a slippage configuration with known and equal variances.
The theoretical stopping rule is designed to ensure PCS for high-confidence levels, i.e., as $\delta\to 0$, but empirically, it tends to be overly conservative for moderate PCS level guarantees. Thus, we implemented a heuristic stopping rule better suited to larger $\delta$. Our results show that this heuristic rule remains conservative, with an estimated PCS exceeding 0.999 in all cases. However, the asymptotically optimal allocation compensates for the conservatism of the stopping rule, achieving smaller sample complexities and higher PCS than competitors. Future work should focus on developing stopping rules that are less conservative for moderate $\delta$. Recent studies, such as Wang et al. (2024) and Tirinzoni and Degenne (2022), propose elimination-based stopping rules that could reduce conservatism and improve computational efficiency. We leave this direction for future research and present this numerical study to illustrate the potential of asymptotically optimized algorithms in large-scale experiments.
5.4.2 Fixed-budget.
In this experiment, we evaluate the performance of the proposed algorithms under configurations studied by Li et al. (2024), including the slippage configuration ($\mu_{0}=0.1$ and $\mu_{i}=0.0$ for $i=2,\dots,K$ with $\sigma_{i}^{2}$ = 1), and the equally-spaced configuration with increasing variances ($\mu_{0}=0.1$ and $\mu_{i}=-i/K$ for $i=2,\dots,K$ with $\sigma_{i}^{2}=1+i/K$). We assume that all system variances are known.
A key insight from Li et al. (2024) is that allocating samples greedily – sampling the system with the highest empirical mean – keeps PCS bounded away from zero as the total sampling budget scales linearly with the number of systems, i.e., $B=cK$ for a fixed constant $c$ as $K\to\infty$. This suggests that a degree of greediness is essential in large-scale R&S to promptly eliminate an empirical best system that is not the true best. They introduced the Explore-First Greedy (EFG) procedure to ensure consistency.
Inspired by their work, we test our proposed algorithms with the EB estimation rule, i.e., the EB-*-IDS algorithms. Figure 2 presents the estimated PCS for the proposed algorithms and benchmarks, including EFG, Greedy, and uniform-allocation procedures. PCS is estimated from $10^{4}$ independent replications, with confidence intervals nearly invisible and thus omitted. As expected, EB-*-IDS algorithms appear to be sample-optimal in the sense described by Hong et al. (2022); Li et al. (2024), as the estimated PCS stabilizes at a positive value as $K\to\infty$. This sample optimality is anticipated since the EB-*-IDS algorithms allocate sufficient samples to the empirical best. We leave a rigorous investigation of this sample optimality for future work. The advantage of our proposed algorithms becomes more evident as the average budget per system increases, demonstrating that a dynamic allocation strategy informed by asymptotic analysis can enhance performance under moderate to low PCS levels.
Figure 2: Estimated PCS are compared for the proposed algorithms and benchmarks. Top row is the slippage configuration with equal variances, bottom row is the equally-spaced configuration with increasing variances.
5.4.3 Computation time.
To compare the computation time of the proposed algorithms, we consider the slippage configuration, where system means are set as $\mu_{0}=0.1$ and $\mu_{i}=0.0$ for $i=2,\dots,K$. The noisy performance of each system is independent and follows Gaussian distributions with a known variance of $\sigma^{2}=1$. We vary the number of systems from $K=10^{1}$ to $10^{4}$. For each instance, we run the algorithm for $100K$ or $1000K$ time steps and perform $10^{3}$ independent repetitions to estimate the sample mean of the running time along with its 95% confidence interval.
Table 4 reports the execution time of our proposed algorithms, measured in average microseconds ($\upmu$s) per iteration. We observe that both TS-KKT-IDS and TS-PPS-IDS are invariant to budget changes, while TTTS-IDS slows down as the time period grows. However, this computational inefficiency in TTTS-IDS is minimal when PCS is below $0.97$. Additionally, the slowdown is less pronounced for larger $K$ as sampling an alternative best system becomes easier with more alternatives available.
The computation times of all three algorithms scale nearly linearly with the number of systems $K$, as expected. For instance, in TS-KKT-IDS, the bottleneck is searching for the minimum among $K-1$ $C_{j}$ functions. Notably, when an algorithm targets constant optimality, such scaling is inevitable – after all, even the computation complexity of calculating the problem complexity lower bound scales linearly with $K$.
TTTS-IDS TS-PPS-IDS TS-KKT-IDS
Budget $100K$ $1000K$ $100K$ $1000K$ $100K$ $1000K$
$K=10^{1}$ $0.66\pm 0.01(0.409)$ $5.93\pm 1.11(0.974)$ $0.84\pm 0.01(0.411)$ $1.04\pm 0.01(0.979)$ $0.35\pm 0.00(0.443)$ $0.39\pm 0.01(0.971)$
$K=10^{2}$ $3.38\pm 0.12(0.161)$ $7.86\pm 0.59(0.991)$ $4.75\pm 0.17(0.179)$ $5.16\pm 0.18(0.980)$ $1.92\pm 0.07(0.204)$ $1.99\pm 0.07(0.988)$
$K=10^{3}$ $22.59\pm 0.98(0.065)$ $25.85\pm 0.99(0.978)$ $32.94\pm 1.43(0.059)$ $33.51\pm 1.44(0.975)$ $13.16\pm 0.57(0.057)$ $13.13\pm 0.57(0.983)$
$K=10^{4}$ $170.13\pm 7.96(0.026)$ $172.56\pm 7.92(0.964)$ $297.96\pm 14.27(0.018)$ $297.52\pm 14.22(0.955)$ $98.79\pm 4.62(0.019)$ $98.62\pm 4.61(0.964)$
Table 4: Computation time for the proposed algorithms, measured in average microseconds ($\upmu$s) per iteration. The $95\%$ confidence interval is provided after the “$\pm$” sign, and the estimated PCS is shown in parentheses.
5.5 Practical considerations in algorithm implementation
Exploration versus exploitation in pure exploration.
The exploration-exploitation trade-off surprisingly arises in pure exploration problems. For example, the estimation rules EB and TS serve to identify the best arm, with EB greedily selecting (i.e., exploiting) the arm with the highest sample mean and TS introducing exploration by sampling from the posterior. While asymptotic performance is largely insensitive to the choice of estimation rule, practical performance under moderate confidence levels or limited budgets can differ significantly. EB offers faster convergence when the greedy guess is correct, as the sampling rule aligns with the correct allocation. However, it risks poor performance when the guess is wrong, as allocations are based on a misspecified problem instance. Conversely, TS reduces the risk of getting stuck with incorrect estimates by exploring more but requires additional samples to compensate for unnecessary exploration when the estimate is already accurate. As a result, EB-*-IDS algorithms [^6] Except for EB-KKT-IDS, where both estimation and detection are greedy without exploration, which result in significant amount of outliers for the slippage configurations. typically achieve better average performance but are prone to extreme outliers caused by highly inaccurate initial sample means, whereas TS-*-IDS algorithms show slightly worse average performance but far fewer outliers. This behavior is discussed in Wu and Zhou (2018a); Jourdan et al. (2022) and conceptually linked to Simchi-Levi et al. (2024). Additionally, as noted by Li et al. (2024), EB generally stabilizes the PCS at positive levels as the number of alternatives grows to infinity, while TS does not.
A similar trade-off arises in the choice of the detection rule (det in Section 3.5.2). While KKT greedily selects the $x$ with the smallest $C_{x}$ as culprit, TS and PPS encourage exploration of culprits. For example, comparing TTTS-IDS with TS-KKT-IDS in Table 1 and Table 3, we see that *-KKT-IDS algorithms achieve better average performance but are prone to extreme outliers (e.g., Table 1, Case 4 and 5, Gaussian). In contrast, *-TS-IDS algorithms exhibit slightly worse average performance but no outliers. However, in most cases, the exploration in the estimation rule alone is sufficient to mitigate outliers.
Common Random Numbers (CRN).
In simulation studies, the performance of different systems is often simulated using common random numbers (CRN), introducing positive correlations between the noisy performances. This typically reduces sampling complexity by decreasing the variances of the differences in sample means, see (Nelson and Matejcik, 1995; Zhong and Hong, 2022).
The main challenge in fully leveraging CRN lies in the need to allow noisy performances of different systems to be mutually correlated. For simplicity of exposition, our framework assumes that the performances are i.i.d. random variables from a single-parameter exponential family. However, our framework can be extended to handle correlated observations.
To demonstrate, consider the case where performances follow a multivariate normal distribution with means $\mu_{i}$, variances $\sigma_{i}^{2}$, and pairwise correlation $\rho$, where both $\sigma_{i}^{2}$ and $\rho$ are known. [^7] Extensions to settings with unknown covariance matrices are possible, as shown in Section 4.2 using the multiparameter exponential family. The generalized Chernoff information is modified to account for the correlation as:
$C_{j}(\bm{p};\bm{\theta})=\frac{(\theta_{I^{*}}-\theta_{j})^{2}}{2\left(\sigma_{I^{*}}^{2}/p_{I^{*}}+\sigma_{j}^{2}/p_{j}-2\rho\sigma_{I^{*}}\sigma_{j}/\sqrt{p_{I^{*}}p_{j}}\right)},\quad\hbox{for }j\in[K]\backslash\{I^{*}\}.$ (17)
It can be verified that (8) holds, along with Theorem 4, enabling the immediate application of an ACC algorithm. For the fixed-confidence setting, the modified $C_{j}$ must also replace the original in the stopping rule. While the universal efficiency of ACC under CRN remains an open question, we present numerical evidence in Table 5 to demonstrate its effectiveness. Establishing formal theoretical guarantees for this extension is a promising direction for future work.
TS-KKT-IDS
Case $K$ $\rho=0.0$ $\rho=0.25$ $\rho=0.5$ $\rho=0.75$
A $10^{2}$ 72$\pm$4(0.1%) 62$\pm$4(1.2%) 52$\pm$3(2.1%) 35$\pm$2(4.3%)
$10^{3}$ 87$\pm$3(0.0%) 80$\pm$3(0.1%) 72$\pm$3(0.3%) 55$\pm$2(1.6%)
$10^{4}$ 112$\pm$1(0.0%) 106$\pm$2(0.0%) 94$\pm$2(0.0%) 75$\pm$2(0.5%)
B $10^{2}$ 616$\pm$16(0.0%) 537$\pm$16(0.2%) 423$\pm$17(0.8%) 254$\pm$10(3.5%)
$10^{3}$ 743$\pm$14(0.0%) 670$\pm$16(0.0%) 554$\pm$17(0.4%) 336$\pm$12(0.8%)
$10^{4}$ 937$\pm$13(0.0%) 832$\pm$17(0.0%) 693$\pm$18(0.0%) 455$\pm$14(0.2%)
Table 5: Sample complexity per arm and estimated PCS are compared for the TS-KKT-IDS algorithm under various choice of the correlation parameter $\rho$. Case A denotes the slippage configuration with equal variances in Section 5.4.1 and Case B denotes the equally-spaced configuration with increasing variances in Section 5.4.2.
6 Conclusion
This paper presents a unified algorithm design framework for general pure-exploration problems, linking asymptotic lower bounds on problem complexity to maximin optimal allocation problems. Through this connection, we derive reformulated Karush-Kuhn-Tucker conditions that incorporate dual variables, offering a new perspective on the top-two algorithm design principle as a method for tracking complementary slackness conditions. Additionally, the stationarity conditions motivate an information-directed selection rule which, when combined with the TTTS algorithm, achieves optimal performance for Gaussian BAI in terms of posterior convergence, fixed-confidence, and static fixed-budget settings. This approach enables straightforward yet optimal algorithm designs that adapt to various pure-exploration problems.
Our framework demonstrates broad applicability, extending optimally to problems such as $\varepsilon$-BAI and thresholding bandits, with the significant advantage of not relying on forced exploration. We propose modifications to Thompson sampling for easy use by practitioners facing pure-exploration problems. Experiments validate our algorithms’ superior performance in terms of fixed-budget probability of incorrect selection and fixed-confidence sample complexity, showing robust performance across varied configurations, including large-scale problems.
This work lays the foundation for simple, adaptable algorithms in pure-exploration problems. Future directions include refining stopping rules to improve practical efficiency, especially in finite-sample or moderate-confidence regimes (Simchowitz et al., 2017), as well as extending theoretical guarantees to broader settings, such as those with heavy-tailed distributions (Agrawal et al., 2020). Another promising avenue is to develop parallelizable versions of the algorithm, which may be well-suited for modern, large-scale simulation applications.
Acknowledgments
Wei You received support from the Hong Kong Research Grants Council [ECS Grant 26212320 and GRF Grant 16212823]. A preliminary version of this paper appeared as an extended abstract in the Proceedings of the 36th Annual Conference on Learning Theory, COLT’23, with the title “Information-Directed Selection for Top-Two Algorithms.”
Appendix
[…]
subsubsection
[…]
Appendix A Proofs for Section 2
A.1 Proof of Theorem 1
Theorem 1 follows directly from the following Lemma 4 and Lemma 1.
Lemma 4.
The optimal allocation problem (5) has a unique solution satisfying (1) and (2). Furthermore,
$\Gamma^{*}_{\bm{\theta}}=\min_{j\neq I^{*}}C_{j}(\bm{p}^{*};\bm{\theta})=\sum_{j\neq I^{*}}\frac{1}{d\left(\left(\theta_{j},\sigma_{j}^{2}\right),\left(\bar{\theta}^{*}_{I^{*},j},\sigma_{j}^{2}\right)\right)}.$ (18)
See 1
The rest of this section devotes to the proof of the two lemmas.
Proof of Lemma 4.
The uniqueness of the solution to (5) is a special case of the uniqueness of optimal allocation for the best-$k$-arm identification problem, established in Proposition 9.
To show that (1) and (2) are necessary and sufficient conditions for (5), we invoke the equivalent KKT conditions for the general pure exploration problem presented in Theorem 4, which specializes to
$p_{I^{*}}=\sum_{j^{\prime}\in[K]\backslash\{I^{*}\}}\mu_{j^{\prime}}h_{I^{*}}^{j^{\prime}},$ (19a)
$p_{j}=\mu_{j}h_{j}^{j},\quad\forall j\neq I^{*},$ (19b)
$\phi=\min_{j^{\prime}\in[K]\backslash\{I^{*}\}}C_{j^{\prime}}(\bm{p}),\quad\text{and}\quad\sum_{j^{\prime}\in[K]\backslash\{I^{*}\}}\mu_{j^{\prime}}(\phi-C_{j^{\prime}}(\bm{p}))=0.$ (19c)
Then, (2) is equivalent to (19a) by solving each $\mu^{*}_{j}$ explicitly from (19b) and $\sum_{i}p^{*}_{i}=1$. Notice that by the positivity of $\bm{p}^{*}$ from Proposition 9 and the fact that $\mu^{*}_{j}=p^{*}_{j}/h^{j}_{j}$, we have $\mu^{*}_{j}>0$ for all $j\neq I^{*}$. This, combined with (19c), is equivalent to (1).
For the alternative expression (18), it can be easily verified that
$\mu^{*}_{j}=\frac{C_{j}(\bm{p}^{*})}{d\left(\left(\theta_{j},\sigma_{j}^{2}\right),\left(\bar{\theta}^{*}_{I^{*},j},\sigma_{j}^{2}\right)\right)},\quad\text{where }\bar{\theta}^{*}_{I^{*},j}\triangleq\frac{p^{*}_{I^{*}}\theta_{I^{*}}/\sigma^{2}_{I^{*}}+p^{*}_{j}\theta_{j}/\sigma^{2}_{j}}{p^{*}_{I^{*}}/\sigma^{2}_{I^{*}}+p^{*}_{j}/\sigma^{2}_{j}}.$
From (19c), we have $C_{j}(\bm{p}^{*})=\Gamma^{*}_{\bm{\theta}}$ for all $j\neq I^{*}$, which implies that
$\mu^{*}_{j}=\frac{\Gamma^{*}_{\bm{\theta}}}{d\left(\left(\theta_{j},\sigma_{j}^{2}\right),\left(\bar{\theta}^{*}_{I^{*},j},\sigma_{j}^{2}\right)\right)}.$
Then (18) follows from the dual feasibility condition $\sum_{j\neq I^{*}}\mu_{j}=1$. ∎
Proof of Lemma 1.
This lemma follows from the proof of Qin et al. (2017, Theorem 2, part 1) by directly replacing the equal variances with heterogeneous variances. ∎
Furthermore, the dual variables provide an alternative interpretation of the optimal allocation problem as a zero-sum game between the DM and an adversarial skeptic.
Remark 9 (Game perspective and dual variables).
Consider a two-player zero-sum game where the DM seeks to gather evidence to assert that $I^{*}(\bm{\theta})$ is the best alternative by allocating measurement effort $\bm{p}$. Meanwhile, a skeptic attempts to challenge the DM by proposing an alternative problem instance $\bm{\vartheta}$ such that $I^{*}(\bm{\vartheta})\neq I^{*}(\bm{\theta})$, aiming to “trick” the DM by being statistically close to $\bm{\theta}$. The DM’s goal is to rule out any $\bm{\vartheta}$ proposed by the skeptic. In this two-player zero-sum game, the Nash equilibrium is characterized by mixed strategies on both sides. The dual variable can be interpreted as the optimal mixed strategy employed by the skeptic. Specifically, the skeptic selects from one of $K-1$ “hard instances,” defined as
$\bar{\theta}^{*}_{I^{*},j}=\frac{p^{*}_{I^{*}}\theta_{I^{*}}/\sigma^{2}_{I^{*}}+p^{*}_{j}\theta_{j}/\sigma^{2}_{j}}{p^{*}_{I^{*}}/\sigma^{2}_{I^{*}}+p^{*}_{j}/\sigma^{2}_{j}},\quad\text{for}\quad j\neq I^{*}(\bm{\theta}).$
The $j$-th hard instance $\bm{\vartheta}^{j}$ corresponds to a parameter such that $\mathcal{I}^{*}(\bm{\vartheta}^{j})=j\neq I^{*}(\bm{\theta})$, and is the statistically hardest instance to distinguish from the true $\bm{\theta}$. The probability that the skeptic raises $\bm{\vartheta}^{j}$ is exactly the dual variable $\mu^{*}_{j}$.
Remark 10 (Lai-Robbins-type formula).
The problem complexity measure exibits a Lai-Robbins-type formula below, resembling the one obtained in Lai and Robbins (1985) for regret minimization:
$\Gamma^{*}_{\bm{\theta}}$ $=\max_{\bm{p}\in\mathcal{S}_{K}}\min_{j\neq I^{*}}C_{j}(\bm{p};\bm{\theta})$
$=\left(\sum_{j\neq I^{*}}\frac{\mu_{j}}{C_{j}(\bm{p}^{*})}\right)^{-1}$
$=\left(\sum_{j\neq I^{*}}\frac{1}{d\left(\left(\theta_{j},\sigma_{j}^{2}\right),\left(\bar{\theta}^{*}_{I^{*},j},\sigma_{j}^{2}\right)\right)}\right)^{-1},$
where we used the fact that $\sum_{j\neq I^{*}}\mu_{j}=1$, the information balance (1), and
$\mu^{*}_{j}=\frac{p_{j}^{*}}{1-h^{j}_{I^{*}}(\bm{p}^{*})}=\frac{\Gamma^{*}_{\bm{\theta}}}{d\left(\left(\theta_{j},\sigma_{j}^{2}\right),\left(\bar{\theta}^{*}_{I^{*},j},\sigma_{j}^{2}\right)\right)},\quad\forall j\neq I^{*}.$
A.2 Proof of Theorem 2
By Theorem 1, it suffices to show that the sample allocation converges almost surely to the unique optimal allocation $\bm{p}^{*}$ to (5), i.e., Lemma 2 restated below. See 2
For simplicity of exposition, we focus on the equal-variance case. The proof in this case follows directly from Corollary 3 and Theorem 5.
The case with heterogeneous variances can be handled with straightforward but tedious modifications to Theorem 5. In particular, in any inequality that involves the variances, $\sigma^{2}$ must be replaced appropriately by $\sigma^{2}_{\min}=\min_{i}\sigma^{2}_{i}$ or $\sigma^{2}_{\max}=\max_{i}\sigma^{2}_{i}$ to relax the inequality.
Appendix B A threshold for the stopping rule and the Proof of Theorem 3
B.1 A threshold for the stopping rule to ensure $\delta$-correctness
Let $h(u)=u-\log u$ and define for any $z\in[1,e]$ and $x\geq 0$
$\tilde{h}_{z}(x)=\begin{cases}e^{1/h^{-1}(x)}h^{-1}(x)&\hbox{if }x\geq h(1/\log z),\\ z(x-\log\log z)&\hbox{otherwise.}\end{cases}$
Furthermore, define
$\mathcal{C}_{\mathrm{exp}}(x)=2\tilde{h}_{3/2}\left(\frac{h^{-1}(1+x)+\log(2\zeta(2))}{2}\right),$
where $\zeta(s)=\sum_{n=1}^{\infty}n^{-s}$ is the Riemann zeta function. One can easily verify that $\mathcal{C}_{\mathrm{exp}}(x)\sim x$ when $x\to\infty$.
Proposition 2 (Kaufmann and Koolen 2021, Proposition 15).
The stopping rule (10) with the sequence of thresholds
$\gamma(t,\delta)=3\sum_{i=1}^{K}\log(1+\log N_{t,i})+K\mathcal{C}_{\mathrm{exp}}\left(\frac{\log(1/\delta)}{K}\right)$ (20)
is such that, for every sampling rule and any $\delta>0$,
$\mathbb{P}_{\bm{\theta}}\left(\tau_{\delta}<\infty,\,\,\mathcal{I}(\bm{\theta}_{\tau_{\delta}})=\mathcal{I}(\bm{\theta})\right)\geq 1-\delta,\quad\hbox{for any problem instance }\bm{\theta}\in\Theta,$
where $\bm{\theta}_{t}$ is the sample mean of the arms at time $t$.
B.2 Proof of Theorem 3
To establish this result, we rely on the following technical finding implied by Proposition 14 in Qin and Russo (2024).
Proposition 3 (Proposition 14 in Qin and Russo 2024).
If a sampling rule $\pi$ satisfies that there exist $\rho>0$, $\alpha>0$ and ${T}\in\mathcal{L}$ such that for any $t\geq{T}$, $\min_{i\in[k]}N_{t,i}\geq\rho\cdot t^{\alpha}$, then $\bm{\theta}_{t}\xrightarrow{\mathcal{L}}\bm{\theta}$.
Proof of Theorem 3.
The first condition in Theorem 3 ensures that the policy $\pi$ is $\delta$-correct (Definition 2), i.e., $\pi\in\mathcal{A}$.
Now we fix $\bm{\theta}\in\Theta$ and write $\bm{p}^{*}=\bm{p}^{*}(\bm{\theta})$ for notational convenience. Since $\bm{p}^{*}>\bm{0}$ component-wisely, the second condition $\bm{p}_{t}\xrightarrow{\mathcal{L}}\bm{p}^{*}$ in Theorem 3 ensures that the condition in Proposition 3 is satisfied, and thus we have $\bm{\theta}_{t}\xrightarrow{\mathcal{L}}\bm{\theta}$. Then by the continuous mapping theorem,
$\min_{x\in\mathcal{X}(\bm{\theta}_{t})}C_{x}(\bm{p}_{t};\bm{\theta}_{t})\xrightarrow{\mathcal{L}}\min_{x\in\mathcal{X}(\bm{\theta})}C_{x}(\bm{p}^{*};\bm{\theta})=\Gamma^{*}_{\bm{\theta}},$
where the equality follows from that $\bm{p}^{*}$ is the maximizer of (9). For any $\epsilon>0$, there exists a random time $T_{\epsilon,1}\in\mathcal{L}$ such that for any $t\geq T_{\epsilon,1}$,
$\min_{x\in\mathcal{X}(\bm{\theta}_{t})}C_{x}(\bm{p}_{t};\bm{\theta}_{t})\geq\Gamma^{*}_{\bm{\theta}}-\epsilon.$
On the other hand, there exists a deterministic time $T_{\epsilon,2}$ such that for any $t\geq T_{\epsilon,2}$,
$\gamma(t,\delta)\leq 3K\log(1+\log t)+K\mathcal{C}_{\mathrm{exp}}\left(\frac{\log(1/\delta)}{K}\right)\leq\epsilon t+K\mathcal{C}_{\mathrm{exp}}\left(\frac{\log(1/\delta)}{K}\right).$
Hence, for any $t\geq T_{\epsilon}\triangleq\max\{T_{\epsilon,1},T_{\epsilon,2}\}\in\mathcal{L}$,
$t\cdot\min_{x\in\mathcal{X}(\bm{\theta}_{t})}C_{x}(\bm{p}_{t};\bm{\theta}_{t})-\gamma(t,\delta)\geq\left(\Gamma^{*}_{\bm{\theta}}-2\epsilon\right)t-K\mathcal{C}_{\mathrm{exp}}\left(\frac{\log(1/\delta)}{K}\right),$
which implies
$\tau_{\delta}\leq\max\left\{T_{\epsilon},\frac{K\mathcal{C}_{\mathrm{exp}}\left(\frac{\log(1/\delta)}{K}\right)}{\Gamma^{*}_{\bm{\theta}}-2\epsilon}\right\}\leq T_{\epsilon}+\frac{K\mathcal{C}_{\mathrm{exp}}\left(\frac{\log(1/\delta)}{K}\right)}{\Gamma^{*}_{\bm{\theta}}-2\epsilon}.$
Since $T_{\epsilon}=\max\{T_{\epsilon,1},T_{\epsilon,2}\}\in\mathcal{L}$, we have $\mathbb{E}[T_{\epsilon}]<\infty$, and then
$\limsup_{\delta\to 0}\frac{\mathbb{E}^{\pi}_{\bm{\theta}}[\tau_{\delta}]}{\log(1/\delta)}\leq\frac{1}{\Gamma^{*}_{\bm{\theta}}-2\epsilon},$
where we also use the fact that $\mathcal{C}_{\mathrm{exp}}(x)\sim x$ when $x\to\infty$. Since $\epsilon>0$ can be arbitrarily small, we have $\limsup_{\delta\to 0}\frac{\mathbb{E}^{\pi}_{\bm{\theta}}[\tau_{\delta}]}{\log(1/\delta)}\leq\frac{1}{\Gamma^{*}_{\bm{\theta}}}.$ We note that a matching lower bound on the sample complexity has been established in Wang et al. (2021, Appendix B), i.e., for any other $\delta$-correct algorithm $\pi^{\prime}$,
$\liminf_{\delta\to 0}\frac{\mathbb{E}^{\pi^{\prime}}_{\bm{\theta}}[\tau_{\delta}]}{\log(1/\delta)}\geq\frac{1}{\Gamma^{*}_{\bm{\theta}}},$
and thus we have
$\limsup_{\delta\to 0}\frac{\mathbb{E}^{\pi}_{\bm{\theta}}[\tau_{\delta}]}{\mathbb{E}^{\pi^{\prime}}_{\bm{\theta}}[\tau_{\delta}]}\leq 1.$
The above inequality holds for any $\bm{\theta}\in\Theta$, which implies that $\pi$ is universal efficient (Definition 3). This completes the proof. ∎
Appendix C Proof of Theorem 4
Proof.
Note that the optimal allocation problem (9) is equivalent to
$\Gamma^{*}=\max_{\phi,\bm{p}}$ $\hskip 10.0pt\phi$ (21a)
$\mathrm{s.t.}$ $\hskip 10.0pt\sum_{i\in[K]}p_{i}-1=0,$ (21b)
$\hskip 10.0ptp_{i}\geq 0,\quad\forall i\in[K],$ (21c)
$\hskip 10.0pt\phi-C_{x}(\bm{p})\leq 0,\quad\forall x\in\mathcal{X}.$ (21d)
The corresponding Lagrangian is given by
$\mathcal{L}(\phi,\bm{p},\lambda,\bm{\mu})=\phi-\lambda\left(\sum_{i\in[K]}p_{i}-1\right)-\sum_{x\in\mathcal{X}}\mu_{x}(\phi-C_{x}(\bm{p}))+\sum_{i\in[K]}\iota_{i}p_{i}.$ (22)
Slater’s condition holds since $\bm{p}=\bm{1}/K$ and $\phi=0$ is a feasible solution such that all inequality constraints hold strictly, where $\bm{1}$ is a vector of $1$’s. Hence, the KKT conditions are necessary and sufficient for global optimality, which are then given by
$-1+\sum_{x\in\mathcal{X}}\mu_{x}=0,$ (23a)
$\lambda-\sum_{x\in\mathcal{X}}\mu_{x}\frac{\partial C_{x}(\bm{p})}{\partial p_{i}}-\iota_{i}=0,\quad\forall i\in[K],$ (23b)
$\sum_{i\in[K]}p_{i}-1=0,\quad p_{i}\geq 0,\quad\forall i\in[K],$ (23c)
$\phi-C_{x}(\bm{p})\leq 0,\quad\forall x\in\mathcal{X},$ (23d)
$\mu_{x}\geq 0,\quad\forall x\in\mathcal{X},\quad\iota_{i}\geq 0\quad\forall i\in[K],$ (23e)
$\sum_{x\in\mathcal{X}}\mu_{x}(\phi-C_{x}(\bm{p}))=0,\quad\sum_{i\in[K]}\iota_{i}p_{i}=0.$ (23f)
We first show that the KKT conditions (23) implies (13). To this end, we reformulate the stationarity condition (23b), leaving the feasibility and complementary slackness conditions unchanged. Define the selection functions $h^{x}_{i}$ as in (12), which we repeat here for convenience:
$h^{x}_{i}\triangleq h^{x}_{i}(\bm{\theta},\bm{p})=\frac{p_{i}d(\theta_{i},\vartheta^{x}_{i})}{C_{x}(\bm{p})},\quad\forall(x,i)\in\mathcal{X}\times[K].$ (24)
Now, we can rewrite the KKT conditions by eliminating the dual variables $\lambda$ and $\iota$.
$\lambda-\sum_{x\in\mathcal{X}}\mu_{x}\frac{\partial C_{x}(\bm{p})}{\partial p_{i}}-\iota_{i}=0,\quad\forall i\in[K],$
[…]
Multiply $\text{Multiply }p_{i}\text{ and use C-S for $\bm{\iota}$}\quad\Rightarrow\quad$ $\frac{\lambda}{\phi}p_{i}-\sum_{x\in\mathcal{X}}\mu_{x}\frac{p_{i}\frac{\partial C_{x}(\bm{p})}{\partial p_{i}}}{C_{x}(\bm{p})}=0,\quad\forall i\in[K],$ (26)
[…]
Sum over $\text{Sum over }i\quad\Rightarrow\quad$ $\frac{\lambda}{\phi}-\sum_{x\in\mathcal{X}}\mu_{x}\sum_{i\in[K]}h^{x}_{i}=0,\quad\forall i\in[K],$
[…]
Substituting $\lambda=\phi$ into (26) yields the desired result. Note that (26) is equivalent to (25) if and only if $\bm{p}^{*}$ is strictly positive.
Next, we show that (13) implies the KKT conditions in (23) when the optimal solution is strictly positive. First, when $\bm{p}^{*}$ is strictly positive, we have $\bm{\iota}=\bm{0}$. We have that (23a), (23d), (23e), and (23f) hold automatically. Let $\lambda=\phi$, we have (23b) using (26), (23f) and $\bm{\iota}=\bm{0}$. Finally, (23c) is obtained by summing (13a) over all $i\in[K]$ and noting that $\bm{\mu}\in\mathcal{S}_{|\mathcal{X}|}$. This completes the proof. ∎
Appendix D Proof of Theorem 5
D.1 Preliminary
D.1.1 Notation
For Gaussian BAI with reward distributions $N(\theta_{i},\sigma^{2})$ for $i\in[K]$, the unique best arm is denoted by $I^{*}=I^{*}(\bm{\theta})=\operatorname*{argmax}_{i\in[K]}\theta_{i}$. The functions $C_{i,j}$ can be calculated as follows,
$C_{i,j}(\bm{p})=\frac{\left(\theta_{i}-\theta_{j}\right)^{2}}{2\sigma^{2}\left(p_{i}^{-1}+p_{j}^{-1}\right)},\quad\forall j\neq i.$ (27)
and the selection functions specialize to
$h_{i,j}=h_{i,j}(\bm{p})\triangleq\frac{p_{i}^{-1}}{p_{i}^{-1}+p_{j}^{-1}}=\frac{p_{j}}{p_{i}+p_{j}},\quad\forall j\neq i.$ (28)
The empirical versions of $C_{i,j}(\cdot)$ and $h_{i,j}$ can be obtained by plugging in sample proportion $\bm{p}_{t}$ and mean estimators $\bm{\theta}_{t}$.
Minimum and maximum gaps.
We define the minimum and maximum values between the expected rewards of two different arms:
$\Delta_{\min}\triangleq\min_{i\neq j}\left\lvert\theta_{i}-\theta_{j}\right\rvert\quad\text{and}\quad\Delta_{\max}\triangleq\max_{i,j\in[K]}\left(\theta_{i}-\theta_{j}\right).$
By the assumption that the arm means are unique, we have $\Delta_{\max}\geq\Delta_{\min}>0$.
Another measure of cumulative effort and average effort.
We have introduced one measure of cumulative effort and average effort: the number of samples and the proportion of total samples allocated allocated to arm $i\in[K]$ before time $t\in\mathbb{N}_{0}$ are denoted as
$N_{t,i}=\sum_{\ell=0}^{t-1}\mathds{1}(I_{\ell}=i)\quad\text{and}\quad p_{t,i}=\frac{N_{t,i}}{t}.$
To analyze randomized algorithms (such as TTTS), we introduce the probability of sampling arm $i$ at time $t$ as
$\psi_{t,i}\triangleq\mathbb{P}(I_{t}=i|\mathcal{H}_{t}),$
where the history $\mathcal{H}_{t}$ is the sigma-algebra generated by the observations $\left(I_{0},Y_{1,I_{0}},\ldots,I_{t-1},Y_{t,I_{t-1}}\right)$ available up to time $t$. Then we define another measure of cumulative effort and average effort:
$\Psi_{t,i}\triangleq\sum_{\ell=0}^{t-1}\psi_{\ell,i}\quad\text{and}\quad w_{t,i}\triangleq\frac{\Psi_{t,i}}{t}.$
Posterior means and variances under independent uninformative priors.
Under the independent uninformative prior, the posterior mean and variance of the unknown mean of arm $i\in[K]$ are
$\theta_{t,i}=\frac{1}{N_{t,i}}\sum_{\ell=0}^{t-1}\mathds{1}(I_{\ell}=i)Y_{\ell+1,I_{\ell}}\quad\text{and}\quad\sigma_{t,i}^{2}=\frac{\sigma^{2}}{N_{t,i}}.$
Posterior probability of being the best arm.
To analyze the TTTS algorithm, we need the posterior probability that arm $i$ emerges as the best arm, defined as
$\alpha_{t,i}\triangleq\mathbb{P}_{\tilde{\bm{\theta}}\sim\Pi_{t}}\left(\tilde{\theta}_{i}=\max_{j\in[k]}\tilde{\theta}_{j}\right).$
D.1.2 Definition and properties of strong convergence
The proof of our Theorem 5 requires bounding time until quantities reaches their limits. Following Qin and Russo (2022), we first introduce a class of random variables that are “light-tailed.” Let the space $\mathcal{L}_{p}$ consist of all measurable $X$ with $\|X\|_{p}<\infty$ where $\|X\|_{p}\triangleq\left(\mathbb{E}\left[|X|^{p}\right]\right)^{1/p}$ is the $p$-norm.
Definition 5 (Definition 1 in Qin and Russo 2022).
For a real valued random variable $X$, we say $X\in\mathbb{M}$ if and only if $\|X\|_{p}<\infty$ for all $p\geq 1$. Equivalently, $\mathbb{M}=\cap_{p\geq 1}\mathcal{L}_{p}$.
Then we introduce the so-called strong convergence for random variables, which is a stronger notion of convergence than almost sure convergence.
Definition 6 ($\mathbb{M}$-convergence, Definition 2 in Qin and Russo 2022).
For a sequence of real valued random variables $\{X_{t}\}_{t\in\mathbb{N}_{0}}$ and a scalar $x\in\mathbb{R}$, we say $X_{t}\xrightarrow{\mathbb{M}}x$ if
for $\text{for all }\epsilon>0\,\text{there exists }\,T\in\mathbb{M}\,\text{ such that for all }t\geq T,\,\,|X_{t}-x|\leq\epsilon.$
We say $X_{t}\xrightarrow{\mathbb{M}}\infty$ if
for $\text{for all }c>0\,\text{there exists }\,T\in\mathbb{M}\,\text{ such that forall }t\geq T,\,\,X_{t}\geq c,$
and similarly, we say $X_{t}\xrightarrow{\mathbb{M}}-\infty$ if $-X_{t}\xrightarrow{\mathbb{M}}\infty$.
For a sequence of random vectors $\{\bm{X}_{t}\}_{t\in\mathbb{N}_{0}}$ taking values in $\mathbb{R}^{d}$ and a vector $\bm{x}\in\mathbb{R}^{d}$, where $\bm{X}_{t}=(X_{t,1},\ldots,X_{t,d})$ and $\bm{x}=(x_{1},\ldots,x_{d})$, we say $\bm{X}_{t}\xrightarrow{\mathbb{M}}\bm{x}$ if $X_{t,i}\xrightarrow{\mathbb{M}}x_{i}$ for all $i\in[d]$.
Corollary 3.
If $\bm{X}_{t}\xrightarrow{\mathbb{M}}\bm{x}$, then $\bm{X}_{t}\xrightarrow{\mathcal{L}}\bm{x}$ and $\bm{X}_{t}\xrightarrow{\mathrm{a.s.}}\bm{x}$.
Note that the sufficient condition for optimality in Theorem 3 requires only $\mathcal{L}$-convergence, which is implied by the stronger $\mathbb{M}$-convergence. In the proofs of the fixed-confidence optimality (Theorem 5 and Theorem 6), we establish the stronger $\mathbb{M}$-convergence.
Qin and Russo (2022) also introduces a number of properties of the class $\mathbb{M}$ (Definition 5) and strong convergence (Definition 6).
Lemma 5 (Lemma 1 in Qin and Russo 2022: Closedness of $\mathbb{M}$).
Let $X$ and $Y$ be non-negative random variables in $\mathbb{M}$.
  1. 1.
    $aX+bY\in\mathbb{M}$ for any scalars $a,b\in\mathbb{R}$.
  2. 2.
    $XY\in\mathbb{M}$.
  3. 3.
    $\max\{X,Y\}\in\mathbb{M}$.
  4. 4.
    $X^{q}\in\mathbb{M}$ for any $q\geq 0$.
  5. 5.
    If $g:\mathbb{R}\to\mathbb{R}$ satisfies $\sup_{x\in-[c,c]}|g(x)|<\infty$ for all $c\geq 0$ and $|g(x)|=O(|x|^{q})$ as $|x|\to\infty$ for some $q\geq 0$, then $g(X)\in\mathbb{M}$.
  6. 6.
    If $g:\mathbb{R}\to\mathbb{R}$ is continuous and $|g(x)|=O(|x|^{q})$ as $|x|\to\infty$ for some $q\geq 0$, then $g(X)\in\mathbb{M}$.
As mentioned in Qin and Russo (2022), many of the above properties can be extended to any finite collection of random variables in $\mathbb{M}$. A notable one is that if $X_{i}\in\mathbb{M}$ for each $i\in[d]$ then
$X_{1}+\cdots+X_{d}\in\mathbb{M}\quad\text{and}\quad\max\{X_{1},\ldots,X_{d}\}\in\mathbb{M}.$
Qin and Russo (2022) also shows an equivalence between pointwise convergence and convergence in the maximum norm and a continuous mapping theorem under strong convergence (Definition 6).
Lemma 6 (Lemma 2 in Qin and Russo 2022).
For a sequence of random vectors $\{X_{n}\}_{n\in\mathbb{N}}$ taking values in $\mathbb{R}^{d}$ and a vector $x\in\mathbb{R}^{d}$, if $X_{n}\xrightarrow{\mathbb{M}}x$, then $\|X_{n}-x\|_{\infty}\xrightarrow{\mathbb{M}}0$. Equivalently, if $X_{n,i}\xrightarrow{\mathbb{M}}x_{i}$ for all $i\in[d]$, then for all $\epsilon>0$, there exists $N\in\mathbb{M}$ such that $n\geq N$ implies $|X_{n,i}-x_{i}|\leq\epsilon$ for all $i\in[d]$.
Lemma 7 (Lemma 3 in Qin and Russo 2022: Continuous Mapping).
For any $x$ taking values in a normed vector space and any random sequence $\{X_{n}\}_{n\in\mathbb{N}}$:
  1. 1.
    If $g$ is continuous at $x$, then $X_{n}\xrightarrow{\mathbb{M}}x$ implies $g(X_{n})\xrightarrow{\mathbb{M}}g(x)$.
  2. 2.
    If the range of function $g$ belongs to $\mathbb{R}$ and $g(y)\to\infty$ as $y\to\infty$, then $X_{n}\xrightarrow{\mathbb{M}}\infty$ implies $g(X_{n})\xrightarrow{\mathbb{M}}\infty$.
D.1.3 Maximal inequalities
Following Qin et al. (2017); Shang et al. (2020), we introduce the following path-dependent random variable to control the impact of observation noises:
$W_{1}=\sup_{(t,i)\in\mathbb{N}_{0}\times[K]}\sqrt{\frac{N_{t,i}+1}{\log(N_{t,i}+e)}}\frac{|\theta_{t,i}-\theta_{i}|}{\sigma}.$ (29)
In addition, to control the impact of algorithmic randomness, we introduce the other path-dependent random variable:
$W_{2}=\sup_{(t,i)\in\mathbb{N}_{0}\times[K]}\frac{|N_{t,i}-\Psi_{t,i}|}{\sqrt{(t+1)\log(t+e^{2})}}.$ (30)
The next lemma ensures that these maximal deviations have light tails.
Lemma 8 (Lemma 6 of Qin et al. 2017 and Lemma 4 of Shang et al. 2020).
For any $\lambda>0$,
$\mathbb{E}[e^{\lambda W_{1}}]<\infty\quad\text{and}\quad\mathbb{E}[e^{\lambda W_{2}}]<\infty.$
The definition of $W_{2}$ in Equation (30) implies the following corollary for sufficiently large $t$.
Corollary 4.
There exists a deterministic $T$ such that for any $t\geq T$,
$|N_{t,i}-\Psi_{t,i}|\leq W_{2}t^{0.6},\quad\forall i\in[K].$
We remark that the exponent $0.6$ is not essential in the sense that it can be any constant in $(0.5,1)$. Corollary 4 implies
$p_{t,i}-w_{t,i}\xrightarrow{\mathbb{M}}0,\quad\forall i\in[K].$
D.2 Proof outline
Stopping rule and $\delta$-correct guarantee.
We adopt the Chernoff stopping rule as described in Appendix B.1, where $\delta$-correct guarantee is provided.
Sufficient conditions for optimality in fixed-confidence setting.
Theorem 3 shows a sufficient condition of sampling rules that ensures the optimality in fixed-confidence setting. It is easier to study the following alternative but equivalent sufficient conditions to prove optimality of sampling rules.
Proposition 4 (Alternative sufficient conditions).
The following conditions are equivalent:
  1. 1.
    $\bm{p}_{t}\xrightarrow{\mathbb{M}}\bm{p}^{*}$;
  2. 2.
    $\frac{p_{t,j}}{p_{t,I^{*}}}\xrightarrow{\mathbb{M}}\frac{p_{j}^{*}}{p_{I^{*}}^{*}}$ for any $j\neq I^{*}$;
  3. 3.
    $\bm{w}_{t}\xrightarrow{\mathbb{M}}\bm{p}^{*}$; and
  4. 4.
    $\frac{w_{t,j}}{w_{t,I^{*}}}\xrightarrow{\mathbb{M}}\frac{p_{j}^{*}}{p_{I^{*}}^{*}}$ for any $j\neq I^{*}$.
Proof.
By Corollary 4, conditions 1 and 3 are equivalent.
Now we are going to show conditions 1 and 2 are equivalent. By Lemma 5, condition 1 implies condition 2. On the other hand, applying Lemma 5 to condition 3 gives
$\frac{1-p_{t,I^{*}}}{p_{t,I^{*}}}=\sum_{j\neq I^{*}}\frac{p_{t,j}}{p_{t,I^{*}}}\xrightarrow{\mathbb{M}}\sum_{j\neq I^{*}}\frac{p^{*}_{j}}{p^{*}_{I^{*}}}=\frac{1-p^{*}_{I^{*}}}{p^{*}_{I^{*}}},$
Using Lemma 5 again yields $p_{t,I^{*}}\xrightarrow{\mathbb{M}}p^{*}_{I^{*}},$ and then condition 2 gives $p_{t,j}\xrightarrow{\mathbb{M}}p_{j}^{*}$ for any $j\neq I^{*}$ Hence, condition 2 implies condition 1, and thus they are equivalent. Similarly, conditions 2 and 4 are equivalent. This completes the proof. ∎
To prove our Theorem 5, we are going to show that TTTS-IDS satisfies the last sufficient condition in Proposition 4:
Proposition 5.
Under TTTS-IDS,
$\frac{w_{t,j}}{w_{t,I^{*}}}\xrightarrow{\mathbb{M}}\frac{p_{j}^{*}}{p_{I^{*}}^{*}},\quad\forall j\neq I^{*}.$
The rest of this appendix is devoted to proving Proposition 5.
D.3 Empirical overall balance under TTTS-IDS
Previous papers (Qin et al., 2017; Shang et al., 2020; Jourdan et al., 2022) establishes $\beta$-optimality by proving either $w_{t,I^{*}}\xrightarrow{\mathbb{M}}\beta$ or $p_{t,I^{*}}\xrightarrow{\mathbb{M}}\beta$ where $\beta$ is a fixed tuning parameter taken as an input to the algorithms, while the optimal $\beta^{*}=$ (i.e., $p^{*}_{I^{*}}$) is unknown a priori. In our information-directed selection rule, at each time step, $h_{t}$ is adaptive to the top-two candidates, so novel proof techniques are required. In this subsection, we first show that the “empirical” version of overall balance (3) holds.
Proposition 6.
Under TTTS-IDS,
$w_{t,I^{*}}^{2}-\sum_{j\neq I^{*}}w_{t,j}^{2}\xrightarrow{\mathbb{M}}0\quad\text{and}\quad p_{t,I^{*}}^{2}-\sum_{j\neq I^{*}}p_{t,j}^{2}\xrightarrow{\mathbb{M}}0.$
To prove this proposition, we need the following supporting result.
Lemma 9.
Under TTTS-IDS, for any $t\in\mathbb{N}_{0}$,
$\alpha_{t,I^{*}}\left(\sum_{\ell\neq I^{*}}\frac{\alpha_{t,\ell}}{1-\alpha_{t,I^{*}}}\frac{p_{t,\ell}}{p_{t,I^{*}}+p_{t,\ell}}\right)\leq\psi_{t,I^{*}}\leq\alpha_{t,I^{*}}\left(\sum_{\ell\neq I^{*}}\frac{\alpha_{t,\ell}}{1-\alpha_{t,I^{*}}}\frac{p_{t,\ell}}{p_{t,I^{*}}+p_{t,\ell}}\right)+(1-\alpha_{t,I^{*}}),$
and for any $j\neq I^{*}$,
$\alpha_{t,I^{*}}\frac{\alpha_{t,j}}{1-\alpha_{t,I^{*}}}\frac{p_{t,I^{*}}}{p_{t,I^{*}}+p_{t,j}}\leq\psi_{t,j}\leq\alpha_{t,I^{*}}\frac{\alpha_{t,j}}{1-\alpha_{t,I^{*}}}\frac{p_{t,I^{*}}}{p_{t,I^{*}}+p_{t,j}}+(1-\alpha_{t,I^{*}}).$
Proof.
Fix $t\in\mathbb{N}_{0}$. Recall that $\alpha_{t,i}$ is the posterior probability at time $t$ that arm $i$ appears as the best arm. For a fixed $j\in[K]$ and any $\ell\neq j$, TTTS obtains the top-two candidates
$\left(I_{t}^{(1)},I_{t}^{(2)}\right)=\begin{cases}(j,\ell),&\text{with probability }\alpha_{t,j}\frac{\alpha_{t,\ell}}{1-\alpha_{t,j}},\\ (\ell,j),&\text{with probability }\alpha_{t,\ell}\frac{\alpha_{t,j}}{1-\alpha_{t,\ell}},\end{cases}$
and when TTTS is applied with IDS,
$\psi_{t,j}=\alpha_{t,j}\left(\sum_{\ell\neq j}\frac{\alpha_{t,\ell}}{1-\alpha_{t,j}}\frac{p_{t,\ell}}{p_{t,j}+p_{t,\ell}}\right)+\sum_{\ell\neq j}\alpha_{t,\ell}\frac{\alpha_{t,j}}{1-\alpha_{t,\ell}}\frac{p_{t,\ell}}{p_{t,j}+p_{t,\ell}}.$
For $j=I^{*}$, we have
$\psi_{t,I^{*}}$ $=\alpha_{t,I^{*}}\left(\sum_{\ell\neq I^{*}}\frac{\alpha_{t,\ell}}{1-\alpha_{t,I^{*}}}\frac{p_{t,\ell}}{p_{t,I^{*}}+p_{t,\ell}}\right)+\sum_{\ell\neq I^{*}}\alpha_{t,\ell}\frac{\alpha_{t,I^{*}}}{1-\alpha_{t,\ell}}\frac{p_{t,\ell}}{p_{t,I^{*}}+p_{t,\ell}},$
so
$\psi_{t,I^{*}}$ $\geq\alpha_{t,I^{*}}\left(\sum_{\ell\neq I^{*}}\frac{\alpha_{t,\ell}}{1-\alpha_{t,I^{*}}}\frac{p_{t,\ell}}{p_{t,I^{*}}+p_{t,\ell}}\right)$
and
$\psi_{t,I^{*}}$ $\leq\alpha_{t,I^{*}}\left(\sum_{\ell\neq I^{*}}\frac{\alpha_{t,\ell}}{1-\alpha_{t,I^{*}}}\frac{p_{t,\ell}}{p_{t,I^{*}}+p_{t,\ell}}\right)+\sum_{\ell\neq I^{*}}\alpha_{t,\ell}$
$=\alpha_{t,I^{*}}\left(\sum_{\ell\neq I^{*}}\frac{\alpha_{t,\ell}}{1-\alpha_{t,I^{*}}}\frac{p_{t,\ell}}{p_{t,I^{*}}+p_{t,\ell}}\right)+(1-\alpha_{t,I^{*}}),$
where the inequality uses that for any $\ell\neq I^{*}$, $\frac{\alpha_{t,I^{*}}}{1-\alpha_{t,\ell}}\frac{p_{t,\ell}}{p_{t,I^{*}}+p_{t,\ell}}\leq\frac{\alpha_{t,I^{*}}}{1-\alpha_{t,\ell}}=\frac{\alpha_{t,I^{*}}}{\sum_{\ell^{\prime}\neq\ell}\alpha_{t,\ell^{\prime}}}\leq 1$.
For $j\neq I^{*}$, we have
$\psi_{t,j}$ $=\sum_{\ell\neq j}\alpha_{t,\ell}\frac{\alpha_{t,j}}{1-\alpha_{t,\ell}}\frac{p_{t,\ell}}{p_{t,j}+p_{t,\ell}}+\alpha_{t,j}\left(\sum_{\ell\neq j}\frac{\alpha_{t,\ell}}{1-\alpha_{t,j}}\frac{p_{t,\ell}}{p_{t,j}+p_{t,\ell}}\right)$
$=\alpha_{t,I^{*}}\frac{\alpha_{t,j}}{1-\alpha_{t,I^{*}}}\frac{p_{t,I^{*}}}{p_{t,j}+p_{t,I^{*}}}+\sum_{\ell\neq j,I^{*}}\alpha_{t,\ell}\frac{\alpha_{t,j}}{1-\alpha_{t,\ell}}\frac{p_{t,\ell}}{p_{t,\ell}+p_{t,j}}+\alpha_{t,j}\left(\sum_{\ell\neq j}\frac{\alpha_{t,\ell}}{1-\alpha_{t,j}}\frac{p_{t,\ell}}{p_{t,j}+p_{t,\ell}}\right),$
so
$\psi_{t,j}$ $\geq\alpha_{t,I^{*}}\frac{\alpha_{t,j}}{1-\alpha_{t,I^{*}}}\frac{p_{t,I^{*}}}{p_{t,j}+p_{t,I^{*}}}$
and
$\psi_{t,j}$ $\leq\alpha_{t,I^{*}}\frac{\alpha_{t,j}}{1-\alpha_{t,I^{*}}}\frac{p_{t,I^{*}}}{p_{t,j}+p_{t,I^{*}}}+\left(\sum_{\ell\neq j,I^{*}}\alpha_{t,\ell}\right)+\alpha_{t,j}$
$=\alpha_{t,I^{*}}\frac{\alpha_{t,j}}{1-\alpha_{t,I^{*}}}\frac{p_{t,I^{*}}}{p_{t,j}+p_{t,I^{*}}}+(1-\alpha_{t,I^{*}}),$
where the inequality uses that
  1. 1.
    for any $\ell\neq j,I^{*}$, $\frac{\alpha_{t,j}}{1-\alpha_{t,\ell}}\frac{p_{t,\ell}}{p_{t,\ell}+p_{t,j}}\leq\frac{\alpha_{t,j}}{1-\alpha_{t,\ell}}=\frac{\alpha_{t,j}}{\sum_{\ell^{\prime}\neq\ell}\alpha_{t,\ell^{\prime}}}\leq 1$; and
  2. 2.
    $\sum_{\ell\neq j}\frac{\alpha_{t,\ell}}{1-\alpha_{t,j}}\frac{p_{t,\ell}}{p_{t,j}+p_{t,\ell}}\leq\sum_{\ell\neq j}\frac{\alpha_{t,\ell}}{1-\alpha_{t,j}}=1$.
This completes the proof. ∎
Now we are ready to complete the proof of Proposition 6.
Proof of Proposition 6.
By Corollary 4, it suffices to only show
$w_{t,I^{*}}^{2}-\sum_{j\neq I^{*}}w_{t,j}^{2}\xrightarrow{\mathbb{M}}0.$
For any $t\in\mathbb{N}_{0}$, we let
$G_{t}\triangleq{\Psi_{t,I^{*}}^{2}}-\sum_{j\neq I^{*}}{\Psi_{t,j}^{2}},$
and we are going to show $\frac{G_{t}}{t^{2}}\xrightarrow{\mathbb{M}}0$.
We argue that it suffices to show that there exists $T\in\mathbb{M}$ such that for any $t\geq T$,
$|G_{t+1}-G_{t}|\leq 2W_{2}t^{0.6}+2.$ (31)
The reason is that given (31), for any $t\geq T$
$|G_{t}|=\left|G_{T}+\sum_{\ell=T}^{t-1}(G_{\ell+1}-G_{\ell})\right|$ $\leq\left|G_{T}\right|+\sum_{\ell=T}^{t-1}|G_{\ell+1}-G_{\ell}|$
$\leq\left|G_{T}\right|+\sum_{\ell=T}^{t-1}\left(2W_{2}\ell^{0.6}+2\right)$
$\leq\left|G_{T}\right|+2(t-T)+2W_{2}\int_{T+1}^{t+1}x^{0.6}\mathrm{d}x$
$=\left|G_{T}\right|+2(t-T)+\frac{2W_{2}}{1.6}\left[(t+1)^{1.6}-(T+1)^{1.6}\right]$
$\leq T^{2}+2(t-T)+\frac{2W_{2}}{1.6}\left[(t+1)^{1.6}-(T+1)^{1.6}\right],$
where in the last inequality, $\left|G_{T}\right|\leq T^{2}$ holds since $G_{T}\geq-\sum_{j\neq I^{*}}\Psi_{T,j}^{2}\geq-\left(\sum_{j\neq I^{*}}\Psi_{T,j}\right)^{2}\geq-T^{2},$ and $G_{T}\leq\Psi_{T,I^{*}}^{2}\leq T^{2}.$ This gives $\frac{|G_{t}|}{t^{2}}\xrightarrow{\mathbb{M}}0.$
Now we are going to prove the sufficient condition (31). We can calculate
$G_{t+1}-G_{t}$ $=\left[\left(\Psi_{t,I^{*}}+\psi_{t,I^{*}}\right)^{2}-\Psi_{t,I^{*}}^{2}\right]-\sum_{j\neq I^{*}}\left[\left(\Psi_{t,j}+\psi_{t,j}\right)^{2}-\Psi_{t,j}^{2}\right]$
$=2\left(\psi_{t,I^{*}}\Psi_{t,I^{*}}-\sum_{j\neq I^{*}}\psi_{t,j}\Psi_{t,j}\right)+\left(\psi_{t,I^{*}}^{2}-\sum_{j\neq I^{*}}\psi_{t,j}^{2}\right).$
Then by Lemma 9, we have
$G_{t+1}-G_{t}$
$\geq$ $1\displaystyle 2\left[\alpha_{t,I^{*}}\sum_{j\neq I^{*}}\frac{\alpha_{t,j}}{1-\alpha_{t,I^{*}}}\frac{p_{t,j}\Psi_{t,I^{*}}}{p_{t,I^{*}}+p_{t,j}}-\alpha_{t,I^{*}}\sum_{j\neq I^{*}}\frac{\alpha_{t,j}}{1-\alpha_{t,I^{*}}}\frac{p_{t,I^{*}}\Psi_{t,j}}{p_{t,I^{*}}+p_{t,j}}-(1-\alpha_{t,I^{*}})\sum_{j\neq I^{*}}\Psi_{t,j}\right]-$ 1
$\geq$ $2\alpha_{t,I^{*}}\sum_{j\neq I^{*}}\frac{\alpha_{t,j}}{1-\alpha_{t,I^{*}}}\frac{p_{t,j}\Psi_{t,I^{*}}-p_{t,I^{*}}\Psi_{t,j}}{p_{t,I^{*}}+p_{t,j}}-2(1-\alpha_{t,I^{*}})t-1,$
where the last inequality uses $\sum_{j\neq I^{*}}\Psi_{t,j}\leq t$, and
$G_{t+1}-G_{t}$
$\leq$ $1\displaystyle 2\left[\alpha_{t,I^{*}}\sum_{j\neq I^{*}}\frac{\alpha_{t,j}}{1-\alpha_{t,I^{*}}}\frac{p_{t,j}\Psi_{t,I^{*}}}{p_{t,I^{*}}+p_{t,j}}+(1-\alpha_{t,I^{*}})\Psi_{t,I^{*}}-\alpha_{t,I^{*}}\sum_{j\neq I^{*}}\frac{\alpha_{t,j}}{1-\alpha_{t,I^{*}}}\frac{p_{t,I^{*}}\Psi_{t,j}}{p_{t,I^{*}}+p_{t,j}}\right]+$ 1
$\leq$ $2\alpha_{t,I^{*}}\left(\sum_{j\neq I^{*}}\frac{\alpha_{t,j}}{1-\alpha_{t,I^{*}}}\frac{p_{t,j}\Psi_{t,I^{*}}-p_{t,I^{*}}\Psi_{t,j}}{p_{t,I^{*}}+p_{t,j}}\right)+2(1-\alpha_{t,I^{*}})t+1,$
where the last inequality uses $\Psi_{t,I^{*}}\leq t$. Hence,
$|G_{t+1}-G_{t}|$ $\leq 2\alpha_{t,I^{*}}\sum_{j\neq I^{*}}\frac{\alpha_{t,j}}{1-\alpha_{t,I^{*}}}\left|\frac{p_{t,j}\Psi_{t,I^{*}}-p_{t,I^{*}}\Psi_{t,j}}{p_{t,I^{*}}+p_{t,j}}\right|+2(1-\alpha_{t,I^{*}})t+1.$
By Lemma 13, there exists $T_{1}\in\mathbb{M}$ such that for any $t\geq T_{1}$, $(1-\alpha_{t,I^{*}})t\leq 1/2$, and thus
$|G_{t+1}-G_{t}|$ $\leq 2\alpha_{t,I^{*}}\sum_{j\neq I^{*}}\frac{\alpha_{t,j}}{1-\alpha_{t,I^{*}}}\max_{j\neq I^{*}}\left|\frac{p_{t,j}\Psi_{t,I^{*}}-p_{t,I^{*}}\Psi_{t,j}}{p_{t,I^{*}}+p_{t,j}}\right|+2$
$\leq 2\max_{j\neq I^{*}}\left|\frac{p_{t,j}\Psi_{t,I^{*}}-p_{t,I^{*}}\Psi_{t,j}}{p_{t,I^{*}}+p_{t,j}}\right|+2$
$=2\max_{j\neq I^{*}}\left|\frac{N_{t,j}\Psi_{t,I^{*}}-N_{t,I^{*}}\Psi_{t,j}}{N_{t,I^{*}}+N_{t,j}}\right|+2,$
where the last inequality applies $\alpha_{t,I^{*}}\leq 1$ and $\sum_{j\neq I^{*}}\frac{\alpha_{t,j}}{1-\alpha_{t,I^{*}}}=1$. By Corollary 4, there exists $T_{2}\in\mathbb{M}$ such that for any $t\geq T_{2}$ and $i\in[K]$, $|N_{t,i}-\Psi_{t,i}|\leq W_{2}t^{0.6}$, and thus for $j\neq I^{*}$
$\left|\frac{N_{t,j}\Psi_{t,I^{*}}-N_{t,I^{*}}\Psi_{t,j}}{N_{t,I^{*}}+N_{t,j}}\right|\leq\left|\frac{N_{t,j}\left(N_{t,I^{*}}+W_{2}t^{0.6}\right)-N_{t,I^{*}}\left(N_{t,j}-W_{2}t^{0.6}\right)}{N_{t,I^{*}}+N_{t,j}}\right|\leq W_{2}t^{0.6}.$
Hence, for $t\geq T\triangleq\max\{T_{1},T_{2}\}$, $|G_{t+1}-G_{t}|\leq 2W_{2}t^{0.6}+2$. Note that $T\in\mathbb{M}$ by Lemma 5. This completes the proof of the sufficient condition (31). ∎
D.3.1 Implication of empirical overall balance
Here we present some results that are implied by empirical overall balance (Proposition 6) and needed for proving the sufficient condition for optimality in fixed-confidence setting (Proposition 5). The first one shows that when $t$ is large enough, two measures of average effort are bounded away from 0 and 1.
Lemma 10.
Let $b_{1}\triangleq\frac{1}{\sqrt{32(K-1)}}$ and $b_{2}\triangleq\frac{3}{4}$. Under TTTS-IDS, there exists a random time $T\in\mathbb{M}$ such that for any $t\geq T$,
$b_{1}\leq w_{t,I^{*}}\leq b_{2}\quad\text{and}\quad b_{1}\leq p_{t,I^{*}}\leq b_{2}.$
Since $w_{t,I^{*}}$ and $p_{t,I^{*}}$ are bounded away from 0 when $t$ is large enough, we obtain the following corollary by dividing the equations in Proposition 6 by $w_{t,I^{*}}$ and $p_{t,I^{*}}$, respectively,
Corollary 5.
Under TTTS-IDS,
$\sum_{j\neq I^{*}}\frac{w_{t,j}^{2}}{w_{t,I^{*}}^{2}}\xrightarrow{\mathbb{M}}1\quad\text{and}\quad\sum_{j\neq I^{*}}\frac{p_{t,j}^{2}}{p_{t,I^{*}}^{2}}\xrightarrow{\mathbb{M}}1.$
Proof of Lemma 10.
We only show the upper and lower bounds for $w_{t,I^{*}}$ since proving those for $p_{t,I^{*}}$ is the same.
By Proposition 6, there exists $T_{1}\in\mathbb{M}$ such that for any $t\geq T_{1}$,
$w_{t,I^{*}}^{2}-\sum_{j\neq I^{*}}w_{t,j}^{2}\leq\frac{1}{2}.$
This implies
$\frac{1}{2}\geq w_{t,I^{*}}^{2}-\left(\sum_{j\neq I^{*}}w_{t,j}\right)^{2}=w_{t,I^{*}}^{2}-\left(1-w_{t,I^{*}}\right)^{2}=2w_{t,I^{*}}-1,$
which gives $w_{t,I^{*}}\leq b_{2}\triangleq\frac{3}{4}$.
Using Proposition 6 again, there exists $T_{2}\in\mathbb{M}$ such that for any $t\geq T_{2}$,
$w_{t,I^{*}}^{2}-\sum_{j\neq I^{*}}w_{t,j}^{2}\geq-\frac{1}{32(K-1)}.$
By Cauchy-Schwarz inequality,
$w_{t,I^{*}}^{2}$ $\geq-\frac{1}{32(K-1)}+\frac{\left(\sum_{j\neq I^{*}}w_{t,j}\right)^{2}}{K-1}$
$=-\frac{1}{32(K-1)}+\frac{\left(1-w_{t,I^{*}}\right)^{2}}{K-1}$
$\geq-\frac{1}{32(K-1)}+\frac{1}{16(K-1)}=\frac{1}{32(K-1)},$
where the last inequality uses $w_{t,I^{*}}\leq b_{2}=\frac{3}{4}$. This gives $w_{t,I^{*}}\geq b_{1}\triangleq\frac{1}{\sqrt{32(K-1)}}$. Taking $T\triangleq\max\{T_{1},T_{2}\}$ completes the proof since $T\in\mathbb{M}$ by Lemma 5. ∎
D.4 Sufficient exploration under TTTS-IDS
In this subsection, we show that applied with information-directed selection, TTTS sufficiently explores all arms:
Proposition 7.
Under TTTS-IDS, there exists $T\in\mathbb{M}$ such that for any $t\geq T$,
$\min_{i\in[K]}N_{t,i}\geq\sqrt{t/K}.$
The previous analyses in Qin et al. (2017); Shang et al. (2020); Jourdan et al. (2022) only work for any sequence of tuning parameters $\{h_{t}\}_{t\in\mathbb{N}_{0}}$ that is uniformly strictly bounded, i.e., there exists $h_{\min}>0$ such that with probability one,
$\inf_{t\in\mathbb{N}_{0}}\min\{h_{t},1-h_{t}\}\geq h_{\min}.$
Our IDS is adaptive to the algorithmic randomness in TTTS in the sense that it depends on the algorithmic randomness in picking top-two candidates, so it does not enjoy the property above. We need some novel analysis to prove the sufficient exploration of TTTS-IDS.
Following Shang et al. (2020) of analyzing TTTS, we define the following auxiliary arms, the two “most promising arms”:
$J_{t}^{(1)}\in\operatorname*{argmax}_{i\in[K]}\alpha_{t,i}\quad\text{and}\quad J_{t}^{(2)}\in\operatorname*{argmax}_{i\neq J_{t}^{(1)}}\alpha_{t,i}.$ (32)
We further define the one that is less sampled as:
$J_{t}\triangleq\operatorname*{argmin}_{i\in\left\{J_{t}^{(1)},J_{t}^{(2)}\right\}}N_{t,i}=\operatorname*{argmin}_{i\in\left\{J_{t}^{(1)},J_{t}^{(2)}\right\}}p_{t,i}.$
Note that the identity of $J_{t}$ can change over time. Next we show that IDS allocates a decent amount of effort to $J_{t}$.
Lemma 11.
Under TTTS-IDS, for any $t\in\mathbb{N}_{0}$,
$\psi_{t,J_{t}}\geq\frac{1}{2K(K-1)}.$
Proof.
Fix $t\in\mathbb{N}_{0}$. We have
$\left(I_{t}^{(1)},I_{t}^{(2)}\right)=\left(J_{t}^{(1)},J_{t}^{(2)}\right)\quad\text{with probability}\quad\alpha_{t,J_{t}^{(1)}}\frac{\alpha_{t,J_{t}^{(2)}}}{1-\alpha_{t,J_{t}^{(1)}}}\geq\frac{1}{K(K-1)},$
where the inequality follows from the definition of $J_{t}^{(1)}$ and $J_{t}^{(2)}$. Note that
$h_{J_{t}^{(1)},J_{t}^{(2)}}(\bm{p}_{t};\bm{\theta}_{t})=\frac{p_{t,J_{t}^{(2)}}}{p_{t,J_{t}^{(1)}}+p_{t,J_{t}^{(2)}}}.$
If $J_{t}=\operatorname*{argmin}_{i\in\left\{J_{t}^{(1)},J_{t}^{(2)}\right\}}p_{t,i}=J_{t}^{(1)}$,
$\psi_{t,J_{t}}\geq\frac{1}{K(K-1)}\frac{p_{t,J_{t}^{(2)}}}{p_{t,J_{t}^{(1)}}+p_{t,J_{t}^{(2)}}}\geq\frac{1}{2K(K-1)}.$
Otherwise, $J_{t}=\operatorname*{argmin}_{i\in\left\{J_{t}^{(1)},J_{t}^{(2)}\right\}}p_{t,i}=J_{t}^{(2)}$ and
$\psi_{t,J_{t}}\geq\frac{1}{K(K-1)}\frac{p_{t,J_{t}^{(1)}}}{p_{t,J_{t}^{(1)}}+p_{t,J_{t}^{(2)}}}\geq\frac{1}{2K(K-1)}.$
This completes the proof. ∎
Following Qin et al. (2017); Shang et al. (2020), we define an insufficiently sampled set for any $t\in\mathbb{N}_{0}$ and $s\geq 0$:
$U_{t}^{s}\triangleq\{i\in[K]\,:\,N_{t,i}<s^{1/2}\}.$
Proposition 7 can be rephrased as that there exists $T\in\mathbb{M}$ such that for all $t\geq T$, $U_{t}^{t/K}$ is empty.
Lemma 11 in Shang et al. (2020) shows a nice property of TTTS: if TTTS with a selection or tuning rule allocates a decent amount of effort to $J_{\ell}$ at any time $\ell$, the insufficiently sampled set $U_{t}^{s}$ is empty when $t$ is large enough.
Lemma 12 (Lemmas 9 and 11 in Shang et al. 2020).
Under TTTS with any selection rule (e.g., IDS) such that
$\exists\psi_{\min}>0\quad\forall t\in\mathbb{N}_{0}\quad\psi_{t,J_{t}}\geq\psi_{\min},$
there exists $S\in\mathbb{M}$ such that for any $s\geq S$, $U^{s}_{\lfloor Ks\rfloor}=\emptyset$.
We remark that, while the original result focuses on $\beta$-tuning, the same proof applies to any selection rule that satisfies the condition.
Now we can complete the proof of Proposition 7.
Proof of Proposition 7.
By Lemma 11, IDS satisfies the condition in Lemma 12. Take the corresponding $S$ in Lemma 12 for IDS, and let $T\triangleq KS$. For any $t\geq T$, we let $s=t/K\geq S$, and then by Lemma 12, we have $U^{s}_{\lfloor Ks\rfloor}=U_{t}^{t/K}$ is empty. ∎
D.4.1 Implication of sufficient exploration
Here we present some results that are implied by sufficient exploration (Proposition 7) and needed for proving the sufficient condition for optimality in fixed-confidence setting (Proposition 5). With sufficient exploration, the posterior means strongly converge to the unknown true means, and the probability of any sub-optimal arm being the best decays exponentially.
Lemma 13 (Lemmas 6 and 12 in Shang et al. 2020).
Under TTTS with any selection rule that ensures sufficient exploration (e.g., IDS), i.e.,
$\exists\tilde{T}\in\mathbb{M}\quad\forall t\geq\tilde{T}\quad\min_{i\in[K]}N_{t,i}\geq\sqrt{t/K},$
the following properties hold:
  1. 1.
    $\bm{\theta}_{t}\xrightarrow{\mathbb{M}}\bm{\theta}$; and
  2. 2.
    there exists $T\in\mathbb{M}$ such that for any $t\geq T$ and $j\neq I^{*}$,
    $\alpha_{t,j}\leq\exp\left(-ct^{1/2}\right),\quad\text{where}\quad c\triangleq\frac{\Delta_{\min}^{2}}{16\sigma^{2}K^{1/2}}.$
We remark that, while the original result focuses on $\beta$-tuning, the same proof applies to any selection rule that ensures sufficient exploration.
D.5 Strong convergence to optimal proportions: completing the proof of Proposition 5
In this subsection, we are going to complete the proof of Proposition 5, the sufficient condition for optimality for fixed-confidence setting, which requires a sequence of results.
Lemma 14.
Under TTTS-IDS, for any arm $j\neq I^{*}$,
$\frac{C_{t,I^{*},j}}{p_{t,I^{*}}\cdot f_{j}\left(\frac{p_{t,j}}{p_{t,I^{*}}}\right)}\xrightarrow{\mathbb{M}}1,\quad\text{where}\quad f_{j}(x)\triangleq\frac{\left(\theta_{I^{*}}-\theta_{j}\right)^{2}}{2\sigma^{2}\left(1+\frac{1}{x}\right)}\quad\text{and}\quad C_{t,I^{*},j}=\frac{\left(\theta_{t,I^{*}}-\theta_{t,j}\right)^{2}}{2\sigma^{2}\left(\frac{1}{p_{t,I^{*}}}+\frac{1}{p_{t,j}}\right)}.$
Proof.
Fix $j\neq I^{*}$. We have
$p_{t,I^{*}}f_{j}\left(\frac{p_{t,j}}{p_{t,I^{*}}}\right)=p_{t,I^{*}}\frac{\left(\theta_{I^{*}}-\theta_{j}\right)^{2}}{2\sigma^{2}\left(1+\frac{p_{t,I^{*}}}{p_{t,j}}\right)}=\frac{\left(\theta_{I^{*}}-\theta_{j}\right)^{2}}{2\sigma^{2}\left(\frac{1}{p_{t,I^{*}}}+\frac{1}{p_{t,j}}\right)}.$
By Proposition 7, TTTS-IDS ensures sufficient exploration, and thus by Lemma 13, $\bm{\theta}_{t}\xrightarrow{\mathbb{M}}\bm{\theta}$. Then applying the continuous mapping theorem (Lemma 7), we have
$\frac{C_{t,I^{*},j}}{p_{t,I^{*}}\cdot f_{j}\left(\frac{p_{t,j}}{p_{t,I^{*}}}\right)}\xrightarrow{\mathbb{M}}1.$
This completes the proof. ∎
Lemma 15.
Under TTTS-IDS, for any $\epsilon>0$, there exists a deterministic constant $c_{\epsilon}>0$ and a random time $T_{\epsilon}\in\mathbb{M}$ such that for any $t\geq T_{\epsilon}$ and $j\neq I^{*}$,
$\frac{w_{t,j}}{w_{t,I^{*}}}>\frac{p_{j}^{*}+\epsilon}{p_{I^{*}}^{*}}\quad\implies\quad\psi_{t,j}\leq\exp\left(-c_{\epsilon}t\right)+(K-1)\exp\left(-ct^{1/2}\right),$
where $c$ is defined in Lemma 13.
Proof.
Fix $\epsilon>0$. It suffices to show that for any $j\neq I^{*}$, there exist a deterministic constant $c^{(j)}_{\epsilon}>0$ and a random time $T^{(j)}_{\epsilon}\in\mathbb{M}$ such that for any $t\geq T^{(j)}_{\epsilon}$,
$\frac{w_{t,j}}{w_{t,I^{*}}}>\frac{p_{j}^{*}+\epsilon}{p_{I^{*}}^{*}}\quad\implies\quad\psi_{t,j}\leq\exp\left(-c_{\epsilon}^{(j)}t\right)+(K-1)\exp\left(-ct^{1/2}\right),$
since this completes the proof by taking $T_{\epsilon}\triangleq\max_{j\neq I^{*}}T^{(j)}_{\epsilon}$ and $c_{\epsilon}\triangleq\min_{j\neq I^{*}}c^{(j)}_{\epsilon}$.
From now on, we fix $j\neq I^{*}$. By Corollary 4, there exists $T_{\epsilon,1}^{(j)}\in\mathbb{M}$ such that for any $t\geq T_{\epsilon,1}^{(j)}$,
$\frac{w_{t,j}}{w_{t,I^{*}}}=\frac{\Psi_{t,j}}{\Psi_{t,I^{*}}}>\frac{p_{j}^{*}+\epsilon}{p_{I^{*}}^{*}}\quad\implies\quad\frac{p_{t,j}}{p_{t,I^{*}}}=\frac{N_{t,j}}{N_{t,I^{*}}}>\frac{p_{j}^{*}+\epsilon/2}{p_{I^{*}}^{*}}.$
Then by Corollary 5, there exists $T_{\epsilon,2}^{(j)}\in\mathbb{M}$ such that for any $t\geq T_{\epsilon,2}^{(j)}$,
$\frac{p_{t,j}}{p_{t,I^{*}}}>\frac{p_{j}^{*}+\epsilon/2}{p_{I^{*}}^{*}}\quad\implies\quad\exists A_{t}\neq I^{*}\,\,:\,\,\frac{p_{t,A_{t}}}{p_{t,I^{*}}}\leq\frac{p_{A_{t}}^{*}}{p_{I^{*}}^{*}}.$
From now on, we consider $t\geq\max\left\{T_{\epsilon,1}^{(j)},T_{\epsilon,2}^{(j)}\right\}$ and we have
$\frac{w_{t,j}}{w_{t,I^{*}}}=\frac{\Psi_{t,j}}{\Psi_{t,I^{*}}}>\frac{p_{j}^{*}+\epsilon}{p_{I^{*}}^{*}}\quad\implies\quad\frac{p_{t,j}}{p_{t,I^{*}}}=\frac{N_{t,j}}{N_{t,I^{*}}}>\frac{p_{j}^{*}+\epsilon/2}{p_{I^{*}}^{*}}\quad\text{and}\quad\exists A_{t}\neq I^{*}\,\,:\,\,\frac{p_{t,A_{t}}}{p_{t,I^{*}}}\leq\frac{p_{A_{t}}^{*}}{p_{I^{*}}^{*}}.$ (33)
By Lemmas 9 and 13,
$\psi_{t,j}\leq\alpha_{t,I^{*}}\frac{\alpha_{t,j}}{1-\alpha_{t,I^{*}}}\frac{p_{t,I^{*}}}{p_{t,I^{*}}+p_{t,j}}+(1-\alpha_{t,I^{*}})\leq\frac{\alpha_{t,j}}{1-\alpha_{t,I^{*}}}+(K-1)\exp\left(-ct^{1/2}\right).$
We can upper bound the numerator $\alpha_{t,j}$ as follows,
$\alpha_{t,j}\leq\mathbb{P}_{t}\left(\tilde{\theta}_{j}\geq\tilde{\theta}_{I^{*}}\right)$ $=\mathbb{P}\left(\frac{\tilde{\theta}_{I^{*}}-\tilde{\theta}_{j}-(\theta_{t,I^{*}}-\theta_{t,j})}{\sigma\sqrt{1/N_{t,I^{*}}+1/N_{t,j}}}\leq\frac{-(\theta_{t,I^{*}}-\theta_{t,j})}{\sigma\sqrt{1/N_{t,I^{*}}+1/N_{t,j}}}\right)$
$=\Phi\left(-\sqrt{2tC_{t,I^{*},j}}\right),$
where $\tilde{\bm{\theta}}=\left(\tilde{\theta}_{1},\ldots,\tilde{\theta}_{K}\right)$ is a sample drawn from the posterior distribution $\Pi_{t}$ and $\Phi(\cdot)$ is Gaussian cumulative distribution function. Similarly, we can lower bound the denominator $1-\alpha_{t,I^{*}}$ as follows,
$1-\alpha_{t,I^{*}}=\mathbb{P}_{t}\left(\exists a\neq I^{*}\,:\,\tilde{\theta}_{a}\geq\tilde{\theta}_{I^{*}}\right)\geq\mathbb{P}_{t}\left(\tilde{\theta}_{A_{t}}\geq\tilde{\theta}_{I^{*}}\right)=\Phi\left(-\sqrt{2tC_{t,I^{*},A_{t}}}\right).$
Hence,
$\psi_{t,j}\leq\frac{\Phi\left(-\sqrt{2tC_{t,I^{*},j}}\right)}{\Phi\left(-\sqrt{2tC_{t,I^{*},A_{t}}}\right)}+(K-1)\exp\left(-ct^{1/2}\right).$
The remaining task is to show that there exist a deterministic constant $c^{(j)}_{\epsilon}>0$ and a random time $T^{(j)}_{\epsilon,3}\in\mathbb{M}$ such that for any $t\geq T^{(j)}_{\epsilon,3}$,
$\frac{\Phi\left(-\sqrt{2tC_{t,I^{*},j}}\right)}{\Phi\left(-\sqrt{2tC_{t,I^{*},A_{t}}}\right)}\leq\exp\left(-c_{\epsilon}^{(j)}t\right).$ (34)
By Lemma 14, for any $\delta>0$,
$\frac{\Phi\left(-\sqrt{2tC_{t,I^{*},j}}\right)}{\Phi\left(-\sqrt{2tC_{t,I^{*},A_{t}}}\right)}$ $\leq\frac{\Phi\left(-\sqrt{2tp_{t,I^{*}}f_{j}\left(\frac{p_{t,j}}{p_{t,I^{*}}}\right)(1-\delta)}\right)}{\Phi\left(-\sqrt{2tp_{t,I^{*}}f_{A_{t}}\left(\frac{p_{t,A_{t}}}{p_{t,I^{*}}}\right)(1+\delta)}\right)}$
$\leq\frac{\Phi\left(-\sqrt{2tp_{t,I^{*}}f_{j}\left(\frac{p^{*}_{j}+\epsilon/2}{p^{*}_{I^{*}}}\right)(1-\delta)}\right)}{\Phi\left(-\sqrt{2tp_{t,I^{*}}f_{A_{t}}\left(\frac{p^{*}_{A_{t}}}{p^{*}_{I^{*}}}\right)(1+\delta)}\right)}$
$=\frac{\Phi\left(-\sqrt{2tp_{t,I^{*}}f_{j}\left(\frac{p^{*}_{j}+\epsilon/2}{p^{*}_{I^{*}}}\right)(1-\delta)}\right)}{\Phi\left(-\sqrt{2tp_{t,I^{*}}f_{j}\left(\frac{p^{*}_{j}}{p^{*}_{I^{*}}}\right)(1+\delta)}\right)},$
where the second inequality follows from (33) the monotonicity of $f_{j}$ and $f_{A_{t}}$ and the final equality uses the information balance (27).
We can pick a sufficiently small $\delta$ as a function of $\epsilon$ such that
$c_{1}\triangleq f_{j}\left(\frac{p^{*}_{j}+\epsilon/2}{p^{*}_{I^{*}}}\right)(1-\delta)>f_{j}\left(\frac{p^{*}_{j}}{p^{*}_{I^{*}}}\right)(1+\delta)\triangleq c_{2}.$
Then (34) follows from Gaussian tail upper and lower bounds (or $\frac{1}{t}\log\Phi\left(-\sqrt{t}x\right)\to-x^{2}/2$ as $t\to\infty$) and that $p_{t,I^{*}}$ is bounded away from 0 and 1 when $t$ is large (Lemma 10). This completes the proof. ∎
Lemma 16.
Under TTTS-IDS, for any $\epsilon>0$, there exists $T_{\epsilon}\in\mathbb{M}$ such that for any $t\geq T_{\epsilon}$,
$\frac{w_{t,j}}{w_{t,I^{*}}}\leq\frac{p_{j}^{*}+\epsilon}{p_{I^{*}}^{*}},\quad\forall j\neq I^{*}.$
Proof.
Fix $\epsilon>0$ and $j\neq I^{*}$. By Lemma 15, there exists $\tilde{T}_{\epsilon}\in\mathbb{M}$ such that for any $t\geq\tilde{T}_{\epsilon}$ and $j\neq I^{*}$,
$\frac{w_{t,j}}{w_{t,I^{*}}}>\frac{p_{j}^{*}+\epsilon/2}{p_{I^{*}}^{*}}\quad\implies\quad\psi_{t,j}\leq\exp\left(-c_{\epsilon}t\right)+(K-1)\exp\left(-ct^{1/2}\right).$
Define
$\kappa_{\epsilon}\triangleq\sum_{\ell=0}^{\infty}\left[\exp\left(-c_{\epsilon}\ell\right)+(K-1)\exp\left(-c\ell^{1/2}\right)\right]<\infty.$
From now on, we consider $t\geq\tilde{T}_{\epsilon}$. We need to consider two cases, and we provide an upper bound on $\frac{w_{t,j}}{w_{t,I^{*}}}$ for each case.
  1. 1.
    The first case supposes that
    $\forall\ell\in\left\{\tilde{T}_{\epsilon},\tilde{T}_{\epsilon}+1,\ldots,t-1\right\}:\quad\frac{w_{\ell,j}}{w_{\ell,I^{*}}}\geq\frac{p_{j}^{*}+\epsilon/2}{p^{*}_{I^{*}}}.$
    We have
    $\Psi_{t,j}$ $=\Psi_{\tilde{T}_{\epsilon},j}+\sum_{\ell=\tilde{T}_{\epsilon}}^{t-1}\psi_{\ell,j}$
    $=\Psi_{\tilde{T}_{\epsilon},j}+\sum_{\ell=\tilde{T}_{\epsilon}}^{t-1}\psi_{\ell,j}\mathds{1}\left(\frac{w_{\ell,j}}{w_{\ell,I^{*}}}\geq\frac{p_{j}^{*}+\epsilon/2}{p^{*}_{I^{*}}}\right)$
    $\leq\Psi_{\tilde{T}_{\epsilon},j}+\sum_{\ell=\tilde{T}_{\epsilon}}^{t-1}\left[\exp\left(-c_{\epsilon}\ell\right)+(K-1)\exp\left(-c\ell^{1/2}\right)\right]$
    $\leq\Psi_{\tilde{T}_{\epsilon},j}+\kappa_{\epsilon}.$
    Take $T\in\mathbb{M}$ in Lemma 10, so for any $t\geq T$, $\Psi_{t,I^{*}}\geq b_{1}t$, and thus
    $\frac{w_{t,j}}{w_{t,I^{*}}}=\frac{\Psi_{t,j}}{\Psi_{t,I^{*}}}\leq\frac{\Psi_{\tilde{T}_{\epsilon},j}+\kappa_{\epsilon}}{\Psi_{t,I^{*}}}\leq\frac{\tilde{T}_{\epsilon}+\kappa_{\epsilon}}{b_{1}t}.$ (35)
  2. 2.
    The alternative case supposes that
    $\exists\ell\in\left\{\tilde{T}_{\epsilon},\tilde{T}_{\epsilon}+1,\ldots,t-1\right\}:\quad\frac{w_{\ell,j}}{w_{\ell,I^{*}}}<\frac{p_{j}^{*}+\epsilon/2}{p^{*}_{I^{*}}}.$
    We define
    $L_{t}\triangleq\max\left\{\ell\in\left\{\tilde{T}_{\epsilon},\tilde{T}_{\epsilon}+1,\ldots,t-1\right\}\,:\,\frac{w_{\ell,j}}{w_{\ell,I^{*}}}<\frac{p_{j}^{*}+\epsilon/2}{p^{*}_{I^{*}}}\right\}.$
    We have
    $\Psi_{t,j}$ $=\Psi_{L_{t},j}+\psi_{L_{t},j}+\sum_{\ell=L_{t}+1}^{t-1}\psi_{\ell,j}$
    $=\Psi_{L_{t},j}+\psi_{L_{t},j}+\sum_{\ell=L_{t}+1}^{t-1}\psi_{\ell,j}\mathds{1}\left(\frac{w_{\ell,j}}{w_{\ell,I^{*}}}\geq\frac{p_{j}^{*}+\epsilon/2}{p^{*}_{I^{*}}}\right)$
    $\leq\frac{p_{j}^{*}+\epsilon/2}{p^{*}_{I^{*}}}\Psi_{L_{t},I^{*}}+1+\sum_{\ell=L_{t}+1}^{t-1}\left[\exp\left(-c_{\epsilon}\ell\right)+(K-1)\exp\left(-c\ell^{1/2}\right)\right]$
    $\leq\frac{p_{j}^{*}+\epsilon/2}{p^{*}_{I^{*}}}\Psi_{L_{t},I^{*}}+\left(\kappa_{\epsilon}+1\right),$
    where the first inequality uses $\frac{\Psi_{L_{t},j}}{\Psi_{L_{t},I^{*}}}=\frac{w_{L_{t},j}}{w_{L_{t},I^{*}}}<\frac{p_{j}^{*}+\epsilon/2}{p^{*}_{I^{*}}}$. Then by Lemma 10, for any $t\geq T$,
    $\frac{w_{t,j}}{w_{t,I^{*}}}\leq\frac{p_{j}^{*}+\epsilon/2}{p^{*}_{I^{*}}}\frac{\Psi_{L_{t},I^{*}}}{\Psi_{t,I^{*}}}+\frac{1+\kappa_{\epsilon}}{\Psi_{t,I^{*}}}\leq\frac{p_{j}^{*}+\epsilon/2}{p^{*}_{I^{*}}}+\frac{1+\kappa_{\epsilon}}{b_{1}t}.$ (36)
Putting (35) and (36) together, we find that for any $t\geq T_{1}$,
$\frac{w_{t,j}}{w_{t,I^{*}}}=\frac{\Psi_{t,j}}{\Psi_{t,I^{*}}}\leq\max\left\{\frac{\tilde{T}_{\epsilon}+\kappa_{\epsilon}}{b_{1}t}\,,\,\frac{p_{j}^{*}+\epsilon/2}{p^{*}_{I^{*}}}+\frac{\kappa_{\epsilon}+1}{b_{1}t}\right\}\leq\frac{p_{j}^{*}+\epsilon/2}{p^{*}_{I^{*}}}+\frac{\tilde{T}_{\epsilon}+2\kappa_{\epsilon}+1}{t}.$
Then we have
$t\geq\frac{2\left(\tilde{T}_{\epsilon}+2\kappa_{\epsilon}+1\right)}{\epsilon}\quad\implies\quad\frac{w_{t,j}}{w_{t,I^{*}}}\leq\frac{p_{j}^{*}+\epsilon}{p^{*}_{I^{*}}}.$
Taking $T_{\epsilon}\triangleq\max\left\{\tilde{T}_{\epsilon},T,\frac{2\left(\tilde{T}_{\epsilon}+2\kappa_{\epsilon}+1\right)}{\epsilon}\right\}$ completes the proof since Lemma 5 implies $T_{\epsilon}\in\mathbb{M}$. ∎
Now we are ready to complete the proof of Proposition 5.
Proof of Proposition 5.
Corollary 5 and Lemma 16 give Proposition 5. ∎
Appendix E Proof of Theorem 6
Define $\Delta_{\min}\triangleq\min_{i\in[K]}|\theta_{i}-\mathfrak{T}|$ and $\Delta_{\max}\triangleq\max_{i\in[K]}|\theta_{i}-\mathfrak{T}|$.
E.1 Sufficient exploration
Proposition 8.
There exists $T\in\mathbb{M}$ such that for any $t\geq T$,
$\min_{i\in[K]}N_{t,i}\geq\sqrt{t/K}.$
Following Qin et al. (2017); Shang et al. (2020), we define the following sets for any $t\in\mathbb{N}_{0}$ and $s\geq 0$:
$U_{t}^{s}\triangleq\{i\in[K]\,:\,N_{t,i}<s^{1/2}\}$
and
$V_{t}^{s}\triangleq\{i\in[K]\,:\,N_{t,i}<s^{3/4}\}.$
We let $\overline{U_{t}^{s}}=[K]\setminus U_{t}^{s}$ and $\overline{V_{t}^{s}}=[K]\setminus V_{t}^{s}$.
Lemma 17.
There exists $S\in\mathbb{M}$ such that for any $s\geq S$, if $U_{t}^{s}$ is nonempty, then $I_{t}\in V_{t}^{s}$.
Proof.
By Lemma 8, there exists $S_{1}\in\mathbb{M}$ such that for any $s\geq S_{1}$ and any $t\in\mathbb{N}_{0}$, if $i\in\overline{V_{t}^{s}}$, then $|\theta_{t,i}-\mathfrak{T}|\geq\Delta_{\min}/2$. On the other hand, if $j\in U_{t}^{s}$, then by the definition of $W_{1}$, $|\theta_{t,j}-\mathfrak{T}|\leq\Delta_{\max}+\sigma W_{1}$. Hence, there exists $S_{2}\in\mathbb{M}$ such that for $s\geq S_{1}$, $p_{t,j}d(\theta_{t,j},\mathfrak{T})<p_{t,i}d(\theta_{t,i},\mathfrak{T})$, which implies $I_{t}\notin\overline{V_{t}^{s}}$, i.e., $I_{t}\in V_{t}^{s}$. ∎
The proof of Proposition 8 follows from the above lemma and a variant of Lemma 12.
E.2 Strong convergence to optimal proportions
For notational convenience, let $w_{t,i}\triangleq\frac{\Psi_{t,i}}{t}$ for any $(t,i)\in\mathbb{N}_{0}\times[K]$. Now we define the over-sampled sets as follows: for any $t\in\mathbb{N}_{0}$ and $\epsilon>0$,
$O_{t,\epsilon}\triangleq\left\{j\in[K]\,:\,w_{t,i}>p_{i}^{*}+\epsilon\right\}.$
Lemma 18.
For any arm $i\in[K]$,
$\frac{C_{t,i}}{D_{t,i}}\xrightarrow{\mathbb{M}}1\quad\text{where}\quad D_{t,i}\triangleq p_{t,i}d(\theta_{i},\mathfrak{T})=\frac{p_{t,i}\left(\theta_{i}-\mathfrak{T}\right)^{2}}{2\sigma^{2}}.$
Lemma 19.
For any $\epsilon>0$, there exists a random time $T_{\epsilon}\in\mathbb{M}$ such that for any $t\geq T_{\epsilon}$ and $i\in[K]$,
$i\in O_{t,\epsilon}\quad\implies\quad\psi_{t,i}=0.$
Proof.
There exists $T_{1}\in\mathbb{M}$ such that for any $t\geq T_{1}$ and $i\in[K]$,
$i\in O_{t,\epsilon}\quad\implies\quad p_{t,i}>p_{i}^{*}+\epsilon/2\quad\text{and}\quad\exists J_{t}\neq I^{*}\,\,\text{s.t.}\,\,p_{t,J_{t}}\geq p_{J_{t}}.$
$\frac{C_{t,i}}{C_{t,J_{t}}}$ $=\frac{p_{t,i}\left(\theta_{t,i}-\mathfrak{T}\right)^{2}}{p_{t,J_{t}}\left(\theta_{t,J_{t}}-\mathfrak{T}\right)^{2}}$
$\geq\frac{p_{t,i}\left(\theta_{i}-\mathfrak{T}\right)^{2}(1-\delta)}{p_{t,J_{t}}\left(\theta_{J_{t}}-\mathfrak{T}\right)^{2}(1+\delta)}$
$\geq\frac{\left(p_{i}+\frac{\epsilon}{2}\right)\left(\theta_{i}-\mathfrak{T}\right)^{2}(1-\delta)}{p_{J_{t}}^{*}\left(\theta_{J_{t}}-\mathfrak{T}\right)^{2}(1+\delta)}$
$>1,$
where the first inequality uses the above lemma and picks a sufficient small $\delta$ as a function of $\epsilon$ such that the last inequality holds. ∎
Lemma 20.
For any $\epsilon>0$, there exists $T_{\epsilon}\in\mathbb{M}$ such that for any $t\geq T_{\epsilon}$,
$\forall i\in[K],\quad i\notin O_{t,\epsilon}\quad\text{i.e.}\quad w_{t,i}\leq p_{i}^{*}+\epsilon.$
Proof.
Since $K$ is finite, it suffices to consider a fixed $i\in[K]$. There exists $\tilde{T}_{\epsilon}\in\mathbb{M}$ such that for any $t\geq\tilde{T}_{\epsilon}$,
$i\in O_{t,\epsilon}\quad\implies\quad\psi_{t,i}=0.$
We need to consider two cases.
  1. 1.
    $\exists\ell\in\left\{\tilde{T}_{\epsilon},\tilde{T}_{\epsilon}+1,\ldots,t-1\right\},w_{\ell,i}\leq p_{i}^{*}+\epsilon$.
    Define
    $L_{t}\triangleq\max\left\{\ell\in\left\{\tilde{T}_{\epsilon},\tilde{T}_{\epsilon}+1,\ldots,t-1\right\}\,:\,w_{\ell,i}\leq p_{i}^{*}+\epsilon\right\}.$
    Then
    $w_{t,i}=\frac{\Psi_{t,i}}{t}=\frac{\Psi_{L_{t},i}}{t}\leq\frac{\Psi_{L_{t},i}}{L_{t}}=w_{L_{t},i}\leq p_{i}^{*}+\epsilon.$
  2. 2.
    $\forall\ell\in\left\{\tilde{T}_{\epsilon},\tilde{T}_{\epsilon}+1,\ldots,t-1\right\},w_{\ell,i}>p_{i}^{*}+\epsilon/2$.
    We have
    $w_{t,i}=\frac{\Psi_{t,i}}{t}=\frac{\Psi_{\tilde{T}_{\epsilon},i}}{t}.$
Letting $T_{\epsilon,i}\triangleq\max\left\{\tilde{T}_{\epsilon},\frac{\Psi_{\tilde{T}_{\epsilon},i}}{p_{i}^{*}+\epsilon}\right\}$ completes the proof. ∎
Appendix F Best-$k$ identification
In this section, we consider a natural extension of best-arm identification, where the objective is to find the exact set of arms with the $k$ highest means
$\mathcal{I}({\bm{\theta}})\triangleq\operatorname*{argmax}_{\mathcal{S}\subseteq[K],|\mathcal{S}|=k}\left\{\sum_{i\in\mathcal{S}}\theta_{i}\right\},$
where $k<K$ is a fixed integer. We refer to arms in $\mathcal{I}(\bm{\theta})$ as the top arms and arms in $\mathcal{I}^{c}(\bm{\theta})=[K]\backslash\mathcal{I}(\bm{\theta})$ as the bottom arms.
For the fixed-confidence best-$k$-arm identification problem with $k>1$, we encounter a more complex exploration query, where each pair of arms $(i,j)\in\mathcal{X}(\bm{\theta})=\mathcal{I}(\bm{\theta})\times\mathcal{I}^{c}(\bm{\theta})$ must be compared. Indeed, the alternative set can be decomposed as
$\mathrm{Alt}(\bm{\theta})=\bigcup_{(i,j)\in\mathcal{I}(\bm{\theta})\times\mathcal{I}^{c}(\bm{\theta})}\mathrm{Alt}_{i,j},\quad\hbox{where}\quad\mathrm{Alt}_{i,j}\triangleq\left\{\bm{\vartheta}:\vartheta_{j}>\vartheta_{i}\right\}.$
For each culprit $(i,j)\in\mathcal{X}(\bm{\theta})$, the function $C_{i,j}$ is given by
$C_{i,j}(\bm{p})$ $\triangleq\inf_{\bm{\vartheta}\in\mathrm{Alt}_{i,j}}\sum_{i}p_{i}d(\theta_{i},\vartheta_{i})$
$=p_{i}d(\theta_{i},\bar{\theta}_{i,j})+p_{j}d(\theta_{j},\bar{\theta}_{i,j}),\quad\hbox{where}\quad\bar{\theta}_{i,j}\triangleq\frac{p_{i}\theta_{i}+p_{j}\theta_{j}}{p_{i}+p_{j}},$ (37)
see Lemma 23 for detailed derivations.
The optimal allocation problem (9) is then given by
$\Gamma^{*}_{\bm{\theta}}=\max_{\bm{p}\in\mathcal{S}_{K}}\min_{(i,j)\in\mathcal{I}\times\mathcal{I}^{c}}C_{i,j}(\bm{p}).$ (38)
We present several properties of the optimal solution to (38). These properties generalize those for BAI in Garivier and Kaufmann (2016). We postpone the proof to Appendix F.4.1. To the best of our knowledge, we provide the first uniqueness result for best-$k$-arm identification problem.
Proposition 9 (Properties of the optimal solution).
The optimal solution $\bm{p}^{*}$ to (38) satisfies the following properties
  1. 1.
    Uniqueness: $\bm{p}^{*}$ is unique.
  2. 2.
    Positivity: $\bm{p}^{*}>\bm{0}$.
  3. 3.
    Monotonicity: for all $i,i^{\prime}\in\mathcal{I}$ we have $\theta_{i}\geq\theta_{i^{\prime}}\implies p^{*}_{i}\leq p^{*}_{i^{\prime}}$; and for all $j,j^{\prime}\in\mathcal{I}^{c}$ we have $\theta_{j}\geq\theta_{j^{\prime}}\implies p^{*}_{j}\geq p^{*}_{j^{\prime}}$.
We now revisit one of the most prominent approaches in solving BAI problems with fixed-confidence to optimality and demonstrate the challenges in generalizing them for the best-$k$-arm identification problem. We note that our ACC algorithm does not face these challenges thanks to the dual variables.
F.1 Optimality conditions that are dependent only on the allocation vector.
As stated in Theorem 1, the BAI problem admits a simple optimality condition that depends only on the allocation vector $\bm{p}$. Given the simplicity and popularity of the conditions in Theorem 1, it may be intriguing to consider a direct generalization of these conditions for $k>1$ and to hope that they remain sufficient, as studied in Gao and Chen (2015, Theorem 2) and Zhang et al. (2023, Theorem 3). Unfortunately, these conditions, albeit necessary, fail to remain sufficient. We first summarize the generalized conditions. For ease of presentation, we show results for Gaussian bandits.
Proposition 10.
For Gaussian bandits, the following conditions are necessary for the optimality of $\bm{p}^{*}$ for (38):
  1. 1.
    Crude information balance:
    $\Gamma_{\bm{\theta}}^{*}=\min_{j^{\prime}\in\mathcal{I}^{c}}C_{i,j^{\prime}}(\bm{p}^{*})\quad\forall i\in\mathcal{I}\quad\text{and}\quad\Gamma_{\bm{\theta}}^{*}=\min_{i^{\prime}\in\mathcal{I}}C_{i^{\prime},j}(\bm{p}^{*})\quad\forall j\in\mathcal{I}^{c}.$ (39)
  2. 2.
    Overall balance:
    $\sum_{i\in\mathcal{I}}(p^{*}_{i})^{2}=\sum_{j\in\mathcal{I}^{c}}(p^{*}_{j})^{2}.$ (40)
The information balance suggests that the optimal allocation should balance the statistical evidence gathered to distinguish any top arm $i$ from its corresponding most confusing bottom arm(s)
$j_{i}\in\operatorname*{argmin}_{j^{\prime}\in\mathcal{I}^{c}}C_{i,j^{\prime}}(\bm{p}^{*}),$
and similarly, any bottom arm from its corresponding most confusing top arm(s). However, it is a priori unclear which pairs of $(i,j)$ are the most confusing under the unknown optimal allocation.
Indeed, these simplistic balancing conditions fail to be sufficient when $k>1$. We present a counterexample using Gaussian bandits in Example 4. This example implies that the crude information balance in Proposition 10 cannot guarantee optimality, even when combined with the correct overall balance.
Example 4 (Failure of sufficiency).
We point out that conditions in Proposition 10 do not guarantee the uniqueness of the solution nor the optimality. We present a counterexample in a Gaussian bandit with $(k,K)=(2,4)$. Consider Gaussian rewards with mean $\bm{\theta}=(1,0.7,0,-0.5)$ and variance $\sigma^{2}=1/4$. The solution $\bm{p}_{1}\approx(0.0482,0.459,0.4603,0.0325)$ satisfies Proposition 10 with an objective value of $\Gamma_{\bm{\theta}}(\bm{p}_{1})\approx 0.0873$. For $\bm{p}_{1}$, we have $C_{1,3}(\bm{p}_{1})=C_{1,4}(\bm{p}_{1})=C_{2,4}(\bm{p}_{1})<C_{2,3}(\bm{p}_{1})$. However, the optimal solution is $\bm{p}^{*}\approx(0.1277,0.3894,0.4016,0.0813)$ with an optimal value of $\Gamma^{*}_{\bm{\theta}}\approx 0.1938>\Gamma_{\bm{\theta}}(\bm{p}_{1})$. For $\bm{p}^{*}$, we have $C_{1,3}(\bm{p}^{*})=C_{2,4}(\bm{p}^{*})=C_{2,3}(\bm{p}^{*})<C_{1,4}(\bm{p}^{*})$.
Our analysis shows that to obtain a sufficient condition, the overall balance property (40) needs to be decomposed into finer balancing conditions among subgroups of arms (Propositions 11 and 14); see Example 5 for an illustration. This result indicates that algorithms that rely only on overall balance is highly likely to fail.
Example 5 (Overall balance in subgroups).
Consider Gaussian bandits with mean rewards
$\bm{\theta}=\{0.51,0.5,0,-0.01,-0.092\},$
variance $\sigma^{2}=1/4$ and $k=2$. The optimal allocation is
$\bm{p}^{*}\approx(0.21851674,0.23713642,0.21851674,0.20260818,0.12322192).$
We have $\Gamma^{*}_{\bm{\theta}}=C_{1,3}(\bm{p}^{*})=C_{2,4}(\bm{p}^{*})=C_{2,5}(\bm{p}^{*})=0.0568362041$. Note that the overall balance can be further decomposed into two subgroups: $\{1,2\}$ and $\{3,4,5\}$, i.e, $(p^{*}_{1})^{2}=(p^{*}_{3})^{2}$ and $(p^{*}_{2})^{2}=(p^{*}_{4})^{2}+(p^{*}_{5})^{2}.$
We now formalize the observations in Example 5. Recall from Proposition 10 that there are multiple $(i,j)$ pairs such that $C_{i,j}(\bm{p}^{*})=\Gamma^{*}_{\bm{\theta}}$. To facilitate discussion of the solution structure, we require the following definition.
Definition 7 (Bipartite graph induced by the optimal solution).
The bipartite graph induced by the optimal solution $\bm{p}^{*}$ is denoted by $\mathcal{G}(\mathcal{I},\mathcal{I}^{c},\mathcal{E})$ where the set of edges $\mathcal{E}$ contains any $(i,j)\in\mathcal{I}\times\mathcal{I}^{c}$ such that $C_{i,j}(\bm{p}^{*})=\Gamma^{*}_{\bm{\theta}}$. Following from the uniqueness of the optimal solution, the bipartite graph is also unique.
Remark 11 (Connected subgraphs).
By (39), for any arm $i\in\mathcal{I}\cup\mathcal{I}^{c}$, there must exist an edge $e\in\mathcal{E}$ such that $i$ is incident to $e$. Hence, the bipartite graph $\mathcal{G}(\mathcal{I},\mathcal{I}^{c},\mathcal{E})$ can be uniquely decomposed into a collection of $L$ connected subgraphs $\left\{\mathcal{G}_{\ell}(\mathcal{I}_{\ell},\mathcal{I}^{c}_{\ell},\mathcal{E}_{\ell})\right\}_{\ell\in[L]}$ such that
  1. 1.
    for any $\ell\in[L]$, $\mathcal{I}_{\ell},\mathcal{I}^{c}_{\ell},\mathcal{E}_{\ell}\neq\varnothing$;
  2. 2.
    {}$\{\mathcal{I}_{\ell}\}_{\ell\in[L]}$ are mutually exclusive with $\bigcup_{\ell\in[L]}\mathcal{I}_{\ell}=\mathcal{I}$; $\{\mathcal{I}^{c}_{\ell}\}_{\ell\in[L]}$ are mutually exclusive with $\bigcup_{\ell\in[L]}\mathcal{I}^{c}_{\ell}=\mathcal{I}^{c}$; {}$\{\mathcal{E}_{\ell}\}_{\ell\in[L]}$ are mutually exclusive with $\bigcup_{\ell\in[L]}\mathcal{E}_{\ell}=\mathcal{E}$.
Example 5, provide an illustration of the connected subgraphs.
Now, we present a necessary condition for Gaussian bandits. In Appendix F.4.3, we collect the proof and provide an extension of Proposition 11 to the general exponential family; see Proposition 14.
Proposition 11 (Overall balance in connected subgraphs – Gaussian bandits).
For Gaussian bandits, each connected subgraph of the bipartite graph induced by the optimal solution balances the sum-of-squared allocations, i.e.,
$\sum_{i\in\mathcal{I}_{\ell}}(p^{*}_{i})^{2}=\sum_{j\in\mathcal{I}^{c}_{\ell}}(p^{*}_{j})^{2},\quad\forall\ell\in[L].$ (41)
We remark that (41) is strictly stronger than (40), as Example 5 shows.
Unfortunately, the optimal decomposition of subgraphs relies on the true $\bm{\theta}$, which is unavailable. Due to the combinatorial number of scenarios, we caution that these conditions are too complicated to yield any meaningful algorithm. Nevertheless, we present a complete case study with $(k,K)=(2,5)$ in Appendix F.3 to illustrate the complexity of the information balance and overall balance properties.
F.2 A (conditional) sufficient condition
To begin, we present a sufficient condition for a large set of problem instances $\bm{\theta}$. For the ease of notation, we re-label the arm with the $\ell$-th largest mean as arm $i_{\ell}$, for $\ell=1,\dots,k$ and re-label the arm with the $(\ell+k)$-th largest mean as arm $j_{\ell}$, for $\ell=1,\dots,K-k$. It may be intriguing to think that the optimal solution should gather equal statistical evidence to distinguish (i) the arm $i_{k}$ from any bottom arm $j$; and (ii) the arm $j_{1}$ from any top arm $i$. The next proposition show that this intuitive balancing rule (together with the balance of sum-of-squared allocation) is optimal under certain condition. To simplify notation, we define $\bar{\theta}^{*}_{i,j}\triangleq\bar{\theta}_{i,j}(\bm{p}^{*})$ for all $(i,j)\in\mathcal{I}\times\mathcal{I}^{c}$.
Proposition 12.
Consider any problem instance such that the optimal solution $\bm{p}^{*}$ satisfies
$01-\sum_{j\in\mathcal{I}^{c}\backslash\{j_{1}\}}\frac{d(\theta_{i_{k}},\bar{\theta}^{*}_{i_{k},j})}{d(\theta_{j},\bar{\theta}^{*}_{i_{k},j})}>$ 0 . (42)
Then $\bm{p}^{*}$ is the optimal solution to (38) if
$\Gamma^{*}_{\bm{\theta}}=C_{i_{k},j}(\bm{p}^{*})=C_{i,j_{1}}(\bm{p}^{*}),\quad\forall i\in\mathcal{I},\forall j\in\mathcal{I}^{c},\quad\text{and}$ (43)
$\frac{d(\theta_{i_{k}},\bar{\theta}^{*}_{i_{k},j_{1}})}{d(\theta_{j_{1}},\bar{\theta}^{*}_{i_{k},j_{1}})}\sum_{i\in\mathcal{I}}\frac{d(\theta_{j_{1}},\bar{\theta}^{*}_{i,j_{1}})}{d(\theta_{i},\bar{\theta}^{*}_{i,j_{1}})}=\sum_{j\in\mathcal{I}^{c}}\frac{d(\theta_{i_{k}},\bar{\theta}^{*}_{i_{k},j})}{d(\theta_{j},\bar{\theta}^{*}_{i_{k},j})}.$ (44)
Complete characterization of necessary and sufficient conditions based only on $\bm{p}$ can be derived following the lines of Proposition 12, see Appendix F.3. However, we caution that such results rely on assumptions similar to (42), which is non-trivial to check. Hence, developing a sampling rule based on tracking the sufficient condition can be challenging. We address this challenge by introducing the dual variable, and analyze the KKT conditions.
Remark 12.
For general bandits, condition (42) always holds when (i) $|\mathcal{I}|\leq 2$ and $|\mathcal{I}^{c}|\leq 2$; and (ii) $|\mathcal{I}|=1$ or $|\mathcal{I}^{c}|=1$. When condition (42) holds, there is only one connected subgraph. For Gaussian bandits with equal variances, the assumption (42) reduces to $(p^{*}_{i_{k}})^{2}>\sum_{j\in\mathcal{I}^{c}\backslash\{j_{1}\}}(p^{*}_{j})^{2}$; and (44) reduces to the balancing of the sum-of-squared allocations. Recall the monotonicity in Proposition 9, we have $p_{j_{1}}\geq p_{j}$ for $j\neq j_{1}$. In general, $p^{*}_{j_{1}}-p^{*}_{j}$ increases as the gap $(\theta_{j_{1}}-\theta_{j})$ increases. Roughly speaking, when $(\theta_{j_{1}}-\theta_{j})$ is large enough with respect to $(\theta_{i_{k}}-\theta_{j_{1}})$, condition (42) will hold. This suggest that problem instances closer to the slippage configuration are the most subtle instances in terms of the structure of the optimal solution.
F.3 A Case study for 5 choose 2: A pathway to complete characterization of optimality based on the allocation vector
In Proposition 12, we characterized the sufficient conditions for optimality under the assumption of (42). We now discuss how to obtain complete characterization.
We assume that the mean rewards are mutually different without loss of generality (for the optimal allocation problem). Otherwise, we can combine arms with the same mean rewards and evenly distribute the allocation among them.
To facilitate our discussion, we present $C_{i,j}(\bm{p}^{*})$ as a matrix in ||×$\mathbb{R}^{|\mathcal{I}|\times|\mathcal{I}^{c}|}$, where the $(i,j)$-th entry is “$=$” if $C_{i,j}(\bm{p}^{*})=\Gamma^{*}_{\bm{\theta}}$ and “$>$” if $C_{i,j}(\bm{p}^{*})>\Gamma^{*}_{\bm{\theta}}$. The following matrix represents the case discussed in Proposition 12.
arms $j_{1}$ $j_{2}$cdots $j_{K-k}$
$i_{1}$ $=$ $>$ $\cdots$ $>$
$i_{2}$ $=$ $>$ $\cdots$ $>$
$\vdots$ $\vdots$ $\vdots$ $\vdots$ $\vdots$
$i_{k-1}$ $=$ $>$ $\cdots$ $>$
$i_{k}$ $=$ $=$ $\cdots$ $=$
Recall the assumption in Proposition 12, i.e.,
$01-\sum_{j\in\mathcal{I}^{c}\backslash\{j_{1}\}}\frac{d(\theta_{i_{k}},\bar{\theta}^{*}_{i_{k},j})}{d(\theta_{j},\bar{\theta}^{*}_{i_{k},j})}>$ 0 .
In general, the $=$’s locations can be quite different from the clean structure in the matrix above.
We illustrate with an example and discuss the sufficient conditions for it. Consider choosing the best-$2$ arms among $5$ competing arms. Without loss of generality, we assume that the arms are sorted in descending order of the mean rewards, i.e., =$\mathcal{I}=\{1,2\}$ and $\mathcal{I}^{c}=\{3,4,5\}$. We may have the following scenario.
arms $3$ $4$ $5$
$1$ $=$ $=$ $>$
$2$ $>$ $=$ $=$
Mirroring the proof of Proposition 12, we have the following sufficient conditions for this case.
Proposition 13.
Consider any problem instance such that the optimal solution $\bm{p}^{*}$ satisfies
$\begin{split}&1-\frac{d(\theta_{2},\bar{\theta}^{*}_{2,5})}{d(\theta_{5},\bar{\theta}^{*}_{2,5})}>0,\\ &1-\frac{d(\theta_{4},\bar{\theta}^{*}_{2,4})}{d(\theta_{2},\bar{\theta}^{*}_{2,4})}+\frac{d(\theta_{4},\bar{\theta}^{*}_{2,4})}{d(\theta_{2},\bar{\theta}^{*}_{2,4})}\frac{d(\theta_{2},\bar{\theta}^{*}_{2,5})}{d(\theta_{5},\bar{\theta}^{*}_{2,5})}>0.\end{split}$ (45)
Then $\bm{p}^{*}$ is the optimal solution to (38) if
$\Gamma^{*}_{\bm{\theta}}=C_{1,3}(\bm{p}^{*})=C_{1,4}(\bm{p}^{*})=C_{2,4}(\bm{p}^{*})=C_{2,5}(\bm{p}^{*})$ (46)
and
$1+\frac{d(\theta_{1},\bar{\theta}^{*}_{1,4})}{d(\theta_{4},\bar{\theta}^{*}_{1,4})}\frac{d(\theta_{4},\bar{\theta}^{*}_{2,4})}{d(\theta_{2},\bar{\theta}^{*}_{2,4})}=\frac{d(\theta_{1},\bar{\theta}^{*}_{1,3})}{d(\theta_{3},\bar{\theta}^{*}_{1,3})}+\frac{d(\theta_{1},\bar{\theta}^{*}_{1,4})}{d(\theta_{4},\bar{\theta}^{*}_{1,4})}+\frac{d(\theta_{1},\bar{\theta}^{*}_{1,4})}{d(\theta_{4},\bar{\theta}^{*}_{1,4})}\frac{d(\theta_{4},\bar{\theta}^{*}_{2,4})}{d(\theta_{2},\bar{\theta}^{*}_{2,4})}\frac{d(\theta_{2},\bar{\theta}^{*}_{2,5})}{d(\theta_{5},\bar{\theta}^{*}_{2,5})}.$ (47)
Remark 13 (Complete characterization of the case with $(K,k)=(5,2)$).
We consider Gaussian bandit with $(K,k)=(5,2)$. For simplicity, we consider the problem instances where the mean rewards are mutually different. Recall that we have the balancing of top and bottom arms in the entire bipartite graph in (40), i.e.,
$(p^{*}_{1})^{2}+(p^{*}_{2})^{2}=(p^{*}_{3})^{2}+(p^{*}_{4})^{2}+(p^{*}_{5})^{2}.$
By the monotonicity in Proposition 9, we must have
$(p^{*}_{2})^{2}>(p^{*}_{5})^{2},$
otherwise, $(p^{*}_{3})^{2}\geq(p^{*}_{4})^{2}\geq(p^{*}_{5})^{2}\geq(p^{*}_{2})^{2}\geq(p^{*}_{1})^{2}$, which contradicts with (40).
Note that the condition (45) is equivalent to
$(p^{*}_{2})^{2}>(p^{*}_{5})^{2}\quad\text{and}\quad(p^{*}_{2})^{2}<(p^{*}_{4})^{2}+(p^{*}_{5})^{2}.$
Since we must have $(p^{*}_{2})^{2}>(p^{*}_{5})^{2}$, condition (45) reduces to
$(p^{*}_{2})^{2}<(p^{*}_{4})^{2}+(p^{*}_{5})^{2}.$
Recall that condition (42) reduces to
$(p^{*}_{2})^{2}>(p^{*}_{4})^{2}+(p^{*}_{5})^{2}.$
Hence the two scenarios are mutually exclusive. Finally, recall that Example 5 corresponds to
$\Gamma^{*}_{\bm{\theta}}=C_{1,3}(p^{*}_{1},p^{*}_{3})=C_{2,4}(p^{*}_{2},p^{*}_{4})=C_{2,5}(p^{*}_{2},p^{*}_{5})$
and
$(p^{*}_{2})^{2}=(p^{*}_{4})^{2}+(p^{*}_{5})^{2}\quad\text{and}\quad(p^{*}_{1})^{2}=(p^{*}_{3})^{2},$ (48)
which corresponds to
arms $3$ $4$ $5$
$1$ $=$ $>$ $>$
$2$ $>$ $=$ $=$
Combining (42), (45) and (48), we have now fully characterized the sufficient conditions for the case of choosing $2$ arms among $5$, which we summarize in Table 6.
Condition $(p^{*}_{2})^{2}>(p^{*}_{4})^{2}+(p^{*}_{5})^{2}$ $(p^{*}_{2})^{2}<(p^{*}_{4})^{2}+(p^{*}_{5})^{2}$ $(p^{*}_{2})^{2}=(p^{*}_{4})^{2}+(p^{*}_{5})^{2}$
Information balance
arms $3$ $4$ $5$
$1$ $=$ $>$ $>$
$2$ $=$ $=$ $=$
arms $3$ $4$ $5$
$1$ $=$ $=$ $>$
$2$ $>$ $=$ $=$
arms $3$ $4$ $5$
$1$ $=$ $>$ $>$
$2$ $>$ $=$ $=$
Overall balance $\sum_{i\in\mathcal{I}}(p^{*}_{i})^{2}=\sum_{j\in\mathcal{I}^{c}}(p^{*}_{j})^{2}$ $\sum_{i\in\mathcal{I}}(p^{*}_{i})^{2}=\sum_{j\in\mathcal{I}^{c}}(p^{*}_{j})^{2}$ $(p^{*}_{1})^{2}=(p^{*}_{3})^{2}$
$(p^{*}_{2})^{2}=(p^{*}_{4})^{2}+(p^{*}_{5})^{2}$
Table 6: Complete characterization of the optimal solution for Gaussian bandits, choose 2 from 5 arms, with mutually distinctive mean rewards.
We omit the discussion for general $(K,k)$ pair and instances with arms sharing the same mean. A similar analysis can be carried out. However, we are cautious that any elegant presentation can be derived due to the combinatorial number of scenarios, even though conditions such as (42) and (45) can significantly reduce the number of scenarios to be discussed.
F.4 Proofs
F.4.1 Proof of Proposition 9
Proof of Proposition 9: Uniqueness.
By Lemma 3, $C_{i,j}(\cdot)$ is a concave function. Hence the optimal allocation problem (38) is a concave program. Consequently, the optimal set is convex, see Boyd and Vandenberghe (2004, Chapter 2).
We prove by contradiction. Suppose there exist two optimal solutions $\bm{p}^{\prime}\neq\bm{p}^{\prime\prime}$, then
$\bm{p}(\varepsilon)=(1-\varepsilon)\bm{p}^{\prime}+\varepsilon\bm{p}^{\prime\prime}$
is also an optimal solution for any $\varepsilon\in(0,1)$. To facilitate discussion, we define the index set where $\bm{p}^{\prime}$ differs from $\bm{p}^{\prime\prime}$
$\mathcal{D}=\{i\in[K]:p_{i}^{\prime}\neq p_{i}^{\prime\prime}\}\neq\varnothing.$
Now, consider the optimal solution $\bm{p}^{\prime}$. By the information balance (39), there exist a set ×$\mathcal{B}\subset\mathcal{I}\times\mathcal{I}^{c}$ such that
$C_{i,j}(\bm{p}^{\prime})=\Gamma_{\bm{\theta}}(\bm{p}^{\prime}),\quad\forall(i,j)\in\mathcal{B},\quad\text{and}\quad C_{i^{\prime},j^{\prime}}(\bm{p}^{\prime})>\Gamma_{\bm{\theta}}(\bm{p}^{\prime}),\quad\forall(i^{\prime},j^{\prime})\notin\mathcal{B}.$
For any collection $\mathcal{C}$ of $(i,j)$ pairs, we define
$\mathcal{C}_{\cap\mathcal{D}}\triangleq\{(i,j)\in\mathcal{C}:\{i,j\}\cap\mathcal{D}\neq\varnothing\}.$
  1. 1.
    For any $(i,j)\in(\mathcal{B}\backslash\mathcal{B}_{\cap\mathcal{D}})$, we have
    $C_{i,j}(\bm{p}(\varepsilon))=C_{i,j}(\bm{p}^{\prime})=\Gamma_{\bm{\theta}}(\bm{p}^{\prime}).$ (49)
  2. 2.
    Next, consider any $(i,j)\in\mathcal{B}^{c}$. Since $|\mathcal{B}^{c}|<\infty$, we can define
    $\Delta^{\Gamma_{\bm{\theta}}(\bm{p}^{\prime})}_{\text{min}}\triangleq\min_{(i,j)\notin\mathcal{B}}C_{i,j}(\bm{p}^{\prime})-\Gamma_{\bm{\theta}}(\bm{p}^{\prime})>0.$
    By the continuity of $C_{i,j}(\cdot)$, there exists a sufficiently small $\varepsilon$ such that
    $\min_{(i,j)\notin\mathcal{B}}C_{i,j}(\bm{p}(\varepsilon))-\Gamma_{\bm{\theta}}(\bm{p}^{\prime})>\Delta^{\Gamma_{\bm{\theta}}(\bm{p}^{\prime})}_{\text{min}}/2>0,$
    then
    $\min_{(i,j)\notin\mathcal{B}}C_{i,j}(\bm{p}(\varepsilon))>\Gamma_{\bm{\theta}}(\bm{p}^{\prime})=\Gamma_{\bm{\theta}}(\bm{p}(\varepsilon)),$ (50)
    where the last equality follows from the fact that both $\bm{p}^{\prime}$ and $\bm{p}(\varepsilon)$ are optimal solutions.
  3. 3.
    Finally, consider any $(i,j)\in\mathcal{B}_{\cap\mathcal{D}}$. Invoking Lemma 24, we conclude that $\varepsilon$ can be selected such that
    $C_{i,j}(\bm{p}(\varepsilon))\neq C_{i,j}(\bm{p}^{\prime}),\quad\forall(i,j)\in\mathcal{B}_{\cap\mathcal{D}}.$ (51)
We have the following contradictions.
  1. 1.
    Suppose that there exists some $(i_{1},j_{1})\in\mathcal{B}_{\cap\mathcal{D}}$ such that
    $C_{i_{1},j_{1}}(\bm{p}(\varepsilon))<C_{i_{1},j_{1}}(\bm{p}^{\prime})=\Gamma_{\bm{\theta}}(\bm{p}^{\prime}).$
    Then
    $\Gamma_{\bm{\theta}}(\bm{p}(\varepsilon))\leq C_{i_{1},j_{1}}(\bm{p}(\varepsilon))<\Gamma_{\bm{\theta}}(\bm{p}^{\prime}).$
    This contradicts with the fact that $\bm{p}(\varepsilon)$ is an optimal solution.
  2. 2.
    Suppose that for all $(i,j)\in\mathcal{B}_{\cap\mathcal{D}}$, we have
    $C_{i_{1},j_{1}}(\bm{p}(\varepsilon))>C_{i_{1},j_{1}}(\bm{p}^{\prime})=\Gamma_{\bm{\theta}}(\bm{p}^{\prime}).$
    Recall that $\mathcal{D}\neq\varnothing$, then there exist $i\in\mathcal{D}$ or $j\in\mathcal{D}$. Suppose $i_{2}\in\mathcal{I}\cap\mathcal{D}\neq\varnothing$. Combining (49),(50) and (51), we have
    $\min_{j\in\mathcal{I}^{c}}C_{i_{2},j}(\bm{p}(\varepsilon))>\Gamma_{\bm{\theta}}(\bm{p}^{\prime}).$
    Invoking the information balance in Proposition 10 we conclude that
    $\Gamma_{\bm{\theta}}(\bm{p}(\varepsilon))>\Gamma_{\bm{\theta}}(\bm{p}^{\prime}).$
    This contradicts with the fact that $\bm{p}^{\prime}$ is an optimal solution. Suppose $j_{2}\in\mathcal{I}^{c}\cap\mathcal{D}\neq\varnothing$, the proof follows similarly.
In summary, we have uniqueness of the optimal solution to (38). ∎
Proof of Proposition 9: Positivity.
Consider any fixed $i\in\mathcal{I}$ and any feasible solution $\bm{p}$. If $p_{i}=0$ for some $i\in\mathcal{I}$, then $C_{i,j}(\bm{p})=p_{i}d(\theta_{i},\bar{\theta}_{i,j})+p_{j}d\left(\theta_{j},\frac{p_{i}\theta_{i}+p_{j}\theta_{j}}{p_{i}+p_{j}}\right)=0$ holds for all $j\in\mathcal{I}^{c}$. Similarly, if $p_{j}=0$ for some $j\in\mathcal{I}^{c}$, then $C_{i,j}=0$ holds for all $i\in\mathcal{I}$. However, there exists $\bm{p}=(1/K,1/K,\dots,1/K)$ such that $C_{i,j}=\frac{1}{K}d\left(\theta_{i},\frac{\theta_{i}+\theta_{j}}{2}\right)+\frac{1}{K}d\left(\theta_{j},\frac{\theta_{i}+\theta_{j}}{2}\right)>0$ holds for all $i\in\mathcal{I}$ and $j\in\mathcal{I}^{c}$. Thus, the optimal solution cannot have $p_{i}=0$ for any $i\in[K]$. ∎
Proof of Proposition 9: Monotonicity.
For the ease of exposition, we introduce the following function
$g_{i,j}(x)=d\left(\theta_{i},\frac{\theta_{i}+x\theta_{j}}{1+x}\right)+xd\left(\theta_{j},\frac{\theta_{i}+x\theta_{j}}{1+x}\right),\quad i\in\mathcal{I},j\in\mathcal{I}^{c}.$
Then $C_{i,j}(\bm{p})$ can be written as
$C_{i,j}(\bm{p})=p_{i}g_{i,j}\left(\frac{p_{j}}{p_{i}}\right).$
At the optimal solution $\bm{p}^{*}$, consider any $j$, $j^{\prime}\in\mathcal{I}^{c}$ such that $\theta_{j^{\prime}}\geq\theta_{j}$. Proposition 10 implies that there exist a $i\in\mathcal{I}$ such that
$C_{i,j}(\bm{p}^{*})=\Gamma^{*}_{\bm{\theta}}\leq C_{i,j^{\prime}}(\bm{p}^{*}).$
Note that we must have $p^{*}_{i}>0$, the inequality above implies that
$g_{i,j}\left(\frac{p^{*}_{j}}{p^{*}_{i}}\right)\leq g_{i,j^{\prime}}\left(\frac{p^{*}_{j^{\prime}}}{p^{*}_{i}}\right).$ (52)
By Lemma 25, for any $i,j$ and $j^{\prime}$ such that $\theta_{i}>\theta_{j^{\prime}}\geq\theta_{j}$, we have
$g_{i,j}(x)\geq g_{i,j^{\prime}}(x)\quad\text{for all }x.$ (53)
Plugging $x=\frac{p^{*}_{j}}{p^{*}_{i}}$ into (53), we have
$g_{i,j}\left(\frac{p^{*}_{j}}{p^{*}_{i}}\right)\geq g_{i,j^{\prime}}\left(\frac{p^{*}_{j}}{p^{*}_{i}}\right).$ (54)
Combining (52), (54) and $g_{i,j^{\prime}}(x)>0$, we conclude that $p_{j^{\prime}}^{*}\geq p_{j}^{*}$. This shows the monotonicity of $\bm{p}^{*}$ within the bottom arms $\mathcal{I}^{c}$.
For the top arms $\mathcal{I}$, we introduce
$g_{j,i}(x)=d\left(\theta_{j},\frac{\theta_{j}+x\theta_{i}}{1+x}\right)+xd\left(\theta_{i},\frac{\theta_{j}+x\theta_{i}}{1+x}\right).$
Then $C_{i,j}$ can be written as
$C_{i,j}(\bm{p})=p_{j}g_{j,i}\left(\frac{p_{i}}{p_{j}}\right).$
From the proof of Lemma 25 (See (77) and replace $\theta_{i}$ with $\theta_{j}$ in $h$), we have
$g_{j,i}^{\prime}(x)=d\left(\theta_{i},\frac{\theta_{j}+x\theta_{i}}{1+x}\right)>0$
and, for any $\theta_{i}\geq\theta_{i^{\prime}}>\theta_{j}$, we have
$g_{j,i}(x)\geq g_{j,i^{\prime}}(x)\quad\text{for all }x.$ (55)
The rest of the proof mirrors the proof of the monotonicity within the bottom arms. At the optimal solution $\bm{p}^{*}$, consider any $i$, $i^{\prime}\in\mathcal{I}$ such that $\theta_{i}\geq\theta_{i^{\prime}}$. Proposition 10 also implies that there exists a $j\in\mathcal{I}^{c}$ such that
$C_{i,j}(\bm{p}^{*})=\Gamma^{*}_{\bm{\theta}}\leq C_{i^{\prime},j}(\bm{p}^{*}).$
Note that we must have $p^{*}_{j}>0$, the inequality above implies that
$g_{j,i}\left(\frac{p^{*}_{i}}{p^{*}_{j}}\right)\leq g_{j,i^{\prime}}\left(\frac{p^{*}_{i^{\prime}}}{p^{*}_{j}}\right).$ (56)
Plugging $x=\frac{p^{*}_{i}}{p^{*}_{j}}$ into (55), we have,
$g_{j,i}\left(\frac{p^{*}_{i}}{p^{*}_{j}}\right)\geq g_{j,i^{\prime}}\left(\frac{p^{*}_{i}}{p^{*}_{j}}\right).$ (57)
Combining (56), (57) and $g_{j,i^{\prime}}(x)>0$, we conclude that $p_{i^{\prime}}^{*}\geq p_{i}^{*}$. This shows the monotonicity of $\bm{p}^{*}$ within the top arms $\mathcal{I}$. The proof is complete. ∎
F.4.2 Proof of Proposition 10
Overall balance.
The overall balance in Proposition 10 is a direct corollary of Proposition 14 presented in Appendix F.4.3.
Crude information balance.
Proof of (39).
Note that for any $(i,j)\in\mathcal{I}\times\mathcal{I}^{c}$, $C_{i,j}(\cdot)$ depends on $\bm{p}$ only through $p_{i},p_{j}$. We write $C_{i,j}(p_{i},p_{j})=C_{i,j}(\bm{p})$ with a slight abuse of notation.
We first prove the second statement of (39), i.e.,
$\Gamma_{\bm{\theta}}^{*}=\min_{i^{\prime}\in\mathcal{I}}C_{i^{\prime},j}(\bm{p}^{*}),\quad\forall j\in\mathcal{I}^{c}.$ (58)
We prove by contradiction. Suppose that (58) does not hold. Let $\mathcal{I}^{c}_{\text{min}}\subset\mathcal{I}^{c}$ be the set of arms $j\in\mathcal{I}^{c}$ such that $\min_{i^{\prime}\in\mathcal{I}}C_{i^{\prime},j}(\bm{p}^{*})=\Gamma^{*}_{\bm{\theta}}$. Then $\mathcal{I}^{c}_{\text{min}}\neq\mathcal{I}^{c}$. By the definition of $\Gamma^{*}_{\bm{\theta}}$, $\mathcal{I}^{c}_{\text{min}}$ is non-empty.
Since $C_{i,j}$’s are continuous functions strictly increasing in $p_{j}$, there exists a sufficiently small $\epsilon>0$ such that
$\min_{i\in\mathcal{I}}C_{i,j}\left(p^{*}_{i},p^{*}_{j}-\epsilon/(|\mathcal{I}^{c}|-|\mathcal{I}^{c}_{\text{min}}|)\right)>\Gamma^{*}_{\bm{\theta}},\quad\forall j\in\mathcal{I}^{c}\backslash\mathcal{I}^{c}_{\text{min}}$
and
$\min_{i\in\mathcal{I}}C_{i,j^{\prime}}(p^{*}_{i},p^{*}_{j^{\prime}}+\epsilon/|\mathcal{I}^{c}_{\text{min}}|)>\min_{i\in\mathcal{I}}C_{i,j^{\prime}}(p^{*}_{i},p^{*}_{j^{\prime}})=\Gamma^{*}_{\bm{\theta}},\quad\forall j^{\prime}\in\mathcal{I}^{c}_{\text{min}}.$
Hence, $\bm{p}^{*}$ cannot be an optimal solution, leading to a contradiction. Consequently, we must have (58).
With similar argument, we can show that
$\Gamma^{*}_{\bm{\theta}}=\min_{j^{\prime}\in\mathcal{I}^{c}}C_{i,j}(\bm{p}^{*}),\quad\forall i\in\mathcal{I},$
using the fact that $C_{i,j}$’s are continuous functions strictly increasing in $p_{i}$. ∎
F.4.3 Overall balance in connected subgraphs: Proof of Proposition 11 and its extension
In this section, we first prove Proposition 11 for Gaussian bandits. Then we present an extention of it to general exponential family distributions.
Proof of Proposition 11.
Consider any connected subgraph $\mathcal{G}_{\ell}=\mathcal{G}(\mathcal{I}_{\ell},\mathcal{I}^{c}_{\ell},\mathcal{E}_{\ell})$ of the bipartite graph induced by an optimal solution $\bm{p}$. We show that if
$\sum_{i\in\mathcal{I}_{\ell}}p_{i}^{2}\neq\sum_{j\in\mathcal{I}^{c}_{\ell}}p_{j}^{2}.$
then we can find another $\bm{p}^{\prime}\neq\bm{p}$ with higher objective value, and hence $\bm{p}$ cannot be an optimal solution.
By Proposition 10, for any arm $i\in[K]$, there must exist an edge $e\in\mathcal{E}$ such that $i$ is incident to $e$. Hence, we have $\mathcal{I}_{\ell}\neq\varnothing$ and $\mathcal{I}^{c}_{\ell}\neq\varnothing$. Take any $i_{1}\in\mathcal{I}_{\ell}$ as an anchor point. By the connectedness of $\mathcal{G}_{\ell}$, for any $j\in\mathcal{I}^{c}_{\ell}$, there exist a path $p_{i_{1},j}=(i_{1},j_{1},i_{2},j_{2},\dots,j_{r}=j)$ connecting $i_{1}$ and $j$. Along the path, we perform Taylor’s expansion
$C_{i,j}(\bm{p}+\bm{\epsilon})=C_{i,j}(\bm{p})+\frac{\partial C_{i,j}(\bm{p})}{\partial p_{i}}\epsilon_{i}+\frac{\partial C_{i,j}(\bm{p})}{\partial p_{j}}\epsilon_{j}+o(\|\bm{\epsilon}\|).$
Starting from the first edge $(i_{1},j_{1})$, we define $\tilde{\epsilon}_{i_{1}}$ and $\tilde{\epsilon}_{j_{1}}$ such that the first order terms in the Taylor’s series vanishes
$\frac{\partial C_{i_{1},j_{1}}(\bm{p})}{\partial p_{i_{1}}}\tilde{\epsilon}_{i_{1}}+\frac{\partial C_{i_{1},j_{1}}(\bm{p})}{\partial p_{j_{1}}}\tilde{\epsilon}_{j_{1}}=0.$ (59)
Recall that, for Gaussian bandits with reward distribution $N(\theta_{i},\sigma^{2})$, we have
$C_{i,j}(\bm{p})=\frac{(\theta_{i}-\theta_{j})^{2}}{2\sigma^{2}}\frac{p_{i}p_{j}}{p_{i}+p_{j}}$
and
$\frac{\partial C_{i,j}(\bm{p})}{\partial p_{i}}=\frac{(\theta_{i}-\theta_{j})^{2}}{2\sigma^{2}}\frac{p_{j}^{2}}{(p_{i}+p_{j})^{2}},\quad\frac{\partial C_{i,j}(\bm{p})}{\partial p_{j}}=\frac{(\theta_{i}-\theta_{j})^{2}}{2\sigma^{2}}\frac{p_{i}^{2}}{(p_{i}+p_{j})^{2}}.$
Hence, we can rewrite (59) as
$\tilde{\epsilon}_{j_{1}}=-\tilde{\epsilon}_{i_{1}}\frac{p^{2}_{j_{1}}}{p^{2}_{i_{1}}}.$
Repeat this procedure for $(j_{1},i_{2})$ on the path, we have
$\tilde{\epsilon}_{i_{2}}=-\tilde{\epsilon}_{j_{1}}\frac{p^{2}_{i_{2}}}{p^{2}_{j_{1}}}=\tilde{\epsilon}_{i_{1}}\frac{p^{2}_{j_{1}}}{p^{2}_{i_{1}}}\frac{p^{2}_{i_{2}}}{p^{2}_{j_{1}}}=\tilde{\epsilon}_{i_{1}}\frac{p^{2}_{i_{2}}}{p^{2}_{i_{1}}}.$
Repeat for all edges on the path, we have
$\tilde{\epsilon}_{j}=-\tilde{\epsilon}_{i_{1}}\frac{p^{2}_{j}}{p^{2}_{i_{1}}},\quad j\in\mathcal{I}^{c}_{\ell}.$
Similarly, for all $i\in\mathcal{I}_{\ell}$, we have
$\tilde{\epsilon}_{i}=\tilde{\epsilon}_{i_{1}}\frac{p^{2}_{i}}{p^{2}_{i_{1}}},\quad i\in\mathcal{I}_{\ell}.$
Now, we claim that if $\sum_{i\in\mathcal{I}_{\ell}}p_{i}^{2}\neq\sum_{j\in\mathcal{I}^{c}_{\ell}}p_{j}^{2}$, then we can find a $\bm{\epsilon}$ such that $\Gamma_{\bm{\theta}}(\bm{p}+\bm{\epsilon})>\Gamma_{\bm{\theta}}(\bm{p})$. Consequently, $\bm{p}$ cannot be an optimal solution.
To this end, let
$\delta\equiv\sum_{i\in\mathcal{I}_{\ell}}\tilde{\epsilon}_{i}+\sum_{j\in\mathcal{I}^{c}_{\ell}}\tilde{\epsilon}_{j}=\tilde{\epsilon}_{i_{1}}\left(\sum_{i\in\mathcal{I}_{\ell}}\frac{p^{2}_{i}}{p^{2}_{i_{1}}}-\sum_{j\in\mathcal{I}^{c}_{\ell}}\frac{p^{2}_{j}}{p^{2}_{i_{1}}}\right).$
We consider two cases.
  1. 1.
    If $\sum_{i\in\mathcal{I}_{\ell}}p_{i}^{2}>\sum_{j\in\mathcal{I}^{c}_{\ell}}p_{j}^{2}$, then let $\epsilon_{i_{1}}<0$ be sufficiently small such that the first order terms dominates the remainder in the Taylor series. We have $\delta<0$. Define
    $\epsilon_{i}$ $=\tilde{\epsilon}_{i}-\frac{\delta}{|\mathcal{I}_{\ell}|}>\tilde{\epsilon}_{i},\quad i\in\mathcal{I}_{\ell},$
    $\epsilon_{j}$ $=\tilde{\epsilon}_{j},\quad i\in\mathcal{I}_{\ell}.$
    Note that $\frac{\partial C_{i,j}(\bm{p})}{\partial p_{i}}$ is strictly positive, then the first order term is strictly positive for any $(i,j)$ pairs, and hence $\Gamma_{\bm{\theta}}(\bm{p}+\bm{\epsilon})>\Gamma_{\bm{\theta}}(\bm{p})$.
  2. 2.
    If $\sum_{i\in\mathcal{I}_{\ell}}p_{i}^{2}<\sum_{j\in\mathcal{I}^{c}_{\ell}}p_{j}^{2}$, then let $\epsilon_{i_{1}}>0$ be sufficiently small such that the first order terms dominates the remainder in the Taylor series. We have $\delta<0$. Define
    $\epsilon_{i}$ $=\tilde{\epsilon}_{i}-\frac{\delta}{|\mathcal{I}_{\ell}|}>\tilde{\epsilon}_{i},\quad i\in\mathcal{I}_{\ell},$
    $\epsilon_{j}$ $=\tilde{\epsilon}_{j},\quad i\in\mathcal{I}_{\ell}.$
    Then the first order term is strictly positive for any $(i,j)$ pairs, and hence $\Gamma_{\bm{\theta}}(\bm{p}+\bm{\epsilon})>\Gamma_{\bm{\theta}}(\bm{p})$.
Finally, we remark that it is straightforward to generalize the above result to Gaussian bandits with heterogeneous variances. To this end, note that
$C_{i,j}(\bm{p})=\frac{(\theta_{i}-\theta_{j})^{2}}{2}\frac{p_{i}p_{j}}{\sigma_{j}^{2}p_{i}+\sigma_{i}^{2}p_{j}}$
and
$\frac{\partial C_{i,j}(\bm{p})}{\partial p_{i}}=\frac{(\theta_{i}-\theta_{j})^{2}}{2}\frac{\sigma_{i}^{2}p_{j}^{2}}{(\sigma_{j}^{2}p_{i}+\sigma_{i}^{2}p_{j})^{2}},\quad\frac{\partial C_{i,j}(\bm{p})}{\partial p_{j}}=\frac{(\theta_{i}-\theta_{j})^{2}}{2}\frac{\sigma_{j}^{2}p_{i}^{2}}{(\sigma_{j}^{2}p_{i}+\sigma_{i}^{2}p_{j})^{2}}.$
Consequently, we have
$\frac{\frac{\partial C_{i,j}(\bm{p})}{\partial p_{i}}}{\frac{\partial C_{i,j}(\bm{p})}{\partial p_{j}}}=\frac{p_{j}^{2}/\sigma_{j}^{2}}{p_{i}^{2}/\sigma_{i}^{2}}.$
The rest follows exactly as the case with equal variances. We omit the details. ∎
In Proposition 11, we show that for Gaussian bandits, the optimal solution must balance the top and bottom arms in all connected subgraphs in terms of the sum-of-squared allocations. For general reward distribution, we have similar results for each connected subgraph. However, the form of this necessary condition cannot be easily simplified.
To derive the general result, we mirror the proof of Proposition 11. Consider any connected subgraph $\mathcal{G}_{\ell}=\mathcal{G}(\mathcal{I}_{\ell},\mathcal{I}^{c}_{\ell},\mathcal{E}_{\ell})$ of the bipartite graph induced by the optimal solution $\bm{p}^{*}$. First, for any arm $i\in\mathcal{I}\cup\mathcal{I}^{c}$, there must exist an edge $e\in\mathcal{E}$ such that $i$ is incident to $e$. Hence, we have $\mathcal{I}_{\ell}\neq\varnothing$ and $\mathcal{I}^{c}_{\ell}\neq\varnothing$.
Take any $i_{1}\in\mathcal{I}_{\ell}$ as an anchor point. By the connectedness of $G_{\ell}$, for any $j\in\mathcal{I}^{c}_{\ell}$, there exists a path $x_{i_{1},j}=(i_{1},j_{1},i_{2},j_{2},\dots,j_{r}=j)$ connecting $i_{1}$ and $j$. Along the path, we perform Taylor’s expansion and derive that
$C_{i,j}(\bm{p}+\bm{\epsilon})=C_{i,j}(\bm{p})+\frac{\partial C_{i,j}(\bm{p})}{\partial p_{i}}\epsilon_{i}+\frac{\partial C_{i,j}(\bm{p})}{\partial p_{j}}\epsilon_{j}+o(\|\bm{\epsilon}\|).$
By Lemma 23, we have
$\frac{\partial C_{i,j}(\bm{p}^{*})}{\partial p_{i}}=d(\theta_{i},\bar{\theta}_{i,j})\quad\text{and}\quad\frac{\partial C_{i,j}(\bm{p}^{*})}{\partial p_{j}}=d(\theta_{j},\bar{\theta}_{i,j}).$
Define $\tilde{\epsilon}_{i_{1}}$ and $\tilde{\epsilon}_{j_{1}}$ such that the first order terms in the Taylor’s series vanish
$d(\theta_{i_{1}},\bar{\theta}_{i_{1},j_{1}})\tilde{\epsilon}_{i_{1}}+d(\theta_{j_{1}},\bar{\theta}_{i_{1},j_{1}})\tilde{\epsilon}_{j_{1}}=0.$
Hence, we can write
$\tilde{\epsilon}_{j_{1}}=-\tilde{\epsilon}_{i_{1}}\frac{d(\theta_{i_{1}},\bar{\theta}_{i_{1},j_{1}})}{d(\theta_{j_{1}},\bar{\theta}_{i_{1},j_{1}})}.$
Repeating this procedure along the path connecting $i_{1}$ and $j$, we have
$\tilde{\epsilon}_{j}=\tilde{\epsilon}_{j_{r}}=-\tilde{\epsilon}_{i_{1}}S_{i_{1}\to j},$
where
$S_{i_{1}\to j}=\frac{d(\theta_{i_{r}},\bar{\theta}_{i_{r},j_{r}})}{d(\theta_{j_{r}},\bar{\theta}_{i_{r},j_{r}})}\cdots\frac{d(\theta_{j_{1}},\bar{\theta}_{i_{2},j_{1}})}{d(\theta_{i_{2}},\bar{\theta}_{i_{2},j_{1}})}\frac{d(\theta_{i_{1}},\bar{\theta}_{i_{1},j_{1}})}{d(\theta_{j_{1}},\bar{\theta}_{i_{1},j_{1}})}.$ (60)
Similarly, we have for any $i\in\mathcal{I}_{\ell}$, there exists a path $p_{i_{1},i}=(i_{1},j^{\prime}_{1},i^{\prime}_{2},j^{\prime}_{2},\dots,i^{\prime}_{r^{\prime}}=i)$ connecting $i_{1}$ and $i$. We can then derive
$\tilde{\epsilon}_{i}=\tilde{\epsilon}_{i^{\prime}_{r^{\prime}}}=\tilde{\epsilon}_{i_{1}}S_{i_{1}\to i},$
where
$S_{i_{1}\to i}=\frac{d(\theta_{j^{\prime}_{r^{\prime}-1}},\bar{\theta}_{i^{\prime}_{r^{\prime}},j^{\prime}_{r^{\prime}-1}})}{d(\theta_{i^{\prime}_{r^{\prime}}},\bar{\theta}_{i^{\prime}_{r^{\prime}},j^{\prime}_{r^{\prime}-1}})}\cdots\frac{d(\theta_{j^{\prime}_{1}},\bar{\theta}_{i^{\prime}_{2},j^{\prime}_{1}})}{d(\theta_{i^{\prime}_{2}},\bar{\theta}_{i^{\prime}_{2},j^{\prime}_{1}})}\frac{d(\theta_{i^{\prime}_{1}},\bar{\theta}_{i^{\prime}_{1},j^{\prime}_{1}})}{d(\theta_{j^{\prime}_{1}},\bar{\theta}_{i^{\prime}_{1},j^{\prime}_{1}})}.$ (61)
We adopt the convention that $S_{i_{1}\to i_{1}}=1$. For Gaussian bandits, this simplifies to
$S_{i_{1}\to j}=\frac{(p^{*}_{j})^{2}}{(p^{*}_{i_{1}})^{2}}\quad\text{and}\quad S_{i_{1}\to i}=\frac{(p^{*}_{i})^{2}}{(p^{*}_{i_{1}})^{2}}.$
For general distribution, we cannot easily simplify $S_{i_{1}\to j}$ and $S_{i_{1}\to i}$. With the same argument as in the proof of Proposition 11, we have the following proposition.
Proposition 14 (Overall balance in connected subgraphs).
Consider a bandit with general exponential family rewards. Let $\mathcal{G}$ be the bipartite graph induced by the optimal solution $\bm{p}^{*}$ and let $\mathcal{G}_{\ell}$ be any connected subgraph of $\mathcal{G}$. For any anchor point $i_{1}\in\mathcal{I}_{\ell}$, $\mathcal{G}_{\ell}$ balances the top and bottom arms
$\sum_{i\in\mathcal{I}_{\ell}}S_{i_{1}\to i}=\sum_{j\in\mathcal{I}^{c}_{\ell}}S_{i_{1}\to j}.$ (62)
F.4.4 Proof of Proposition 12
Proof.
We specialize the KKT conditions for general pure exploration problems in (23) here for the best-$k$-arm identification problem as follows
$-1+\sum_{i\in\mathcal{I}}\sum_{j\in\mathcal{I}^{c}}\mu_{i,j}=0,$ (63a)
$\lambda-\sum_{j\in\mathcal{I}^{c}}\mu_{i,j}d(\theta_{i},\bar{\theta}_{i,j})=0,\quad\forall i\in\mathcal{I},$ (63b)
$\lambda-\sum_{i\in\mathcal{I}}\mu_{i,j}d(\theta_{j},\bar{\theta}_{i,j})=0,\quad\forall j\in\mathcal{I}^{c},$ (63c)
$\sum_{i\in[K]}p_{i}-1=0,\quad p_{i}\geq 0,\quad\forall i\in[K],$ (63d)
$\phi-C_{i,j}(\bm{p})\leq 0,\quad\forall i\in\mathcal{I},j\in\mathcal{I}^{c},$ (63e)
$\mu_{i,j}\geq 0,\quad\forall i\in\mathcal{I},j\in\mathcal{I}^{c},$ (63f)
$\sum_{i\in\mathcal{I}}\sum_{j\in\mathcal{I}^{c}}\mu_{i,j}(\phi-C_{i,j}(\bm{p}))=0.$ (63g)
Let $\bm{p}$ be the optimal primal solution so that (63d) holds automatically. We aim to find a KKT solution such that
$\mu_{i,j}=0,\quad\text{if $i\neq i_{k}$ or $j\neq j_{1}$}.$ (64)
In particular, under (64), we find that the solution to (63b) and (63c) satisfies
$\mu_{i,j}$ $=0,\quad\text{if $i\neq i_{k}$ or $j\neq j_{1}$},$
$\mu_{i_{k},j}$ $=\frac{\lambda}{d(\theta_{j},\bar{\theta}_{i_{k},j})},\quad\text{for $j\neq j_{1}$},$
[…]
$\mu_{i_{k},j_{1}}$ $=\lambda\left(1-\sum_{j\in\mathcal{I}^{c}\backslash\{j_{1}\}}\frac{d(\theta_{i_{k}},\bar{\theta}_{i_{k},j})}{d(\theta_{j},\bar{\theta}_{i_{k},j})}\right)\bigg{/}d(\theta_{i_{k}},\bar{\theta}_{i_{k},j_{1}})$
$=\lambda\left(1-\sum_{i\in\mathcal{I}\backslash\{i_{k}\}}\frac{d(\theta_{j_{1}},\bar{\theta}_{i,j_{1}})}{d(\theta_{i},\bar{\theta}_{i,j_{1}})}\right)\bigg{/}d(\theta_{j_{1}},\bar{\theta}_{i_{k},j_{1}}).$ (65)
where the multiplier $\lambda$ is determined by (63a), i.e.,
$\sum_{j\neq j_{1}}\mu_{i_{k},j}+\sum_{i\neq i_{k}}\mu_{i,j_{1}}+\mu_{i_{k},j_{1}}=1.$ (66)
For the dual feasibility condition (63f) to hold, we require condition (42) to hold, i.e.,
$1-\sum_{j\in\mathcal{I}^{c}\backslash\{j_{1}\}}\frac{d(\theta_{i_{k}},\bar{\theta}_{i_{k},j})}{d(\theta_{j},\bar{\theta}_{i_{k},j})}>0,$ (67)
For (65) to hold, we require (44), i.e.,
$\frac{d(\theta_{i_{k}},\bar{\theta}_{i_{k},j_{1}})}{d(\theta_{j_{1}},\bar{\theta}_{i_{k},j_{1}})}\sum_{i\in\mathcal{I}}\frac{d(\theta_{j_{1}},\bar{\theta}_{i,j_{1}})}{d(\theta_{i},\bar{\theta}_{i,j_{1}})}=\sum_{j\in\mathcal{I}^{c}}\frac{d(\theta_{i_{k}},\bar{\theta}_{i_{k},j})}{d(\theta_{j},\bar{\theta}_{i_{k},j})}.$ (68)
We have now checked that equations (63a), (63b), (63c), (63d) and (63f) hold.
For the primal variable (equivalent to objective value) $\phi$, we let
$\phi=C_{i_{k},j}(\bm{p})=C_{i,j_{1}}(\bm{p}),\quad\forall(i,j)\in\mathcal{I}\times\mathcal{I}^{c}.$ (69)
By noting that $\mu_{i,j}=0$ for $i\neq i_{k}$ and $j\neq j_{1}$, we have (63g). At last, when $\Gamma^{*}_{\bm{\theta}}=\phi$, the inequality in (63e) also holds automatically.
Hence, all the conditions in (63) holds. The proof is complete. ∎
Appendix G Examples of algorithms for pure-exploration problems
In this section, we collect examples of pure-exploration problems. For each pure-exploration problem, we explicitly derive the following components:
  1. 1.
    Correct answer $\mathcal{I}(\bm{\theta})$.
  2. 2.
    The alternative set $\mathrm{Alt}(\bm{\theta})$, the culprit set $\mathcal{X}(\bm{\theta})$, and the decomposition of $\mathrm{Alt}(\bm{\theta})$ that satisfies Assumption 1.
  3. 3.
    The unique solution $\bm{\vartheta}^{x}$ in Lemma 3.
  4. 4.
    The generalized Chernoff information $C_{x}(\bm{p})$.
  5. 5.
    The partial derivative of $C_{x}(\bm{p})$, from which the selection functions $h_{i}^{x}(\bm{p})$ can be derived.
With explicit formula for these components, the KKT detection rule (Algorithm 6) and the IDS selection rule (Algorithm 4) can be immediately implemented. To implement the TS detection rule (Algorithm 7), we need to sample from the posterior distribution, which is usually straightforward in the case of single-parameter exponential family. To implememnt the PPS detection rule (Algorithm 8), we need to calculate the posterior probability of the alternative set, which is also made possible by the explicit formula for the alternative sets.
G.1 BAI or IZ-free R&S
  1. 1.
    Correct answer: $\mathcal{I}(\bm{\theta})=\{I^{*}\}$, where $I^{*}=\operatorname*{argmax}_{i}\{\theta_{i}\}$.
  2. 2.
    $\mathrm{Alt}(\bm{\theta})=\cup_{j\in\mathcal{X}}\mathrm{Alt}_{j}$ with $\mathcal{X}=[K]\backslash\{I^{*}\}$ and $\mathrm{Alt}_{j}=\{\bm{\theta}:\theta_{j}>\theta_{I^{*}}\}$.
  3. 3.
    $\bm{\vartheta}^{j}$ is given by $\vartheta^{j}_{I^{*}}=\vartheta^{j}_{j}=\frac{p_{I^{*}}\theta_{I^{*}}+p_{j}\theta_{j}}{p_{I^{*}}+p_{j}}$ and $\vartheta^{j}_{k}=\theta_{k}$ for all $j\in\mathcal{X}$ and $k\neq I^{*},j$.
  4. 4.
    $C_{j}(\bm{p})=p_{I^{*}}d(\theta_{I^{*}},\vartheta^{j}_{I^{*}})+p_{j}d(\theta_{j},\vartheta^{j}_{j})$ for all $j\in\mathcal{X}$.
  5. 5.
    $\frac{\partial C_{j}}{\partial p_{i}}=d(\theta_{i},\vartheta^{j}_{i})$ for all $j\in\mathcal{X}$, most of them are zero, except for $i=I^{*}$ or $j$.
G.2 Best-$k$ identification
  1. 1.
    Correct answer: best-$k$ set $\mathcal{I}(\bm{\theta})=\operatorname*{argmax}_{\mathcal{S}\subseteq[K],|\mathcal{S}|=k}\left\{\sum_{i\in\mathcal{S}}\theta_{i}\right\}$.
  2. 2.
    $\mathrm{Alt}(\bm{\theta})=\cup_{(i,j)\in\mathcal{X}}\mathrm{Alt}_{i,j}$ with $\mathcal{X}=\mathcal{I}\times\mathcal{I}^{c}$ and $\mathrm{Alt}_{i,j}=\{\bm{\theta}:\theta_{j}>\theta_{i}\}$.
  3. 3.
    $\bm{\vartheta}^{(i,j)}$ is given by $\vartheta^{(i,j)}_{i}=\vartheta^{(i,j)}_{j}=\frac{p_{i}\theta_{i}+p_{j}\theta_{j}}{p_{i}+p_{j}}$ and $\vartheta^{(i,j)}_{k}=\theta_{k}$ for all $(i,j)\in\mathcal{X}$ and $k\neq i,j$.
  4. 4.
    $C_{(i,j)}(\bm{p})=p_{i}d(\theta_{i},\vartheta^{(i,j)}_{i})+p_{j}d(\theta_{j},\vartheta^{(i,j)}_{j})$ for all $(i,j)\in\mathcal{X}$.
  5. 5.
    $\frac{\partial C_{(i,j)}}{\partial p_{k}}=d(\theta_{k},\vartheta^{(i,j)}_{k})$ for all $(i,j)\in\mathcal{X}$, most of them are zero, except for $k=i$ or $j$.
G.3 Thresholding bandits
Consider a set of $K$ arms. Each arm $i$ has Gaussian reward with mean $\theta_{i}$. A threshold $S\in\mathbb{R}$ is given. The goal is to select the set of arms whose mean rewards are above a given threshold $\mathfrak{T}$.
  1. 1.
    Correct answer: $\mathcal{I}(\bm{\theta})=\{k\in[K]:\theta_{k}>\mathfrak{T}\}$.
  2. 2.
    $\mathrm{Alt}(\bm{\theta})=\cup_{j\in\mathcal{X}}\mathrm{Alt}_{j}$ where $\mathcal{X}=[K]$ and $\mathrm{Alt}_{j}=\{\bm{\theta}^{\prime}:(S-\theta^{\prime}_{j})(S-\theta_{j})<0\}$.
  3. 3.
    $\bm{\vartheta}^{j}$ is given by $\vartheta^{j}_{j}=\mathfrak{T}$ and $\vartheta^{j}_{k}=\theta_{k}$ for all $j\in\mathcal{X}$ and $k\neq j$.
  4. 4.
    $C_{j}(\bm{p})=\sum_{i\in[K]}p_{i}d(\theta_{i},\vartheta^{j}_{i})=p_{j}d(\theta_{i},\mathfrak{T})$.
  5. 5.
    $\frac{\partial C_{j}}{\partial p_{i}}=d(\theta_{i},\vartheta^{j}_{i})$ for all $j\in\mathcal{X}$, most of them are zero, except for $i=j$.
G.4 Thresholding bandit (selecting one arm closest to a threshold)
Consider a set of $K$ arms. Each arm $i$ has Gaussian reward with mean $\theta_{i}$. A threshold $\mathfrak{T}\in\mathbb{R}$ is given. The goal is to select the arm whose mean reward is the closest to the threshold $\mathfrak{T}$.
  1. 1.
    Correct answer: $\mathcal{I}(\bm{\theta})=I^{*}=\operatorname*{argmin}_{i\in[K]}|\theta_{i}-\mathfrak{T}|$.
  2. 2.
    $\mathrm{Alt}(\bm{\theta})=\cup_{j\in\mathcal{X}}\mathrm{Alt}_{j}$, where $\mathcal{X}=[K]\backslash{I^{*}}$ and $\mathrm{Alt}_{j}=\{\bm{\theta}:|\theta_{j}-\mathfrak{T}|<|\theta_{I^{*}}-\mathfrak{T}|\}$.
  3. 3.
    $\bm{\vartheta}=\{\bm{\vartheta}^{j}:j\in\mathcal{X}\}$ and
    $\bm{\vartheta}^{j}\in\operatorname*{argmin}_{\bm{\theta}^{\prime}\in\{\bm{\theta}:|\theta_{j}-\mathfrak{T}|<|\theta_{I^{*}}-\mathfrak{T}|\}}\sum_{k=1}^{K}p_{k}d(\theta_{k},\theta^{\prime}_{k}).$
    To solve this minimization problem in closed form, we note that
    $\min_{\bm{\theta}^{\prime}\in\{\bm{\theta}:|\theta_{j}-\mathfrak{T}|<|\theta_{I^{*}}-\mathfrak{T}|\}}\sum_{k=1}^{K}p_{k}d(\theta_{k},\theta^{\prime}_{k})=\min_{|\theta^{\prime}_{j}-\mathfrak{T}|<|\theta^{\prime}_{I^{*}}-\mathfrak{T}|}p_{I^{*}}d(\theta_{I^{*}},\theta^{\prime}_{I^{*}})+p_{j}d(\theta_{j},\theta^{\prime}_{j}).$
    That is to say $\vartheta^{j}_{k}=0$ for $k\neq I^{*},j$. Moreover, at the minimum, $|\vartheta^{j}_{j}-\mathfrak{T}|=|\vartheta^{j}_{I^{*}}-\mathfrak{T}|$ holds. Let $x=\theta^{\prime}_{j}-\mathfrak{T}$. Then the minimization problem reduces to
    $\min_{|\theta^{\prime}_{j}-\mathfrak{T}|<|\theta^{\prime}_{I^{*}}-\mathfrak{T}|}p_{I^{*}}d(\theta_{I^{*}},\theta^{\prime}_{I^{*}})+p_{j}d(\theta_{j},\theta^{\prime}_{j})$
    $=\min\left\{\min_{x}p_{I^{*}}d(\theta_{I^{*}},\mathfrak{T}+x)+p_{j}d(\theta_{j},\mathfrak{T}+x),\min_{x}p_{I^{*}}d(\theta_{I^{*}},\mathfrak{T}-x)+p_{j}d(\theta_{j},\mathfrak{T}+x)\right\}$
    $=\min\Bigg{\{}p_{I^{*}}d\left(\theta_{I^{*}},\frac{p_{I^{*}}\theta_{I^{*}}+p_{j}\theta_{j}}{p_{I^{*}}+p_{j}}\right)+p_{j}d\left(\theta_{j},\frac{p_{I^{*}}\theta_{I^{*}}+p_{j}\theta_{j}}{p_{I^{*}}+p_{j}}\right),$
    $\hskip 40.0ptp_{I^{*}}d\left(\theta_{I^{*}},\frac{\theta_{I^{*}}p_{I^{*}}-\theta_{j}p_{j}+2p_{j}\mathfrak{T}}{p_{I^{*}}+p_{j}}\right)+p_{j}d\left(\theta_{j},\frac{-\theta_{I^{*}}p_{I^{*}}+\theta_{j}p_{j}+2p_{I^{*}}\mathfrak{T}}{p_{I^{*}}+p_{j}}\right)\Bigg{\}}$
    $=\min\left\{\frac{p_{I^{*}}p_{j}}{p_{I^{*}}+p_{j}}\frac{(\theta_{I^{*}}-\theta_{j})^{2}}{2},\frac{p_{I^{*}}p_{j}}{p_{I^{*}}+p_{j}}\frac{(2\mathfrak{T}-\theta_{I^{*}}-\theta_{j})^{2}}{2}\right\}$
    This implies that, if $(\theta_{I^{*}}-\theta_{j})^{2}>(2\mathfrak{T}-\theta_{I^{*}}-\theta_{j})^{2}$, then
    $\vartheta^{j}_{I^{*}}=\frac{\theta_{I^{*}}p_{I^{*}}-\theta_{j}p_{j}+2p_{j}\mathfrak{T}}{p_{I^{*}}+p_{j}},\quad\vartheta^{j}_{j}=\frac{-\theta_{I^{*}}p_{I^{*}}+\theta_{j}p_{j}+2p_{I^{*}}\mathfrak{T}}{p_{I^{*}}+p_{j}};$
    if $(\theta_{I^{*}}-\theta_{j})^{2}<(2\mathfrak{T}-\theta_{I^{*}}-\theta_{j})^{2}$, then
    $\vartheta^{j}_{I^{*}}=\vartheta^{j}_{j}=\frac{p_{I^{*}}\theta_{I^{*}}+p_{j}\theta_{j}}{p_{I^{*}}+p_{j}}.$
  4. 4.
    $C_{j}(\bm{p})=\sum_{i\in[K]}p_{i}d(\theta_{i},\vartheta^{j}_{i})$ for all $j\in\mathcal{X}$.
  5. 5.
    $\frac{\partial C_{j}}{\partial p_{i}}=d(\theta_{i},\vartheta^{j}_{i})$ for all $j\in\mathcal{X}$, most of them are zero, except for $i=I^{*}$ and $j$.
G.5 All-$\varepsilon$-best arm identification
Find all arms with mean larger than $\max_{j\in[K]}\theta_{j}-\varepsilon$, i.e., the set $\mathcal{I}_{\varepsilon}(\bm{\theta})$. See Mason et al. (2020).
  1. 1.
    Correct answer: $\mathcal{I}(\bm{\theta})=\{i:\theta_{i}\geq\max_{j\in[K]}\theta_{j}-\varepsilon\}$.
  2. 2.
    $\mathrm{Alt}(\bm{\theta})=\cup_{(i,j)\in\mathcal{X}}\mathrm{Alt}_{i,j}$, where $\mathcal{X}=\left\{(i,j)\in[K]\times[K]:i\in\mathcal{I},j\neq i\right\}$, and for any $(i,j)\in\mathcal{X}$, $\mathrm{Alt}_{i,j}=\{\bm{\vartheta}:\vartheta_{j}-\varepsilon>\vartheta_{i}\}$. To interpret, if there exists an arm $j\neq i$ with $\theta_{j}-\theta_{i}>\epsilon$, then arm $i$ must not be in the $\varepsilon$-best set.
  3. 3.
    For any $(i,j)\in\mathcal{X}$, $C_{i,j}=p_{i}d(\theta_{i},\bar{\theta}_{i,j,\varepsilon}-\varepsilon)+p_{j}d(\theta_{j},\bar{\theta}_{i,j,\varepsilon}),$ where $\bar{\theta}_{i,j,\varepsilon}=\frac{p_{i}(\theta_{i}+\varepsilon)+p_{j}\theta_{j}}{p_{i}+p_{j}}$.
  4. 4.
    The partial derivatives can be derived accordingly.
Appendix H Existing algorithms in the literature
H.1 Fixed-confidence algorithms
KL-LUCB (Kaufmann and Kalyanakrishnan, 2013).
At each round $t$, the algorithm computes the upper confidence bound $U_{i}(t)$ and lower confidence bound $L_{i}(t)$ of the mean reward for each arm $i$:
$U_{i}(t)$ $=\max\{q\in[\theta_{t,i},\infty):T_{t,i}d(\theta_{t,i},q)\leq\beta_{t,\delta}\},$ (70)
$L_{i}(t)$ $=\min\{q\in(-\infty,\theta_{t,i}]:T_{t,i}d(\theta_{t,i},q)\leq\beta_{t,\delta}\},$
where $\beta_{t,\delta}$ is the exploration rate we specify to be $\log\left((\log t+1)/\delta\right)$. Then it samples two critical arms $u_{t}$ and $l_{t}$:
$u_{t}=\operatorname*{argmax}_{i\notin\widehat{\mathcal{I}}_{t}}U_{i}(t)\quad\text{and}\quad l_{t}=\operatorname*{argmin}_{i\in\widehat{\mathcal{I}}_{t}}L_{i}(t),$
The algorithm stops when $U_{u_{t}}(t)>L_{l_{t}}(t)$.
UGapE (Gabillon et al., 2012).
At each round $t$, the algorithm first compute the same $U_{i}(t)$ and $L_{i}(t)$ as (70). Then it computes the gap index $B_{i}(t)=\max_{i^{\prime}\neq i}^{k}U_{i^{\prime}}(t)-L_{i}(t)$ for each arm $i$, where $\max^{k}$ denotes the operator returning the $k$-th largest value. Based on $B_{i}(t)$, it evaluates a set of $k$ arms: $J(t)=\{j:\operatorname*{argmin}^{j}B_{j}(t)\}$. From $J(t)$, it computes two critical arms $u_{t}$ and $l_{t}$:
$u_{t}=\operatorname*{argmax}_{i\notin J_{t}}U_{i}(t)\quad\text{and}\quad l_{t}=\operatorname*{argmin}_{i\in J_{t}}L_{i}(t),$
and samples one with larger $\beta_{t,\cdot}$. The algorithm stops when $\max_{i\in J(t)}B_{i}(t)>0$.
$m$-LinGapE (Réda et al., 2021a).
At each round $t$, the algorithm first computes the same $U_{i}(t)$ and $L_{i}(t)$ as (70). Then it computes the gap index $B_{i,j}(t)=U_{i}(t)-L_{j}(t)$ for each arm $i$. It chooses $J(t)$ as the set of $k$ arms with largest empirical means. From $J(t)$, it computes two critical arms $u_{t}$ and $l_{t}$:
$u_{t}=\operatorname*{argmax}_{j\in J_{t}}\max_{i\notin J_{t}}B_{i,j}(t)\quad\text{and}\quad l_{t}=\operatorname*{argmax}_{i\notin J_{t}}B_{i,u_{t}}(t),$
and samples one which would reduce the variance of the empirical mean difference between them. (This is the so-called “greedy selection rule” in Réda et al. (2021a).) The algorithm stops when
$B_{l_{t},u_{t}}(t)>0.$
We remark that, in Réda et al. (2021a), the authors introduce a meta algorithm called “Gap-Index-Focused” Algorithm (GIFA), which includes KL-LUCB, UGapE and $m$-LinGapE in its framework. Refer to Algorithm 1 and Table 1 in Réda et al. (2021a).
MisLid (Réda et al., 2021b).
The algorithm is designed from the lower bound maximin problem and achieves the asymptotic optimality. It maintains a learner over the probability simplex of number of arms. At each round $t$, it computes a gain function and updates the learner. Then it perform tracking-rule to sample an arm. We adopt AdaHedge as the learner and use empirical gain function $g_{t}$ as
$\operatorname*{argmin}_{\bm{\lambda}\in\text{Alt}(\bm{\mu}^{t})}\sum_{i=1}^{K}\bm{p}^{t}_{i}(\bm{\mu}^{t}_{i}-\bm{\lambda})^{2}$
where $\bm{p}^{t}$ is probability distribution returned by AdaHedge, $\bm{\mu}_{t}$ is the empirical mean estimate and $\text{Alt}(\bm{\mu}^{t})$ defines the set of bandit model returning the top-$k$ arm set different from $(\bm{\mu}^{t})$. As suggested by Réda et al. (2021b), the empirical gain function reduced the sample complexity practically compared with the optimistic gain function suggested by theory. For the tracking rule, we adopt C-Tracking in our experiments. For the stopping rule, we adopt the same General Likelihood Ratio Test statistic with threshold $\log\left((\log(t)+1)/\delta\right)$.
FWS (Wang et al., 2021).
The algorithm track the lower bound maximin problem. At each round $t$, it performs forced exploration or Frank-Wolfe update. Due to the non-smoothness of objective function, it construct an $r_{t}$-sub-differential subspaces and then solve a zero-sum game. Note $r_{t}$ is a tuning parameter. We choose $r_{t}=t^{-0.9}/K$ as suggested in Wang et al. (2021). For the stopping rule, we adopt the same General Likelihood Ratio Test statistic with threshold $\log\left((\log(t)+1)/\delta\right)$.
Lazy-Mirror-Ascent (Ménard, 2019).
The algorithm treats the lower bound problem as
$\max_{\bm{w}}F(\bm{w}).$
It uses a sampling rule based on an online lazy mirror ascent. At each round $t$, it tracks the allocation weights $\bm{w}(t)$ and updates it by
$\tilde{\bm{w}}(t+1)=\operatorname*{argmax}_{\bm{w}}\eta_{t+1}\sum_{s=K}^{t}\bm{w}\cdot\nabla F(\tilde{\bm{w}}(s);\bm{\mu}(s))-\text{kl}(\bm{w},\bm{\pi});$
${\bm{w}}(t+1)=(1-\gamma_{t})\tilde{\bm{w}}(t+1)+\gamma_{t}\bm{\pi},$
where $\bm{\mu}(s)$ is the empirical mean estimates and $\bm{\pi}$ is the uniform distribution. There exists two tuning parameter $\eta_{t}$ and $\gamma_{t}$. Based on the code released by Wouter M. Koolen (https://bitbucket.org/wmkoolen/tidnabbil/src/master/) and the suggestion in Ménard (2019), we choose $\eta_{t}={\sqrt{\log K/t}}/{F^{*}}$ and $\gamma_{t}=1/(4\sqrt{t})$. Wouter M. Koolen comments that, although setting $F^{*}$ relies on the knowledge of true arm means, it is critical for the performance. In appendix of Wang et al. (2021), the authors also comment that this method is very sensitive to the learning rate $\eta_{t}$, the current choice is only for experimental comparison and cannot be used in real-world scenarios.
H.2 Fixed-budget algorithms
In the fixed-budget setting, some fixed-confidence algorithms can be adjusted for use. Meanwhile, another branch of algorithms following the Optimal Computing Budget Allocation (OCBA) framework is popular in the simulation optimization community. We review some of them below.
UGapE (Gabillon et al., 2012).
The UGapE provide a variant running in the fixed-budget setting. At each round $t$, the algorithm also identifies the arm with largest gap index $B_{i}(t)=\max_{i^{\prime}\neq i}^{k}U_{i^{\prime}}(t)-L_{i}(t)$. In the fixed-budget setting, $U_{i}(t)$ and $L_{i}(t)$ are defined by
$U_{i}(t)=\theta_{t,i}+\sqrt{\frac{a\times\texttt{budget}}{4HN_{t,i}}},\quad L_{i}(t)=\theta_{t,i}-\sqrt{\frac{a\times\texttt{budget}}{4HN_{t,i}}}.$ (71)
A tuning parameter $a$ for exploration is needed. Furthermore, the algorithm needs to estimate the sample complexity $H$ of the problem, which is defined by $H=\left(\sum_{i=1}^{K}\frac{1}{\Delta_{i}}\right)$ where the gap $\Delta_{i}$ for each arm is defined as
$\Delta_{i}=\begin{cases}\theta_{i}-\theta_{k+1}\hskip 10.0pt\text{if $i\leq k$,}\\ \theta_{k}-\theta_{i}\hskip 20.0pt\text{if $i>k$.}\end{cases}$
In our implementation, we give advantage to UGapE by selecting the best problem-dependent tuning parameter and pass-through the true sample complexity $H$, both of which are unknown beforehand.
SAR (Bubeck et al., 2013).
The SAR maintains an active set $\mathcal{A}$ and a selected set $\mathcal{S}$, which are initialized to be $\mathcal{A}=[K]$ and $\mathcal{S}=\emptyset$ The time horizon (budget) $T$ is divided into $K-1$ phases. At the beginning, the algorithm computes the following quantities:
$\overline{\log}(K)=\frac{1}{2}+\sum_{i=2}^{K}\frac{1}{i},\ n_{0}=0,\ \text{and}\ n_{p}=\left\lceil{\frac{1}{\overline{\log}(K)}\frac{T-K}{K+1-p}}\right\rceil\text{ for $p$ in $1,\cdots,K-1$}.$
At each phase $p$, the algorithm samples each active arm $i\in\mathcal{A}$ for $n_{p}-n_{p-1}$ times. Sort the arms in $\mathcal{A}$ in descending order, say $\theta_{t,1}\geq\theta_{t,2}\geq\cdots$. Define the following empirical gap
$\widehat{\Delta}_{i}=\begin{cases}\theta_{t,i}-\theta_{t,k-|S|+1}\hskip 10.0pt\text{if $i\leq k-|\mathcal{S}|$,}\\ \theta_{t,k-|S|}-\theta_{t,i}\hskip 20.0pt\text{if $i>k-|\mathcal{S}|$.}\end{cases}$
Let $i_{p}=\operatorname*{argmax}\widehat{\Delta}_{i}$ and $\mathcal{A}=\mathcal{A}\backslash\{i_{p}\}$. If $\theta_{t,i_{p}}>\theta_{t,k-|\mathcal{S}|+1}$, then $\mathcal{S}=\mathcal{S}\cup\{i_{p}\}$.
OCBAss (Gao and Chen, 2015).
For the Gaussian bandits with equal known variances, OCBAss can be described as follows. At the beginning, the algorithm samples each arm for $\Delta_{0}$ times. Then at time $t$, for all $i\in{\mathcal{I}}_{t}$ and $j\notin{\mathcal{I}}_{t}$, the algorithm compute $C_{i,j}$ defined as follows:
$C_{i,j}=\frac{N_{t,i}}{t}d\left(\theta_{t,i},\frac{N_{t,i}\theta_{t,i}+N_{t,j}\theta_{t,j}}{N_{t,i}+N_{t,j}}\right)+\frac{N_{t,j}}{t}d\left(\theta_{t,j},\frac{N_{t,i}\theta_{t,i}+N_{t,j}\theta_{t,j}}{N_{t,i}+N_{t,j}}\right).$ (72)
Then it identifies the $(i_{t},j_{t})$-pair that minimizes $C_{i,j}$ as candidate arms. If $\sum_{i\in{\mathcal{I}}_{t}}N_{t,i}^{2}<\sum_{j\notin{\mathcal{I}}_{t}}N_{t,j}^{2}$, then sample $i_{t}$ for $n$ times. Otherwise, sample $j_{t}$ for $n$ times. In the experiment, we choose the parameter $\Delta_{0}=n=10$ suggested by Gao and Chen (2015).
At first glance, the OCBAss algorithm appears very similar to the TS-KKT(0) algorithm, as they both identify the $(i,j)$-pair that minimizes $C_{i,j}$ as candidate arms. However, there is a key difference between them: when choosing from the $(i,j)$-pair, OCBAss balances the sum-of-squared allocation from the entire bipartite graph, while TS-KKT tries to balance the sum-of-squared allocation in each and every connected subgraph. Despite this difference, OCBAss still demonstrates impressive empirical performance and is optimal at least for a large problem instance. It is also worth noting that OCBAss discovered this sum-of-squared allocation structure some time ago, earlier than the BAI paper (Garivier and Kaufmann, 2016).
OCBASS (Gao and Chen, 2016).
At the beginning, the algorithm samples each arm for $\Delta_{0}$ times. Then at time $t$, the algorithm computes $C_{i,j}$ defined by (72). Then it identifies the $(i_{t},j_{t})$-pair that minimizes $C_{i,j}$ as candidate arms. Choose $i_{t}$ or $j_{t}$ arbitrarily and sample it for $n$ times. In the experiment, we choose the parameter $\Delta_{0}=n=10$ suggested by Gao and Chen (2016).
The OCBASS algorithm is very similar to the OCBA-ss algorithm. In particular, they both identify the $(i,j)$-pair that minimizes $C_{i,j}$ as candidate arms. In each iteration of the algorithm, there will be multiple arms that achieves the minimum value of $C_{i,j}$. However, the important step of balancing sum-of-squared allocation is missing, which implies that OCBASS cannot be optimal.
Appendix I Additional numerical experiments
I.1 Fixed-budget experiments
In this section, we test proposed algorithms in the fixed-budget setting. We consider the best-$5$-arm identification problem for a Gaussian bandit with $\theta_{i}=1-0.05*(i-1)$ and $\sigma_{i}^{2}=1/4$ for $i=[20]$.
Figure 3: Fix-budget performance estimated over $10^{7}$ independent replications.
In Figure 3, we present the fixed-budget performance of our proposed TS-KKT-IDS algorithm, along with benchmark methods reviewed in Appendix H.2. In the subplot of Figure 3, we provide a zoomed-out illustration on a logarithmic scale, which suggests a superior convergence rate in the probability of incorrect selection (PICS) for our algorithm compared to existing methods.
Figure 4: Compare the optimal allocation rate with the empirical allocation rate of the algorithms.
In Figure 4, we compare the optimal allocation rate with the empirical allocation rate of the algorithms, given a budget of $10^{6}$. We observe that both TS-KKT-IDS and OCBAss (omitted) allocate nearly exactly at the optimal rate. In this simpler case, OCBAss satisfies the correct sufficient condition and thus also converges to the optimal rate. In contrast, OCBASS, lacking the key step of balancing the sum-of-squared allocations, does not converge to the optimal allocation rate, though it remains sufficiently close. The UGapE and SAR algorithms deviate more significantly from the optimal allocation rate, although their empirical performance in terms of PICS is still respectable.
Figure 5: The probability of false selection for varying values of $k$ with a fixed budget of $1500$, estimated over $30000$ independent replications.
In Figure 5, we consider the case where the number of best arms to be identified varies from $k=1$ to $19$, displaying the PICS for each $k$ with a fixed budget of $1500$ samples. We observed the our TS-KKT-IDS algorithm consistently outperforms the alternatives across the range of $k$.
Appendix J Miscellaneous results and proofs
J.1 Facts about the exponential family
We restrict our attention to reward distribution following an one-dimensional natural exponential family. In particular, a natural exponential family has identity natural statistic in canonical form, i.e.,
$p(y|\eta)=b(y)\exp\{\eta y-A(\eta)\},\quad\eta\in\mathcal{T}=\left\{\eta:\int b(y)e^{\eta y}\mathrm{d}y<\infty\right\}$ (73)
[…]
$\theta(\eta)=\mathbb{E}_{\eta}[Y_{t,i}]=A^{\prime}(\eta),\quad\text{and}\quad A^{\prime\prime}(\eta)=\text{Var}_{\eta}(Y_{t,i})>0.$
We immediately have the following lemma.
Lemma 21.
At any $\eta\in\mathcal{T}^{o}$, $\theta(\eta)$ is strictly increasing and $A(\eta)$ is strictly convex.
Lemma 21 implies that $\theta(\cdot)$ is a one-to-one mapping. Consequently, we can use $\theta$ to denote the distribution $p(\cdot|\eta)$ for convenience. Let $\theta_{1}$ and $\theta_{2}$ be the mean reward of the distribution $p(\cdot|\eta_{1})$ and $p(\cdot|\eta_{2})$. Denote by $d(\theta_{1},\theta_{2})$ the Kullback-Leibler (KL) divergence from $p(\cdot|\eta_{1})$ to $p(\cdot|\eta_{2})$. We have convenient closed-form expressions of $d(\cdot,\cdot)$ and its partial derivative.
Lemma 22.
The partial derivatives of the KL divergense of a one-dimensional natural exponential family satisfies
$d(\theta_{1},\theta_{2})=A(\eta_{2})-A(\eta_{1})-\theta_{1}(\eta_{2}-\eta_{1}),$ (74)
$\frac{\partial d(\theta_{1},\theta_{2})}{\partial\theta_{1}}=\eta_{1}-\eta_{2},\quad\frac{\partial d(\theta_{1},\theta_{2})}{\partial\theta_{2}}=(\theta_{2}-\theta_{1})\frac{\mathrm{d}\eta(\theta_{2})}{\mathrm{d}\theta_{2}}.$ (75)
J.2 Proof of Lemma 3
For easy reference, we restate Lemma 3 here. See 3
Proof.
When the reward distribution follows a single-parameter exponential family, the KL divergence $d(m,m^{\prime})$ is continuous in $(m,m^{\prime})$ and strictly convex in $m^{\prime}$. A direct application of Wang et al. (2021, Theorem 4 and Lemma 7) yields a unique $\bm{\vartheta}^{x}$ that achieves the infimum in (7), as well as the differentiability of $C_{x}$. For the concavity, note that $C_{x}$ is the infimum over a family of linear functions in $\bm{p}$. This directly implies the concavity of $C_{x}$ in $\bm{p}$; see Boyd and Vandenberghe (2004, Chapter 3.2).
We say a function $f$ is homogeneous of degree one if it satisfies $f(s\bm{p})=sf(\bm{p})$ for any $s\geq 0$. The fact that $C_{x}(\bm{p})$ is homogeneous of degree one follows directly from its definition in (7). Combined with differentiability, the PDE characterization follows from Euler’s theorem for homogeneous functions; see Apostol (1969, p. 287) for example. ∎
J.3 Properties of $C_{x}(\cdot)$ in BAI and best-$k$-arm identification
We provide useful properties of the functions $C_{j}$ for BAI and the functions $C_{i,j}$ for best-$k$-arm identification.
Lemma 23.
For BAI, the minimizer $\bm{\vartheta}^{j}$ for (7) is given by
$\bm{\theta}^{j}=\bar{\theta}_{I^{*},j}(\bm{p})\triangleq\frac{p_{I^{*}}\theta_{I^{*}}+p_{j}\theta_{j}}{p_{I^{*}}+p_{j}},\quad\forall j\neq I^{*}.$
In particular, $C_{j}(\bm{p})$ is concave in $\bm{p}$, strictly increasing in $p_{I^{*}}$ and $p_{j}$, and
$\frac{\partial C_{j}(\bm{p})}{\partial p_{I^{*}}}=d(\theta_{I^{*}},\bar{\theta}_{I^{*},j})\quad\text{and}\quad\frac{\partial C_{j}(\bm{p})}{\partial p_{j}}=d(\theta_{j},\bar{\theta}_{I^{*},j}).$
For best-$k$-arm identification, the minimizer $\bm{\vartheta}^{(i,j)}$ for (7) is given by
$\bm{\theta}^{(i,j)}=\bar{\theta}_{i,j}(\bm{p})\triangleq\frac{p_{i}\theta_{i}+p_{j}\theta_{j}}{p_{i}+p_{j}},\quad\forall(i,j)\in\mathcal{I}\times\mathcal{I}^{c}.$
In particular, $C_{i,j}(\bm{p})$ is concave in $\bm{p}$, strictly increasing in $p_{i}$ and $p_{j}$, and
$\frac{\partial C_{i,j}(\bm{p})}{\partial p_{i}}=d(\theta_{i},\bar{\theta}_{i,j})\quad\text{and}\quad\frac{\partial C_{i,j}(\bm{p})}{\partial p_{j}}=d(\theta_{j},\bar{\theta}_{i,j}).$
Proof.
Note that $C_{j}$ for BAI is equivalent to $C_{I^{*},j}$ for best-$k$-arm identification, where $I^{*}$ is the best arm. To avoid redundancy, we focus on the functions $C_{i,j}$. To find the minimizer, we note that
$C_{i,j}(\bm{p})$ $\triangleq\inf_{\bm{\vartheta}\in\mathrm{Alt}_{i,j}}\sum_{i}p_{i}d(\theta_{i},\vartheta_{i})$
$=\inf_{\vartheta_{j}>\vartheta_{i}}\left\{p_{i}d(\theta_{i},\vartheta_{i})+p_{j}d(\theta_{j},\vartheta_{j})\right\}$
$=\inf_{\tilde{\theta}}\left\{p_{i}d(\theta_{i},\tilde{\theta})+p_{j}d(\theta_{j},\tilde{\theta})\right\}$
$=p_{i}d(\theta_{i},\bar{\theta}_{i,j})+p_{j}d(\theta_{j},\bar{\theta}_{i,j}),\quad\hbox{where}\quad\bar{\theta}_{i,j}\triangleq\frac{p_{i}\theta_{i}+p_{j}\theta_{j}}{p_{i}+p_{j}}.$ (76)
In the second line, we set $\vartheta_{k}=\theta_{k}$ for all $k\neq i,j$; the third line follows from the monotonicity of the KL divergence, and the last equality results from analytically solving the one-dimensional convex optimization.
The partial derivatives then follows directly from Lemma 3.
Alternatively, one may calculate the derivatives directly using Lemma 22 and the definition of $C_{i,j}$ in (76), as follows
$\frac{\partial C_{i,j}(\bm{p})}{\partial p_{i}}$ $=d(\theta_{i},\bar{\theta}_{i,j})+p_{i}\frac{\partial d(\theta_{i},\bar{\theta})}{\partial p_{i}}+p_{j}\frac{\partial d(\theta_{j},\bar{\theta})}{\partial p_{i}}$
$=d(\theta_{i},\bar{\theta}_{i,j})+\frac{\partial\bar{\theta}}{\partial p_{i}}\cdot\left[p_{i}\frac{\partial d(\theta_{i},\bar{\theta})}{\partial\bar{\theta}}+p_{j}\frac{\partial d(\theta_{j},\bar{\theta})}{\partial\bar{\theta}}\right]$
$=d(\theta_{i},\bar{\theta}_{i,j}).$
The last equality follows from the first-order optimality condition for solving the convex optimization problem in the third line of (76). The calculation of $\frac{\partial C_{i,j}(\bm{p})}{\partial p_{j}}$ proceeds similarly. ∎
Lemma 24.
For any $(i,j)\in\mathcal{I}\times\mathcal{I}^{c}$, the contour of $C_{i,j}(\cdot)$ as a bivariate function of $(p_{i},p_{j})$ does not contain any line segment.
Proof.
To prove the result, we need some basic facts.
  1. 1.
    We first give the Hessian matrix of $C_{i,j}(\bm{p})$ as follows
    $H_{i,j}=\begin{bmatrix}\frac{\partial^{2}C_{i,j}(\bm{p})}{\partial^{2}p_{i}}&\frac{\partial^{2}C_{i,j}(\bm{p})}{\partial p_{i}\partial p_{j}}\\ \frac{\partial^{2}C_{i,j}(\bm{p})}{\partial p_{j}\partial p_{i}}&\frac{\partial^{2}C_{i,j}(\bm{p})}{\partial^{2}p_{j}}\end{bmatrix}=\frac{(\theta_{i}-\theta_{j})^{2}}{(p_{i}+p_{j})^{3}}\frac{\mathrm{d}\eta(\bar{\theta}_{i,j})}{\mathrm{d}\theta}\begin{bmatrix}-p_{j}^{2}&p_{i}p_{j}\\ p_{i}p_{j}&-p_{i}^{2}\end{bmatrix},$
    where $\eta$ is the natural parameter of the exponential family. By Lemma 21, $\frac{\mathrm{d}\eta(\bar{\theta})}{\mathrm{d}\theta}$ is strictly positive. Note that $H_{i,j}$ has an eigenvalue 00. The other eigenvalue can be calculated correspondingly,
    $\lambda_{\min}(H_{i,j})=-\frac{p_{i}^{2}+p_{j}^{2}}{(p_{i}+p_{j})^{3}}\frac{\mathrm{d}\eta(\bar{\theta}_{i,j})}{\mathrm{d}\theta}(\theta_{i}-\theta_{j})^{2}.$
    Note that $\lambda_{\min}(H_{i,j})<0$ whenever $p_{i}$ or $p_{j}$ is greater than 00.
  2. 2.
    Recall that $C_{i,j}$ satisfies the partial differential equation in (8), which has a general solution in the form of
    $C_{i,j}(\bm{p})=p_{i}g_{i,j}\left(\frac{p_{j}}{p_{i}}\right),$
    where $g_{i,j}(\cdot)$ is any differentiable function. This implies that the graph of $C_{i,j}(\bm{p})$ is the boundary of a cone. The graph of a function $f(x)$ is defined as $G(f)=\{(x,f(x)):x\in D\}$, where $D$ is the domain of $f(x)$. In particular, for any point on the surface, the line segment connecting the point and the origin belongs to the graph of $C_{i,j}(\bm{p})$.
We prove the result by contradiction. Suppose that there exist a line segment in the contour of $C_{i,j}(p_{i},p_{j})$. Then the triangle specified by the two end points of the line segment and the origin belongs to the graph of $C_{i,j}$. This implies that at any point in the interior of this triangle, $C_{i,j}$ has a Hessian of $\bm{0}$. This contradicts with the fact that the Hessian of $C_{i,j}$ at any point $(p_{i},p_{j})$ such that $p_{i}$ or $p_{j}$ is greater than 00 is semi-negative definite, with a eigenvalue strictly smaller than 00. ∎
Lemma 25.
For any $(i,j)\in\mathcal{I}\times\mathcal{I}^{c}$, $C_{i,j}(\bm{p})$ can be written as
$C_{i,j}(\bm{p})=p_{i}g_{i,j}\left(\frac{p_{j}}{p_{i}}\right),$
where
$g_{i,j}(x)=d\left(\theta_{i},\frac{\theta_{i}+x\theta_{j}}{1+x}\right)+xd\left(\theta_{j},\frac{\theta_{i}+x\theta_{j}}{1+x}\right).$
Furthermore, we have
$g^{\prime}_{i,j}(x)=d\left(\theta_{j},\frac{\theta_{i}+x\theta_{j}}{1+x}\right)>0.$
Moreover, for any $i,j$ and $j^{\prime}$ such that $\theta_{i}>\theta_{j^{\prime}}\geq\theta_{j}$, we have
$g_{i,j}(x)\geq g_{i,j^{\prime}}(x)\quad\text{for all }x.$
Proof.
We start by identifying that
$C_{i,j}(p_{i},p_{j})=p_{i}g_{i,j}\left(\frac{p_{j}}{p_{i}}\right).$
Fix $p_{i}$ and let $x=\frac{p_{j}}{p_{i}}$,
$\frac{\partial C_{i,j}}{\partial p_{j}}=p_{i}g_{i,j}^{\prime}(x)\frac{\partial x}{\partial p_{j}}=g_{i,j}^{\prime}(x).$
By Lemma 23, we have
$\frac{\partial C_{i,j}}{\partial p_{j}}=d\left(\theta_{j},\frac{p_{i}\theta_{i}+p_{j}\theta_{j}}{p_{i}+p_{j}}\right)=d\left(\theta_{j},\frac{\theta_{i}+x\theta_{j}}{1+x}\right).$
Hence,
$g_{i,j}^{\prime}(x)=d\left(\theta_{j},\frac{\theta_{i}+x\theta_{j}}{1+x}\right)>0.$
To this end, we fix $\theta_{i}$ and $x$ and define
$h(\alpha)=d\left(\theta_{i},\frac{\theta_{i}+x\alpha}{1+x}\right)+xd\left(\alpha,\frac{\theta_{i}+x\alpha}{1+x}\right).$
It follows that $h(\theta_{j})=g_{i,j}(x)$. To complete the proof, we need to show that, for any $\theta_{i}>\theta_{j^{\prime}}\geq\theta_{j}$, $g_{i,j}(x)\geq g_{i,j^{\prime}}(x)$. It suffices to verify that $h^{\prime}(\alpha)<0$ when $\theta_{i}>\alpha$. To see this, let
$m=\frac{\theta_{i}+x\alpha}{1+x}.$
We recall the partial derivative of $d(\cdot,\cdot)$ in (75) and calculate $h^{\prime}(\alpha)$ as follows.
$h^{\prime}(\alpha)$ $=\frac{\partial d(\theta_{i},m)}{\partial m}\frac{x}{1+x}+x\left[\frac{\partial d(\alpha,m)}{\partial\alpha}+\frac{\partial d(\alpha,m)}{\partial m}\frac{x}{1+x}\right]$
$=(m-\theta_{i})\frac{\mathrm{d}\eta}{\mathrm{d}m}\frac{x}{1+x}+x\left[\eta(\alpha)-\eta(m)+(m-\alpha)\frac{\mathrm{d}\eta}{\mathrm{d}m}\frac{x}{1+x}\right]$
$=x(\eta(\alpha)-\eta(m))+((m-\theta_{i})+x(m-\alpha))\frac{\mathrm{d}\eta}{\mathrm{d}m}\frac{x}{1+x}$
$=x(\eta(\alpha)-\eta(m)).$
Noticing that $\eta^{\prime}(\theta)=1/\theta^{\prime}(\eta)>0$ and $m-\alpha=(\theta_{i}-\alpha)/(1+x)$, we have
$h^{\prime}(\alpha)\begin{cases}<0,\quad\text{if }\theta_{i}>\alpha,\\ >0,\quad\text{if }\theta_{i}<\alpha.\end{cases}$ (77)
This completes the proof. ∎
J.4 Proof of Corollary 2
Assuming strictly positiveness of $\bm{p}^{*}$, recall from (23b) that
$\lambda-\sum_{x\in\mathcal{X}}\mu_{x}\frac{\partial C_{x}(\bm{p})}{\partial p_{i}}=0,\quad\forall i\in[K],$
where we used the fact that $\iota_{i}=0$ due to complementary slackness. By the definition of
$\mathcal{X}_{i}\triangleq\left\{x\in\mathcal{X}(\bm{\theta}):\frac{\partial C_{x}(\bm{p})}{\partial p_{i}}>0\right\},$
we have
$\lambda-\sum_{x\in\mathcal{X}_{i}}\mu_{x}\frac{\partial C_{x}(\bm{p})}{\partial p_{i}}=0,\quad\forall i\in[K].$ (78)
Notice that $\lambda=\phi=\Gamma^{*}_{\bm{\theta}}>0$ from the proof of Theorem 4 and that $\frac{\partial C_{x}(\bm{p})}{\partial p_{i}}>0$, there must exist a $x_{0}\in\mathcal{X}_{i}$ such that $\mu_{x}>0$. The complementary-slackness condition then implies that $C_{x_{0}}(\bm{p}^{*})=\phi=\Gamma^{*}_{\bm{\theta}}$. Consequently,
$\Gamma_{\bm{\theta}}^{*}=\min_{x\in\mathcal{X}_{i}}C_{x}(\bm{p}^{*}),\quad\hbox{for all }i\in[K].$