### v3

# Dual-Directed Algorithm Design for Efficient Pure Exploration

Abstract
While experimental design often focuses on selecting the single best alternative from a finite set (e.g., in ranking and selection or best-arm identification), many pure-exploration problems pursue richer goals. Given a specific goal, adaptive experimentation aims to achieve it by strategically allocating sampling effort, with the underlying sample complexity characterized by a maximin optimization problem. By introducing dual variables, we derive necessary and sufficient conditions for an optimal allocation, yielding a unified algorithm design principle that extends the top-two approach beyond best-arm identification. This principle gives rise to Information-Directed Selection, a hyperparameter-free rule that dynamically evaluates and chooses among candidates based on their current informational value. We prove that, when combined with Information-Directed Selection, top-two Thompson sampling attains asymptotic optimality for Gaussian best-arm identification, resolving a notable open question in the pure-exploration literature. Furthermore, our framework produces asymptotically optimal algorithms for pure-exploration thresholding bandits and $\varepsilon$-best-arm identification (i.e., ranking and selection with probability-of-good-selection guarantees), and more generally establishes a recipe for adapting Thompson sampling across a broad class of pure-exploration problems. Extensive numerical experiments highlight the efficiency of our proposed algorithms compared to existing methods.
1 Introduction
A common goal in experimental design is to identify the best-performing option among a finite set of alternatives. Adaptive experiments aim to achieve this goal efficiently by reducing measurement effort on poorly performing options and concentrating resources on more promising ones. This problem is widely known as ranking and selection (R&S) in the operations research literature, and as best-arm identification (BAI) in the machine learning literature. The R&S problem serves as a crucial abstraction for many applications. For example, it is used to determine optimal hyperparameters for algorithms (Shang et al. 2019). In the context of large language models, it helps in identifying the most effective beam search direction for automatic prompt optimization (Pryzant et al. 2023). More recently, the approach has been applied in brain-computer interface systems to interpret neural signals under a variety of stimulus paradigms (Zhou et al. 2024). For further details on the relationship between BAI and R&S, see Glynn and Juneja (2018).
A distinguishing characteristic of R&S and BAI problems is their emphasis on end outcomes. In these settings, the decision-maker (DM) is not penalized for losses incurred during the adaptive phase. Instead, the primary focus is on the accuracy of the final decision, which can have substantial real-world consequences. For instance, the final decision might result in the deployment of a “winning” feature, the construction of infrastructure, or the approval of a new drug. The goal is to accurately find the best alternative by the conclusion of the experiment, particularly since resources are committed to the selected option. In contrast to scenarios that seek immediate gains through exploitation of high-performing alternatives, these problems prioritize exploration to secure a well-informed final decision. Consequently, such sequential decision problems are often referred to as pure-exploration problems.
The study of pure-exploration problems dates back to the 1950s in the context of R&S (Bechhofer 1954). Since then, a wide range of algorithms have been developed to address R&S in both the simulation and statistical learning communities. Examples include elimination algorithms (Kim and Nelson 2001, Fan et al. 2016, Bubeck et al. 2013), Optimal Computing Budget Allocation algorithms (Chen et al. 2000, Gao and Chen 2015, Wu and Zhou 2018b), knowledge-gradient algorithms (Frazier et al. 2008), and UCB-type algorithms (Kalyanakrishnan et al. 2012, Kaufmann and Kalyanakrishnan 2013). For comprehensive surveys of R&S, we refer readers to Fu and Henderson (2017), Hong et al. (2021), Fan et al. (2024).
1.1 Motivations
A unified approach to general pure-exploration problems.
While identifying the best alternative (e.g., in R&S) is a common goal, adaptive experimentation is also well-suited for pure exploration problems that aim to achieve broader goals beyond best-arm identification. For instance, the DM may want to identify alternatives sufficiently close to the best-performing one, often referred to as R&S with guarantees on the probability of good selection (PGS) or $\varepsilon$-BAI in the machine learning literature. Other objectives include identifying all alternatives exceeding a specified performance threshold (Locatelli et al. 2016), selecting the best-$k$ alternatives (Chen et al. 2008), or determining whether the mean performance lies within a convex feasible set (Qiao and Tewari 2024).
Despite various scattered studies in the literature addressing specific exploration tasks, developing a unified and practical algorithmic framework capable of handling these diverse queries remains a significant challenge. To the best of our knowledge, general-purpose algorithms have been discussed in a limited number of works, such as Wang et al. (2021), Ménard (2019); however, these studies approach the problem as a generic optimization problem, whereas we explicitly leverage structural properties specific to pure-exploration problems.
Generalizing top-two algorithms.
A recent class of algorithms, known as top-two algorithms, initially introduced by Russo (2020) for BAI, has gathered increasing attention. These algorithms employ a two-phase approach for selecting the next alternatives to sample. In the first phase, a pair of top-two candidates—referred to as leader and challenger—are identified. The second phase, called selection step, involves sampling one of these candidates, often based on a potentially biased coin toss. [^1] We use the term “sampling” specifically to refer to the act of picking an alternative to draw a sample from, and reserve the term “selection” to describe the step of determining which candidate should be sampled. The simplicity and effectiveness of top-two algorithms have led to their widespread adoption, with notable contributions from Qin et al. (2017), Shang et al. (2020), and Jourdan et al. (2022).
However, a notable limitation of these algorithms is their dependence on tuning parameters to achieve optimality. The existence of simple, parameter-free algorithms that are provably optimal remains an open question (see Section 8, Russo 2020). Furthermore, the existing literature on top-two algorithms has predominantly focused on BAI, leaving broader pure-exploration tasks largely unexplored.
Generalizing Thompson sampling.
Thompson sampling (TS) has achieved notable success in minimizing regret for multi-armed bandits, attracting attention from both academia and industry (Russo et al. 2018). However, its inherent focus on maximizing cumulative rewards leads to excessive exploitation of the estimated best alternative, resulting in insufficient exploration of perceived suboptimal options to confidently verify whether the perceived best alternative is indeed the best. Consequently, modifications to TS are necessary for its effective application to pure exploration. Adapting TS for pure exploration could have an immediate impact in areas where TS is already widely used, such as A/B testing (Graepel et al. 2010), news article recommendation (Chapelle and Li 2011), and revenue management (Ferreira et al. 2018). Several studies have proposed TS variants for pure exploration, beginning with the top-two Thompson sampling algorithm (Russo 2020). Other adaptations include the Murphy sampling (Kaufmann et al. 2018), which modifies classical TS to test whether the lowest mean falls below a given threshold.
However, there is a lack of unified guidelines on systematically adapting TS for general pure-exploration queries. We address this gap with the introduction of Algorithm 7.
A unified approach to different problem formulations.
Various formulations have been developed to evaluate algorithm performance in pure exploration. The fixed-budget problem is often framed as a stochastic dynamic programming problem. However, the complexity of solving such problems exactly has led to a focus on approximate solutions, such as restricting analysis to procedures based on fixed static allocations. Building on OCBA (Chen et al. 2000), Glynn and Juneja (2004) provide a large-deviation characterization of the probability of incorrect selection (PICS) for static-allocation procedures, identifying an optimal static allocation that maximizes the exponential decay rate of PICS. This approach introduces a notion of problem complexity as the optimal decay exponent. Subsequent works (Chen and Ryzhov 2023, Avci et al. 2023) build upon this framework by designing dynamic procedures that converge to the optimal static allocation.
The fixed-precision formulation, also referred to as the fixed-confidence setting in the bandit literature, traces its origins to Chernoff (1952)’s seminal work on sequential binary hypothesis testing. R&S and BAI extend this to multiple hypotheses with sequentially chosen experiments. In this setting, the DM determines when to terminate the experiment and propose an estimated top performer. The goal is to minimize the expected sample size while guaranteeing a pre-specified probability of correct selection (PCS). Garivier and Kaufmann (2016) provided matching lower and upper bounds on the sample complexity, framing it as an optimization problem that prescribes the optimal sample allocation ratio.
The posterior convergence rate setting (Russo 2020) maintains a posterior distribution over the unknown parameters and evaluates the algorithms by the posterior PICS. This approach operates within a frequentist framework, where the problem instance is fixed, yet it adopts a Bayesian objective to evaluate performance. Russo (2020) identifies the optimal decay exponent for the posterior PICS as the complexity measure.
Remarkably, for Gaussian models, the problem complexity across the three formulations are identical, see Section 2.5.2. Despite extensive research on individual formulations of R&S and BAI, a unified algorithmic framework capable of addressing all three major formulations simultaneously remains an open challenge.
1.2 Contributions and Organization
This paper proposes a unified algorithm design framework that accommodates diverse problem formulations and a wide range of pure-exploration queries. Our main contributions are summarized as follows.
  1. 1.
    We identify common structural elements across various problem formulations and exploration queries, formulating the problem complexity into a distinctive non-smooth maximin convex optimization problem. This formulation characterizes the inherent difficulty of each query and serves as the foundation for our unified algorithmic framework. Guided by the shared structure of these problem complexity optimization problems, our framework seamlessly adapts to diverse pure-exploration tasks.
  2. 2.
    We establish necessary and sufficient conditions for the optimality of problem complexity optimization using dual variables, offering a fresh perspective on existing algorithms and enabling generalizations. Notably, our use of dual variables uncovers the foundation of top-two algorithms and TS variants in the complementary slackness condition of the Karush-Kuhn-Tucker conditions. This provides a natural generalization of the top-two principle to diverse pure-exploration problems and introduces an adaptive, provably optimal selection step, addressing an open question in Russo (2020).
  3. 3.
    Our design principle is highly flexible and easily adaptable to practical settings with minimal modifications. We demonstrate the versatility of our algorithm framework through its application to Gaussian noises with unequal and unknown variances, correlated rewards using common random numbers, and general pure-exploration problems. Additionally, we offer a systematic and accessible approach to tailoring Thompson sampling for general pure-exploration problems in Algorithm 7.
  4. 4.
    We provide theoretical guarantees for the optimality of our algorithms in various formulations and exploration queries. In particular, we show that our algorithm is optimal for R&S problems under Gaussian models under all three major formulations—fixed-budget, fixed-precision, and posterior convergence rate. We further establish optimality of our framework in pure-exploration thresholding bandits and $\varepsilon$-BAI.
  5. 5.
    We conduct extensive numerical studies to showcase the compelling performance of our unified algorithm across various settings. These experiments encompass a wide range of scenarios, including the fixed-confidence and fixed-budget settings, large-scale problems, applications with common random numbers, and general exploration tasks beyond conventional R&S and BAI. This comprehensive evaluation highlights the performance and robustness of our unified algorithm.
Organization.
The paper is organized as follows. In Section 2, we present an illustrative study on indifference-zone-free R&S problems in Gaussian models, demonstrating its optimality in posterior convergence rate and static fixed-budget settings. Section 3 introduces the problem formulation for general pure exploration and provides detailed explanations of our unified algorithm design principle, where we also present optimality for Gaussian BAI in the fixed-confidence setting. In Section 4, we discuss extensions beyond the base cases, including exploration queries with multiple correct answers and Gaussian best-arm identification with unknown variances. Finally, in Section 5, we conduct an extensive numerical study and discuss practical considerations for algorithm implementation.
2 An Illustrative Example: Indifference-Zone-Free Ranking and Selection in Gaussian Models
To illustrate our algorithm design, we study the canonical indifference-zone-free R&S problems in Gaussian models with known but potentially unequal variances, commonly known as BAI in the machine learning literature. Our analysis leads to optimal algorithms aimed at maximizing the posterior convergence rate (Russo 2020). The unknown-variance case and additional extensions are discussed further in Section 4.
2.1 Problem Formulation
Consider a set of $K$ systems, each with an unknown mean performance. For each system $i\in[K]\triangleq\{1,2,\dots,K\}$, the mean performance is characterized by an unknown parameter $\theta_{i}\in\mathbb{R}$. At each time step $t\in\mathbb{N}_{0}\triangleq\{0,1,\dots\}$, a decision-maker (DM) selects a system $I_{t}\in[K]$ and observes its random performance: $Y_{t+1,I_{t}}=\theta_{I_{t}}+W_{t,I_{t}},$ with $W_{t,I_{t}}\sim N(0,\sigma_{I_{t}}^{2}),$ where $W_{t,I_{t}}$ is a Gaussian noise with known variance $\sigma_{I_{t}}^{2}$, independent across time and systems. The true mean vector $\bm{\theta}\triangleq(\theta_{1},\dots,\theta_{K})$ is fixed but unknown. We assume there is a unique best system, denoted by $I^{*}(\bm{\theta})\triangleq\operatorname*{argmax}_{i\in[K]}\theta_{i}$, or simply $I^{*}$ when the dependence on $\bm{\theta}$ is clear. The DM’s goal is to identify this unknown best system $I^{*}$ by sequentially allocating a finite sampling budget $T$ among the $K$ systems.
Although this problem is set in a frequentist framework, with both the mean vector $\bm{\theta}$ and best system $I^{*}(\bm{\theta})$ fixed but unknown, we focus on minimizing the posterior probability that a suboptimal system is mistakenly identified as the best, a performance criteria popularized by Russo (2020).
Our performance criteria utilizes a prior distribution $\Pi_{0}$ with density $\pi_{0}$. We focus on uninformative priors for the mean vector, which are mutually independent Gaussian distributions with infinite variance. Given the history $\mathcal{H}_{t}\triangleq(I_{0},Y_{1,I_{0}},\dots,I_{t-1},Y_{t,I_{t-1}})$ collected by any allocation rule, the posterior distribution of each mean $\theta_{i}$ is normal, $N(\theta_{t,i},\sigma^{2}_{t,i})$, with posterior mean and variance given by: $\theta_{t,i}\triangleq\frac{1}{N_{t,i}}\sum_{\ell=0}^{t-1}{\bm{1}(I_{\ell}=i)Y_{\ell+1,I_{\ell}}}$ and $\sigma^{2}_{t,i}\triangleq\frac{\sigma_{i}^{2}}{N_{t,i}},$ where $N_{t,i}\triangleq\sum_{\ell=0}^{t-1}\bm{1}(I_{\ell}=i)$ is the number of samples allocated to system $i$ up to time $t$.
Define the set of mean vectors for which system $I^{*}=I^{*}(\bm{\theta})$ is uniquely the best as:
$\Theta_{I^{*}}\triangleq\bigl{\{}(\vartheta_{1},\ldots,\vartheta_{K})\in\mathbb{R}^{K}:\vartheta_{I^{*}}>\vartheta_{j},\,\,\forall j\neq I^{*}\bigr{\}}.$
Our goal is to minimize the posterior probability $\Pi_{t}(\Theta^{c}_{I^{*}})$, the probability of misidentifying the best system, where $\Theta^{c}_{I^{*}}$ is the complement of $\Theta_{I^{*}}$. Typically, this probability decays exponentially fast as $t\to\infty$, motivating the asymptotic objective to maximize the exponential decay rate
$\liminf_{t\to\infty}-\frac{1}{t}\log\Pi_{t}(\Theta^{c}_{I^{*}}).$
This criterion leads us to define optimality for any fixed problem instance $\bm{\theta}$ as follows.
Definition 1 (Optimal allocation rules).
An allocation rule $\textup{{A}}^{*}$ is said to be asymptotically optimal for $\bm{\theta}$ if, for every other non-anticipating allocation rule A,
$\liminf_{t\to\infty}\frac{-\log\Pi_{t}^{\textup{{A}}^{*}}\!(\Theta^{c}_{I^{*}})}{-\log\Pi_{t}^{\textup{{A}}}(\Theta^{c}_{I^{*}})}\geq 1\quad\text{with probability one,}$,
where the superscript indicates dependence on the allocation rule that determines data collection.
We now present a sufficient condition for asymptotic optimality in terms of long-run allocation proportions. Let $\bm{p}_{t}=(p_{t,1},\dots,p_{t,K})$ denote the empirical allocation vector, with $p_{t,i}$ being the fraction of samples allocated to system $i$ by time $t$. The proof of Theorem 1 is collected in Appendix B.
Theorem 1 (A sufficient condition for optimality).
An allocation rule is asymptotically optimal for $\bm{\theta}$ if its empirical allocation vector $\bm{p}_{t}$ converges almost surely to the unique component-wise strictly positive probability vector $\bm{p}^{*}=(p^{*}_{1},\dots,p^{*}_{K})$ satisfying:
  1. 1.
    Information balance:
    $\frac{(\theta_{I^{*}}-\theta_{i})^{2}}{\sigma_{I^{*}}^{2}/p^{*}_{I^{*}}+\sigma_{i}^{2}/p^{*}_{i}}=\frac{(\theta_{I^{*}}-\theta_{j})^{2}}{\sigma_{I^{*}}^{2}/p^{*}_{I^{*}}+\sigma_{j}^{2}/p^{*}_{j}},\quad\forall i,j\neq I^{*}.$ (1)
  2. 2.
    Stationarity:
    $p^{*}_{I^{*}}=\sum_{j\neq I^{*}}h^{j}_{I^{*}}(\bm{p}^{*})\cdot\frac{p_{j}^{*}}{1-h^{j}_{I^{*}}(\bm{p}^{*})},\quad\text{where}\quad h^{j}_{I^{*}}(\bm{p}^{*})\triangleq\frac{\sigma_{I^{*}}^{2}/p^{*}_{I^{*}}}{\sigma_{I^{*}}^{2}/p^{*}_{I^{*}}+\sigma_{j}^{2}/p^{*}_{j}},\quad\forall j\neq I^{*}.$ (2)
Remark 1 (Equivalent form of stationarity condition).
Substituting $\bigl{\{}h^{j}_{I^{*}}(\bm{p}^{*})\bigr{\}}_{j\neq I^{*}}$ into the stationarity condition in (2) recovers the conventional balance-of-sum-of-squares condition:
$p^{*}_{I^{*}}=\frac{\sigma^{2}_{I^{*}}}{p^{*}_{I^{*}}}\sum_{j\neq I^{*}}\frac{(p^{*}_{j})^{2}}{\sigma^{2}_{j}}\quad\iff\quad\frac{(p^{*}_{I^{*}})^{2}}{\sigma^{2}_{I^{*}}}=\sum_{j\neq I^{*}}\frac{(p^{*}_{j})^{2}}{\sigma^{2}_{j}}.$ (3)
This condition has been previously observed in the static fixed-budget setting by Glynn and Juneja (2004, Theorem 1) and Chen and Ryzhov (2023, Lemma 1), and in the fixed-confidence setting by Garivier and Kaufmann (2016, Theorem 5) and Bandyopadhyay et al. (2024, Proposition 2.1). Although these two conditions are mathematically equivalent, the stationarity condition in (2) explicitly guides our algorithm design principle; see Section 2.3.
2.2 Top-Two Design Principle as Information Balance Condition
We demonstrate that the information balance condition in (1) naturally leads to the well-known top-two design principle introduced by Russo (2020), where one identifies the two most promising candidates and then carefully selects from these two options. To illustrate this connection, we analyze the top-two Thompson sampling (TTTS) algorithm proposed in their work. This condition also underpins various related “top-two” algorithms (Qin et al. 2017, Jourdan et al. 2022, Bandyopadhyay et al. 2024).
2.2.1 Top-Two Thompson Sampling
The TTTS algorithm starts with uninformative priors and maintains a posterior distribution $\Pi_{t}$ over the true mean vector $\bm{\theta}$. In each iteration $t$, TTTS first generates a sample $\widetilde{\bm{\theta}}=(\widetilde{\theta}_{1},\ldots,\widetilde{\theta}_{K})$ from the posterior distribution $\Pi_{t}$ and identifies a best system under this sample, denoted by $I_{t}^{(1)}\in\operatorname*{argmax}_{i\in[K]}\widetilde{\theta}_{i}$. The system $I_{t}^{(1)}$ is referred to as leader. TTTS then repeatedly samples $\widetilde{\bm{\theta}}^{\prime}=(\widetilde{\theta}^{\prime}_{1},\ldots,\widetilde{\theta}^{\prime}_{K})$ from $\Pi_{t}$ until another system $I_{t}^{(2)}\in\operatorname*{argmax}_{i\in[K]}\widetilde{\theta}^{\prime}_{i}$ that differs from $I_{t}^{(1)}$ is identified. We refer to this system $I_{t}^{(2)}$ as challenger, and the pair $(I_{t}^{(1)},I_{t}^{(2)})$ is referred to as top-two candidates.
Algorithm 1 Top-two Thompson sampling (Russo 2020)
1:Posterior distribution $\Pi_{t}$.
2:Sample $\widetilde{\bm{\theta}}=\bigl{(}\widetilde{\theta}_{1},\ldots,\widetilde{\theta}_{K}\bigr{)}\sim\Pi_{t}$ and define $I_{t}^{(1)}\in\operatorname*{argmax}_{i\in[K]}\widetilde{\theta}_{i}$.
3:Repeatedly sample $\widetilde{\bm{\theta}}^{\prime}=\bigl{(}\widetilde{\theta}^{\prime}_{1},\ldots,\widetilde{\theta}^{\prime}_{K}\bigr{)}\sim\Pi_{t}$ until $I_{t}^{(2)}\in\operatorname*{argmax}_{i\in[K]}\widetilde{\theta}^{\prime}_{i}$ differs from $I_{t}^{(1)}$.
4:return $\bigl{(}I_{t}^{(1)},I_{t}^{(2)}\bigr{)}$.
Given the top-two candidates $(I_{t}^{(1)},I_{t}^{(2)})$, TTTS selects exactly one candidate to measure using a potentially biased coin toss. Specifically, TTTS measures system $I_{t}^{(1)}$ with probability $\beta$ and system $I_{t}^{(2)}$ with probability $1-\beta$, where $\beta$ represents the probability of selecting the leader. A common choice is $\beta=0.5$. The detailed discussion regarding an optimal selection of the coin bias $\beta$ is provided in Section 2.3.
2.2.2 Top-Two Algorithms and Information Balance Condition
The idea of TTTS and other top-two algorithms arises naturally from (1). Recall that the key step in TTTS is sampling repeatedly from the posterior to obtain a challenger. We motivate TTTS by heuristically arguing that a system $j$ appears as a challenger, given that a system $I^{*}$ is the leader, asymptotically only when the following is minimized
$C_{j}(\bm{p})=C_{j}(\bm{p};\bm{\theta})\triangleq\frac{(\theta_{I^{*}}-\theta_{j})^{2}}{2\bigl{(}\sigma_{I^{*}}^{2}/p_{I^{*}}+\sigma_{j}^{2}/p_{j}\bigr{)}},$ (4)
where $\bm{p}=(p_{1},\ldots,p_{K})$ is the proportion of samples allocated among the systems. When the context is clear, we will omit the dependence of $C_{j}(\bm{p};\bm{\theta})$ on the mean vector $\bm{\theta}$.
To see this, suppose $i$ is the leader, then $j$ appears as the challenger only when the posterior sample satisfies $\widetilde{\theta}^{\prime}_{j}>\widetilde{\theta}^{\prime}_{i}$. By Russo (2020, Proposition 5), this occurs under the posterior distribution with an asymptotic probability proportional to $\exp\left(-tC_{j}(\bm{p})\right)$. The challenger is the system with the highest posterior sample, which corresponds to the system with the smallest $\exp\left(-tC_{j}(\bm{p})\right)$ when $t$ is large. Thus, TTTS asymptotically chooses the $j$ with the smallest $C_{j}(\bm{p})$ as the challenger. Since $C_{j}(\bm{p})$ is a monotonically increasing function in $\bm{p}$, TTTS satisfies the information balance condition by maintaining equality among the $C_{j}(\bm{p})$’s.
2.3 Information-Directed Selection among Top-Two via Stationarity Condition
More importantly, we demonstrate how the stationarity condition in (2) allows us to address a significant open problem identified in Russo (2020): achieving exact optimality rather than merely $\beta$-optimality. Specifically, the original TTTS algorithm adopts a non-adaptive tuning parameter $\beta$ to decide which of the top-two candidates to select in each round. An additional tuning step is required to guarantee that $\beta$ converges to the unknown optimal $\beta^{*}$, which depends on the unknown mean vector $\bm{\theta}$.
We propose the information-directed selection (IDS) for top-two algorithms under Gaussian distributions. IDS takes the current sample allocation $\bm{p}_{t}$ and the current top-two candidates $(I_{t}^{(1)},I_{t}^{(2)})$ as input and returns a selected candidate. IDS is hyperparameter-free and guarantees asymptotic optimality.
Algorithm 2 Information-directed selection
1:Sample allocation $\bm{p}$ and top-two candidates $\left(i,j\right)$.
2:return $i$ with probability $h_{i}^{j}(\bm{p})=\frac{\sigma_{i}^{2}/p_{i}}{\sigma_{i}^{2}/p_{i}+\sigma_{j}^{2}/p_{j}}$ and $j$ with probability $h_{j}^{j}(\bm{p})=\frac{\sigma_{j}^{2}/p_{j}}{\sigma_{i}^{2}/p_{i}+\sigma_{j}^{2}/p_{j}}$.
We now argue that IDS naturally arises from the stationarity condition in (2). To see this, note that TTTS selects a system using a two-step procedure: it first identifies a pair of candidates and then measures one from the pair. This procedure naturally provides a law of total probability formula by conditioning on the choice of the challenger. In particular, the left-hand side of (2), representing the proportion of samples allocated to the best system $I^{*}$, can be decomposed into a summation of terms $h^{j}_{I^{*}}(\bm{p}^{*})\cdot\frac{p_{j}^{*}}{1-h^{j}_{I^{*}}(\bm{p}^{*})}$. By IDS, $h^{j}_{I^{*}}(\bm{p}^{*})$ is interpreted as the probability that system $I^{*}$ is sampled, given that $j$ is proposed as the challenger. Let $\mu^{*}_{j}$ denote the probability that $j$ is chosen as the challenger. We argue that
$\mu^{*}_{j}=\frac{p_{j}^{*}}{1-h^{j}_{I^{*}}(\bm{p}^{*})},\quad\forall j\neq I^{*}.$ (5)
Specifically, system $j$ is sampled only if it appears as the challenger (with probability $\mu^{*}_{j}$) and is selected by IDS (with probability $1-h^{j}_{I^{*}}(\bm{p}^{*})$). Therefore, the relationship $p_{j}^{*}=\mu^{*}_{j}(1-h^{j}_{I^{*}}(\bm{p}^{*}))$ must hold. As a sanity check, condition (3) ensures that $\bm{\mu}$ is a proper probability vector: $\mu_{j}^{*}>0$ for all $j\neq I^{*}$ and
$\sum_{j\neq I^{*}}\mu^{*}_{j}=\sum_{j\neq I^{*}}\frac{p^{*}_{j}}{1-h^{j}_{I^{*}}(\bm{p}^{*})}=\sum_{j\neq I^{*}}\frac{p^{*}_{j}}{\frac{\sigma_{j}^{2}/p^{*}_{j}}{\sigma_{i}^{2}/p^{*}_{i}+\sigma_{j}^{2}/p^{*}_{j}}}=\sum_{j\neq I^{*}}p^{*}_{j}+p^{*}_{I^{*}}\cdot\sum_{j\neq I^{*}}\frac{(p^{*}_{j})^{2}/\sigma^{2}_{j}}{(p^{*}_{I^{*}})^{2}/\sigma^{2}_{I^{*}}}=\sum_{i\in[K]}p_{i}=1,$
where the second-to-last inequality follows from (3).
Remark 2 (Selection versus tuning).
Our IDS algorithm fundamentally differs from the standard tuning rules presented in Russo (2020). In their tuning rules, the same tuning parameter is applied irrespective of the proposed top-two candidates, given the history. In contrast, our IDS method explicitly selects between proposed candidates using a candidate-specific selection function $h^{I_{t}^{(2)}}_{I_{t}^{(1)}}(\bm{p}_{t})$, computed based on the information collected from these candidates.
2.4 Optimality of Top-Two Algorithms with Information-Directed Selection
Among the top-two algorithms, the analysis of TTTS (Algorithm 1) is arguably the most complex, as both the leader and the challenger are random variables that depend on the history. Therefore, in this section, we focus on the most challenging TTTS-IDS algorithm. Specifically, we apply Algorithm 2 to select one of the top-two candidates proposed by Algorithm 1. We prove that TTTS-IDS is asymptotically optimal, as defined in Definition 1, by verifying that it satisfies the sufficient conditions established in Theorem 1. This resolves an open problem in Russo (2020, Section 8).
Theorem 2.
Suppose that $\bm{\theta}$ satisfies $\theta_{i}\neq\theta_{j}$ for all $i\neq j$. Then, TTTS-IDS ensures that $\bm{p}_{t}$ converges almost surely to the optimal allocation $\bm{p}^{*}$, which verifies the sufficient condition for optimality in Theorem 1.
This result is a direct consequence of Theorem 5 in Section 3.6, which establishes convergence in a stronger sense than almost sure convergence.
2.5 Discussions
The primary novelty of our algorithm lies in the direct incorporation of dual variables, enabling significant generalizability. Furthermore, our optimality results extend beyond the posterior convergence rate setting.
2.5.1 Novel Optimality Conditions by Incorporating Dual Variables
The literature has explored optimality conditions such as (1) and (3); see, for instance, Glynn and Juneja (2004, Theorem 1) and Garivier and Kaufmann (2016, Theorem 5). Specifically, these results highlight that the unique optimal allocation vector $\bm{p}^{*}$ is indeed the solution to
$\Gamma^{*}_{\bm{\theta}}\triangleq\max_{\bm{p}\in\mathcal{S}_{K}}\min_{j\neq I^{*}}C_{j}(\bm{p}),$ (6)
where $\mathcal{S}_{K}$ denote the probability simplex of dimension $K-1$.
Our approach advocates using the condition (2) rather than the equivalent condition (3), as it goes a step further by directly incorporating dual variables into the optimality condition. The variables $\mu^{*}_{j}$’s defined in (5) turn out to be the optimal dual variables that correspond to the inequality constraints (7) in the following reformulation of the optimal allocation problem (6):
$\Gamma^{*}_{\bm{\theta}}=\max_{\phi,\bm{p}\in\mathcal{S}_{K}}$ $\hskip 10.0pt\phi,$
$\mathrm{s.t.}$ $\hskip 10.0pt\phi-C_{j}(\bm{p})\leq 0,\quad\forall j\neq I^{*}.$ (7)
The conditions in Theorem 1 are derived directly from the Karush-Kuhn-Tucker (KKT) conditions, where the information balance condition (1) corresponds to the complementary slackness condition, and (2) represents the stationarity condition.
The novelty of incorporating dual variables lies in their significant potential for generalization, as we shall demonstrate in Section 3. This is particularly important because generalizing existing approaches (e.g., Chen and Ryzhov 2023, Bandyopadhyay et al. 2024) for general pure-exploration problems is challenging. For example, Zhang et al. (2023, Table 1) notes that the information balance conditions for identifying the best-$k$ alternatives involve a complex combinatorial structure, making the explicit generalization of (1) nearly infeasible. You et al. (2023) further illustrates these complexities through a case study of selecting the best two among five alternatives. However, by leveraging dual variables, we can circumvent these combinatorial difficulties by implicitly enforcing information balance through the complementary slackness condi‘tion:
$\mu^{*}_{j}\cdot\Bigl{(}\min_{j\neq I^{*}}C_{j}(\bm{p}^{*})-C_{j}(\bm{p}^{*})\Bigr{)}=0,\quad\forall j\neq I^{*}.$
This condition applies to a broad range of problems beyond the canonical R&S; see Theorem 4.
2.5.2 Optimality in Alternative Performance Criteria
The same optimal allocation $\bm{p}^{*}$ arises in other performance criteria studied in the literature, and thus our algorithm is also optimal in such settings.
Static-allocation fixed-budget setting.
Consider the frequentist PCS, defined as $\mathbb{P}_{\bm{\theta}}\bigl{(}\widehat{I}_{T}^{*}=I^{*}\bigr{)}$ for a fixed budget $T$, where the probability is taken over the randomness in both observations and the algorithm and $\widehat{I}_{T}^{*}$ is the output of the algorithm at the end of horizon. Due to the dynamic nature of typical algorithms—where the decision at later stages depends on earlier observations—analyzing PCS in the fixed-budget setting is challenging (Hong et al. 2021, Komiyama 2024). To address this, a common framework in the simulation literature, initiated by Glynn and Juneja (2004) and followed by Chen and Ryzhov (2023), focuses on PCS under a static algorithm. These static algorithms allocate samples proportionally to a fixed probability vector $\bm{p}\in\mathcal{S}_{K}$. The framework characterizes the large deviation rate of the probability of incorrect selection under a static algorithm as follows: $-\lim_{T\to\infty}\frac{1}{T}\log\bigl{(}1-\mathbb{P}_{\bm{p},\bm{\theta}}(\widehat{I}_{T}^{*}=I^{*})\bigr{)}=\widetilde{\Gamma}_{\bm{\theta}}(\bm{p})$, for some large deviation rate function $\widetilde{\Gamma}_{\bm{\theta}}(\bm{p})$, where $\mathbb{P}_{\bm{p},\bm{\theta}}$ denotes the probability under system parameter $\bm{\theta}$ and static allocation $\bm{p}$. For Gaussian models, the large deviation rate function $\widetilde{\Gamma}_{\bm{\theta}}(\bm{p})=\Gamma_{\bm{\theta}}(\bm{p})$. Glynn and Juneja (2004) proceeds to maximize the large deviation rate, leading to the same optimal allocation problem as in (6). [^2] For general distributions beyond Gaussian, the expression of $C_{j}$ in the optimal allocation for static fixed-budget and posterior convergence rate settings differs. However, the maximin structure remains unchanged, and our analysis extends to these cases. A common goal of fixed-budget algorithms is to ensure that the dynamic allocation converges almost surely to the optimal static allocation $\bm{p}^{*}$. Our algorithm achieves this goal, as guaranteed by Theorem 2.
Fixed-confidence setting.
The fixed-confidence formulation for BAI, which leads to the same optimal allocation $\bm{p}^{*}$, has been studied in Garivier and Kaufmann (2016, Lemma 4 and Theorem 5). In Section 3, we extend this formulation to general pure-exploration problems and prove that our algorithm achieves optimality for R&S in Gaussian models under the fixed-confidence setting.
3 General Pure-Exploration Problems
We now broaden our focus from traditional R&S problems to address a wider range of pure-exploration queries relevant to decision-making. To align with the common terminology in machine learning literature on pure exploration, we adopt terms such as “arm” instead of “system” and “reward” instead of “performance.” We will show that many formulations commonly studied in simulation, such as the PGS and the indifference-zone (IZ) approach, have clear counterparts within the pure-exploration framework.
We consider a multi-armed bandit problem with $K$ arms, denoted by $i\in[K]$. The rewards from arm $i$ are independent and identically distributed random variables following a distribution $F_{\theta_{i}}$, parameterized by the mean $\theta_{i}$ and belonging to a single-parameter exponential family. We assume that the rewards from different arms are mutually independent. A problem instance is defined by a fixed but unknown vector $\bm{\theta}\in\mathbb{R}^{K}$. The DM’s goal is to answer an exploration query regarding $\bm{\theta}$ by adaptively allocating the sampling budget across the arms and observing noisy rewards.
Exploration query.
An exploration query specifies a question regarding the unknown mean vector $\bm{\theta}$. We assume that the query has a unique answer, denoted by $\mathcal{I}(\bm{\theta})$. [^3] Exploration queries with multiple correct answers require more care, see Section 4.1 for an example. Several notable exploration queries have been explored in the literature. For the canonical R&S or the BAI problem (Section 2), the query identifies the unique best arm $\mathcal{I}({\bm{\theta}})=\operatorname*{argmax}_{i\in[K]}\theta_{i}$. Alternatively, one might seek to identify the exact set of best-$k$ arms, a problem referred to as best-$k$ identification, where $\mathcal{I}({\bm{\theta}})=\operatorname*{argmax}_{\mathcal{K}\subseteq[K],|\mathcal{K}|=k}\left\{\sum_{i\in\mathcal{K}}\theta_{i}\right\}$. Other exploration queries include R&S under the indifference-zone (IZ) assumption (Bechhofer 1954), or identifying all $\varepsilon$-good arms (Mason et al. 2020), where $\mathcal{I}(\bm{\theta})=\{i\in[K]:\theta_{i}\geq\max_{i\in[K]}\theta_{i}-\varepsilon\}$. Another example is the pure-exploration thresholding bandit problem (Lai and Liao 2012, Locatelli et al. 2016), where the goal is to identify all arms with mean rewards exceeding a threshold $\mathfrak{T}$, represented by $\mathcal{I}({\bm{\theta}})=\{i\in[K]:\theta_{i}>\mathfrak{T}\}$.
Fixed-confidence performance criteria.
We focus on the fixed-confidence setting, which aims to identify the correct answer with a PCS guarantee of at least $1-\delta$ for a fixed $\delta\in(0,1)$, [^4] In simulation literature, the notation $\delta$ commonly represents the indifference-zone parameter. In contrast, our work considers an IZ-free formulation, and $1-\delta$ denotes the target PCS level. while minimizing the sampling budget. In addition to the allocation rule discussed in Section 2, an algorithm requires both stopping and decision rules. At each time step $t\in\mathbb{N}_{0}$, the DM first determines whether to stop according to a stopping rule $\tau_{\delta}$. If stopping occurs at $t=\tau_{\delta}$, the DM provides an estimate ^$\widehat{\mathcal{I}}_{\tau_{\delta}}$ of the answer $\mathcal{I}(\bm{\theta})$ based on a decision rule. If not, she selects an arm $I_{t}$ according to an allocation rule and observes the reward from arm $I_{t}$, denoted by $Y_{t+1,I_{t}}$. This procedure is summarized in Algorithm 3 where the information that the DM has at the beginning of time $t$ is denoted by $\mathcal{H}_{t}=(I_{0},Y_{1,I_{0}},\dots,I_{t-1},Y_{t,I_{t-1}})$.
Algorithm 3 A template for anytime pure-exploration algorithms
1:Stopping rule (stop), decision rule (decision), allocation rule (allocate), and $\mathcal{H}_{0}=\varnothing$.
2:for $t=0,1,\dots$ do
3:     if $\textup{{stop}}(\mathcal{H}_{t})$ then
4:         return ^$\widehat{\mathcal{I}}=\textup{{decision}}(\mathcal{H}_{t})$.
5:     else
6:         Run allocation rule $I_{t}=\textup{{allocate}}(\mathcal{H}_{t})$.
7:         Select arm $I_{t}$, observe reward $Y_{t+1,I_{t}}$, and update history $\mathcal{H}_{t+1}=\mathcal{H}_{t}\cup\{I_{t},Y_{t+1,I_{t}}\}$.
8:     end if
9:end for
We consider problem instances that satisfy a mild assumption. Let $\overline{\Theta}$ denote parameters yielding a unique answer to the exploration query, and $\mathfrak{I}$ be the set of all possible correct answers. For each $\mathcal{I}\in\mathfrak{I}$, let $\Theta_{\mathcal{I}}$ denote the set of parameters with $\mathcal{I}$ as the correct answer, i.e., $\Theta_{\mathcal{I}}\triangleq\{\bm{\vartheta}\in\overline{\Theta}:\mathcal{I}(\bm{\vartheta})=\mathcal{I}\}.$ We assume that each $\Theta_{\mathcal{I}}$ is an open set. This assumption rules out statistically unidentifiable scenarios. Our focus is on the parameter space $\Theta\triangleq\cup_{\mathcal{I}\in\mathfrak{I}}\Theta_{\mathcal{I}}$. We note that this assumption is generally unrestrictive; for instance, in R&S, we have $\overline{\Theta}=\Theta$, i.e., we consider all problem instances with a unique solution.
Definition 2 ($\delta$-correct).
A policy is said to be $\delta$-correct if
$\mathbb{P}_{\bm{\theta}}\bigl{(}\tau_{\delta}<\infty,\widehat{\mathcal{I}}_{\tau_{\delta}}=\mathcal{I}(\bm{\theta})\bigr{)}\geq 1-\delta,\quad\hbox{for any }\bm{\theta}\in\Theta.$
Let $\mathcal{A}$ denote the class of policies that are $\delta$-correct for any $\delta>0$. The goal is to design a policy $\pi$ that minimizes the expected sample complexity, $\mathbb{E}^{\pi}_{\bm{\theta}}[\tau_{\delta}]$ over $\mathcal{A}$, where $\tau_{\delta}\triangleq\min\{t\geq 0:\textup{{stop}}(\mathcal{H}_{t})=\textup{{True}}\}$ denotes the stopping time when the policy halts. A policy is deemed optimal in the fixed-confidence setting if it achieves universal efficiency, as defined below.
Definition 3 (Universal efficiency in fixed confidence).
A policy $\pi^{*}$ is universally efficient if
  1. 1.
    $\pi^{*}$ is $\delta$-correct (i.e., $\pi^{*}\in\mathcal{A}$), and
  2. 2.
    for any other $\pi\in\mathcal{A}$, we have
    $\limsup_{\delta\to 0}\frac{\mathbb{E}^{\pi^{*}}_{\bm{\theta}}\![\tau_{\delta}]}{\mathbb{E}^{\pi}_{\bm{\theta}}[\tau_{\delta}]}\leq 1,\quad\hbox{for any }\bm{\theta}\in\Theta.$
Compared with the “pointwise” optimality in Definition 1, Definition 3 is stronger, as it requires optimality to hold universally for all $\bm{\theta}$ under a $\delta$-correctness constraint for all $\delta$ and $\bm{\theta}\in\Theta$.
3.1 Preliminary
We now introduce concepts and notation that lay the foundation of our algorithm design principle.
The pitfalls.
For an algorithm to correctly answer a pure-exploration query, it must distinguish between problem instances that yield different answers. For a given problem instance $\bm{\theta}\in\Theta$, the intrinsic complexity of the pure-exploration query in the fixed-confidence setting is closely related to the alternative set, defined as the set of parameters that yield an answer different from $\mathcal{I}(\bm{\theta})$:
$\mathrm{Alt}(\bm{\theta})\triangleq\bigl{\{}\bm{\vartheta}\in\Theta:\mathcal{I}(\bm{\vartheta})\neq\mathcal{I}(\bm{\theta})\bigr{\}}.$
To formulate a maximin problem as in (6), we rely on the following assumption.
Assumption 1 (Wang et al. 2021).
For any fixed problem instance $\bm{\theta}$, the alternative set $\mathrm{Alt}(\bm{\theta})$ is a finite union of convex sets. Namely, there exists a finite set $\mathcal{X}(\bm{\theta})$ and a collection of convex sets $\{\mathrm{Alt}_{x}(\bm{\theta}):x\in\mathcal{X}(\bm{\theta})\}$ such that $\mathrm{Alt}(\bm{\theta})=\cup_{x\in\mathcal{X}(\bm{\theta})}\mathrm{Alt}_{x}(\bm{\theta})$.
This condition is indeed satisfied by commonly encountered exploration queries; see Appendix A. We refer to the collection $\mathcal{X}(\bm{\theta})$ as the set of pitfalls. These pitfalls represent the fundamental causes that lead an alternative problem instance $\bm{\vartheta}$ to yield an answer different from $\mathcal{I}(\bm{\theta})$. Thus, the pitfall set is determined by both the exploration query and the problem instance $\bm{\theta}$.
Example 1 (BAI).
For BAI, the pitfall set is defined as $\mathcal{X}(\bm{\theta})=[K]\backslash\{I^{*}(\bm{\theta})\}$, where each $j\in\mathcal{X}(\bm{\theta})$ represents a scenario in which, under an alternative problem instance $\bm{\vartheta}\in\Theta$, system $j$ outperforms the best system $I^{*}(\bm{\theta})$ identified under $\bm{\theta}$. In such scenarios, $I^{*}(\bm{\theta})$ is an incorrect selection due to the existence of the pitfall $j$. The corresponding alternative set for each pitfall $j\in\mathcal{X}(\bm{\theta})$ is $\mathrm{Alt}_{j}(\bm{\theta})=\bigl{\{}\bm{\vartheta}\in\Theta:\vartheta_{j}\geq\vartheta_{I^{*}(\bm{\theta})}\bigr{\}}$.
Example 2 (Best-$k$ identification).
The goal is to identify the set of $k$ arms with the largest means. The exploration query can be broken down into verifying whether any top arm $i\in\mathcal{I}(\bm{\theta})$ has a higher mean than any bottom arm $j\in\mathcal{I}^{c}(\bm{\theta})$. If there exist pairs $(i,j)\in\mathcal{X}(\bm{\theta})\triangleq\mathcal{I}(\bm{\theta})\times\mathcal{I}^{c}(\bm{\theta})$ such that $\vartheta_{j}>\vartheta_{i}$, then either $i\notin\mathcal{I}(\bm{\vartheta})$ or $j\in\mathcal{I}(\bm{\vartheta})$, leading to $\mathcal{I}(\bm{\vartheta})\neq\mathcal{I}(\bm{\theta})$. A pitfall in this context is specified by a pair of arms $(i,j)\in\mathcal{X}(\bm{\theta})$, with the alternative set $\mathrm{Alt}_{i,j}(\bm{\theta})=\{\bm{\vartheta}\in\Theta:\vartheta_{j}>\vartheta_{i}\}$.
Example 3 (Pure-exploration thresholding bandits).
The goal is to identify the set of arms with means above a threshold $\mathfrak{T}$, i.e., $\mathcal{I}(\bm{\theta})\triangleq\{i\in[K]:\theta_{i}>\mathfrak{T}\}$. The exploration query decomposes into sub-queries comparing each arm $i\in[K]$ with the threshold $\mathfrak{T}$. A pitfall is specified by an arm $i\in\mathcal{X}(\bm{\theta})\triangleq[K]$ with the corresponding alternative set $\mathrm{Alt}_{i}(\bm{\theta})=\{\bm{\vartheta}\in\Theta:(\mathfrak{T}-\vartheta_{i})(\mathfrak{T}-\theta_{i})<0\}$. This indicates that arm $i$ is on opposite sides of $\mathfrak{T}$ under $\bm{\vartheta}$ and $\bm{\theta}$, resulting in an incorrect answer.
The generalized Chernoff information $C_{x}$.
Let $\theta_{t,i}\triangleq\frac{1}{N_{t,i}}\sum_{\ell=0}^{t-1}{\bm{1}(I_{\ell}=i)Y_{\ell+1,I_{\ell}}}$ denotes the sample mean of arm $i$ at time $t$ and $\bm{\theta}_{t}=(\theta_{t,1},\dots,\theta_{t,K})$. To correctly answer the pure-exploration query, we must overcome every potential pitfall. From a hypothesis testing perspective, this involves testing the hypotheses $H_{0,x}:\bm{\theta}\in\mathrm{Alt}_{x}(\bm{\theta}_{t})$ versus $H_{1,x}:\bm{\theta}\notin\mathrm{Alt}_{x}(\bm{\theta}_{t})$, for all $x\in\mathcal{X}(\bm{\theta}_{t})$, where $H_{1,x}$ corresponds to the scenario in which we overcome pitfall $x$. For instance, in R&S, rejecting $H_{0,x}$ indicates sufficient evidence to support the desired ordering $\theta_{I^{*}(\bm{\theta}_{t})}>\theta_{x}$ for any $x\neq I^{*}(\bm{\theta}_{t})$.
To this end, we utilize the generalized log-likelihood ratio test (GLRT) statistic, as proposed in Kaufmann and Koolen (2021). Let $\mathcal{S}_{K}$ denote the probability simplex of dimension $K-1$ and let $\bm{p}\in\mathcal{S}_{K}$ represent a generic allocation vector, indicating the proportion of samples allocated to the $K$ arms. If an algorithm allocates according to $\bm{p}_{t}$ at time $t$, then the GLRT statistic given by
$\log\frac{\sup_{\bm{\vartheta}\in\Theta}L(\bm{\vartheta})}{\sup_{\bm{\vartheta}\in\mathrm{Alt}_{x}(\bm{\theta}_{t})}L(\bm{\vartheta})}=\inf_{\bm{\vartheta}\in\mathrm{Alt}_{x}(\bm{\theta}_{t})}\log\frac{L(\bm{\theta}_{t})}{L(\bm{\vartheta})}=t\cdot\inf_{\bm{\vartheta}\in\mathrm{Alt}_{x}(\bm{\theta}_{t})}\sum_{i\in[K]}p_{i}d(\theta_{t,i},\vartheta_{i}),$
where $d(\cdot,\cdot)$ is the Kullback-Leibler (KL) divergence and we use the fact that the unconstrained maximum likelihood estimator for the mean of a single-parameter exponential family is $\bm{\theta}_{t}$.
We extend $C_{j}$ in (4) as the population version of the GLRT statistic normalized by the sample size $t$: for all $x\in\mathcal{X}(\bm{\theta})$,
$C_{x}(\bm{p})=C_{x}(\bm{p};\bm{\theta})\triangleq\inf_{\bm{\vartheta}\in\mathrm{Alt}_{x}(\bm{\theta})}\sum_{i\in[K]}p_{i}d(\theta_{i},\vartheta_{i}).$ (8)
We refer to $C_{x}$ as the generalized Chernoff information, extending a classical concept in hypothesis testing (Cover and Thomas 2006). When the context is clear, we will omit the dependence of $C_{x}(\bm{p};\bm{\theta})$ on the problem instance $\bm{\theta}$. The following result establishes the basic properties of $C_{x}$.
Lemma 1.
Under Assumption 1, for each $x\in\mathcal{X}$ and component-wise strictly positive $\bm{p}\in\mathcal{S}_{K}$,
  1. 1.
    There exists a unique $\bm{\vartheta}^{x}\in\mathrm{Alt}_{x}(\bm{\theta})$ that achieves the infimum in (8).
  2. 2.
    The functions $C_{x}(\bm{p})$ are differentiable and concave, with $\frac{\partial C_{x}(\bm{p})}{\partial p_{i}}=d(\theta_{i},\vartheta^{x}_{i})$ for all $i\in[K].$
  3. 3.
    The functions $C_{x}(\bm{p})$ satisfy the following partial differential equation
    $C_{x}(\bm{p})=\sum_{i\in[K]}p_{i}\frac{\partial C_{x}(\bm{p})}{\partial p_{i}}.$ (9)
When the sampling effort is distributed according to $\bm{p}$, the function $C_{x}$ quantifies the average information gain per sample needed to assert $H_{1,x}$, i.e., to overcome pitfall $x$. The partial differential equation (PDE) in (9) shows that this information gain can be effectively decomposed into contributions from the sample allocations to individual arms; see the discussion in Remark 5.
The optimal allocation problem.
Recall that $H_{0,x}$ must be rejected for all pitfalls $x\in\mathcal{X}(\bm{\theta})$ to correctly answer the pure-exploration query. Furthermore, $C_{x}(\bm{p})$ represents the information gathered to reject $H_{0,x}$ under sample allocation $\bm{p}$—the larger $C_{x}(\bm{p})$ is, the easier it is to reject $H_{0,x}$. By taking the minimum over all pitfalls $x\in\mathcal{X}(\bm{\theta})$, we define
$\Gamma_{\bm{\theta}}(\bm{p})\triangleq\min_{x\in\mathcal{X}(\bm{\theta})}C_{x}(\bm{p}),$
which corresponds to the GLRT statistic for the principal pitfall, i.e., the pitfall that is hardest to overcome. To maximize information gathering, we should follow an allocation $\bm{p}$ that maximizes $\Gamma_{\bm{\theta}}(\bm{p})$. This leads to the following optimal allocation problem:
$\Gamma^{*}_{\bm{\theta}}\triangleq\max_{\bm{p}\in\mathcal{S}_{K}}\Gamma_{\bm{\theta}}(\bm{p})=\max_{\bm{p}\in\mathcal{S}_{K}}\min_{x\in\mathcal{X}(\bm{\theta})}C_{x}(\bm{p}),$ (10)
which generalizes (6) in the context of R&S problems in Gaussian models.
The stopping and decision rules.
A fixed-confidence algorithm requires three components: stopping, decision, and allocation rules. We follow standard practices in the literature for the stopping and decision rules, allowing us to focus on designing optimal allocation rules.
For the stopping rule, Kaufmann and Koolen (2021) provide the Chernoff stopping rule based on the GLRT for exponential family bandits:
$\tau_{\delta}=\inf\bigl{\{}t\in\mathbb{N}_{0}:t\cdot\Gamma_{\bm{\theta}_{t}}(\bm{p}_{t})>\gamma(t,\delta)\bigr{\}},$ (11)
where $\gamma(t,\delta)$ is a threshold function. Kaufmann and Koolen (2021, Proposition 15) provide a $\gamma(t,\delta)$ ensuring that, combined with any allocation rule, the stopping rule in (11) is $\delta$-correct for any pure-exploration query with a unique correct answer. For completeness, we rephrase their results in Appendix C.1.
We adopt a specific decision rule
$\textup{{decision}}(\mathcal{H_{\tau_{\delta}}})=\mathcal{I}(\bm{\theta}_{\tau_{\delta}}),$ (12)
where $\mathcal{I}(\bm{\theta}_{\tau_{\delta}})$ is the correct answer based on the sample mean $\bm{\theta}_{\tau_{\delta}}$ upon stopping. [^5] In cases where multiple correct answers exist under $\bm{\theta}_{\tau_{\delta}}$, ties are broken arbitrarily. For simplicity, we abuse the definition $\mathcal{I}(\bm{\theta}_{\tau_{\delta}})$ to represent one of the correct answers.
3.2 A Sufficient Condition of Allocation Rules for Fixed-Confidence Optimality
We now present a sufficient condition for an allocation rule to achieve universal efficiency in the fixed-confidence setting. This condition requires the allocation of sampling effort to converge “rapidly” to the optimal solution of (10). We refer to this mode of convergence as $\mathcal{L}$-convergence, which is stronger than almost sure convergence. Specifically, $\mathcal{L}$-convergence ensures that for any fixed neighborhood of the target, the sequence remains within that neighborhood indefinitely after a random time with finite expectation. Let the space $\mathcal{L}$ consist of all real-valued random variables $T$ with $\|T\|_{1}<\infty$ where $\|T\|_{1}\triangleq\mathbb{E}\left[|T|\right]$.
Definition 4 ($\mathcal{L}$-convergence).
For a sequence of real-valued random variables $\{X_{t}\}_{t\in\mathbb{N}_{0}}$ and a scalar $x\in\mathbb{R}$, we say $X_{t}\xrightarrow{\mathcal{L}}x$ if
for $\text{for all }\epsilon>0,\text{ there exists }T\in\mathcal{L}\text{ such that for all }t\geq T,|X_{t}-x|\leq\epsilon.$
For a sequence of random vectors $\{\bm{X}_{t}\}_{t\in\mathbb{N}_{0}}$ taking values in $\mathbb{R}^{d}$ and a vector $\bm{x}\in\mathbb{R}^{d}$ where $\bm{X}_{t}=(X_{t,1},\ldots,X_{t,d})$ and $\bm{x}=(x_{1},\ldots,x_{d})$, we say $\bm{X}_{t}\xrightarrow{\mathcal{L}}\bm{x}$ if $X_{t,i}\xrightarrow{\mathcal{L}}x_{i}$ for all $i\in[d]$.
It is straightforward to observe that $\mathcal{L}$-convergence implies almost sure convergence.
Lemma 2.
If $\bm{X}_{t}\xrightarrow{\mathcal{L}}\bm{x}$, then $\bm{X}_{t}\xrightarrow{\mathrm{a.s.}}\bm{x}$.
Universal efficiency requires controlling the expectation of the stopping time $\tau_{\delta}$ in (11). This is achieved by establishing a random time $T_{\varepsilon}\in\mathcal{L}$ such that $|\Gamma_{\bm{\theta}_{t}}(\bm{p}_{t})-\Gamma^{*}|<\epsilon$ for all $t>T_{\epsilon}$. Since $T_{\varepsilon}\in\mathcal{L}$, the expectation $\mathbb{E}^{\pi}_{\bm{\theta}}[\tau_{\delta}]$ is well-controlled. Notably, $\Gamma_{\bm{\theta}_{t}}(\bm{p}_{t})\xrightarrow{\mathcal{L}}\Gamma^{*}$ is implied by $\bm{p}_{t}\xrightarrow{\mathcal{L}}\bm{p}^{*}$. We formalize this sufficient condition below, with the proof deferred to Appendix C. This condition generalizes the result in Qin et al. (2017), which was specifically tailored to BAI.
Theorem 3 (A sufficient condition for optimality).
Suppose that for any $\bm{\theta}\in\Theta$, the optimal allocation problem (10) admits a unique, component-wise strictly positive optimal allocation $\bm{p}^{*}$. A policy $\pi$ is universally efficient (Definition 3) if the following sufficient conditions are satisfied:
  1. 1.
    the stopping rule (11) and decision rule (12) are used, and
  2. 2.
    its allocation rule satisfies that
    $\quad\bm{p}_{t}\xrightarrow{\mathcal{L}}\bm{p}^{*},\quad\text{for all }\bm{\theta}\in\Theta.$
In particular,
$\quad\lim_{\delta\to 0}\frac{\mathbb{E}^{\pi}_{\bm{\theta}}[\tau_{\delta}]}{\log(1/\delta)}=(\Gamma^{*}_{\bm{\theta}})^{-1},\quad\text{for all }\bm{\theta}\in\Theta.$
Remark 3 (Uniqueness and positivity of optimal solution).
For BAI, Gaussian $\varepsilon$-BAI, pure-exploration thresholding bandits, and best-$k$ identification with single-parameter exponential family distributions, the existence of a unique and componentwise strictly positive optimal solution $\bm{p}^{*}(\bm{\theta})>0$ for any $\bm{\theta}\in\Theta$ is established in Garivier and Kaufmann (2016), Jourdan et al. (2023b), Ménard (2019), and You et al. (2023), respectively.
Remark 4 (Other performance criteria).
We discussed two alternative performance criteria in Theorem 1 and Section 2.5.2, both focusing on the almost sure convergence of the empirical allocation $\bm{p}_{t}\xrightarrow{\mathrm{a.s.}}\bm{p}^{*}$. For the posterior convergence rate setting (Russo 2020), the optimal allocation aligns with (10), even for other single-parameter exponential family distributions. For the static-allocation fixed-budget setting (Glynn and Juneja 2004) under Gaussian distributions, the optimal allocation also coincides with (10) due to the symmetry of the KL divergence. For other distributions, swapping the parameters in the KL divergence in (8) is necessary, but the properties of $C_{x}$ in (9) and the maximin structure of (10) remain valid, leading to a similar sufficient condition. Our sufficient condition in Theorem 3, which requires the $\mathcal{L}$-convergence of the sample allocation, is stronger than almost sure convergence. Consequently, this condition also implies optimality in the posterior convergence rate and the large deviations rate of static allocations.
3.3 Optimality Conditions for the Optimal Allocation Problem
We define selection functions that generalize $h^{j}_{i}(\bm{p})$ in Algorithm 2 to general pure-exploration problems: for $\ x\in\mathcal{X}(\bm{\theta})\ \hbox{and}\ i\in[K]$,
$h_{i}^{x}(\bm{p})=h_{i}^{x}(\bm{p};\bm{\theta})\triangleq\frac{p_{i}\frac{\partial C_{x}(\bm{p})}{\partial p_{i}}}{C_{x}(\bm{p})}=\frac{p_{i}d(\theta_{i},\vartheta^{x}_{i})}{\sum_{k\in[K]}p_{k}d(\theta_{k},\vartheta^{x}_{k})}.$ (13)
Following from the PDE characterization of $C_{x}$ in Lemma 1, $\bm{h}^{x}(\bm{p})=\{h_{i}^{x}(\bm{p}):i\in[K]\}\in\mathcal{S}_{K}$ forms a valid probability vector for any $x\in\mathcal{X}(\bm{\theta})$. In fact, $\bm{h}^{x}(\bm{p})$ can be interpreted as the probability distribution over the arms to be sampled when $x$ is identified as the principal pitfall.
We derive the following optimality conditions for (10) in the cases where the optimal solution $\bm{p}^{*}$ is component-wise strictly positive. These conditions extend those presented in Theorem 1. The detailed proof can be found in Appendix D.
Theorem 4.
Assume that $\bm{p}^{*}$ is component-wise strictly positive. A feasible solution $(\phi,\bm{p})$ to (10) is optimal if and only if $\phi=\min_{x\in\mathcal{X}(\bm{\theta})}C_{x}(\bm{p})$ and there exist dual variables $\bm{\mu}\in\mathcal{S}_{|\mathcal{X}(\bm{\theta})|}$ such that
$p_{i}=\sum_{x\in\mathcal{X}(\bm{\theta})}\mu_{x}h_{i}^{x},\quad\forall i\in[K],$ (14a)
$\mu_{x}(\phi-C_{x}(\bm{p}))=0,\quad\forall x\in\mathcal{X}(\bm{\theta}).$ (14b)
The condition (14) is an equivalent form of the KKT condition for optimality, where (14a) is the stationarity condition and (14b) is the complementary slackness condition.
The following result extends the information balance for BAI in (1). It is a direct corollary of (14) and the assumption that $\bm{p}^{*}$ is component-wise strictly positive. We defer the proof to Appendix G.4.
Corollary 1 (Crude information balance).
Assume that $\bm{p}^{*}$ is component-wise strictly positive, then
$\Gamma_{\bm{\theta}}^{*}=\min_{x\in\mathcal{X}_{i}}C_{x}(\bm{p}^{*}),\quad\hbox{for all }i\in[K],\quad\text{where}\quad\mathcal{X}_{i}\triangleq\Bigl{\{}x\in\mathcal{X}(\bm{\theta}):\frac{\partial C_{x}(\bm{p})}{\partial p_{i}}>0\Bigr{\}}.$
Surprisingly, this information balance condition proves too crude to guarantee optimality in pure-exploration problems beyond BAI. Even when combined with the correct overall balance condition, it may fail; see You et al. (2023, Example 4) for a counterexample. To address this challenge, we introduce dual variables that implicitly encode a more refined information balance.
3.4 Dual-Directed Algorithm Design
Based on Theorem 4, an allocation rule can be immediately derived by tracking the two sets of conditions in (14). Our general algorithmic design framework follows the same idea as that discussed in Section 2.2 and Section 2.3. In particular, our algorithm takes two steps: it first identifies a principal pitfall; and then selects an arm that is guided by the information gain relevant to the detected principal pitfall.
We first discuss how the principal pitfall is detected. Recall that the smaller the value of $C_{x}$, the less information we have to rule out $x$ as the root cause of a potential incorrect answer. Additionally, the dual variable $\mu_{x}$ can be interpreted as the proportion of time that $x$ is detected as the principal pitfall. In our algorithm, the pitfall $x$ with the smallest (empirical estimation of) $C_{x}$ is detected as the principal pitfall $x_{t}$. This step is guided by the complementary slackness conditions in (14b), which indicate that only the pitfall $x$ with the smallest $C_{x}$ has $\mu_{x}>0$ and therefore requires attention.
Once $x_{t}$ has been detected as the principal pitfall, the algorithm employs an information-directed selection (IDS) rule to select an arm for sampling, specified by $\bm{h}^{x_{t}}$ defined in (13). This procedure, formalized in Algorithm 4, extends the IDS rule for BAI (Algorithm 2) to general pure-exploration problems. The IDS rule is designed to depend on the detected principal pitfall $x_{t}$, aiming to guide the average information gain required to overcome the principal pitfall. In particular, this process satisfies the stationarity conditions in (14a), where the sampling probability of arm $i$ is decomposed into a law-of-total-probability formula. Here, $h^{x}_{i}$ specifies the probability of sampling $i$ given that $x$ is detected as the principal pitfall.
Algorithm 4 Information-directed selection for general pure exploration
1:Sample allocation $\bm{p}$, estimate $\widehat{\bm{\theta}}$ and the detected principal pitfall $x$.
2:return $I\in[K]$ sampled from the distribution $\bm{h}^{x}(\bm{p};\widehat{\bm{\theta}})$.
Remark 5 (Information-directed selection).
$C_{x}(\bm{p})$ is defined as the population version of the GLRT statistic, normalized by the sample size $t$, for testing $H_{0,x}:\bm{\theta}\in\mathrm{Alt}_{x}(\bm{\theta}_{t})$ versus $H_{1,x}:\bm{\theta}\notin\mathrm{Alt}_{x}(\bm{\theta}_{t})$. Thus, it quantifies the information gathered per unit sample under the allocation vector $\bm{p}$ to determine whether the pitfall $x$ is responsible for an incorrect answer. The PDE characterization of $C_{x}(\bm{p})$ in (9) reveals a decomposition of this information. To interpret, $\frac{\partial C_{x}(\bm{p})}{\partial p_{i}}$ represents the information gain per unit allocation to arm $i$, while $p_{i}\frac{\partial C_{x}(\bm{p})}{\partial p_{i}}$ quantifies the information gain per unit sample from allocating a proportion $p_{i}$ of samples to arm $i$. The selection function $h^{x}_{i}(\bm{p})=\frac{p_{i}\frac{\partial C_{x}(\bm{p})}{\partial p_{i}}}{C_{x}(\bm{p})}$ suggests that sampling should be done proportionally to the information gain contributed by each arm when determining the allocation among the candidates.
Moreover, the selection functions $h^{x}_{i}(\bm{p})$ identify a set of candidates whose allocation probabilities are positive for a given pitfall $x$ under allocation $\bm{p}$.
Definition 5 (Active candidate set).
The active candidate set for a pitfall $x$ under allocation $\bm{p}$ is
$\mathfrak{C}_{x}\triangleq\Bigl{\{}i\in[K]:\frac{\partial C_{x}(\bm{p})}{\partial p_{i}}>0\Bigr{\}}.$
The active candidate set is essential, as only arms in $\mathfrak{C}_{x}$ contribute to the information gain that is critical for overcoming pitfall $x$. For the specific case of best-arm identification, the cardinality of $\mathfrak{C}_{x}$ is always two, leading to the well-known class of top-“two” algorithms. However, for general pure-exploration tasks, the cardinality of the candidate set may depend on the pitfall $x$ and can differ from two; see examples in Appendix A.
Remark 6 (The role of the dual variables).
The novelty of our algorithm design framework lies in the incorporation of dual variables. Existing constant-optimal algorithms typically fall into two categories. The first solves the optimal allocation problem using plug-in estimators at each iteration and tracks the resulting solutions, as pioneered by Garivier and Kaufmann (2016) and extended in works such as Kato and Ariu (2024), Al Marjani et al. (2022). The second tracks optimality conditions solely based on the allocation vector $\bm{p}$ (e.g., (1) and (2)), as demonstrated in Chen and Ryzhov (2023), Avci et al. (2023), Bandyopadhyay et al. (2024). Our approach is more fundamental and flexible due to the introduction of dual variables, which encode the information balance structure described in Corollary 1. This enables us to avoid explicit characterization of the optimal information balance structure, which is often impractical without knowledge of the true parameter $\bm{\theta}$ (Zhang et al. 2023). Furthermore, dual variables facilitate a law-of-total-probability-type stationarity condition (14a), naturally leading to the IDS rule that achieves optimality across various pure-exploration tasks.
3.5 A Modular Template for Proposed Algorithms
Our algorithm is designed to nominate relevant candidates tailored to the detected principal pitfall (as defined in Definition 5), from which we extract the information necessary to overcome that pitfall effectively. Accordingly, we name it the Pitfall-Adapted Nomination (PAN) algorithm.
Our primary focus is on PAN allocation rules, which determine the arm to be sampled at each time step based on the observed history. These rules are independent of the total budget of samples $T$ or the confidence parameter $\delta$, making them anytime. In the fixed-confidence setting, a PAN allocation rule is coupled with a stopping rule to ensure $\delta$-correctness. As detailed in Section 3.1, we adopt standard stopping and decision rules from the literature for this setting. In contrast, the fixed-budget and posterior convergence rate settings terminate at a pre-specified budget $T$ and are evaluated based on their respective PCS measures. This versatility demonstrates the adaptability of the PAN framework to various performance criteria.
We propose PAN allocation rules that follow three steps: first, estimate (est) the unknown parameters using $\widehat{\bm{\theta}}$, which also provides the most plausible answer $\mathcal{I}(\widehat{\bm{\theta}})$ to the exploration query; next, detect (det) the principal pitfall $x$ that could make $\mathcal{I}(\widehat{\bm{\theta}})$ incorrect; and finally, select (sel) one arm to sample based on the detected principal pitfall $x$. A template for our allocation rule is summarized in Algorithm 5.
Algorithm 5 Pitfall-Adapted Nomination (PAN) allocation rule for general pure exploration
1:Current history $\mathcal{H}$, estimation rule (est), detection rule (det), and selection rule (sel).
2:Run estimation rule to obtain $\widehat{\bm{\theta}}=\textup{{est}}(\mathcal{H})$.
3:Run detection rule to obtain the principal pitfall $x=\textup{{det}}(\widehat{\bm{\theta}},\mathcal{H})$.
4:return $I=\textup{{sel}}\bigl{(}x,\widehat{\bm{\theta}},\mathcal{H}\bigr{)}$.
We will follow the naming convention of est-det-sel for our proposed PAN algorithms. Below, we elaborate on the options for the three subroutines required by PAN. We provide explicit formulas of $\mathcal{X}(\bm{\theta}),C_{x}(\bm{p})$ and $\bm{h}^{x}(\bm{p})$ for a wide range of pure-exploration problems in Appendix A.
3.5.1 Estimation (est) Rules
We consider two estimation rules. The empirical-best (EB) rule estimates the means using the empirical mean $\bm{\theta}_{t}$ based on the history $\mathcal{H}_{t}$ at time $t$. The Thompson-sampling (TS) rule estimates the means by drawing a sample $\widetilde{\bm{\theta}}_{t}$ from the posterior distribution $\Pi_{t}$ at time $t$.
The estimate $\widehat{\bm{\theta}}$ returned by the estimation rule determines the perceived best answer $\mathcal{I}(\widehat{\bm{\theta}})$, the pitfall set $\mathcal{X}(\widehat{\bm{\theta}})$, and the decomposed alternative set $\mathrm{Alt}_{x}(\widehat{\bm{\theta}})$. Along with the sample allocation $\bm{p}$, this determines the estimated generalized Chernoff information $C_{x}(\bm{p},\widehat{\bm{\theta}})$.
3.5.2 Detection (det) Rules
We propose three detection rules.
  1. 1.
    The KKT detection rule (Algorithm 6) explicitly follows the complementary slackness conditions (14b), and identifies the principal pitfall as the $x$ that minimizes $C_{x}(\bm{p},\widehat{\bm{\theta}})$.
    Algorithm 6 KKT detection rule for general pure exploration
    1:Current sample allocation $\bm{p}$ and an estimate $\widehat{\bm{\theta}}$.
    2:return principal pitfall $x=\operatorname*{argmin}_{x\in\mathcal{X}(\widehat{\bm{\theta}})}C_{x}(\bm{p},\widehat{\bm{\theta}})$, breaking ties arbitrarily.
  2. 2.
    The TS detection rule (Algorithm 7) generalizes the TTTS algorithm (Algorithm 1) for BAI to more general pure-exploration problems. It identifies an alternative problem instance with a different answer (and hence, a pitfall) that is most likely to emerge from the posterior distribution. This framework offers a systematic approach to adapting the classical TS algorithm for pure-exploration problems.
    Algorithm 7 TS detection rule for general pure exploration
    1:Estimation $\widehat{\bm{\theta}}$ and posterior distribution $\Pi$.
    2:Repeatedly sample $\widetilde{\bm{\theta}}\sim\Pi$ until $\widetilde{\bm{\theta}}\in\mathrm{Alt}(\widehat{\bm{\theta}})$.
    3:return principal pitfall $x\in\bigl{\{}x^{\prime}\in\mathcal{X}(\widehat{\bm{\theta}}):\widetilde{\bm{\theta}}\in\mathrm{Alt}_{x^{\prime}}(\widehat{\bm{\theta}})\bigr{\}}$, breaking ties arbitrarily.
  3. 3.
    The Posterior-Probability-Sampling (PPS) detection rule (Algorithm 8) samples the principal pitfall $x$ with probability proportional to $\Pi(\mathrm{Alt}_{x}(\widehat{\bm{\theta}}))$, where $\Pi$ denotes the current posterior distribution.
    Algorithm 8 PPS detection rule for general pure exploration
    1:Estimation $\widehat{\bm{\theta}}$ and posterior distribution $\Pi$.
    2:return principal pitfall $x\in\mathcal{X}(\widehat{\bm{\theta}})$ sampled with probability proportional to $\Pi\bigl{(}\mathrm{Alt}_{x}(\widehat{\bm{\theta}})\bigr{)}$.
Following the same rationale as in Section 2.2.2, Russo (2020, Proposition 5) establish that the TS and PPS detection rules are asymptotically equivalent to the KKT rule, as $t\to\infty$.
3.5.3 Selection (sel) Rule
For the selection rule, we follow the IDS rule (Algorithm 4).
Remark 7 (When does the design framework work?).
The derivation of the PAN algorithm relies only on two ingredients that lead directly to the KKT conditions in Theorem 4: (i) Assumption 1 induces the maximin convex optimization problem central to our framework, which is satisfied by many common exploration queries, as demonstrated in Appendix A; and (ii) by definition (8) using the GLRT, the generalized Chernoff information is homogeneous of degree one, satisfying $C_{x}(t\cdot\bm{p})=t\cdot C_{x}(\bm{p})$ for any $t>0$, which is equivalent to the PDE characterization in (9) by Euler’s theorem for homogeneous functions; see Apostol (1969, p. 287) for example.
Remark 8 (Extensions).
Our analysis extends beyond the base case to accommodate: (i) reward distributions outside the single-parameter exponential family, and (ii) exploration queries with multiple correct answers. We present these extensions in Section 4. Specifically, in Section 4.1, we consider the problem of identifying one of the $\varepsilon$-good alternatives, known as $\varepsilon$-best-arm identification ($\varepsilon$-BAI, Eckman and Henderson 2018, Jourdan et al. 2023b). This problem aligns with R&S under a PGS guarantee. The reason $\varepsilon$-BAI is compatible with our framework is that it can be reformulated as a modified pure-exploration problem with a unique answer. In Section 4.2, we examine the Gaussian case with unknown variances—a scenario in which the reward distribution no longer belongs to a single-parameter exponential family. Similar analyses are possible for other distributions; for instance, Agrawal et al. (2020) provides an analysis for heavy-tailed distributions. Importantly, our algorithm can be readily extended to accommodate these cases.
3.6 Universal Efficiency
In this section, we apply Algorithm 5 to Gaussian BAI (Example 1) and to pure-exploration thresholding bandits (Example 3), demonstrating that two specialized versions of Algorithm 5 are universally optimal for each problem, respectively.
3.6.1 Gaussian BAI
For BAI presented in Example 1 with Gaussian rewards, we focus on the TTTS sub-routine (Algorithm 1) combined with the information-directed selection sub-routine (Algorithm 2). Under the naming convention of Section 3.5, this is the TS-TS-IDS algorithm. To honor the original naming convention proposed in Russo (2020), we refer to this algorithm as TTTS-IDS. As a theoretical guarantee, we prove that for Gaussian BAI, TTTS-IDS satisfies the sufficient condition for optimal allocation rules in Theorem 3. This resolves an open problem in the pure-exploration literature raised by Russo (2020, Section 8). The proof is deferred to Appendix E.
Theorem 5.
TTTS-IDS ensures that, for any $\bm{\theta}$ satisfying $\theta_{i}\neq\theta_{j}$ for all $i\neq j$, we have $\bm{p}_{t}\xrightarrow{\mathcal{L}}\bm{p}^{*},$ where the optimal allocation $\bm{p}^{*}$ is identified in Theorem 1, which verifies the sufficient condition for optimal allocation rules in Theorem 3.
In conjunction with Theorem 2 on the fixed-budget posterior convergence rate and the discussion of the static fixed-budget setting in Remark 4, the TTTS-IDS is optimal across all three settings.
Remark 9 (Fixed-confidence optimality of other top-two algorithms).
The analysis of TTTS-IDS is arguably the most complex among the top-two algorithms, as both the leader and challenger are random variables conditioned on the history. Our proof strategy can be extended to other top-two algorithms combined with IDS, such as TS-KKT-IDS, TS-PPS-IDS, and those presented in Russo (2020), Qin et al. (2017), Shang et al. (2020), Jourdan et al. (2022).
3.6.2 Pure-Exploration Thresholding Bandits
For pure-exploration thresholding bandits discussed in Example 3, we prove the universal efficiency of the EB-KKT-IDS algorithm. For this problem, we have $\mathcal{X}=[K]$ and the optimal allocation problem simplifies to
$\Gamma^{*}_{\bm{\theta}}$ $=\max_{\bm{p}\in\mathcal{S}_{K}}\min_{i\in\mathcal{X}}\inf_{\bm{\vartheta}\in\mathrm{Alt}_{i}}\sum_{i}p_{i}d(\theta_{i},\vartheta_{i})=\max_{\bm{p}\in\mathcal{S}_{K}}\min_{i\in\mathcal{X}}\inf_{(\mathfrak{T}-\vartheta_{i})(\mathfrak{T}-\theta_{i})<0}p_{i}d(\theta_{i},\vartheta_{i})=\max_{\bm{p}\in\mathcal{S}_{K}}\min_{i\in[K]}p_{i}d(\theta_{i},\mathfrak{T}),$
where the second equality follows from the fact that we can always set $\vartheta_{j}=\theta_{j}$ for $j\neq i$, and the last equality follows from the monotonicity of the KL divergence. We apply Lemma 1 to obtain that $\frac{\partial C_{i}}{\partial p_{i}}=d(\theta_{i},\mathfrak{T})$ and $\frac{\partial C_{i}}{\partial p_{j}}=0$ for any $j\neq i$. Hence, we have $h^{i}_{i}(\bm{p})=\frac{p_{i}d(\theta_{i},\mathfrak{T})}{C_{i}(\bm{p})}=1$, and $h^{i}_{j}(\bm{p})=0$, for all $j\neq i$ and $i\in[K]$.
With the explicit forms of $\mathcal{X}$, $C_{x}$, and $\bm{h}^{x}$ derived, we can immediately implement EB-KKT-IDS algorithm. Theorem 6 establishes that EB-KKT-IDS is universally efficient. The proof deferred to Appendix F.
Theorem 6.
EB-KKT-IDS ensures that for any $\bm{\theta}\in\Theta$, $\bm{p}_{t}\xrightarrow{\mathcal{L}}\bm{p}^{*}$.
4 Extensions
In this section, we briefly discuss several important problems that the main framework introduced in Section 3 does not cover. The first is related to alternative performance criteria, specifically the probability of good selection (PGS), also known as $(\epsilon,\delta)$-PAC (probably approximately correct) in the learning literature. The second concerns a practical scenario in which Gaussian distributions have unknown variances.
4.1 Optimal Probability of Good Selection Guarantee: Gaussian $\varepsilon$-Best-Arm Identification
This section demonstrates that our framework can be extended to develop optimal algorithms for a variant of the R&S problem, specifically R&S with a PGS guarantee. This approach is related to, but distinct from, the PCS guarantee in the indifference-zone (IZ) formulation (Perng 1969, Jennison et al. 1982). For a detailed discussion, see Eckman and Henderson (2018) and Hong et al. (2021). In the bandit literature, this problem is referred to as the $\varepsilon$-best-arm identification problem (Garivier and Kaufmann 2021, Jourdan et al. 2023b), or $(\epsilon,\delta)$-PAC learning.
Consider the same $K$-armed bandits. Let $\mathcal{I}_{\varepsilon}(\bm{\theta})\triangleq\{i\in[K]:\theta_{i}\geq\theta_{I^{*}}-\varepsilon\}$ denote the set of “good enough arms”, whose gap from the best arm $I^{*}=\operatorname*{argmax}_{i\in[K]}\theta_{i}$ (assuming uniqueness) is within a specified tolerance $\varepsilon$. The $\varepsilon$-BAI problem aims to find any arm $i\in\mathcal{I}_{\varepsilon}(\bm{\theta})$.
To apply our framework, the first step is to establish a lower bound on the problem complexity. Notice that the $\varepsilon$-BAI problem differs significantly from the BAI problem in that the correct answer is non-unique; any arm $i\in\mathcal{I}_{\varepsilon}$ can serve as a correct answer. When $\varepsilon$ exceeds the gap between the best and second-best arm, there may be multiple correct answers. For a discussion on pure-exploration problems with multiple correct answers, see Degenne and Koolen (2019).
In this section, we restrict our focus to Gaussian bandits with a homogeneous variance of $\sigma^{2}=1$. For Gaussian bandits, Jourdan et al. (2023b, Lemma 9) establishes the lower bound
$\lim_{\delta\to 0}\frac{\mathbb{E}_{\bm{\theta}}[\tau_{\delta}]}{\log(1/\delta)}\geq(\Gamma^{*}_{\bm{\theta},\varepsilon})^{-1},\text{ where }\Gamma^{*}_{\bm{\theta},\varepsilon}=\max_{i\in\mathcal{I}_{\varepsilon}(\bm{\theta})}\max_{\bm{p}\in\mathcal{S}_{K}}\min_{j\neq i}\frac{(\theta_{i}-\theta_{j}+\varepsilon)^{2}}{2\left(\frac{1}{p_{i}}+\frac{1}{p_{j}}\right)}=\max_{\bm{p}\in\mathcal{S}_{K}}\min_{j\neq I^{*}}\frac{(\theta_{I^{*}}-\theta_{j}+\varepsilon)^{2}}{2\left(\frac{1}{p_{I^{*}}}+\frac{1}{p_{j}}\right)}.$
To interpret the complexity lower bound, note that finding any $\varepsilon$-good arm can be achieved by asking, for each arm $i$, the question: Is arm $i$ an $\varepsilon$-good arm? This question can be further decomposed into smaller hypothesis tests $H_{0,i,j}:\theta_{j}-\varepsilon\geq\theta_{i}$ versus $H_{1,i,j}:\theta_{j}-\varepsilon<\theta_{i}$ for any $j\neq i$. To see this, if there exists another arm $i\neq j$ for which $H_{0,i,j}$ holds, then arm $i$ cannot be $\varepsilon$-good, as $\theta_{j}-\theta_{i}>\varepsilon$. Therefore, to assert that arm $i$ is $\varepsilon$-good, we need to reject $H_{0,i,j}$ for all $j\neq i$. The term $\frac{(\theta_{i}-\theta_{j}+\varepsilon)^{2}}{2\left(1/p_{i}+1/p_{j}\right)}$ represents the test statistic, where a higher value increases the likelihood of rejection. Consequently, the quantity $\min_{j\neq i}\frac{(\theta_{i}-\theta_{j}+\varepsilon)^{2}}{2\left(1/p_{i}+1/p_{j}\right)}$ corresponds to the information needed to assert that $i$ is $\varepsilon$-good. The problem complexity depends on maximizing this information across all arms under their optimal allocations. For the last equality, we observe that the best arm $I^{*}$ is always the easiest to confirm as $\varepsilon$-good.
With the maximin complexity lower bound, the $\varepsilon$-BAI problem falls within the scope of our framework. Specifically, we can adapt the KKT detection rule (Algorithm 6) by substituting in
$C_{j}(\bm{p};\bm{\theta})=\frac{(\theta_{I^{*}(\bm{\theta})}-\theta_{j}+\varepsilon)^{2}}{2\bigl{(}1/p_{I^{*}(\bm{\theta})}+1/p_{j}\bigr{)}}\quad\text{and}\quad h^{j}_{i}(\bm{p})=\frac{p_{j}}{p_{i}+p_{j}},\quad h^{j}_{j}(\bm{p})=\frac{p_{i}}{p_{i}+p_{j}}\quad\text{for all}\quad j\neq I^{*}(\bm{\theta}).$
This approach yields the EB-TCε algorithm studied in Jourdan et al. (2023b). Following our proof techniques for Theorem 5, Jourdan et al. (2023b) establishes the fixed-confidence optimality of EB-TCε, when combined with IDS and recommending the empirical best arm upon stopping.
Proposition 1 (Lemma 24 in Jourdan et al. 2023b).
Under EB-TCε-IDS, we have $\bm{p}_{t}\xrightarrow{\mathcal{L}}\bm{p}^{*}.$
4.2 Dealing with Unknown Variances
Gaussian distributions with unknown variances no longer belong to the single-parameter exponential family. Fortunately, our framework can still be applied by making appropriate modifications to the GLRT statistic.
Let $(\bm{\theta},\bm{\sigma}^{2})$ denote the unknown mean and variance vectors. The alternative set is given by $\mathrm{Alt}(\bm{\theta},\bm{\sigma}^{2})=\{(\bm{\vartheta},\bm{\varsigma}^{2})\in\Theta\times\mathbb{R}_{+}^{K}:I^{*}(\bm{\vartheta})\neq I^{*}(\bm{\theta})\},$ where $I^{*}(\bm{\theta})$ is the index of the best arm under $\bm{\theta}$. This set can be decomposed into a union of convex sets, satisfying Assumption 1, $\mathrm{Alt}(\bm{\theta},\bm{\sigma}^{2})=\cup_{x\in\mathcal{X}(\bm{\theta},\bm{\sigma}^{2})}\mathrm{Alt}_{x}(\bm{\theta},\bm{\sigma}^{2}),$ where $\mathcal{X}(\bm{\theta},\bm{\sigma}^{2})\triangleq[K]\backslash\{I^{*}(\bm{\theta})\}$ and $\mathrm{Alt}_{x}(\bm{\theta},\bm{\sigma}^{2})=\{(\bm{\vartheta},\bm{\varsigma}^{2})\in\Theta\times\mathbb{R}_{+}^{K}:\vartheta_{x}>\vartheta_{I^{*}(\bm{\theta})}\}.$ The generalized Chernoff information is similarly defined as the population version of the GLRT statistic, with the key distinction that the likelihood ratio now involves the KL divergence for a Gaussian distribution with two parameters: for all $x\in\mathcal{X}(\bm{\theta},\bm{\sigma}^{2})$,
$C_{x}(\bm{p};\bm{\theta},\bm{\sigma}^{2})\triangleq\inf_{(\bm{\vartheta},\bm{\varsigma}^{2})\in\mathrm{Alt}_{x}(\bm{\theta},\bm{\sigma}^{2})}\sum_{i\in[K]}p_{i}d\bigl{(}(\theta_{i},\sigma^{2}_{i}),(\vartheta_{i},\varsigma^{2}_{i})\bigr{)},$ (15)
where $d\bigl{(}(\theta,\sigma^{2}),(\vartheta,\varsigma^{2})\bigr{)}=\frac{1}{2}\bigl{[}(\theta-\vartheta)^{2}/\varsigma^{2}+\sigma^{2}/\varsigma^{2}-1-\log(\sigma^{2}/\varsigma^{2})\bigr{]}.$ Following Lemma 11 of Jourdan et al. (2023a),
$C_{x}(\bm{p};\bm{\theta},\bm{\sigma}^{2})=\inf_{\vartheta\in\left[\theta_{x},\theta_{I^{*}(\bm{\theta})}\right]}\frac{1}{2}\left[{p_{I^{*}(\bm{\theta})}}\log\left(1+\frac{(\theta_{I^{*}(\bm{\theta})}-\vartheta)^{2}}{\sigma^{2}_{I^{*}(\bm{\theta})}}\right)+{p_{x}}\log\left(1+\frac{(\theta_{x}-\vartheta)^{2}}{\sigma^{2}_{x}}\right)\right].$
Added complexity arises because the minimizer $\vartheta$ cannot be determined in closed form. However, it can be efficiently computed using a bisection search.
Compared to the known-variance case, the generalized Chernoff information takes a distinct form, highlighting a fundamental difference between the two problems. Notably, the generalized Chernoff information under unknown variances is strictly smaller than that for known variances, and there are problem instances where the ratio between the two can be arbitrarily small (Jourdan et al. 2023a, Lemma 4 and 12). Crucially, substituting empirical variances into the complexity measures derived for known-variance cases, such as those in Glynn and Juneja (2004), Garivier and Kaufmann (2016), does not produce optimal results.
Fortunately, the key components of our algorithmic framework remain intact. The first component is the maximin formulation of the optimal allocation problem, which now reads:
$\Gamma_{\bm{\theta},\bm{\sigma}^{2}}^{*}=\max_{\bm{p}\in\mathcal{S}_{k}}\min_{x\in\mathcal{X}(\bm{\theta},\bm{\sigma}^{2})}C_{x}(\bm{p};\bm{\theta},\bm{\sigma}^{2}).$
[…]
$h_{i}^{x}(\bm{p};\bm{\theta},\bm{\sigma}^{2})\triangleq\frac{p_{i}\frac{\partial}{\partial p_{i}}C_{x}(\bm{p};\bm{\theta},\bm{\sigma}^{2})}{C_{x}(\bm{p};\bm{\theta},\bm{\sigma}^{2})},\quad\text{for all }x\in\mathcal{X}(\bm{\theta})\text{ and }i\in[K],$ (16)
is still a valid probability vector since $C_{x}(\bm{p};\bm{\theta},\bm{\sigma}^{2})$ is homogeneous of degree one by definition. The KKT condition retains the same form as in Theorem 4, with appropriate substitutions to account for the unknown variances. Specifically, $C_{x}(\bm{p};\bm{\theta},\bm{\sigma}^{2})$ defined in (15) replaces $C_{x}(\bm{p};\bm{\theta})$, and $\mathcal{X}(\bm{\theta},\bm{\sigma}^{2})$ replaces $\mathcal{X}(\bm{\theta})$. Additionally, the new IDS rule is applied, using $h_{i}^{x}(\bm{p};\bm{\theta},\bm{\sigma}^{2})$ as defined in (16). With these specifications, a KKT detection rule can be implemented directly. Similarly, Bayesian detection rules, such as TS and PPS, are straightforward to implement when paired with an appropriate conjugate prior.
5 Numerical Experiments
For the numerical study, we focus on the following problem instances of the best-$k$ identification problem. For each case, we examine both Bernoulli and Gaussian bandits (with $\sigma^{2}=1$) and adopt an uninformative prior for our algorithms. Additional numerical experiments are collected in Appendix I. Implementation of our algorithms is available at https://github.com/cnyouwei/PE.
Case ID $K,k$ $\bm{\theta}$
1 $5,2$ $(0.1,0.2,0.3,0.4,0.5)$
2 $20,5$ $\theta_{i}=0.05*i,\forall i\in[20]$
3 $15,1$ $\theta_{i}=0.3,\forall i\in[14];\theta_{15}=0.7$
4 $100,10$ $\theta_{i}=0.3,\forall i\in[90];\theta_{i}=0.7,\forall i\in[100]\backslash[90]$
5 $50,25$ $\theta_{i}=0.2,\forall i\in[10];\theta_{i}=0.5,\forall i\in[25]\backslash[10];\theta_{i}=0.8,\forall i\in[50]\backslash[25]$
5.1 Convergence
We established in Theorem 5 that the empirical allocation under our proposed algorithms converges strongly to the optimal allocation for Gaussian BAI. In Figure 1, we present the empirical convergence rate of $\mathcal{O}(1/\sqrt{t})$ to the optimal value for our algorithms in Case 1, where $\bm{\theta}=(0.5,0.4,0.3,0.2,0.1)$, for identifying the best-$2$ under both Gaussian and Bernoulli reward distributions. These results suggest that this convergence behavior extends to more general settings. We record the trajectories of $\Gamma_{\bm{\theta}}(\bm{p}_{t})$ over $10^{3}$ independent replications and plot the mean trajectories, with shaded areas representing the first and third quartiles.
Figure 1: $\mathcal{O}(t^{-1/2})$ convergence to the optimal value, with shaded areas representing the first and third quartiles.
5.2 Fixed-Confidence Setting
For benchmarks, we compare our proposed algorithms with existing ones, which are briefly summarized in Appendix H. The $m$-LinGapE and MisLid algorithms, implemented by Tirinzoni and Degenne (2022), are designed for linear bandits and therefore are not implemented for Bernoulli bandits. To ensure fair comparisons, we set the exploration rate as $\beta({t,\delta})=\log\bigl{(}(\log(t)+1)/\delta\bigr{)}$ in UGapE and KL-LUCB, along with their respective stopping rules. Table 1 presents the fixed-confidence sample complexity, highlighting the superior performance of our proposed algorithms. The FWS algorithm also performs competitively, though with significantly higher computational complexity. Table 1 also reports computation times for selected cases, demonstrating our algorithm’s efficiency. Notably, computation time is primarily constrained by the stopping rule calculation, which scales as $\mathcal{O}(k(K-k))$; see Section 5.4.3 for further discussion.
Case $\delta$ LB TTTS-IDS TS-KKT-IDS TS-PPS-IDS FWS $m$-LinGapE MisLid LMA KL-LUCB UGapE Uniform
Bernoulli 1 0.1 487 838$\pm$12 763$\pm$14 857$\pm$13 852$\pm$14 1131$\pm$13 1643$\pm$26 1639$\pm$25 1322$\pm$13
0.01 974 1349$\pm$19 1287$\pm$19 1414$\pm$20 1323$\pm$20 1683$\pm$18 2687$\pm$38 2666$\pm$38 2330$\pm$17
2 0.1 1637 3417$\pm$27 2779$\pm$25 3462$\pm$26 2964$\pm$25 6329$\pm$25 5453$\pm$48 5726$\pm$50 18465$\pm$42
0.01 3275 5007$\pm$38 4406$\pm$37 5067$\pm$38 4565$\pm$36 9314$\pm$39 8856$\pm$74 8851$\pm$73 31236$\pm$60
3 0.1 155 350$\pm$1 282$\pm$1 354$\pm$1 284$\pm$1 517$\pm$1 533$\pm$4 550$\pm$2 610$\pm$1
0.01 311 504$\pm$2 435$\pm$2 506$\pm$2 443$\pm$1 725$\pm$1 843$\pm$6 901$\pm$2 899$\pm$1
4 0.1 1114 3536$\pm$2 2755$\pm$1 3668$\pm$1 2834$\pm$3 7350$\pm$1 3859$\pm$3 4562$\pm$2 8992$\pm$1
0.01 2228 4644$\pm$2 3522$\pm$2 4716$\pm$2 3868$\pm$3 9772$\pm$1 6048$\pm$5 7161$\pm$2 11607$\pm$1
5 0.1 934 2868$\pm$2 2121$\pm$2 2951$\pm$2 2536$\pm$5 4513$\pm$2 3870$\pm$48 3419$\pm$4 6703$\pm$2
0.01 1869 3759$\pm$3 2869$\pm$3 3820$\pm$3 3341$\pm$5 6053$\pm$3 5203$\pm$5 5566$\pm$22 8745$\pm$2
Gaussian 1 0.1 2159 3919$\pm$58 3667$\pm$61 3991$\pm$61 3684$\pm$60 4327$\pm$58 4411$\pm$65 5809$\pm$63 7775$\pm$121 7618$\pm$117 5975$\pm$58
0.01 4318 6230$\pm$85 5949$\pm$89 6169$\pm$85 5974$\pm$89 6717$\pm$84 6786$\pm$91 8431$\pm$78 11932$\pm$168 11956$\pm$167 10162$\pm$80
2 0.1 9459 19424$\pm$151 16371$\pm$146 19282$\pm$148 17385$\pm$153 24526$\pm$162 24534$\pm$182 34280$\pm$152 32650$\pm$284 32828$\pm$284 101990$\pm$225
0.01 18918 29214$\pm$218 26274$\pm$215 29685$\pm$225 27227$\pm$227 33563$\pm$215 36828$\pm$270 46383$\pm$220 52323$\pm$436 50977$\pm$417 178697$\pm$322
3 0.1 647 1484$\pm$5 1181$\pm$5 1455$\pm$5 1240$\pm$7 1720$\pm$4 1964$\pm$7 1959$\pm$5 2497$\pm$44 2331$\pm$7 2617$\pm$5
0.01 1294 2152$\pm$7 1827$\pm$7 2110$\pm$7 1904$\pm$7 2477$\pm$5 2883$\pm$9 2768$\pm$6 3627$\pm$25 3757$\pm$10 3933$\pm$6
4 0.1 4605 14756$\pm$8 19101$\pm$161* 15035$\pm$6 17251$\pm$52 19585$\pm$5 25815$\pm$15 24967$\pm$5 53498$\pm$441* 66248$\pm$463* 38579$\pm$4
0.01 9210 19543$\pm$12 20279$\pm$87* 19742$\pm$8 21669$\pm$53 25595$\pm$6 32946$\pm$16 32672$\pm$6 29321$\pm$128* 31197$\pm$76* 49312$\pm$5
5 0.1 4177 12994$\pm$14 16492$\pm$191* 13113$\pm$11 19577$\pm$102 15660$\pm$11 20502$\pm$20 18676$\pm$10 55535$\pm$552* 56074$\pm$480* 31350$\pm$8
0.01 8355 17226$\pm$15 17340$\pm$111* 17291$\pm$14 21875$\pm$96 20333$\pm$14 27519$\pm$28 25050$\pm$13 25373$\pm$138* 25896$\pm$88* 40921$\pm$9
Case $\delta$ ($K,k$) Execution time per iteration in microseconds ($\upmu$s)
Gaussian 1 0.1 (5,2) 90.3 8.5 7.5 405.8 12.3 17.5 12.9 1.2 3.6 3.2
0.01 (5,2) 738.7 8.8 8.0 385.6 12.2 16.1 12.2 1.1 3.9 3.1
4 0.1 (100,10) 279.4 657.5 530.3 31295.7 12998.1 9844.7 4440.2 10.4 153.9 217.1
0.01 (100,10) 591.4 680.4 531.2 32563.8 13660.1 10323.8 4555.7 10.3 167.6 218.9
Table 1: Fixed-confidence sample complexity, averaged over $1000$ replications. Cases where the algorithm did not stop within $10^{6}$ steps for some replications are marked with *, with problem complexity recorded as $10^{6}$ in these cases. The $95\%$ confidence interval is provided after the “$\pm$” sign.
5.3 Advantage of IDS over $\beta$-Tuning
To isolate the effect of the selection rule, we implement PAN in Algorithm 5 with the Thompson sampling estimation rule and the PPS detection rule, comparing the IDS selection rule in Algorithm 2 with $0.5$-tuning, i.e., set $h^{j}_{i}=h^{j}_{j}=0.5$. We consider the slippage configuration where $\bm{\theta}$ is given by $\theta_{i}=0.75,i\in[k]$, and $\theta_{j}=0.5,j\in[K]\backslash[k]$ for both Gaussian and Bernoulli bandits. Instances are labeled by $(K,k)$-distribution. Table 2 reports the significant reduction in sample complexity achieved by using IDS.
$(K,k)$-dist $(500,1)$-G $(500,2)$-G $(500,3)$-G $(500,5)$-G $(500,1)$-B $(500,2)$-B $(500,3)$-B $(500,5)$-B
TS-PPS-0.5 285823 285591 283401 283180 62596 61914 62660 62734
TS-PPS-IDS 199374 216668 223860 241869 44640 47044 50436 52767
Reduction 30.2% 24.1% 21.0% 14.6% 28.7% 24.0% 19.5% 15.9%
Table 2: Fixed-confidence sample complexity with $\delta=0.001$, averaged over 100 replications.
5.4 Large-Scale Experiments
The simulation literature has recently seen a surge of interest in large-scale R&S problems, which involve a substantial number of alternative systems. For those interested, we recommend the recent survey by Fan et al. (2024), which provides a comprehensive overview of large-scale simulation optimization. In this section, we evaluate the performance of our proposed algorithms on several large-scale problem instances, considering both fixed-confidence and fixed-budget settings.
5.4.1 Fixed-Confidence Setting
In this experiment, we consider the large-scale slippage configuration studied in Zhong and Hong (2022), where the means of the systems are set as $\mu_{0}=0.1$ and $\mu_{i}=0.0$ for $i=2,\dots,K$. The noisy performances of the systems are independent and follows Gaussian distributions with a known variance of $\sigma^{2}=1$. We set the desired confidence level to $\delta=0.05$.
In Table 3, we compare the fixed-confidence performance of Procedure $\mathcal{KT}_{0}$ (Zhong and Hong 2022), Procedure $\mathcal{KN}$ (Kim and Nelson 2001), and two proposed algorithms, TS-KKT-IDS and TTTS-IDS. The simulation results for Procedures $\mathcal{KT}_{0}$ and $\mathcal{KN}$ are taken from Zhong and Hong (2022, Table 2). Both $\mathcal{KT}_{0}$ and $\mathcal{KN}$ require the specification of an indifference zone parameter, which is set to the (unknown) true gap of $0.1$ in this experiment. In contrast, our algorithms do not require parameter tuning and are therefore entirely parameter-free. For our fixed-confidence algorithms, the stopping rule (11) is applied using a heuristic threshold $\gamma(t,\delta)=\Phi^{-1}(1-\delta)$, i.e., the $1-\delta$ quantile of the standard normal distribution. All results are averaged over $1000$ independent replications.
TS-KKT-IDS TTTS-IDS EB-TS-IDS Procedure $\mathcal{KT}_{0}$ Procedure $\mathcal{KN}$
$K$ PCS $\mathbb{E}[\tau_{\delta}]/K$ PCS $\mathbb{E}[\tau_{\delta}]/K$ PCS $\mathbb{E}[\tau_{\delta}]/K$ PCS $\mathbb{E}[\tau_{\delta}]/K$ PCS $\mathbb{E}[\tau_{\delta}]/K$
$10^{2}$ $0.999$ $616\pm 16$ $1.0$ $758\pm 14$ $0.999$ $727\pm 12$ $0.963$ $1462\pm 3$ $0.978$ $884\pm 8$
$10^{3}$ $1.0$ $764\pm 14$ $1.0$ $864\pm 12$ $1.0$ $788\pm 7$ $0.965$ $1512\pm 1$ $0.989$ $1165\pm 8$
$10^{4}$ $1.0$ $937\pm 13$ $1.0$ $1015\pm 12$ $0.999$ $898\pm 5$ $0.956$ $1527\pm 1$ $0.994$ $1419\pm 9$
$10^{5}$ $1.0$ $1069\pm 14$ $1.0$ $1154\pm 13$ $1.0$ $1024\pm 3$ $0.955$ $1528\pm 1$ $0.994$ $1686\pm 8$
Table 3: Sample complexity and estimated PCS are compared for the proposed algorithms, Procedure $\mathcal{KT}_{0}$ and Procedure $\mathcal{KN}$ across different numbers of systems in a slippage configuration with known and equal variances. The $95\%$ confidence interval is provided after the “$\pm$” sign, and the estimated PCS is shown in parentheses.
The theoretical stopping rule is designed to ensure PCS for high-confidence levels, i.e., as $\delta\to 0$, but empirically, it tends to be overly conservative for moderate PCS level guarantees. Thus, we implemented a heuristic stopping rule better suited to larger $\delta$. Our results show that this heuristic rule remains conservative, with an estimated PCS exceeding 0.999 in all cases. However, the asymptotically optimal allocation compensates for the conservatism of the stopping rule, achieving smaller sample complexities and higher PCS than competitors. Future work should focus on developing stopping rules that are less conservative for moderate $\delta$. Recent studies, such as Wang et al. (2024) and Tirinzoni and Degenne (2022), propose elimination-based stopping rules that could reduce conservatism and improve computational efficiency. We leave this direction for future research and present this numerical study to illustrate the potential of asymptotically optimized algorithms in large-scale experiments.
5.4.2 Fixed-Budget Setting
In this experiment, we evaluate the performance of the proposed algorithms under configurations studied by Li et al. (2025), including the slippage configuration ($\mu_{0}=0.1$ and $\mu_{i}=0.0$ for $i=2,\dots,K$ with $\sigma_{i}^{2}$ = 1), and the equally-spaced configuration with increasing variances ($\mu_{0}=0.1$ and $\mu_{i}=-i/K$ for $i=2,\dots,K$ with $\sigma_{i}^{2}=1+i/K$). We assume that system variances are known.
A key insight from Li et al. (2025) is that allocating samples greedily—sampling the system with the highest empirical mean—keeps PCS bounded away from zero as the total sampling budget scales linearly with the number of systems, i.e., $B=cK$ for a fixed constant $c$ as $K\to\infty$. This suggests that a degree of greediness is essential in large-scale R&S to promptly eliminate an empirical best system that is not the true best. They introduced the Explore-First Greedy (EFG) procedure to ensure consistency.
Inspired by their work, we test our proposed algorithms with the EB estimation rule, i.e., the EB-*-IDS algorithms. Figure 2 presents the estimated PCS for the proposed algorithms and benchmarks, including EFG, Greedy, and uniform-allocation procedures. PCS is estimated from $10^{4}$ independent replications, with confidence intervals nearly invisible and thus omitted. As expected, EB-*-IDS algorithms appear to be sample-optimal in the sense described by Hong et al. (2022), Li et al. (2025), as the estimated PCS stabilizes at a positive value as $K\to\infty$. This sample optimality is anticipated since the EB-*-IDS algorithms allocate sufficient samples to the empirical best. We leave a rigorous investigation of this sample optimality for future work. The advantage of our proposed algorithms becomes more evident as the average budget per system increases, demonstrating that a dynamic allocation strategy informed by asymptotic analysis can enhance performance under moderate to low PCS levels.
Figure 2: Estimated PCS are compared for the proposed algorithms and benchmarks. Top row is the slippage configuration with equal variances, bottom row is the equally-spaced configuration with increasing variances.
5.4.3 Computation Time
To compare the computation time of the proposed algorithms, we consider the slippage configuration, where system means are set as $\mu_{0}=0.1$ and $\mu_{i}=0.0$ for $i=2,\dots,K$. The noisy performance of each system is independent and follows Gaussian distributions with a known variance of $\sigma^{2}=1$. We vary the number of systems from $K=10^{1}$ to $10^{4}$. For each instance, we run the algorithm for $100K$ or $1000K$ time steps and perform $10^{3}$ independent repetitions to estimate the sample mean of the running time along with its 95% confidence interval.
Table 4 reports the execution time of our proposed algorithms, measured in average microseconds ($\upmu$s) per iteration. We observe that both TS-KKT-IDS and TS-PPS-IDS are invariant to budget changes, while TTTS-IDS slows down as the time period grows. However, this computational inefficiency in TTTS-IDS is minimal when PCS is below $0.97$. Additionally, the slowdown is less pronounced for larger $K$ as sampling an alternative best system becomes easier with more alternatives available.
All three algorithms scale nearly linearly with the number of systems $K$, as expected. For instance, in TS-KKT-IDS, the bottleneck is searching for the minimum among the $K-1$ number of $C_{j}$ functions. Notably, when an algorithm targets constant optimality, such scaling is inevitable—after all, even the computation complexity of calculating the problem complexity lower bound scales linearly with $K$.
TTTS-IDS TS-PPS-IDS TS-KKT-IDS
Budget $100K$ $1000K$ $100K$ $1000K$ $100K$ $1000K$
$K=10^{1}$ $0.66\pm 0.01(0.409)$ $5.93\pm 1.11(0.974)$ $0.84\pm 0.01(0.411)$ $1.04\pm 0.01(0.979)$ $0.35\pm 0.00(0.443)$ $0.39\pm 0.01(0.971)$
$K=10^{2}$ $3.38\pm 0.12(0.161)$ $7.86\pm 0.59(0.991)$ $4.75\pm 0.17(0.179)$ $5.16\pm 0.18(0.980)$ $1.92\pm 0.07(0.204)$ $1.99\pm 0.07(0.988)$
$K=10^{3}$ $22.59\pm 0.98(0.065)$ $25.85\pm 0.99(0.978)$ $32.94\pm 1.43(0.059)$ $33.51\pm 1.44(0.975)$ $13.16\pm 0.57(0.057)$ $13.13\pm 0.57(0.983)$
$K=10^{4}$ $170.13\pm 7.96(0.026)$ $172.56\pm 7.92(0.964)$ $297.96\pm 14.27(0.018)$ $297.52\pm 14.22(0.955)$ $98.79\pm 4.62(0.019)$ $98.62\pm 4.61(0.964)$
Table 4: Computation time for the proposed algorithms, measured in average microseconds ($\upmu$s) per iteration. The $95\%$ confidence interval is provided after the “$\pm$” sign, and the estimated PCS is shown in parentheses.
5.5 Practical Considerations in Algorithm Implementation
Exploration versus exploitation in pure exploration.
The exploration-exploitation trade-off arises even in pure-exploration problems. For instance, the estimation rules EB and TS aim to identify the best arm. EB exploits by selecting the arm with the highest sample mean, while TS explores by sampling from the posterior. Asymptotic performance is largely unaffected by the choice of estimation rule, but practical performance can differ under moderate confidence levels or limited budgets. EB converges faster when the greedy guess is correct, but it risks poor performance when the guess is wrong, as allocations are based on a misspecified problem instance. In contrast, TS explores more to avoid getting stuck with incorrect estimates but requires additional samples when the estimate is already accurate. Thus, EB-*-IDS algorithms [^6] Except for EB-KKT-IDS, where both estimation and detection are greedy, leading to significant outliers in slippage configurations. typically achieve better average performance but are more susceptible to extreme outliers due to inaccurate initial estimates. In contrast, TS-*-IDS algorithms generally exhibit slightly lower average performance but fewer outliers. This behavior is discussed in Wu and Zhou (2018a), Jourdan et al. (2022) and conceptually linked to Simchi-Levi et al. (2024). As a side note, Li et al. (2025) observes that EB generally stabilizes the PCS at positive levels as the number of alternatives grows, while TS does not.
A similar trade-off exists in the choice of the detection rule (det in Section 3.5.2). While KKT greedily selects the $x$ with the smallest $C_{x}$ as the principal pitfall, TS and PPS encourage exploration of other pitfalls. For instance, comparing TTTS-IDS with TS-KKT-IDS in Table 1 and Table 3, we find that *-KKT-IDS algorithms perform better on average but are more prone to extreme outliers (e.g., Table 1, Case 4 and 5, Gaussian). In contrast, *-TS-IDS algorithms exhibit slightly worse average performance but no outliers. However, in most cases, exploration in the estimation rule alone suffices to mitigate outliers.
Common Random Numbers (CRN).
In simulation studies, the performances of different systems are often simulated using CRN, introducing positive correlations between the noisy performances. This typically reduces sampling complexity by decreasing the variances of the differences in sample means, see Nelson and Matejcik (1995), Zhong and Hong (2022).
The main challenge in fully leveraging CRN lies in the need to allow noisy performances of different systems to be mutually correlated. For simplicity of exposition, we introduced our framework under the assumption that the rewards are i.i.d. random variables from a single-parameter exponential family. However, our framework can be extended to handle correlated observations.
To demonstrate, consider the case where performances follow a multivariate normal distribution with means $\mu_{i}$, variances $\sigma_{i}^{2}$, and pairwise correlation $\rho$, where both $\sigma_{i}^{2}$ and $\rho$ are known. [^7] Extensions to unknown covariance matrices are possible using the multiparameter exponential family, as in Section 4.2. The generalized Chernoff information is modified to account for the correlation as: for $j\in[K]\backslash\{I^{*}\}$,
$C_{j}(\bm{p};\bm{\theta})=\frac{(\theta_{I^{*}}-\theta_{j})^{2}}{2\Bigl{(}\frac{\sigma_{I^{*}}^{2}}{p_{I^{*}}}+\frac{\sigma_{j}^{2}}{p_{j}}-\frac{2\rho\sigma_{I^{*}}\sigma_{j}}{\sqrt{p_{I^{*}}p_{j}}}\Bigr{)}}.$
It can be verified that (9) holds, along with Theorem 4, allowing the immediate application of an PAN algorithm. For the fixed-confidence setting, the modified $C_{j}$ must replace the original in the stopping rule. Although the universal efficiency of PAN under CRN remains an open question, numerical evidence in Table 5 demonstrates its effectiveness.
TS-KKT-IDS
Case $K$ $\rho=0.0$ $\rho=0.25$ $\rho=0.5$ $\rho=0.75$
A $10^{2}$ 72$\pm$4(0.999) 62$\pm$4(0.988) 52$\pm$3(0.979) 35$\pm$2(0.957)
$10^{3}$ 87$\pm$3(1.000) 80$\pm$3(0.999) 72$\pm$3(0.997) 55$\pm$2(0.984)
$10^{4}$ 112$\pm$1(1.000) 106$\pm$2(1.000) 94$\pm$2(1.000) 75$\pm$2(0.995)
B $10^{2}$ 616$\pm$16(1.000) 537$\pm$16(0.998) 423$\pm$17(0.992) 254$\pm$10(0.965)
$10^{3}$ 743$\pm$14(1.000) 670$\pm$16(1.000) 554$\pm$17(0.996) 336$\pm$12(0.992)
$10^{4}$ 937$\pm$13(1.000) 832$\pm$17(1.000) 693$\pm$18(1.000) 455$\pm$14(0.998)
Table 5: Sample complexity per arm and estimated PCS are compared for the TS-KKT-IDS algorithm with varying $\rho$. Case A corresponds to the slippage configuration with equal variances (Section 5.4.1), and Case B to the equally-spaced configuration with increasing variances (Section 5.4.2). The $95\%$ confidence interval is shown after “$\pm$” signs, followed by the estimated PCS.
6 Conclusion
This paper introduces a unified algorithm design principle for general pure-exploration problems by linking asymptotic lower bounds on problem complexities to maximin optimal allocation problems. This connection yields reformulated Karush-Kuhn-Tucker conditions involving dual variables, offering a novel interpretation of the top-two algorithm principle as a method for tracking complementary slackness conditions. Furthermore, the derived stationarity conditions motivate the information-directed selection rule, which, when integrated with TTTS algorithm, achieves optimal performance for Gaussian BAI in posterior convergence, fixed-confidence, and fixed-budget settings.
Our framework demonstrates broad applicability, optimally extending to settings such as $\varepsilon$-BAI and pure-exploration thresholding bandits, notably without relying on forced exploration. We propose practical modifications to Thompson sampling, simplifying its implementation for practitioners addressing various pure-exploration problems. Experimental results confirm our algorithms’ superior performance, highlighting improvements in both fixed-budget probability of incorrect selection and fixed-confidence sample complexity, with robust results across diverse and large-scale problem configurations.
Future research directions include refining stopping rules to enhance efficiency in finite-sample settings or under moderate confidence requirements (Simchowitz et al. 2017), and extending theoretical results to broader settings, such as heavy-tailed distributions (Agrawal et al. 2020). Another promising avenue is developing parallelizable algorithm variants suited for large-scale simulations (Avci et al. 2023) and incorporating elimination procedures to reduce parallelization overhead (Zhong and Hong 2022).
Acknowledgments
We express our sincere gratitude to the anonymous reviewers and the Associate Editor for their valuable feedback, which significantly improved the quality of the paper. We are also grateful to Daniel Russo, Sandeep Juneja, Po-An Wang, Shane Henderson, Xiaowei Zhang, Jun Luo, the participants of the 2023 INFORMS Applied Probability Society Conference, the 2023 INFORMS Annual Meeting, and the 2024 INFORMS MSOM Conference for their insightful comments on this work. A preliminary version of this paper appeared as an extended abstract in the Proceedings of the 36th Annual Conference on Learning Theory, COLT’23, with the title “Information-Directed Selection for Top-Two Algorithms.”
Appendix
[…]
subsubsection
[…]
Appendix A Proposed Algorithms for Selected Pure-Exploration Queries
In this section, we collect a wide variety of examples of pure-exploration queries that fall within our algorithmic framework and provide explicit formulas for implementing our algorithms.
To implement the KKT detection rule (Algorithm 6) and the IDS selection rule (Algorithm 4), we require expressions for the pitfall set $\mathcal{X}(\bm{\theta})$, the generalized Chernoff information $C_{x}(\bm{p})$, and the selection functions $h_{i}^{x}(\bm{p})$. To implement the TS detection rule (Algorithm 7), we require an explicit formula for the pitfall set $\mathcal{X}(\bm{\theta})$, the alternative set $\mathrm{Alt}(\bm{\theta})$, and an oracle to sample from the posterior distribution (this is usually straightforward in the case of a single-parameter exponential family). Moreover, to implement the PPS detection rule (Algorithm 8), we need to calculate the posterior probability of the alternative set, which is also made possible by the explicit formulas for the alternative sets.
To enable direct implementation of our algorithms, we explicitly derive the following components for each pure-exploration problem:
  1. 1.
    Correct answer $\mathcal{I}(\bm{\theta})$.
  2. 2.
    The alternative set $\mathrm{Alt}(\bm{\theta})$, the pitfall set $\mathcal{X}(\bm{\theta})$, and the decomposition of $\mathrm{Alt}(\bm{\theta})$.
  3. 3.
    The unique solution $\bm{\vartheta}^{x}$ in Lemma 1.
  4. 4.
    The generalized Chernoff information $C_{x}(\bm{p})$.
  5. 5.
    The partial derivative of $C_{x}(\bm{p})$, from which the selection functions $h_{i}^{x}(\bm{p})$ can be derived.
A.1 Variants of BAI
BAI or IZ-free R&S.
This is studied in detail in Section 2. The goal is to find the arm with the highest mean.
  1. 1.
    Correct answer: $\mathcal{I}(\bm{\theta})=\{I^{*}(\bm{\theta})\}$, where $I^{*}(\bm{\theta})=\operatorname*{argmax}_{i}\{\theta_{i}\}$.
  2. 2.
    $\mathrm{Alt}(\bm{\theta})=\cup_{j\in\mathcal{X}(\bm{\theta})}\mathrm{Alt}_{j}(\bm{\theta})$ with $\mathcal{X}(\bm{\theta})=[K]\backslash\mathcal{I}(\bm{\theta})$ and $\mathrm{Alt}_{j}(\bm{\theta})=\{\bm{\vartheta}:\vartheta_{j}>\vartheta_{I^{*}(\bm{\theta})}\}$.
  3. 3.
    $\bm{\vartheta}^{j}$ is given by $\vartheta^{j}_{I^{*}(\bm{\theta})}=\vartheta^{j}_{j}=\frac{p_{I^{*}(\bm{\theta})}\theta_{I^{*}(\bm{\theta})}+p_{j}\theta_{j}}{p_{I^{*}(\bm{\theta})}+p_{j}}$ and $\vartheta^{j}_{k}=\theta_{k}$ for all $j\in\mathcal{X}(\bm{\theta})$ and $k\neq I^{*}(\bm{\theta}),j$888 Our algorithm relies on the vector $\bm{\vartheta}^{j}$ as defined in the function $\bm{h}^{x}(\bm{p})$ used in IDS. Note that $\bm{\vartheta}^{j}$ is undefined when $p_{I^{*}}=p_{j}=0$. In these cases, we assign $\bm{h}^{x}(\bm{p})=\frac{1}{K}\bm{1},$ which corresponds to the uniform vector. This choice is made without loss of generality, since beginning with one allocation per arm ensures that the empirical allocations will never become exactly zero, no matter how small the allocation may get. A similar argument applies to the best-$k$ identification, pure-exploration thresholding bandit (selecting the arm closest to a threshold), and all-$\varepsilon$-best-arm identification scenarios, so we do not repeat this discussion for those cases.
  4. 4.
    $C_{j}(\bm{p})=p_{I^{*}(\bm{\theta})}d(\theta_{I^{*}(\bm{\theta})},\vartheta^{j}_{I^{*}(\bm{\theta})})+p_{j}d(\theta_{j},\vartheta^{j}_{j})$ for all $j\in\mathcal{X}(\bm{\theta})$.
  5. 5.
    $\frac{\partial C_{j}(\bm{p})}{\partial p_{i}}=d(\theta_{i},\vartheta^{j}_{i})$ for all $j\in\mathcal{X}(\bm{\theta})$, most of them are zero, except for $i=I^{*}(\bm{\theta})$ or $j$.
Best-$k$ Identification.
The goal is to find $k$ arms with the highest means.
  1. 1.
    Correct answer: $\mathcal{I}(\bm{\theta})=\operatorname*{argmax}_{\mathcal{S}\subseteq[K],|\mathcal{S}|=k}\left\{\sum_{i\in\mathcal{S}}\theta_{i}\right\}$.
  2. 2.
    $\mathrm{Alt}(\bm{\theta})=\cup_{(i,j)\in\mathcal{X}(\bm{\theta})}\mathrm{Alt}_{i,j}(\bm{\theta})$ with $\mathcal{X}(\bm{\theta})=\mathcal{I}(\bm{\theta})\times\mathcal{I}(\bm{\theta})^{c}$ and $\mathrm{Alt}_{i,j}(\bm{\theta})=\{\bm{\vartheta}:\vartheta_{j}>\vartheta_{i}\}$.
  3. 3.
    $\bm{\vartheta}^{(i,j)}$ is given by $\vartheta^{(i,j)}_{i}=\vartheta^{(i,j)}_{j}=\frac{p_{i}\theta_{i}+p_{j}\theta_{j}}{p_{i}+p_{j}}$ and $\vartheta^{(i,j)}_{k}=\theta_{k}$ for all $(i,j)\in\mathcal{X}(\bm{\theta})$ and $k\neq i,j$.
  4. 4.
    $C_{i,j}(\bm{p})=p_{i}d(\theta_{i},\vartheta^{(i,j)}_{i})+p_{j}d(\theta_{j},\vartheta^{(i,j)}_{j})$ for all $(i,j)\in\mathcal{X}(\bm{\theta})$.
  5. 5.
    $\frac{\partial C_{i,j}(\bm{p})}{\partial p_{k}}=d(\theta_{k},\vartheta^{(i,j)}_{k})$ for all $(i,j)\in\mathcal{X}(\bm{\theta})$, most of them are zero, except for $k=i$ or $j$.
Pairwise Pure Exploration.
The BAI and best-$k$ identification problems can be further generalized to the so-called pairwise pure exploration [Wu et al., 2025]. Let $\mathcal{R}\subset\{(i,j):i\neq j\}$ be any subset of (ordered) pairs. [^9] The set $\mathcal{R}$ must be admissible in the sense that $(i,j)$ and $(j,i)$ are not simultaneously in $\mathcal{R}$. A pairwise pure exploration query partitions the permutation group of the arm set $[K]$ into mutually exclusive and collectively exhaustive subsets $\{\mathcal{R}_{1},\dots,\mathcal{R}_{M}\}$, and the goal is to determine which of the subsets $\mathcal{R}_{m}$ is consistent with the true ranking of $\bm{\theta}$.
  1. 1.
    Correct answer: $\mathcal{I}(\bm{\theta})=\{m^{*}(\bm{\theta})\}$, where $m^{*}(\bm{\theta})$ is the unique index such that $\bm{\theta}$ satisfies all constraints in $\mathcal{R}_{m^{*}(\bm{\theta})}$.
  2. 2.
    $\mathrm{Alt}(\bm{\theta})=\cup_{(i,j)\in\mathcal{X}(\bm{\theta})}\mathrm{Alt}_{i,j}(\bm{\theta})$ with $\mathcal{X}(\bm{\theta})=\mathcal{R}_{m^{*}(\bm{\theta})}$ and $\mathrm{Alt}_{i,j}(\bm{\theta})=\{\bm{\vartheta}:\vartheta_{j}>\vartheta_{i}\}$.
  3. 3.
    $\bm{\vartheta}^{(i,j)}$ is given by $\vartheta^{(i,j)}_{i}=\vartheta^{(i,j)}_{j}=\frac{p_{i}\theta_{i}+p_{j}\theta_{j}}{p_{i}+p_{j}}$ and $\vartheta^{(i,j)}_{k}=\theta_{k}$ for all $(i,j)\in\mathcal{X}(\bm{\theta})$ and $k\neq i,j$.
  4. 4.
    $C_{i,j}(\bm{p})=p_{i}d(\theta_{i},\vartheta^{(i,j)}_{i})+p_{j}d(\theta_{j},\vartheta^{(i,j)}_{j})$ for all $(i,j)\in\mathcal{X}(\bm{\theta})$.
  5. 5.
    $\frac{\partial C_{i,j}(\bm{p})}{\partial p_{k}}=d(\theta_{k},\vartheta^{(i,j)}_{k})$ for all $(i,j)\in\mathcal{X}(\bm{\theta})$, most of them are zero, except for $k=i$ or $j$.
All-$\varepsilon$-Best-Arm Identification.
The goal is to find the set of all arms with means larger than $\max_{j\in[K]}\theta_{j}-\varepsilon$; see Mason et al. [2020].
  1. 1.
    Correct answer: $\mathcal{I}(\bm{\theta})=\{i:\theta_{i}\geq\max_{j\in[K]}\theta_{j}-\varepsilon\}$.
  2. 2.
    $\mathrm{Alt}(\bm{\theta})=\cup_{(i,j)\in\mathcal{X}}\mathrm{Alt}_{i,j}(\bm{\theta})$, where $\mathcal{X}(\bm{\theta})=\left\{(i,j)\in[K]\times[K]:i\in\mathcal{I}(\bm{\theta}),j\neq i\right\}$, and for any $(i,j)\in\mathcal{X}(\bm{\theta})$, $\mathrm{Alt}_{i,j}(\bm{\theta})=\{\bm{\vartheta}:\vartheta_{j}-\varepsilon>\vartheta_{i}\}$. To interpret, if there exists an arm $j\neq i$ with $\theta_{j}-\theta_{i}>\epsilon$, then arm $i$ cannot be $\varepsilon$-good. Notice the subtle difference between the pitfall set $\mathcal{X}(\bm{\theta})$ here and that for the best-$k$ identification problem.
  3. 3.
    $\bm{\vartheta}^{(i,j)}$ is given by $\vartheta^{(i,j)}_{i}=\frac{p_{i}(\theta_{i}+\varepsilon)+p_{j}\theta_{j}}{p_{i}+p_{j}}-\varepsilon$, $\vartheta^{(i,j)}_{j}=\frac{p_{i}(\theta_{i}+\varepsilon)+p_{j}\theta_{j}}{p_{i}+p_{j}}$ and $\vartheta^{(i,j)}_{k}=\theta_{k}$ for all $(i,j)\in\mathcal{X}(\bm{\theta})$ and $k\neq i,j$.
  4. 4.
    $C_{i,j}(\bm{p})=p_{i}d(\theta_{i},\vartheta^{(i,j)}_{i})+p_{j}d(\theta_{j},\vartheta^{(i,j)}_{j})$ for all $(i,j)\in\mathcal{X}(\bm{\theta})$.
  5. 5.
    $\frac{\partial C_{i,j}(\bm{p})}{\partial p_{k}}=d(\theta_{k},\vartheta^{(i,j)}_{k})$ for all $(i,j)\in\mathcal{X}(\bm{\theta})$, most of them are zero, except for $k=i$ or $j$.
A.2 Variants of Pure-Exploration Thresholding Bandits
Pure-Exploration Thresholding Bandits.
This is studied in Section 3.6.2. The goal is to find the set of all arms whose mean rewards are above a given threshold $\mathfrak{T}\in\mathbb{R}$.
  1. 1.
    Correct answer: $\mathcal{I}(\bm{\theta})=\{k\in[K]:\theta_{k}>\mathfrak{T}\}$.
  2. 2.
    $\mathrm{Alt}(\bm{\theta})=\cup_{j\in\mathcal{X}}\mathrm{Alt}_{j}(\bm{\theta})$ where $\mathcal{X}=[K]$ and $\mathrm{Alt}_{j}(\bm{\theta})=\{\bm{\vartheta}:(\mathfrak{T}-\vartheta_{j})(\mathfrak{T}-\theta_{j})<0\}$.
  3. 3.
    $\bm{\vartheta}^{j}$ is given by $\vartheta^{j}_{j}=\mathfrak{T}$ and $\vartheta^{j}_{k}=\theta_{k}$ for all $j\in\mathcal{X}$ and $k\neq j$.
  4. 4.
    $C_{j}(\bm{p})=\sum_{i\in[K]}p_{i}d(\theta_{i},\vartheta^{j}_{i})=p_{j}d(\theta_{j},\mathfrak{T})$.
  5. 5.
    $\frac{\partial C_{j}(\bm{p})}{\partial p_{i}}=d(\theta_{i},\vartheta^{j}_{i})$ for all $j\in\mathcal{X}$, most of them are zero, except for $i=j$.
Pure-Exploration Thresholding Bandit (Selecting One Arm Closest to a Threshold).
The goal is to find the arm whose mean reward is the closest to a given threshold $\mathfrak{T}\in\mathbb{R}$.
  1. 1.
    Correct answer: $\mathcal{I}(\bm{\theta})=\{I^{*}(\bm{\theta})\}$, where $I^{*}(\bm{\theta})=\operatorname*{argmin}_{i\in[K]}|\theta_{i}-\mathfrak{T}|$.
  2. 2.
    $\mathrm{Alt}(\bm{\theta})=\cup_{j\in\mathcal{X}}\mathrm{Alt}_{j}(\bm{\theta})$, where $\mathcal{X}(\bm{\theta})=[K]\backslash\mathcal{I}(\bm{\theta})$ and $\mathrm{Alt}_{j}(\bm{\theta})=\{\bm{\theta}:|\theta_{j}-\mathfrak{T}|<|\theta_{I^{*}(\bm{\theta})}-\mathfrak{T}|\}$.
  3. 3.
    $\bm{\vartheta}^{j}$ is given by $\vartheta^{j}_{I^{*}(\bm{\theta})}=\vartheta^{j}_{j}=\frac{p_{I^{*}(\bm{\theta})}\theta_{I^{*}(\bm{\theta})}+p_{j}\theta_{j}}{p_{I^{*}(\bm{\theta})}+p_{j}}$, and $\vartheta^{j}_{k}=\theta_{k}$ for all $k\neq I^{*}(\bm{\theta}),j.$
  4. 4.
    $C_{j}(\bm{p})=\sum_{i\in[K]}p_{i}d(\theta_{i},\vartheta^{j}_{i})$ for all $j\in\mathcal{X}$.
  5. 5.
    $\frac{\partial C_{j}(\bm{p})}{\partial p_{i}}=d(\theta_{i},\vartheta^{j}_{i})$ for all $j\in\mathcal{X}$, most of them are zero, except for $i=I^{*}(\bm{\theta})$ and $j$.
Signed Bandits.
This is an variant of pure-exploration thresholding bandit problem. The goal is to identified the whether the arm means are all above or all below a given threshold $\mathfrak{T}\in\mathbb{R}$. We assume that either all arm means are above the threshold $\mathfrak{T}$ or below it—let this restricted parameter space be denoted as $\widetilde{\Theta}$; see Ménard [2019].
  1. 1.
    Correct answers: $\mathcal{I}(\bm{\theta})=\{+\}$ if $\min_{i\in[K]}\theta_{i}>\mathfrak{T}$ and $\mathcal{I}(\bm{\theta})=\{-\}$ if $\max_{i\in[K]}\theta_{i}<\mathfrak{T}$.
  2. 2.
    $\mathrm{Alt}(\bm{\theta})=\{\bm{\vartheta}\in\widetilde{\Theta}:\max_{i\in[K]}\vartheta_{i}<\mathfrak{T}\}$ if $\mathcal{I}(\bm{\theta})=\{+\}$ and $\mathrm{Alt}(\bm{\theta})=\{\bm{\vartheta}\in\widetilde{\Theta}:\min_{i\in[K]}\vartheta_{i}>\mathfrak{T}\}$ if $\mathcal{I}(\bm{\theta})=\{-\}$. In both cases, the pitfall sets are singleton sets and no decomposition is needed. We remark that this satisfies Assumption 1.
  3. 3.
    For both $\mathcal{I}(\bm{\theta})=\{+\}$ and $\mathcal{I}(\bm{\theta})=\{-\}$, we have $\bm{\vartheta}=\mathfrak{T}\bm{1}$, i.e., a constant vector of $\mathfrak{T}$.
  4. 4.
    $C(\bm{p})=\sum_{i\in[K]}p_{i}d(\theta_{i},\mathfrak{T})$.
  5. 5.
    $\frac{\partial C(\bm{p})}{\partial p_{i}}=d(\theta_{i},\mathfrak{T})$ for all $i\in[K]$.
Murphy Sampling.
Consider yet another variant of pure-exploration thresholding bandit, where the goal is to determine whether the smallest mean reward is lower than a given threshold $\mathfrak{T}$ or not. If the smallest mean is lower than $\mathfrak{T}$, we say that the answer is “feasible,” and we say that the answer is “infeasible” if otherwise; see Kaufmann et al. [2018], Juneja and Krishnasamy [2019].
  1. 1.
    Correct answers: $\mathcal{I}(\bm{\theta})=\{\text{feasible}\}$ if $\min_{i\in[K]}\theta_{i}<\mathfrak{T}$ and $\mathcal{I}(\bm{\theta})=\{\text{infeasible}\}$ if $\min_{i\in[K]}\theta_{i}>\mathfrak{T}$.
  2. 2.
    $\mathrm{Alt}(\bm{\theta})=\{\bm{\vartheta}\in\widetilde{\Theta}:\min_{i\in[K]}\vartheta_{i}>\mathfrak{T}\}$ if $\mathcal{I}(\bm{\theta})=\{\text{feasible}\}$ (so that the pitfall set is a singleton set and no decomposition is needed) and $\mathrm{Alt}(\bm{\theta})=\{\bm{\vartheta}\in\widetilde{\Theta}:\min_{i\in[K]}\vartheta_{i}<\mathfrak{T}\}=\cup_{j\in\mathcal{X}}\mathrm{Alt}_{j}(\bm{\theta})=\cup_{j\in\mathcal{X}}\{\bm{\vartheta}:\vartheta_{j}<\mathfrak{T}\}$ if $\mathcal{I}(\bm{\theta})=\{\text{infeasible}\}$, where $\mathcal{X}(\bm{\theta})=[K]$.
  3. 3.
    If $\mathcal{I}(\bm{\theta})=\{\text{feasible}\}$, then $\bm{\vartheta}=\mathfrak{T}\bm{1}$, i.e., a constant vector of $\mathfrak{T}$. If $\mathcal{I}(\bm{\theta})=\{\text{infeasible}\}$, then $\bm{\vartheta}^{j}$ is given by $\vartheta^{j}_{j}=\mathfrak{T}$ and $\vartheta^{j}_{k}=\theta_{k}$ for all $j\in\mathcal{X}$ and $k\neq j$.
  4. 4.
    $C(\bm{p})=\sum_{i\in[K]}p_{i}d(\theta_{i},\mathfrak{T})$ if $\mathcal{I}(\bm{\theta})=\{\text{feasible}\}$, and $C_{j}(\bm{p})=\sum_{i\in[K]}p_{i}d(\theta_{i},\vartheta^{j}_{i})=p_{j}d(\theta_{j},\mathfrak{T})$ if $\mathcal{I}(\bm{\theta})=\{\text{infeasible}\}$.
  5. 5.
    $\frac{\partial C(\bm{p})}{\partial p_{i}}=d(\theta_{i},\mathfrak{T})$ for all $i\in[K]$, if $\mathcal{I}(\bm{\theta})=\{\text{feasible}\}$. $\frac{\partial C_{j}(\bm{p})}{\partial p_{i}}=d(\theta_{i},\vartheta^{j}_{i})$ for all $j\in\mathcal{X}$, if $\mathcal{I}(\bm{\theta})=\{\text{infeasible}\}$.
A.3 Partition Identification.
In Juneja and Krishnasamy [2019], the authors identify a class of pure-exploration problems, termed partition identification, which generalizes the variants of pure-exploration thresholding bandits discussed above. These problems, where the goal is to decide in which partition of the parameter space the true mean vector $\bm{\theta}$ lies, fall naturally within our framework.
Half-Space Problem.
The goal is to check whether the mean vector lies on one side or the other of the half-space, i.e., $\{\bm{\theta}\in\mathbb{R}^{K}:\sum_{i=1}^{K}a_{i}\theta_{i}>b\}$ or $\{\bm{\theta}\in\mathbb{R}^{K}:\sum_{i=1}^{K}a_{i}\theta_{i}<b\},$ for a specified $(a_{1},\dots,a_{K},b)\in\mathbb{R}^{K+1}$; see Juneja and Krishnasamy [2019, Theorem 5].
  1. 1.
    Correct answers: $\mathcal{I}(\bm{\theta})=\{+\}$ if $\sum_{i=1}^{K}a_{i}\theta_{i}>b$ and $\mathcal{I}(\bm{\theta})=\{-\}$ if $\sum_{i=1}^{K}a_{i}\theta_{i}<b$.
  2. 2.
    If $\mathcal{I}(\bm{\theta})=\{+\}$, then $\mathrm{Alt}(\bm{\theta})=\{\bm{\vartheta}\in\mathbb{R}^{K}:\sum_{i=1}^{K}a_{i}\vartheta_{i}<b\}.$ If $\mathcal{I}(\bm{\theta})=\{-\}$, then $\mathrm{Alt}(\bm{\theta})=\{\bm{\vartheta}\in\mathbb{R}^{K}:\sum_{i=1}^{K}a_{i}\vartheta_{i}>b\}.$ In both cases, the pitfall sets are singleton and no decomposition is needed.
  3. 3.
    For both cases, define the unique optimizer $\bm{\vartheta}^{*}=\operatorname*{argmin}_{\bm{\vartheta}\in\mathbb{R}^{K}:\sum_{i=1}^{K}a_{i}\vartheta_{i}=b}\sum_{i=1}^{K}p_{i}d(\theta_{i},\vartheta_{i})$. The first-order optimality conditions read $\frac{\partial}{\partial\vartheta_{i}}(p_{i}d(\theta_{i},\vartheta^{*}_{i}))=-\lambda a_{i},$ for $i=1,\dots,K,$ with the Lagrange multiplier $\lambda$ chosen such that $\sum_{i=1}^{K}a_{i}\vartheta^{*}_{i}=b.$
  4. 4.
    For both cases, $C(\bm{p})=\sum_{i=1}^{K}p_{i}d(\theta_{i},\vartheta^{*}_{i}).$
  5. 5.
    For both cases, $\frac{\partial C(\bm{p})}{\partial p_{i}}=d(\theta_{i},\vartheta^{*}_{i})$ for all $i\in[K].$
Convex-Set Feasibility Problem.
We note that the half-space problem naturally extends to verifying whether the mean vector $\bm{\theta}$ lies within a given convex feasible set; see, e.g., Juneja and Krishnasamy [2019, Section 3.3]. In particular, if the complement of the convex feasible set can be expressed as a finite union of convex sets, then our Assumption 1 is satisfied. For example, when the feasibility set is a convex polytope [Gangrade et al., 2024], its complement can be written as a finite union of half-spaces, and hence the assumption holds. Due to the wide variety of potential feasible sets and the associated technical complexity, we omit detailed derivations.
Appendix B Proof of Theorem 1
We first formally establish the existence and uniqueness of a probability vector that satisfies the information balance condition (1) and stationarity condition (2) in Theorem 1, as well as its properties, which will be needed in the proof of the sufficient condition in Theorem 1.
Lemma 3 (Existence and properties of $\bm{p}^{*}$).
There exists a unique and component-wise strictly positive probability vector $\bm{p}^{*}$ satisfying information balance condition (1) and stationarity condition (2). Furthermore, $\bm{p}^{*}$ is the unique solution to the maximin optimization problem in (6), i.e.,
$\bm{p}^{*}=\operatorname*{argmax}_{\bm{p}\in\mathcal{S}_{K}}\min_{j\neq I^{*}}C_{j}(\bm{p}).$
Proof.
The proof of this result directly follows from the proofs of Garivier and Kaufmann [2016, Lemma 4 and Theorem 5]. Although the statements of Garivier and Kaufmann [2016, Lemma 4 and Theorem 5] are given for the case of equal variances, their proofs can be extended straightforwardly to the case of unequal variances, establishing that:
  1. 1.
    the solution to maximin optimization problem in (6) is unique, and
  2. 2.
    it is the only probability vector that satisfies both the information balance condition (1) and the sum-of-squares balance condition (3).
As shown in Remark 1, the sum-of-squares balance condition (3) is equivalent to the stationarity condition (2). Hence, we have $\bm{p}^{*}=\operatorname*{argmax}_{\bm{p}\in\mathcal{S}_{K}}\min_{j\neq I^{*}}C_{j}(\bm{p})$. The strict positivity of each component of $\bm{p}^{*}$ follows from Garivier and Kaufmann [2016, Theorem 5] and its proof. ∎
The sufficient condition in Theorem 1 is directly led by the next result, which shows that
  1. 1.
    the optimal value of the maximin optimization problem in (6), denoted by $\Gamma_{\bm{\theta}}^{*}$, represents the optimal rate of posterior convergence, and
  2. 2.
    an allocation rule with asymptotic allocation converging to the unique solution of this maximin optimization problem, denoted by $\bm{p}^{*}$, is indeed optimal.
Theorem 7 (Optimal posterior convergence rate—A stronger result that implies Theorem 1).
For any allocation rule A, with probability one,
$\limsup_{t\to\infty}-\frac{1}{t}\log\Pi_{t}^{\textup{{A}}}(\Theta^{c}_{I^{*}})\leq\Gamma_{\bm{\theta}}^{*},$
where $\Gamma_{\bm{\theta}}^{*}$ is the optimal value of the maximin optimization problem in (6).
Furthermore, if an allocation rule $\textup{{A}}^{*}$ ensures $\bm{p}_{t}\xrightarrow{\mathrm{a.s.}}\bm{p}^{*}$ for the unique $\bm{p}^{*}$ defined in Theorem 1, then with probability one,
$\lim_{n\to\infty}-\frac{1}{t}\log\Pi_{t}^{\textup{{A}}^{*}}\!(\Theta^{c}_{I^{*}})=\Gamma_{\bm{\theta}}^{*}.$
Proof.
The proof of this result directly follows from that of Qin et al. [2017, Theorem 2]. Although the statement of Qin et al. [2017, Theorem 2] is given for the case of equal variances, its proof can be extended straightforwardly to the case of unequal variances. ∎
Appendix C A Threshold for the Stopping Rule and the Proof of Theorem 3
C.1 A Threshold for the Stopping Rule to Ensure $\delta$-Correctness
Let $h(u)=u-\log u$ and define for any $z\in[1,e]$ and $x\geq 0$
$\widetilde{h}_{z}(x)=\begin{cases}e^{1/h^{-1}(x)}h^{-1}(x),&\hbox{if }x\geq h(1/\log z),\\ z(x-\log\log z),&\hbox{otherwise.}\end{cases}$
Furthermore, define
$\mathcal{C}_{\mathrm{exp}}(x)=2\widetilde{h}_{3/2}\left(\frac{h^{-1}(1+x)+\log(2\zeta(2))}{2}\right),$
where $\zeta(s)=\sum_{n=1}^{\infty}n^{-s}$ is the Riemann zeta function. One can easily verify that $\mathcal{C}_{\mathrm{exp}}(x)\sim x$ when $x\to\infty$.
Proposition 2 (Kaufmann and Koolen 2021, Proposition 15).
The stopping rule (11) with the sequence of thresholds
$\gamma(t,\delta)=3\sum_{i=1}^{K}\log(1+\log N_{t,i})+K\mathcal{C}_{\mathrm{exp}}\left(\frac{\log(1/\delta)}{K}\right)$
is such that, for every allocation rule and any $\delta>0$,
$\mathbb{P}_{\bm{\theta}}\left(\tau_{\delta}<\infty,\mathcal{I}(\bm{\theta}_{\tau_{\delta}})=\mathcal{I}(\bm{\theta})\right)\geq 1-\delta,\quad\hbox{for any problem instance }\bm{\theta}\in\Theta,$
where $\bm{\theta}_{t}$ is the sample mean of the arms at time $t$.
C.2 Proof of Theorem 3
We rely on the following technical finding given by Proposition 14 in Qin and Russo [2024].
Proposition 3 (Proposition 14 in Qin and Russo 2024).
If an allocation rule satisfies that there exist $\rho>0$, $\alpha>0$ and ${T}\in\mathcal{L}$ such that for any $t\geq{T}$, $\min_{i\in[K]}N_{t,i}\geq\rho\cdot t^{\alpha}$, then $\bm{\theta}_{t}\xrightarrow{\mathcal{L}}\bm{\theta}$.
Proof of Theorem 3.
The first condition in Theorem 3 ensures that the policy $\pi$ is $\delta$-correct (Definition 2).
Now we fix $\bm{\theta}\in\Theta$ and write $\bm{p}^{*}=\bm{p}^{*}(\bm{\theta})$ for notational convenience. Since $\bm{p}^{*}>\bm{0}$ component-wisely, the second condition $\bm{p}_{t}\xrightarrow{\mathcal{L}}\bm{p}^{*}$ in Theorem 3 ensures that the condition in Proposition 3 is satisfied, and thus we have $\bm{\theta}_{t}\xrightarrow{\mathcal{L}}\bm{\theta}$. Then by the continuous mapping theorem,
$\min_{x\in\mathcal{X}(\bm{\theta}_{t})}C_{x}(\bm{p}_{t};\bm{\theta}_{t})\xrightarrow{\mathcal{L}}\min_{x\in\mathcal{X}(\bm{\theta})}C_{x}(\bm{p}^{*};\bm{\theta})=\Gamma^{*}_{\bm{\theta}},$
where the equality follows from that $\bm{p}^{*}$ is the maximizer of (10). For any $\epsilon>0$, there exists a random time $T_{\epsilon,1}\in\mathcal{L}$ such that for any $t\geq T_{\epsilon,1}$,
$\min_{x\in\mathcal{X}(\bm{\theta}_{t})}C_{x}(\bm{p}_{t};\bm{\theta}_{t})\geq\Gamma^{*}_{\bm{\theta}}-\epsilon.$
On the other hand, there exists a deterministic time $T_{\epsilon,2}$ such that for any $t\geq T_{\epsilon,2}$,
$\gamma(t,\delta)\leq 3K\log(1+\log t)+K\mathcal{C}_{\mathrm{exp}}\left(\frac{\log(1/\delta)}{K}\right)\leq\epsilon t+K\mathcal{C}_{\mathrm{exp}}\left(\frac{\log(1/\delta)}{K}\right).$
Hence, for any $t\geq T_{\epsilon}\triangleq\max\{T_{\epsilon,1},T_{\epsilon,2}\}\in\mathcal{L}$,
$t\cdot\min_{x\in\mathcal{X}(\bm{\theta}_{t})}C_{x}(\bm{p}_{t};\bm{\theta}_{t})-\gamma(t,\delta)\geq\left(\Gamma^{*}_{\bm{\theta}}-2\epsilon\right)t-K\mathcal{C}_{\mathrm{exp}}\left(\frac{\log(1/\delta)}{K}\right),$
which implies
$\tau_{\delta}\leq\max\left\{T_{\epsilon},\frac{K\mathcal{C}_{\mathrm{exp}}\left(\frac{\log(1/\delta)}{K}\right)}{\Gamma^{*}_{\bm{\theta}}-2\epsilon}\right\}\leq T_{\epsilon}+\frac{K\mathcal{C}_{\mathrm{exp}}\left(\frac{\log(1/\delta)}{K}\right)}{\Gamma^{*}_{\bm{\theta}}-2\epsilon}.$
Since $T_{\epsilon}=\max\{T_{\epsilon,1},T_{\epsilon,2}\}\in\mathcal{L}$, we have $\mathbb{E}[T_{\epsilon}]<\infty$, and then
$\limsup_{\delta\to 0}\frac{\mathbb{E}^{\pi}_{\bm{\theta}}[\tau_{\delta}]}{\log(1/\delta)}\leq\frac{1}{\Gamma^{*}_{\bm{\theta}}-2\epsilon},$
where we also use the fact that $\mathcal{C}_{\mathrm{exp}}(x)\sim x$ when $x\to\infty$. Since $\epsilon>0$ can be arbitrarily small, we have $\limsup_{\delta\to 0}\frac{\mathbb{E}^{\pi}_{\bm{\theta}}[\tau_{\delta}]}{\log(1/\delta)}\leq\frac{1}{\Gamma^{*}_{\bm{\theta}}}.$ We note that a matching lower bound on the sample complexity has been established in Wang et al. [2021, Appendix B], i.e., for any other $\delta$-correct algorithm $\pi^{\prime}$,
$\liminf_{\delta\to 0}\frac{\mathbb{E}^{\pi^{\prime}}_{\bm{\theta}}[\tau_{\delta}]}{\log(1/\delta)}\geq\frac{1}{\Gamma^{*}_{\bm{\theta}}},\quad\text{andthus we have}\quad\limsup_{\delta\to 0}\frac{\mathbb{E}^{\pi}_{\bm{\theta}}[\tau_{\delta}]}{\mathbb{E}^{\pi^{\prime}}_{\bm{\theta}}[\tau_{\delta}]}\leq 1.$
The above inequality holds for any $\bm{\theta}\in\Theta$, which implies that $\pi$ is universal efficient (Definition 3). ∎
Appendix D Proof of Theorem 4
Proof.
Note that the optimal allocation problem (10) is equivalent to
$\Gamma^{*}=\max_{\phi,\bm{p}}$ $\hskip 10.0pt\phi$
$\mathrm{s.t.}$ $\hskip 10.0pt\sum_{i\in[K]}p_{i}-1=0,$
$\hskip 10.0ptp_{i}\geq 0,\quad\forall i\in[K],$
$\hskip 10.0pt\phi-C_{x}(\bm{p})\leq 0,\quad\forall x\in\mathcal{X}.$
The corresponding Lagrangian is given by
$\mathcal{L}(\phi,\bm{p},\lambda,\bm{\mu})=\phi-\lambda\left(\sum_{i\in[K]}p_{i}-1\right)-\sum_{x\in\mathcal{X}}\mu_{x}(\phi-C_{x}(\bm{p}))+\sum_{i\in[K]}\iota_{i}p_{i}.$
Slater’s condition holds since $\bm{p}=\bm{1}/K$ and $\phi=0$ is a feasible solution such that all inequality constraints hold strictly, where $\bm{1}$ is a vector of $1$’s. Hence, the KKT conditions are necessary and sufficient for global optimality, which are then given by
$-1+\sum_{x\in\mathcal{X}}\mu_{x}=0,$ (18a)
$\lambda-\sum_{x\in\mathcal{X}}\mu_{x}\frac{\partial C_{x}(\bm{p})}{\partial p_{i}}-\iota_{i}=0,\quad\forall i\in[K],$ (18b)
$\sum_{i\in[K]}p_{i}-1=0,\quad p_{i}\geq 0,\quad\forall i\in[K],$ (18c)
$\phi-C_{x}(\bm{p})\leq 0,\quad\forall x\in\mathcal{X},$ (18d)
$\mu_{x}\geq 0,\quad\forall x\in\mathcal{X},\quad\iota_{i}\geq 0\quad\forall i\in[K],$ (18e)
$\sum_{x\in\mathcal{X}}\mu_{x}(\phi-C_{x}(\bm{p}))=0,\quad\sum_{i\in[K]}\iota_{i}p_{i}=0.$ (18f)
We first show that the KKT conditions (18) implies (14). To this end, we reformulate the stationarity condition (18b), leaving the feasibility and complementary slackness conditions unchanged. Define the selection functions $h^{x}_{i}$ as in (13), which we repeat here for convenience:
$h^{x}_{i}\triangleq h^{x}_{i}(\bm{\theta},\bm{p})=\frac{p_{i}d(\theta_{i},\vartheta^{x}_{i})}{C_{x}(\bm{p})},\quad\forall(x,i)\in\mathcal{X}\times[K].$
Now, we can rewrite the KKT conditions by eliminating the dual variables $\lambda$ and $\iota$.
$\lambda-\sum_{x\in\mathcal{X}}\mu_{x}\frac{\partial C_{x}(\bm{p})}{\partial p_{i}}-\iota_{i}=0,\quad\forall i\in[K],$
[…]
Multiply $\text{Multiply }p_{i}\text{ and use C-S for $\bm{\iota}$}\quad\Rightarrow\quad$ $\frac{\lambda}{\phi}p_{i}-\sum_{x\in\mathcal{X}}\mu_{x}\frac{p_{i}\frac{\partial C_{x}(\bm{p})}{\partial p_{i}}}{C_{x}(\bm{p})}=0,\quad\forall i\in[K],$ (20)
[…]
Sum over $\text{Sum over }i\quad\Rightarrow\quad$ $\frac{\lambda}{\phi}-\sum_{x\in\mathcal{X}}\mu_{x}\sum_{i\in[K]}h^{x}_{i}=0,\quad\forall i\in[K],$
[…]
Substituting $\lambda=\phi$ into (20) yields the desired result. Note that (20) is equivalent to (19) if and only if $\bm{p}^{*}$ is strictly positive.
Next, we show that (14) implies the KKT conditions in (18) when the optimal solution is strictly positive. First, when $\bm{p}^{*}$ is strictly positive, we have $\bm{\iota}=\bm{0}$, then (18a), (18d), (18e), and (18f) hold automatically. Let $\lambda=\phi$, we have (18b) using (20), (18f) and $\bm{\iota}=\bm{0}$. Finally, (18c) is obtained by summing (14a) over all $i\in[K]$ and noting that $\bm{\mu}\in\mathcal{S}_{|\mathcal{X}|}$. This completes the proof. ∎
Appendix E Proof of Theorem 5
For clarity of exposition, our analysis focuses on the case of a common variance. The arguments readily generalize to the heterogeneous variance setting by substituting the common variance $\sigma^{2}$ with either $\sigma^{2}_{\min}=\min_{i\in[K]}\sigma^{2}_{i}$ or $\sigma^{2}_{\max}=\max_{i\in[K]}\sigma^{2}_{i}$, depending on the direction of the inequality in question.
E.1 Preliminary
E.1.1 Notation
Minimum and maximum gaps.
We define the minimum and maximum values between the expected rewards of two different arms:
$\Delta_{\min}\triangleq\min_{i\neq j}\left\lvert\theta_{i}-\theta_{j}\right\rvert\quad\text{and}\quad\Delta_{\max}\triangleq\max_{i,j\in[K]}\left(\theta_{i}-\theta_{j}\right).$
Since the arm means are unique, we have $\Delta_{\max}\geq\Delta_{\min}>0$.
Another measure of cumulative and average effort.
We have defined one measure of cumulative and average effort: the number of samples and the proportion of total samples allocated to arm $i\in[K]$ before time $t\in\mathbb{N}_{0}$ are given by
$N_{t,i}=\sum_{\ell=0}^{t-1}\mathds{1}(I_{\ell}=i)\quad\text{and}\quad p_{t,i}=\frac{N_{t,i}}{t}.$
To analyze randomized algorithms (such as TTTS), we introduce an alternative notion based on sampling probabilities. Specifically, let
$\psi_{t,i}\triangleq\mathbb{P}(I_{t}=i|\mathcal{H}_{t}),$
denote the probability of sampling arm $i$ at time $t$, conditional on the history
$\mathcal{H}_{t}=\left(I_{0},Y_{1,I_{0}},\ldots,I_{t-1},Y_{t,I_{t-1}}\right).$
We then define the corresponding cumulative and average effort as
$\Psi_{t,i}\triangleq\sum_{\ell=0}^{t-1}\psi_{\ell,i}\quad\text{and}\quad w_{t,i}\triangleq\frac{\Psi_{t,i}}{t}.$
Posterior means and variances under independent uninformative priors.
Under the independent uninformative prior, the posterior mean and variance of arm $i$’s unknown mean at time $t$ are given by:
$\theta_{t,i}=\frac{1}{N_{t,i}}\sum_{\ell=0}^{t-1}\mathds{1}(I_{\ell}=i)Y_{\ell+1,I_{\ell}}\quad\text{and}\quad\sigma_{t,i}^{2}=\frac{\sigma^{2}}{N_{t,i}}.$
Posterior probability of being the best arm.
To analyze the TTTS algorithm, we consider the posterior probability at time $t$ that arm $i$ is optimal, defined as
$\alpha_{t,i}\triangleq\mathbb{P}_{\widetilde{\bm{\theta}}\sim\Pi_{t}}\!\left(\widetilde{\theta}_{i}=\max_{j\in[K]}\widetilde{\theta}_{j}\right).$
E.1.2 Definition and Properties of Strong Convergence
The proof of our Theorem 5 involves bounding the time it takes for certain quantities to approach their limit values. Following Qin and Russo [2022], we first introduce a class of random variables that are “light-tailed.” Let the space $\mathcal{L}_{p}$ consist of all measurable $T$ with $\|T\|_{p}<\infty$ where $\|T\|_{p}\triangleq\left(\mathbb{E}\left[|T|^{p}\right]\right)^{1/p}$ is the $p$-norm.
Definition 6 (Definition 1 in Qin and Russo 2022).
For a real valued random variable $T$, we say $T\in\mathbb{M}$ if and only if $\|T\|_{p}<\infty$ for all $p\geq 1$. Equivalently, $\mathbb{M}=\cap_{p\geq 1}\mathcal{L}_{p}$.
We then introduce the concept of strong convergence for random variables—a stronger notion than $\mathcal{L}$-convergence (Definition 4) and almost sure convergence. While the sufficient condition for allocation rules to be optimal in Theorem 3 requires only $\mathcal{L}$-convergence, the proof of Theorem 5 establishes this stronger $\mathbb{M}$-convergence for our allocation rule.
Definition 7 (Definition 2 in Qin and Russo 2022: $\mathbb{M}$-convergence).
For a sequence of real valued random variables $\{X_{t}\}_{t\in\mathbb{N}_{0}}$ and a scalar $x\in\mathbb{R}$, we say $X_{t}\xrightarrow{\mathbb{M}}x$ if
for $\text{for all }\epsilon>0\text{ there exists }T\in\mathbb{M}\text{ such that for all }t\geq T,|X_{t}-x|\leq\epsilon.$
We say $X_{t}\xrightarrow{\mathbb{M}}\infty$ if
for $\text{for all }c>0\text{ there exists }T\in\mathbb{M}\text{ such that for all }t\geq T,X_{t}\geq c,$
and similarly, we say $X_{t}\xrightarrow{\mathbb{M}}-\infty$ if $-X_{t}\xrightarrow{\mathbb{M}}\infty$.
For a sequence of random vectors $\{\bm{X}_{t}\}_{t\in\mathbb{N}_{0}}$ taking values in $\mathbb{R}^{d}$ and a vector $\bm{x}\in\mathbb{R}^{d}$ where $\bm{X}_{t}=(X_{t,1},\ldots,X_{t,d})$ and $\bm{x}=(x_{1},\ldots,x_{d})$, we say $\bm{X}_{t}\xrightarrow{\mathbb{M}}\bm{x}$ if $X_{t,i}\xrightarrow{\mathbb{M}}x_{i}$ for all $i\in[d]$.
Corollary 2.
If $\bm{X}_{t}\xrightarrow{\mathbb{M}}\bm{x}$, then $\bm{X}_{t}\xrightarrow{\mathcal{L}}\bm{x}$ and $\bm{X}_{t}\xrightarrow{\mathrm{a.s.}}\bm{x}$.
Qin and Russo [2022] introduces a number of properties of the class $\mathbb{M}$ (Definition 6) and strong convergence (Definition 7).
Lemma 4 (Lemma 1 in Qin and Russo 2022: Closedness of $\mathbb{M}$).
Let $X$ and $Y$ be non-negative random variables in $\mathbb{M}$.
  1. 1.
    $aX+bY\in\mathbb{M}$ for any scalars $a,b\in\mathbb{R}$.
  2. 2.
    $XY\in\mathbb{M}$.
  3. 3.
    $\max\{X,Y\}\in\mathbb{M}$.
  4. 4.
    $X^{q}\in\mathbb{M}$, for any $q\geq 0$.
  5. 5.
    If $g:\mathbb{R}\to\mathbb{R}$ satisfies $\sup_{x\in-[c,c]}|g(x)|<\infty$ for all $c\geq 0$ and $|g(x)|=\mathcal{O}(|x|^{q})$ as $|x|\to\infty$ for some $q\geq 0$, then $g(X)\in\mathbb{M}$.
  6. 6.
    If $g:\mathbb{R}\to\mathbb{R}$ is continuous and $|g(x)|=\mathcal{O}(|x|^{q})$ as $|x|\to\infty$ for some $q\geq 0$, then $g(X)\in\mathbb{M}$.
As mentioned in Qin and Russo [2022], many of the above properties can be extended to any finite collection of random variables in $\mathbb{M}$. A notable one is that if $X_{i}\in\mathbb{M}$ for each $i\in[d]$ then
$X_{1}+\cdots+X_{d}\in\mathbb{M}\quad\text{and}\quad\max\{X_{1},\ldots,X_{d}\}\in\mathbb{M}.$
Qin and Russo [2022] also shows an equivalence between pointwise convergence and convergence in the maximum norm and a continuous mapping theorem under strong convergence (Definition 7).
Lemma 5 (Lemma 2 in Qin and Russo 2022).
For a sequence of random vectors $\{X_{n}\}_{n\in\mathbb{N}}$ taking values in $\mathbb{R}^{d}$ and a vector $x\in\mathbb{R}^{d}$, if $X_{n}\xrightarrow{\mathbb{M}}x$, then $\|X_{n}-x\|_{\infty}\xrightarrow{\mathbb{M}}0$. Equivalently, if $X_{n,i}\xrightarrow{\mathbb{M}}x_{i}$ for all $i\in[d]$, then for all $\epsilon>0$, there exists $N\in\mathbb{M}$ such that $n\geq N$ implies $|X_{n,i}-x_{i}|\leq\epsilon$ for all $i\in[d]$.
Lemma 6 (Lemma 3 in Qin and Russo 2022: Continuous Mapping).
For any $x$ taking values in a normed vector space and any random sequence $\{X_{n}\}_{n\in\mathbb{N}}$:
  1. 1.
    If $g$ is continuous at $x$, then $X_{n}\xrightarrow{\mathbb{M}}x$ implies $g(X_{n})\xrightarrow{\mathbb{M}}g(x)$.
  2. 2.
    If the range of function $g$ belongs to $\mathbb{R}$ and $g(y)\to\infty$ as $y\to\infty$, then $X_{n}\xrightarrow{\mathbb{M}}\infty$ implies $g(X_{n})\xrightarrow{\mathbb{M}}\infty$.
E.1.3 Maximal Inequalities
Following Qin et al. [2017] and Shang et al. [2020], we define
$W_{1}=\sup_{(t,i)\in\mathbb{N}_{0}\times[K]}\sqrt{\frac{N_{t,i}+1}{\log(N_{t,i}+e)}}\cdot\frac{|\theta_{t,i}-\theta_{i}|}{\sigma},$
which is a path-dependent random variable that controls the impact of observation noises. In addition, to control the impact of algorithmic randomness, we introduce the other path-dependent random variable:
$W_{2}=\sup_{(t,i)\in\mathbb{N}_{0}\times[K]}\frac{|N_{t,i}-\Psi_{t,i}|}{\sqrt{(t+1)\log(t+e^{2})}}.$ (21)
The next lemma ensures that these maximal deviations have light tails.
Lemma 7 (Lemma 6 of Qin et al. 2017 and Lemma 4 of Shang et al. 2020).
For any $\lambda>0$,
$\mathbb{E}[e^{\lambda W_{1}}]<\infty\quad\text{and}\quad\mathbb{E}[e^{\lambda W_{2}}]<\infty.$
The definition of $W_{2}$ in Equation (21) implies the following corollary for sufficiently large $t$.
Corollary 3.
There exists a deterministic $T$ such that for any $t\geq T$,
$|N_{t,i}-\Psi_{t,i}|\leq W_{2}t^{0.6},\quad\forall i\in[K].$
Corollary 3 implies that
$p_{t,i}-w_{t,i}\xrightarrow{\mathbb{M}}0,\quad\forall i\in[K].$
Note that the exponent $0.6$ in Corollary 3 is not essential—it can be replaced by any constant in the interval $(0.5,1)$.
E.2 A Restatement of Theorem 5
Alternative sufficient conditions for optimality.
Theorem 3 provides a sufficient condition for allocation rules to be optimal. Below, we present alternative but equivalent sufficient conditions.
Proposition 4 (Alternative sufficient conditions).
The following conditions are equivalent:
  1. 1.
    $\bm{p}_{t}\xrightarrow{\mathbb{M}}\bm{p}^{*}$;
  2. 2.
    $\frac{p_{t,j}}{p_{t,I^{*}}}\xrightarrow{\mathbb{M}}\frac{p_{j}^{*}}{p_{I^{*}}^{*}}$ for any $j\neq I^{*}$;
  3. 3.
    $\bm{w}_{t}\xrightarrow{\mathbb{M}}\bm{p}^{*}$; and
  4. 4.
    $\frac{w_{t,j}}{w_{t,I^{*}}}\xrightarrow{\mathbb{M}}\frac{p_{j}^{*}}{p_{I^{*}}^{*}}$ for any $j\neq I^{*}$.
Proof.
First, by Corollary 3, conditions 1 and 3 are equivalent.
We now show that conditions 1 and 2 are equivalent. By Lemma 4, condition 1 implies condition 2. Conversely, applying Lemma 4 to condition 2 yields
$\frac{1-p_{t,I^{*}}}{p_{t,I^{*}}}=\sum_{j\neq I^{*}}\frac{p_{t,j}}{p_{t,I^{*}}}\xrightarrow{\mathbb{M}}\sum_{j\neq I^{*}}\frac{p^{*}_{j}}{p^{*}_{I^{*}}}=\frac{1-p^{*}_{I^{*}}}{p^{*}_{I^{*}}},$
Using Lemma 4 again, we obtain $p_{t,I^{*}}\xrightarrow{\mathbb{M}}p^{*}_{I^{*}}.$ Then, by condition 2, it follows that $p_{t,j}\xrightarrow{\mathbb{M}}p_{j}^{*}$ for any $j\neq I^{*}$ Therefore, condition 2 implies condition 1, establishing their equivalence. By a similar argument, conditions 2 and 4 are also equivalent. This completes the proof. ∎
To prove Theorem 5, it suffices to show that TTTS-IDS satisfies condition 4 in Proposition 4:
Proposition 5 (A restatement of Theorem 5).
Under TTTS-IDS,
$\frac{w_{t,j}}{w_{t,I^{*}}}\xrightarrow{\mathbb{M}}\frac{p_{j}^{*}}{p_{I^{*}}^{*}},\quad\forall j\neq I^{*}.$
The rest of this appendix is devoted to proving Proposition 5.
E.3 Empirical Overall Balance under TTTS-IDS
Previous papers [Qin et al., 2017, Shang et al., 2020, Jourdan et al., 2022] establishes $\beta$-optimality by proving either $w_{t,I^{*}}\xrightarrow{\mathbb{M}}\beta$ or $p_{t,I^{*}}\xrightarrow{\mathbb{M}}\beta$ where $\beta$ is a fixed tuning parameter taken as an input to the algorithms, while the optimal parameter $\beta^{*}$ (i.e., $p^{*}_{I^{*}}$) is unknown a priori. Our information-directed selection rule adapts to the top-two candidates at each time step, which necessitates novel proof techniques. To address this, we first establish that the “empirical” version of the overall balance condition (3) holds.
Proposition 6.
Under TTTS-IDS,
$w_{t,I^{*}}^{2}-\sum_{j\neq I^{*}}w_{t,j}^{2}\xrightarrow{\mathbb{M}}0\quad\text{and}\quad p_{t,I^{*}}^{2}-\sum_{j\neq I^{*}}p_{t,j}^{2}\xrightarrow{\mathbb{M}}0.$
To prove this proposition, we need the following supporting result.
Lemma 8.
Under TTTS-IDS, for any $t\in\mathbb{N}_{0}$,
$\alpha_{t,I^{*}}\left(\sum_{\ell\neq I^{*}}\frac{\alpha_{t,\ell}}{1-\alpha_{t,I^{*}}}\frac{p_{t,\ell}}{p_{t,I^{*}}+p_{t,\ell}}\right)\leq\psi_{t,I^{*}}\leq\alpha_{t,I^{*}}\left(\sum_{\ell\neq I^{*}}\frac{\alpha_{t,\ell}}{1-\alpha_{t,I^{*}}}\frac{p_{t,\ell}}{p_{t,I^{*}}+p_{t,\ell}}\right)+(1-\alpha_{t,I^{*}}),$
and for any $j\neq I^{*}$,
$\alpha_{t,I^{*}}\frac{\alpha_{t,j}}{1-\alpha_{t,I^{*}}}\frac{p_{t,I^{*}}}{p_{t,I^{*}}+p_{t,j}}\leq\psi_{t,j}\leq\alpha_{t,I^{*}}\frac{\alpha_{t,j}}{1-\alpha_{t,I^{*}}}\frac{p_{t,I^{*}}}{p_{t,I^{*}}+p_{t,j}}+(1-\alpha_{t,I^{*}}).$
Proof.
Fix $t\in\mathbb{N}_{0}$. Recall that $\alpha_{t,i}$ is the posterior probability at time $t$ that arm $i$ appears as the best arm. For a fixed $j\in[K]$ and any $\ell\neq j$, TTTS obtains the top-two candidates
$\left(I_{t}^{(1)},I_{t}^{(2)}\right)=\begin{cases}(j,\ell),&\text{with probability }\alpha_{t,j}\frac{\alpha_{t,\ell}}{1-\alpha_{t,j}},\\ (\ell,j),&\text{with probability }\alpha_{t,\ell}\frac{\alpha_{t,j}}{1-\alpha_{t,\ell}},\end{cases}$
and when TTTS is applied with IDS,
$\psi_{t,j}=\alpha_{t,j}\left(\sum_{\ell\neq j}\frac{\alpha_{t,\ell}}{1-\alpha_{t,j}}\frac{p_{t,\ell}}{p_{t,j}+p_{t,\ell}}\right)+\sum_{\ell\neq j}\alpha_{t,\ell}\frac{\alpha_{t,j}}{1-\alpha_{t,\ell}}\frac{p_{t,\ell}}{p_{t,j}+p_{t,\ell}}.$
For $j=I^{*}$, we have
$\psi_{t,I^{*}}$ $=\alpha_{t,I^{*}}\left(\sum_{\ell\neq I^{*}}\frac{\alpha_{t,\ell}}{1-\alpha_{t,I^{*}}}\frac{p_{t,\ell}}{p_{t,I^{*}}+p_{t,\ell}}\right)+\sum_{\ell\neq I^{*}}\alpha_{t,\ell}\frac{\alpha_{t,I^{*}}}{1-\alpha_{t,\ell}}\frac{p_{t,\ell}}{p_{t,I^{*}}+p_{t,\ell}},$
so
$\psi_{t,I^{*}}$ $\geq\alpha_{t,I^{*}}\left(\sum_{\ell\neq I^{*}}\frac{\alpha_{t,\ell}}{1-\alpha_{t,I^{*}}}\frac{p_{t,\ell}}{p_{t,I^{*}}+p_{t,\ell}}\right)$
and
$\psi_{t,I^{*}}$ $\leq\alpha_{t,I^{*}}\left(\sum_{\ell\neq I^{*}}\frac{\alpha_{t,\ell}}{1-\alpha_{t,I^{*}}}\frac{p_{t,\ell}}{p_{t,I^{*}}+p_{t,\ell}}\right)+\sum_{\ell\neq I^{*}}\alpha_{t,\ell}$
$=\alpha_{t,I^{*}}\left(\sum_{\ell\neq I^{*}}\frac{\alpha_{t,\ell}}{1-\alpha_{t,I^{*}}}\frac{p_{t,\ell}}{p_{t,I^{*}}+p_{t,\ell}}\right)+(1-\alpha_{t,I^{*}}),$
where the inequality uses that for any $\ell\neq I^{*}$, $\frac{\alpha_{t,I^{*}}}{1-\alpha_{t,\ell}}\frac{p_{t,\ell}}{p_{t,I^{*}}+p_{t,\ell}}\leq\frac{\alpha_{t,I^{*}}}{1-\alpha_{t,\ell}}=\frac{\alpha_{t,I^{*}}}{\sum_{\ell^{\prime}\neq\ell}\alpha_{t,\ell^{\prime}}}\leq 1$.
For $j\neq I^{*}$, we have
$\psi_{t,j}$ $=\sum_{\ell\neq j}\alpha_{t,\ell}\frac{\alpha_{t,j}}{1-\alpha_{t,\ell}}\frac{p_{t,\ell}}{p_{t,j}+p_{t,\ell}}+\alpha_{t,j}\left(\sum_{\ell\neq j}\frac{\alpha_{t,\ell}}{1-\alpha_{t,j}}\frac{p_{t,\ell}}{p_{t,j}+p_{t,\ell}}\right)$
$=\alpha_{t,I^{*}}\frac{\alpha_{t,j}}{1-\alpha_{t,I^{*}}}\frac{p_{t,I^{*}}}{p_{t,j}+p_{t,I^{*}}}+\sum_{\ell\neq j,I^{*}}\alpha_{t,\ell}\frac{\alpha_{t,j}}{1-\alpha_{t,\ell}}\frac{p_{t,\ell}}{p_{t,\ell}+p_{t,j}}+\alpha_{t,j}\left(\sum_{\ell\neq j}\frac{\alpha_{t,\ell}}{1-\alpha_{t,j}}\frac{p_{t,\ell}}{p_{t,j}+p_{t,\ell}}\right),$
so
$\psi_{t,j}$ $\geq\alpha_{t,I^{*}}\frac{\alpha_{t,j}}{1-\alpha_{t,I^{*}}}\frac{p_{t,I^{*}}}{p_{t,j}+p_{t,I^{*}}}$
and
$\psi_{t,j}$ $\leq\alpha_{t,I^{*}}\frac{\alpha_{t,j}}{1-\alpha_{t,I^{*}}}\frac{p_{t,I^{*}}}{p_{t,j}+p_{t,I^{*}}}+\left(\sum_{\ell\neq j,I^{*}}\alpha_{t,\ell}\right)+\alpha_{t,j}=\alpha_{t,I^{*}}\frac{\alpha_{t,j}}{1-\alpha_{t,I^{*}}}\frac{p_{t,I^{*}}}{p_{t,j}+p_{t,I^{*}}}+(1-\alpha_{t,I^{*}}),$
where the inequality uses that
  1. 1.
    for any $\ell\neq j,I^{*}$, $\frac{\alpha_{t,j}}{1-\alpha_{t,\ell}}\frac{p_{t,\ell}}{p_{t,\ell}+p_{t,j}}\leq\frac{\alpha_{t,j}}{1-\alpha_{t,\ell}}=\frac{\alpha_{t,j}}{\sum_{\ell^{\prime}\neq\ell}\alpha_{t,\ell^{\prime}}}\leq 1$; and
  2. 2.
    $\sum_{\ell\neq j}\frac{\alpha_{t,\ell}}{1-\alpha_{t,j}}\frac{p_{t,\ell}}{p_{t,j}+p_{t,\ell}}\leq\sum_{\ell\neq j}\frac{\alpha_{t,\ell}}{1-\alpha_{t,j}}=1$.
This completes the proof. ∎
Now we are ready to complete the proof of Proposition 6.
Proof of Proposition 6.
By Corollary 3, it suffices to only show
$w_{t,I^{*}}^{2}-\sum_{j\neq I^{*}}w_{t,j}^{2}\xrightarrow{\mathbb{M}}0.$
For any $t\in\mathbb{N}_{0}$, we let
$G_{t}\triangleq{\Psi_{t,I^{*}}^{2}}-\sum_{j\neq I^{*}}{\Psi_{t,j}^{2}},$
and we are going to show $\frac{G_{t}}{t^{2}}\xrightarrow{\mathbb{M}}0$.
We argue that it suffices to show that there exists $T\in\mathbb{M}$ such that for any $t\geq T$,
$|G_{t+1}-G_{t}|\leq 2W_{2}t^{0.6}+2.$ (22)
The reason is that given (22), for any $t\geq T$
$|G_{t}|=\left|G_{T}+\sum_{\ell=T}^{t-1}(G_{\ell+1}-G_{\ell})\right|$ $\leq\left|G_{T}\right|+\sum_{\ell=T}^{t-1}|G_{\ell+1}-G_{\ell}|$
$\leq\left|G_{T}\right|+\sum_{\ell=T}^{t-1}\left(2W_{2}\ell^{0.6}+2\right)$
$\leq\left|G_{T}\right|+2(t-T)+2W_{2}\int_{T+1}^{t+1}x^{0.6}\mathrm{d}x$
$=\left|G_{T}\right|+2(t-T)+\frac{2W_{2}}{1.6}\left[(t+1)^{1.6}-(T+1)^{1.6}\right]$
$\leq T^{2}+2(t-T)+\frac{2W_{2}}{1.6}\left[(t+1)^{1.6}-(T+1)^{1.6}\right],$
where in the last inequality, $\left|G_{T}\right|\leq T^{2}$ holds since $G_{T}\geq-\sum_{j\neq I^{*}}\Psi_{T,j}^{2}\geq-\left(\sum_{j\neq I^{*}}\Psi_{T,j}\right)^{2}\geq-T^{2},$ and $G_{T}\leq\Psi_{T,I^{*}}^{2}\leq T^{2}.$ This gives $\frac{|G_{t}|}{t^{2}}\xrightarrow{\mathbb{M}}0.$
Now we are going to prove the sufficient condition (22). We can calculate
$G_{t+1}-G_{t}$ $=\left[\left(\Psi_{t,I^{*}}+\psi_{t,I^{*}}\right)^{2}-\Psi_{t,I^{*}}^{2}\right]-\sum_{j\neq I^{*}}\left[\left(\Psi_{t,j}+\psi_{t,j}\right)^{2}-\Psi_{t,j}^{2}\right]$
$=2\left(\psi_{t,I^{*}}\Psi_{t,I^{*}}-\sum_{j\neq I^{*}}\psi_{t,j}\Psi_{t,j}\right)+\left(\psi_{t,I^{*}}^{2}-\sum_{j\neq I^{*}}\psi_{t,j}^{2}\right).$
Then by Lemma 8, we have
$G_{t+1}-G_{t}$
$\geq$ $1\displaystyle 2\left[\alpha_{t,I^{*}}\sum_{j\neq I^{*}}\frac{\alpha_{t,j}}{1-\alpha_{t,I^{*}}}\frac{p_{t,j}\Psi_{t,I^{*}}}{p_{t,I^{*}}+p_{t,j}}-\alpha_{t,I^{*}}\sum_{j\neq I^{*}}\frac{\alpha_{t,j}}{1-\alpha_{t,I^{*}}}\frac{p_{t,I^{*}}\Psi_{t,j}}{p_{t,I^{*}}+p_{t,j}}-(1-\alpha_{t,I^{*}})\sum_{j\neq I^{*}}\Psi_{t,j}\right]-$ 1
$\geq$ $2\alpha_{t,I^{*}}\sum_{j\neq I^{*}}\frac{\alpha_{t,j}}{1-\alpha_{t,I^{*}}}\frac{p_{t,j}\Psi_{t,I^{*}}-p_{t,I^{*}}\Psi_{t,j}}{p_{t,I^{*}}+p_{t,j}}-2(1-\alpha_{t,I^{*}})t-1,$
where the last inequality uses $\sum_{j\neq I^{*}}\Psi_{t,j}\leq t$, and
$G_{t+1}-G_{t}$
$\leq$ $1\displaystyle 2\left[\alpha_{t,I^{*}}\sum_{j\neq I^{*}}\frac{\alpha_{t,j}}{1-\alpha_{t,I^{*}}}\frac{p_{t,j}\Psi_{t,I^{*}}}{p_{t,I^{*}}+p_{t,j}}+(1-\alpha_{t,I^{*}})\Psi_{t,I^{*}}-\alpha_{t,I^{*}}\sum_{j\neq I^{*}}\frac{\alpha_{t,j}}{1-\alpha_{t,I^{*}}}\frac{p_{t,I^{*}}\Psi_{t,j}}{p_{t,I^{*}}+p_{t,j}}\right]+$ 1
$\leq$ $2\alpha_{t,I^{*}}\left(\sum_{j\neq I^{*}}\frac{\alpha_{t,j}}{1-\alpha_{t,I^{*}}}\frac{p_{t,j}\Psi_{t,I^{*}}-p_{t,I^{*}}\Psi_{t,j}}{p_{t,I^{*}}+p_{t,j}}\right)+2(1-\alpha_{t,I^{*}})t+1,$
where the last inequality uses $\Psi_{t,I^{*}}\leq t$. Hence,
$|G_{t+1}-G_{t}|$ $\leq 2\alpha_{t,I^{*}}\sum_{j\neq I^{*}}\frac{\alpha_{t,j}}{1-\alpha_{t,I^{*}}}\left|\frac{p_{t,j}\Psi_{t,I^{*}}-p_{t,I^{*}}\Psi_{t,j}}{p_{t,I^{*}}+p_{t,j}}\right|+2(1-\alpha_{t,I^{*}})t+1.$
By Lemma 12, there exists $T_{1}\in\mathbb{M}$ such that for any $t\geq T_{1}$, $(1-\alpha_{t,I^{*}})t\leq 1/2$, and thus
$|G_{t+1}-G_{t}|$ $\leq 2\alpha_{t,I^{*}}\sum_{j\neq I^{*}}\frac{\alpha_{t,j}}{1-\alpha_{t,I^{*}}}\max_{j\neq I^{*}}\left|\frac{p_{t,j}\Psi_{t,I^{*}}-p_{t,I^{*}}\Psi_{t,j}}{p_{t,I^{*}}+p_{t,j}}\right|+2$
$\leq 2\max_{j\neq I^{*}}\left|\frac{p_{t,j}\Psi_{t,I^{*}}-p_{t,I^{*}}\Psi_{t,j}}{p_{t,I^{*}}+p_{t,j}}\right|+2=2\max_{j\neq I^{*}}\left|\frac{N_{t,j}\Psi_{t,I^{*}}-N_{t,I^{*}}\Psi_{t,j}}{N_{t,I^{*}}+N_{t,j}}\right|+2,$
where the last inequality applies $\alpha_{t,I^{*}}\leq 1$ and $\sum_{j\neq I^{*}}\frac{\alpha_{t,j}}{1-\alpha_{t,I^{*}}}=1$. By Corollary 3, there exists $T_{2}\in\mathbb{M}$ such that for any $t\geq T_{2}$ and $i\in[K]$, $|N_{t,i}-\Psi_{t,i}|\leq W_{2}t^{0.6}$, and thus for $j\neq I^{*}$
$\left|\frac{N_{t,j}\Psi_{t,I^{*}}-N_{t,I^{*}}\Psi_{t,j}}{N_{t,I^{*}}+N_{t,j}}\right|\leq\left|\frac{N_{t,j}\left(N_{t,I^{*}}+W_{2}t^{0.6}\right)-N_{t,I^{*}}\left(N_{t,j}-W_{2}t^{0.6}\right)}{N_{t,I^{*}}+N_{t,j}}\right|\leq W_{2}t^{0.6}.$
Hence, for $t\geq T\triangleq\max\{T_{1},T_{2}\}$, $|G_{t+1}-G_{t}|\leq 2W_{2}t^{0.6}+2$. Note that $T\in\mathbb{M}$ by Lemma 4. This completes the proof of the sufficient condition (22). ∎
E.3.1 Implication of Empirical Overall Balance
Here, we present a result implied by the empirical overall balance (Proposition 6) that is needed to prove the sufficient condition in Proposition 5: for sufficiently large $t$, two measures of average effort allocated to the best arm remain bounded away from 0 and 1.
Lemma 9.
Let $b_{1}\triangleq\frac{1}{\sqrt{32(K-1)}}$ and $b_{2}\triangleq\frac{3}{4}$. Under TTTS-IDS, there exists a random time $T\in\mathbb{M}$ such that for any $t\geq T$,
$b_{1}\leq w_{t,I^{*}}\leq b_{2}\quad\text{and}\quad b_{1}\leq p_{t,I^{*}}\leq b_{2}.$
Since $w_{t,I^{*}}$ and $p_{t,I^{*}}$ are bounded away from 0 for sufficiently large $t$, we derive the following alternative forms of the empirical overall balance (Proposition 6).
Corollary 4.
Under TTTS-IDS,
$\sum_{j\neq I^{*}}\frac{w_{t,j}^{2}}{w_{t,I^{*}}^{2}}\xrightarrow{\mathbb{M}}1\quad\text{and}\quad\sum_{j\neq I^{*}}\frac{p_{t,j}^{2}}{p_{t,I^{*}}^{2}}\xrightarrow{\mathbb{M}}1.$
We now proceed to prove Lemma 9 as follows.
Proof of Lemma 9.
We only show the upper and lower bounds for $w_{t,I^{*}}$ since proving those for $p_{t,I^{*}}$ is the same.
By Proposition 6, there exists $T_{1}\in\mathbb{M}$ such that for any $t\geq T_{1}$,
$w_{t,I^{*}}^{2}-\sum_{j\neq I^{*}}w_{t,j}^{2}\leq\frac{1}{2}.$
This implies
$\frac{1}{2}\geq w_{t,I^{*}}^{2}-\left(\sum_{j\neq I^{*}}w_{t,j}\right)^{2}=w_{t,I^{*}}^{2}-\left(1-w_{t,I^{*}}\right)^{2}=2w_{t,I^{*}}-1,$
which gives $w_{t,I^{*}}\leq b_{2}\triangleq\frac{3}{4}$.
Using Proposition 6 again, there exists $T_{2}\in\mathbb{M}$ such that for any $t\geq T_{2}$,
$w_{t,I^{*}}^{2}-\sum_{j\neq I^{*}}w_{t,j}^{2}\geq-\frac{1}{32(K-1)}.$
By Cauchy-Schwarz inequality,
$w_{t,I^{*}}^{2}\geq-\frac{1}{32(K-1)}+\frac{\left(\sum_{j\neq I^{*}}w_{t,j}\right)^{2}}{K-1}$ $=-\frac{1}{32(K-1)}+\frac{\left(1-w_{t,I^{*}}\right)^{2}}{K-1}$
$\geq-\frac{1}{32(K-1)}+\frac{1}{16(K-1)}=\frac{1}{32(K-1)},$
where the last inequality uses $w_{t,I^{*}}\leq b_{2}=\frac{3}{4}$. This gives $w_{t,I^{*}}\geq b_{1}\triangleq\frac{1}{\sqrt{32(K-1)}}$. Taking $T\triangleq\max\{T_{1},T_{2}\}$ completes the proof since $T\in\mathbb{M}$ by Lemma 4. ∎
E.4 Sufficient Exploration under TTTS-IDS
In this subsection, we show that TTTS-IDS sufficiently explores all arms:
Proposition 7.
Under TTTS-IDS, there exists $T\in\mathbb{M}$ such that for any $t\geq T$,
$\min_{i\in[K]}N_{t,i}\geq\sqrt{t/K}.$
The previous analyses in Qin et al. [2017], Shang et al. [2020], Jourdan et al. [2022] only work for any sequence of tuning parameters $\{h_{t}\}_{t\in\mathbb{N}_{0}}$ that is uniformly strictly bounded, i.e., there exists $h_{\min}>0$ such that with probability one,
$\inf_{t\in\mathbb{N}_{0}}\min\{h_{t},1-h_{t}\}\geq h_{\min}.$
Our IDS is adaptive to the algorithmic randomness in TTTS in the sense that it depends on the algorithmic randomness in picking top-two candidates, so it does not enjoy the property above. We need some novel analysis to prove the sufficient exploration of TTTS-IDS.
Following the analysis of TTTS by Shang et al. [2020], we define the following two “most promising arms”:
$J_{t}^{(1)}\in\operatorname*{argmax}_{i\in[K]}\alpha_{t,i}\quad\text{and}\quad J_{t}^{(2)}\in\operatorname*{argmax}_{i\neq J_{t}^{(1)}}\alpha_{t,i}.$
We further define the one that is less sampled as:
$J_{t}\triangleq\operatorname*{argmin}_{i\in\{J_{t}^{(1)},J_{t}^{(2)}\}}N_{t,i}=\operatorname*{argmin}_{i\in\{J_{t}^{(1)},J_{t}^{(2)}\}}p_{t,i}.$
Even though $J_{t}$ may change over time, we show that IDS allocates a decent effort to $J_{t}$.
Lemma 10.
Under TTTS-IDS, for any $t\in\mathbb{N}_{0}$,
$\psi_{t,J_{t}}\geq\frac{1}{2K(K-1)}.$
Proof.
Fix $t\in\mathbb{N}_{0}$. We have
$\left(I_{t}^{(1)},I_{t}^{(2)}\right)=\left(J_{t}^{(1)},J_{t}^{(2)}\right)\quad\text{with probability}\quad\alpha_{t,J_{t}^{(1)}}\frac{\alpha_{t,J_{t}^{(2)}}}{1-\alpha_{t,J_{t}^{(1)}}}\geq\frac{1}{K(K-1)},$
where the inequality follows from the definition of $J_{t}^{(1)}$ and $J_{t}^{(2)}$. Note that
$h_{J_{t}^{(1)}}^{J_{t}^{(2)}}(\bm{p}_{t};\bm{\theta}_{t})=\frac{p_{t,J_{t}^{(2)}}}{p_{t,J_{t}^{(1)}}+p_{t,J_{t}^{(2)}}}\quad\text{and}\quad h_{J_{t}^{(2)}}^{J_{t}^{(2)}}(\bm{p}_{t};\bm{\theta}_{t})=\frac{p_{t,J_{t}^{(1)}}}{p_{t,J_{t}^{(1)}}+p_{t,J_{t}^{(2)}}}.$
If $J_{t}=\operatorname*{argmin}_{i\in\left\{J_{t}^{(1)},J_{t}^{(2)}\right\}}p_{t,i}=J_{t}^{(1)}$,
$\psi_{t,J_{t}}\geq\frac{1}{K(K-1)}\frac{p_{t,J_{t}^{(2)}}}{p_{t,J_{t}^{(1)}}+p_{t,J_{t}^{(2)}}}\geq\frac{1}{2K(K-1)}.$
Otherwise, $J_{t}=\operatorname*{argmin}_{i\in\left\{J_{t}^{(1)},J_{t}^{(2)}\right\}}p_{t,i}=J_{t}^{(2)}$ and
$\psi_{t,J_{t}}\geq\frac{1}{K(K-1)}\frac{p_{t,J_{t}^{(1)}}}{p_{t,J_{t}^{(1)}}+p_{t,J_{t}^{(2)}}}\geq\frac{1}{2K(K-1)}.$
This completes the proof. ∎
Following Qin et al. [2017], we define an insufficiently sampled set for any $t\in\mathbb{N}_{0}$ and $s\geq 0$:
$U_{t}^{s}\triangleq\{i\in[K]:N_{t,i}<s^{1/2}\}.$
Lemmas 9 and 11 in Shang et al. [2020] establish a key property of TTTS:
Lemma 11.
Under TTTS with any selection rule (e.g., IDS) such that
$\exists\psi_{\min}>0\quad\forall t\in\mathbb{N}_{0}\quad\psi_{t,J_{t}}\geq\psi_{\min},$
there exists $S\in\mathbb{M}$ such that for any $s\geq S$, $U^{s}_{\lfloor Ks\rfloor}=\emptyset$.
We remark that, while the original results focus on $\beta$-tuning, the same proofs extend to any selection rule that satisfies the stated condition.
We are now ready to complete the proof of Proposition 7.
Proof of Proposition 7.
By Lemma 10, IDS satisfies the condition in Lemma 11. Take the corresponding $S$ in Lemma 11 for IDS, and let $T\triangleq KS$. For any $t\geq T$, we let $s=t/K\geq S$, and then by Lemma 11, we have $U^{s}_{\lfloor Ks\rfloor}=U_{t}^{t/K}$ is empty. ∎
E.4.1 Implication of Sufficient Exploration
Here we present some results that are implied by sufficient exploration (Proposition 7) and needed for proving the sufficient condition for optimality in fixed-confidence setting (Proposition 5). With sufficient exploration, the posterior means strongly converge to the unknown true means, and the probability of any sub-optimal arm being the best decays exponentially.
Lemma 12 (Lemmas 6 and 12 in Shang et al. 2020).
Under TTTS with any selection rule that ensures sufficient exploration (e.g., IDS), i.e.,
$\exists\widetilde{T}\in\mathbb{M}\text{ such that for all }t\geq\widetilde{T}\text{ we have }\min_{i\in[K]}N_{t,i}\geq\sqrt{t/K},$
the following properties hold:
  1. 1.
    $\bm{\theta}_{t}\xrightarrow{\mathbb{M}}\bm{\theta}$, and
  2. 2.
    there exists $T\in\mathbb{M}$ such that for any $t\geq T$ and $j\neq I^{*}$,
    $\alpha_{t,j}\leq\exp\left(-ct^{1/2}\right),\quad\text{where}\quad c\triangleq\frac{\Delta_{\min}^{2}}{16\sigma^{2}K^{1/2}}.$
We remark that, while the original results focus on $\beta$-tuning, the same proofs extend to any selection rule that ensures sufficient exploration.
E.5 Strong Convergence to Optimal Proportions: Completing the Proof of Proposition 5
In this subsection, we complete the proof of Proposition 5 by establishing a sequence of supporting results.
Lemma 13.
Under TTTS-IDS, for any arm $j\neq I^{*}$,
$\frac{C_{t,j}}{p_{t,I^{*}}\cdot f_{j}\left(\frac{p_{t,j}}{p_{t,I^{*}}}\right)}\xrightarrow{\mathbb{M}}1,\quad\text{where}\quad f_{j}(x)\triangleq\frac{\left(\theta_{I^{*}}-\theta_{j}\right)^{2}}{2\sigma^{2}\left(1+\frac{1}{x}\right)}\quad\text{and}\quad C_{t,j}=\frac{\left(\theta_{t,I^{*}}-\theta_{t,j}\right)^{2}}{2\sigma^{2}\left(\frac{1}{p_{t,I^{*}}}+\frac{1}{p_{t,j}}\right)}.$
Proof.
Fix $j\neq I^{*}$. We have
$p_{t,I^{*}}f_{j}\left(\frac{p_{t,j}}{p_{t,I^{*}}}\right)=p_{t,I^{*}}\frac{\left(\theta_{I^{*}}-\theta_{j}\right)^{2}}{2\sigma^{2}\left(1+\frac{p_{t,I^{*}}}{p_{t,j}}\right)}=\frac{\left(\theta_{I^{*}}-\theta_{j}\right)^{2}}{2\sigma^{2}\left(\frac{1}{p_{t,I^{*}}}+\frac{1}{p_{t,j}}\right)}.$
By Proposition 7, TTTS-IDS guarantees sufficient exploration. Consequently, Lemma 12 implies that $\bm{\theta}_{t}\xrightarrow{\mathbb{M}}\bm{\theta}$. Applying the continuous mapping theorem (Lemma 6) then yields the desired convergence. ∎
Lemma 14.
Under TTTS-IDS, for any $\epsilon>0$, there exists a deterministic constant $c_{\epsilon}>0$ and a random time $T_{\epsilon}\in\mathbb{M}$ such that for any $t\geq T_{\epsilon}$ and $j\neq I^{*}$,
$\frac{w_{t,j}}{w_{t,I^{*}}}>\frac{p_{j}^{*}+\epsilon}{p_{I^{*}}^{*}}\quad\implies\quad\psi_{t,j}\leq\exp\left(-c_{\epsilon}t\right)+(K-1)\exp\left(-ct^{1/2}\right),$
where $c$ is defined in Lemma 12.
Proof.
Fix $\epsilon>0$. It suffices to show that for any $j\neq I^{*}$, there exist a deterministic constant $c^{(j)}_{\epsilon}>0$ and a random time $T^{(j)}_{\epsilon}\in\mathbb{M}$ such that for any $t\geq T^{(j)}_{\epsilon}$,
$\frac{w_{t,j}}{w_{t,I^{*}}}>\frac{p_{j}^{*}+\epsilon}{p_{I^{*}}^{*}}\quad\implies\quad\psi_{t,j}\leq\exp\left(-c_{\epsilon}^{(j)}t\right)+(K-1)\exp\left(-ct^{1/2}\right),$
since this completes the proof by taking $T_{\epsilon}\triangleq\max_{j\neq I^{*}}T^{(j)}_{\epsilon}$ and $c_{\epsilon}\triangleq\min_{j\neq I^{*}}c^{(j)}_{\epsilon}$.
From now on, we fix $j\neq I^{*}$. By Corollary 3, there exists $T_{\epsilon,1}^{(j)}\in\mathbb{M}$ such that for any $t\geq T_{\epsilon,1}^{(j)}$,
$\frac{w_{t,j}}{w_{t,I^{*}}}=\frac{\Psi_{t,j}}{\Psi_{t,I^{*}}}>\frac{p_{j}^{*}+\epsilon}{p_{I^{*}}^{*}}\quad\implies\quad\frac{p_{t,j}}{p_{t,I^{*}}}=\frac{N_{t,j}}{N_{t,I^{*}}}>\frac{p_{j}^{*}+\epsilon/2}{p_{I^{*}}^{*}}.$
Then by Corollary 4, there exists $T_{\epsilon,2}^{(j)}\in\mathbb{M}$ such that for any $t\geq T_{\epsilon,2}^{(j)}$,
$\frac{p_{t,j}}{p_{t,I^{*}}}>\frac{p_{j}^{*}+\epsilon/2}{p_{I^{*}}^{*}}\quad\implies\quad\exists A_{t}\neq I^{*}:\frac{p_{t,A_{t}}}{p_{t,I^{*}}}\leq\frac{p_{A_{t}}^{*}}{p_{I^{*}}^{*}}.$
From now on, we consider $t\geq\max\left\{T_{\epsilon,1}^{(j)},T_{\epsilon,2}^{(j)}\right\}$ and we have
$\frac{w_{t,j}}{w_{t,I^{*}}}=\frac{\Psi_{t,j}}{\Psi_{t,I^{*}}}>\frac{p_{j}^{*}+\epsilon}{p_{I^{*}}^{*}}\implies\frac{p_{t,j}}{p_{t,I^{*}}}=\frac{N_{t,j}}{N_{t,I^{*}}}>\frac{p_{j}^{*}+\epsilon/2}{p_{I^{*}}^{*}}\text{ and }\exists A_{t}\neq I^{*}:\frac{p_{t,A_{t}}}{p_{t,I^{*}}}\leq\frac{p_{A_{t}}^{*}}{p_{I^{*}}^{*}}.$ (23)
By Lemmas 8 and 12,
$\psi_{t,j}\leq\alpha_{t,I^{*}}\frac{\alpha_{t,j}}{1-\alpha_{t,I^{*}}}\frac{p_{t,I^{*}}}{p_{t,I^{*}}+p_{t,j}}+(1-\alpha_{t,I^{*}})\leq\frac{\alpha_{t,j}}{1-\alpha_{t,I^{*}}}+(K-1)\exp\left(-ct^{1/2}\right).$
We can upper bound the numerator $\alpha_{t,j}$ as follows,
$\alpha_{t,j}\leq\mathbb{P}_{t}\left(\widetilde{\theta}_{j}\geq\widetilde{\theta}_{I^{*}}\right)$ $=\mathbb{P}\left(\frac{\widetilde{\theta}_{I^{*}}-\widetilde{\theta}_{j}-(\theta_{t,I^{*}}-\theta_{t,j})}{\sigma\sqrt{1/N_{t,I^{*}}+1/N_{t,j}}}\leq\frac{-(\theta_{t,I^{*}}-\theta_{t,j})}{\sigma\sqrt{1/N_{t,I^{*}}+1/N_{t,j}}}\right)=\Phi\left(-\sqrt{2tC_{t,j}}\right),$
where $\widetilde{\bm{\theta}}=\bigl{(}\widetilde{\theta}_{1},\ldots,\widetilde{\theta}_{K}\bigr{)}$ is a sample drawn from the posterior distribution $\Pi_{t}$ and $\Phi(\cdot)$ is Gaussian cumulative distribution function. Similarly, we can lower bound the denominator $1-\alpha_{t,I^{*}}$ as follows,
$1-\alpha_{t,I^{*}}=\mathbb{P}_{t}\left(\exists a\neq I^{*}:\widetilde{\theta}_{a}\geq\widetilde{\theta}_{I^{*}}\right)\geq\mathbb{P}_{t}\left(\widetilde{\theta}_{A_{t}}\geq\widetilde{\theta}_{I^{*}}\right)=\Phi\left(-\sqrt{2tC_{t,A_{t}}}\right).$
Hence,
$\psi_{t,j}\leq\frac{\Phi\left(-\sqrt{2tC_{t,j}}\right)}{\Phi\left(-\sqrt{2tC_{t,A_{t}}}\right)}+(K-1)\exp\left(-ct^{1/2}\right).$
The remaining task is to show that there exist a deterministic constant $c^{(j)}_{\epsilon}>0$ and a random time $T^{(j)}_{\epsilon,3}\in\mathbb{M}$ such that for any $t\geq T^{(j)}_{\epsilon,3}$,
$\frac{\Phi\left(-\sqrt{2tC_{t,j}}\right)}{\Phi\left(-\sqrt{2tC_{t,A_{t}}}\right)}\leq\exp\left(-c_{\epsilon}^{(j)}t\right).$ (24)
By Lemma 13, for any $\delta>0$,
$\frac{\Phi\left(-\sqrt{2tC_{t,j}}\right)}{\Phi\left(-\sqrt{2tC_{t,A_{t}}}\right)}\leq\frac{\Phi\left(-\sqrt{2tp_{t,I^{*}}f_{j}\left(\frac{p_{t,j}}{p_{t,I^{*}}}\right)(1-\delta)}\right)}{\Phi\left(-\sqrt{2tp_{t,I^{*}}f_{A_{t}}\left(\frac{p_{t,A_{t}}}{p_{t,I^{*}}}\right)(1+\delta)}\right)}$ $\leq\frac{\Phi\left(-\sqrt{2tp_{t,I^{*}}f_{j}\left(\frac{p^{*}_{j}+\epsilon/2}{p^{*}_{I^{*}}}\right)(1-\delta)}\right)}{\Phi\left(-\sqrt{2tp_{t,I^{*}}f_{A_{t}}\left(\frac{p^{*}_{A_{t}}}{p^{*}_{I^{*}}}\right)(1+\delta)}\right)}$
$=\frac{\Phi\left(-\sqrt{2tp_{t,I^{*}}f_{j}\left(\frac{p^{*}_{j}+\epsilon/2}{p^{*}_{I^{*}}}\right)(1-\delta)}\right)}{\Phi\left(-\sqrt{2tp_{t,I^{*}}f_{j}\left(\frac{p^{*}_{j}}{p^{*}_{I^{*}}}\right)(1+\delta)}\right)},$
where the second inequality follows from (23) and the monotonicity of $f_{j}$ and $f_{A_{t}}$, while the final equality uses the information balance (1).
We can pick a sufficiently small $\delta$ as a function of $\epsilon$ such that
$c_{1}\triangleq f_{j}\left(\frac{p^{*}_{j}+\epsilon/2}{p^{*}_{I^{*}}}\right)(1-\delta)>f_{j}\left(\frac{p^{*}_{j}}{p^{*}_{I^{*}}}\right)(1+\delta)\triangleq c_{2}.$
Then (24) follows from Gaussian tail upper and lower bounds (or $\frac{1}{t}\log\Phi\left(-\sqrt{t}x\right)\to-x^{2}/2$ as $t\to\infty$) and that $p_{t,I^{*}}$ is bounded away from 0 and 1 for sufficiently large $t$ (Lemma 9). This completes the proof. ∎
The result above suggests that the proportion ratios tend to self-correct over time. The following result formalizes this intuition.
Lemma 15.
Under TTTS-IDS, for any $\epsilon>0$, there exists $T_{\epsilon}\in\mathbb{M}$ such that for any $t\geq T_{\epsilon}$,
$\frac{w_{t,j}}{w_{t,I^{*}}}\leq\frac{p_{j}^{*}+\epsilon}{p_{I^{*}}^{*}},\quad\forall j\neq I^{*}.$
Proof.
Fix $\epsilon>0$ and $j\neq I^{*}$. By Lemma 14, there exists $\widetilde{T}_{\epsilon}\in\mathbb{M}$ such that for any $t\geq\widetilde{T}_{\epsilon}$ and $j\neq I^{*}$,
$\frac{w_{t,j}}{w_{t,I^{*}}}>\frac{p_{j}^{*}+\epsilon/2}{p_{I^{*}}^{*}}\quad\implies\quad\psi_{t,j}\leq\exp\left(-c_{\epsilon}t\right)+(K-1)\exp\left(-ct^{1/2}\right).$
Define
$\kappa_{\epsilon}\triangleq\sum_{\ell=0}^{\infty}\left[\exp\left(-c_{\epsilon}\ell\right)+(K-1)\exp\left(-c\ell^{1/2}\right)\right]<\infty.$
From now on, we consider $t\geq\widetilde{T}_{\epsilon}$. We consider two cases, and provide upper bounds on $\frac{w_{t,j}}{w_{t,I^{*}}}$ for each case.
  1. 1.
    The first case supposes that
    $\forall\ell\in\left\{\widetilde{T}_{\epsilon},\widetilde{T}_{\epsilon}+1,\ldots,t-1\right\}:\quad\frac{w_{\ell,j}}{w_{\ell,I^{*}}}\geq\frac{p_{j}^{*}+\epsilon/2}{p^{*}_{I^{*}}}.$
    We have
    $\Psi_{t,j}=\Psi_{\widetilde{T}_{\epsilon},j}+\sum_{\ell=\widetilde{T}_{\epsilon}}^{t-1}\psi_{\ell,j}$ $=\Psi_{\widetilde{T}_{\epsilon},j}+\sum_{\ell=\widetilde{T}_{\epsilon}}^{t-1}\psi_{\ell,j}\mathbf{1}\left(\frac{w_{\ell,j}}{w_{\ell,I^{*}}}\geq\frac{p_{j}^{*}+\epsilon/2}{p^{*}_{I^{*}}}\right)$
    $\leq\Psi_{\widetilde{T}_{\epsilon},j}+\sum_{\ell=\widetilde{T}_{\epsilon}}^{t-1}\left[\exp\left(-c_{\epsilon}\ell\right)+(K-1)\exp\left(-c\ell^{1/2}\right)\right]\leq\Psi_{\widetilde{T}_{\epsilon},j}+\kappa_{\epsilon}.$
    Take $T\in\mathbb{M}$ in Lemma 9, so for any $t\geq T$, $\Psi_{t,I^{*}}\geq b_{1}t$, and thus
    $\frac{w_{t,j}}{w_{t,I^{*}}}=\frac{\Psi_{t,j}}{\Psi_{t,I^{*}}}\leq\frac{\Psi_{\widetilde{T}_{\epsilon},j}+\kappa_{\epsilon}}{\Psi_{t,I^{*}}}\leq\frac{\widetilde{T}_{\epsilon}+\kappa_{\epsilon}}{b_{1}t}.$ (25)
  2. 2.
    The alternative case supposes that
    $\exists\ell\in\left\{\widetilde{T}_{\epsilon},\widetilde{T}_{\epsilon}+1,\ldots,t-1\right\}:\quad\frac{w_{\ell,j}}{w_{\ell,I^{*}}}<\frac{p_{j}^{*}+\epsilon/2}{p^{*}_{I^{*}}}.$
    We define
    $L_{t}\triangleq\max\left\{\ell\in\left\{\widetilde{T}_{\epsilon},\widetilde{T}_{\epsilon}+1,\ldots,t-1\right\}:\frac{w_{\ell,j}}{w_{\ell,I^{*}}}<\frac{p_{j}^{*}+\epsilon/2}{p^{*}_{I^{*}}}\right\}.$
    We have
    $\Psi_{t,j}$ $=\Psi_{L_{t},j}+\psi_{L_{t},j}+\sum_{\ell=L_{t}+1}^{t-1}\psi_{\ell,j}=\Psi_{L_{t},j}+\psi_{L_{t},j}+\sum_{\ell=L_{t}+1}^{t-1}\psi_{\ell,j}\mathbf{1}\left(\frac{w_{\ell,j}}{w_{\ell,I^{*}}}\geq\frac{p_{j}^{*}+\epsilon/2}{p^{*}_{I^{*}}}\right)$
    $\leq\frac{p_{j}^{*}+\epsilon/2}{p^{*}_{I^{*}}}\Psi_{L_{t},I^{*}}+1+\sum_{\ell=L_{t}+1}^{t-1}\left[\exp\left(-c_{\epsilon}\ell\right)+(K-1)\exp\left(-c\ell^{1/2}\right)\right]$
    $\leq\frac{p_{j}^{*}+\epsilon/2}{p^{*}_{I^{*}}}\Psi_{L_{t},I^{*}}+\left(\kappa_{\epsilon}+1\right),$
    where the first inequality uses $\frac{\Psi_{L_{t},j}}{\Psi_{L_{t},I^{*}}}=\frac{w_{L_{t},j}}{w_{L_{t},I^{*}}}<\frac{p_{j}^{*}+\epsilon/2}{p^{*}_{I^{*}}}$. Then by Lemma 9, for any $t\geq T$,
    $\frac{w_{t,j}}{w_{t,I^{*}}}\leq\frac{p_{j}^{*}+\epsilon/2}{p^{*}_{I^{*}}}\frac{\Psi_{L_{t},I^{*}}}{\Psi_{t,I^{*}}}+\frac{1+\kappa_{\epsilon}}{\Psi_{t,I^{*}}}\leq\frac{p_{j}^{*}+\epsilon/2}{p^{*}_{I^{*}}}+\frac{1+\kappa_{\epsilon}}{b_{1}t}.$ (26)
Putting (25) and (26) together, we find that for any $t\geq T_{1}$,
$\frac{w_{t,j}}{w_{t,I^{*}}}=\frac{\Psi_{t,j}}{\Psi_{t,I^{*}}}\leq\max\left\{\frac{\widetilde{T}_{\epsilon}+\kappa_{\epsilon}}{b_{1}t},\frac{p_{j}^{*}+\epsilon/2}{p^{*}_{I^{*}}}+\frac{\kappa_{\epsilon}+1}{b_{1}t}\right\}\leq\frac{p_{j}^{*}+\epsilon/2}{p^{*}_{I^{*}}}+\frac{\widetilde{T}_{\epsilon}+2\kappa_{\epsilon}+1}{t}.$
Then we have
$t\geq\frac{2\left(\widetilde{T}_{\epsilon}+2\kappa_{\epsilon}+1\right)}{\epsilon}\quad\implies\quad\frac{w_{t,j}}{w_{t,I^{*}}}\leq\frac{p_{j}^{*}+\epsilon}{p^{*}_{I^{*}}}.$
Taking $T_{\epsilon}\triangleq\max\left\{\widetilde{T}_{\epsilon},T,\frac{2\left(\widetilde{T}_{\epsilon}+2\kappa_{\epsilon}+1\right)}{\epsilon}\right\}$ completes the proof since Lemma 4 implies $T_{\epsilon}\in\mathbb{M}$. ∎
Now we are ready to complete the proof of Proposition 5.
Proof of Proposition 5.
Proposition 5 follows directly from Lemma 15 and Corollary 4. ∎
Appendix F Proof of Theorem 6
We overload the notation $\Delta_{\min}$ and $\Delta_{\max}$ for pure-exploration thresholding bandits as follows,
$\Delta_{\min}\triangleq\min_{i\in[K]}|\theta_{i}-\mathfrak{T}|\quad\text{and}\quad\Delta_{\max}\triangleq\max_{i\in[K]}|\theta_{i}-\mathfrak{T}|.$
The identification assumption ensures that $\Delta_{\min}>0$.
F.1 Sufficient Exploration
We first show that our algorithm sufficiently explore all arms.
Proposition 8.
There exists $T\in\mathbb{M}$ such that for any $t\geq T$, $\min_{i\in[K]}N_{t,i}\geq\sqrt{t/K}.$
Following Qin et al. [2017] and Shang et al. [2020], we define the following sets for any $t\in\mathbb{N}_{0}$ and $s\geq 0$:
$U_{t}^{s}\triangleq\{i\in[K]:N_{t,i}<s^{1/2}\}\quad\text{and}\quad V_{t}^{s}\triangleq\{i\in[K]:N_{t,i}<s^{3/4}\}.$
We let $\overline{U_{t}^{s}}=[K]\setminus U_{t}^{s}$ and $\overline{V_{t}^{s}}=[K]\setminus V_{t}^{s}$.
The following auxiliary result is needed to establish Proposition 8.
Lemma 16.
There exists $S\in\mathbb{M}$ such that for any $s\geq S$, if $U_{t}^{s}$ is nonempty, then $I_{t}\in V_{t}^{s}$.
Proof.
By the definition of $W_{1}$ and Lemma 7, there exists $S_{1}\in\mathbb{M}$ such that for any $s\geq S_{1}$ and any $t\in\mathbb{N}_{0}$, the condition $i\in\overline{V_{t}^{s}}$ implies $|\theta_{t,i}-\mathfrak{T}|\geq\Delta_{\min}/2$. On the other hand, by the definition of $W_{1}$, the condition $j\in U_{t}^{s}$ implies $|\theta_{t,j}-\mathfrak{T}|\leq\Delta_{\max}+\sigma W_{1}$. Hence, there exists $S_{2}\in\mathbb{M}$ such that for $s\geq S_{1}$, $p_{t,j}d(\theta_{t,j},\mathfrak{T})<p_{t,i}d(\theta_{t,i},\mathfrak{T})$, which implies $I_{t}\notin\overline{V_{t}^{s}}$. This completes the proof. ∎
We now complete the proof of Proposition 8.
Proof of Proposition 8.
Proposition 8 follows from the property of our algorithm, established in the preceding lemma, together with the arguments used in the proof of Lemma 11 in Shang et al. [2020]. ∎
F.2 Strong Convergence to Optimal Proportions: Completing the Proof of Theorem 6
Given the sufficient exploration ensured by our algorithm, we can establish the following auxiliary result.
Lemma 17.
For any arm $i\in[K]$,
$\frac{C_{t,i}}{D_{t,i}}\xrightarrow{\mathbb{M}}1\quad\text{where}\quad C_{t,i}=p_{t,i}d(\theta_{t,i},\mathfrak{T})\text{ and }D_{t,i}\triangleq p_{t,i}d(\theta_{i},\mathfrak{T}).$
Proof.
This result follows directly from the same arguments used in the proof of Lemma 13. ∎
To complete the proof of Theorem 6, we define the over-sampled set at time $t\in\mathbb{N}_{0}$ with $\epsilon>0$ as
$O_{t,\epsilon}\triangleq\left\{j\in[K]:w_{t,i}>p_{i}^{*}+\epsilon\right\},$
where $w_{t,i}=\frac{\Psi_{t,i}}{t}$ denotes the average sampling effort, based on the sampling probabilities. The next result establishes that if a suboptimal arm is over-sampled at some sufficiently large time step, it will have zero probability of being sampled at that time.
Lemma 18.
For any $\epsilon>0$, there exists a random time $T_{\epsilon}\in\mathbb{M}$ such that for any $t\geq T_{\epsilon}$ and $i\in[K]$,
$i\in O_{t,\epsilon}\quad\implies\quad\psi_{t,i}=0.$
Proof.
There exists $T_{1}\in\mathbb{M}$ such that for any $t\geq T_{1}$ and $i\in[K]$,
$i\in O_{t,\epsilon}\quad\implies\quad p_{t,i}>p_{i}^{*}+\epsilon/2\quad\text{and thus}\quad\exists J_{t}\in[K]\text{ s.t. }p_{t,J_{t}}\leq p^{*}_{J_{t}}.$
Hence,
$\frac{C_{t,i}}{C_{t,J_{t}}}=\frac{p_{t,i}\left(\theta_{t,i}-\mathfrak{T}\right)^{2}}{p_{t,J_{t}}\left(\theta_{t,J_{t}}-\mathfrak{T}\right)^{2}}$ $\geq\frac{p_{t,i}\left(\theta_{i}-\mathfrak{T}\right)^{2}(1-\delta)}{p_{t,J_{t}}\left(\theta_{J_{t}}-\mathfrak{T}\right)^{2}(1+\delta)}\geq\frac{\left(p_{i}^{*}+\frac{\epsilon}{2}\right)\left(\theta_{i}-\mathfrak{T}\right)^{2}(1-\delta)}{p_{J_{t}}^{*}\left(\theta_{J_{t}}-\mathfrak{T}\right)^{2}(1+\delta)}>1,$
where the first inequality applies the preceding lemma, along with choosing a sufficient small $\delta$ as a function of $\epsilon$, so that the last inequality holds. ∎
The result above suggests that the allocation tends to self-correct over time. The following result formalizes this intuition.
Lemma 19.
For any $\epsilon>0$, there exists $T_{\epsilon}\in\mathbb{M}$ such that for any $t\geq T_{\epsilon}$,
$w_{t,i}\leq p_{i}^{*}+\epsilon,\quad\forall i\in[K].$
Proof.
The preceding result establishes that there exists $\widetilde{T}_{\epsilon}\in\mathbb{M}$ such that for any $t\geq\widetilde{T}_{\epsilon}$ and $i\in[K]$,
$i\in O_{t,{\epsilon}}\quad\implies\quad\psi_{t,i}=0.$
Fix an arm $i$. There are two possible cases.
  1. 1.
    $\exists\ell\in\left\{\widetilde{T}_{\epsilon},\widetilde{T}_{\epsilon}+1,\ldots,t-1\right\},w_{\ell,i}\leq p_{i}^{*}+\epsilon$.
    Define
    $L_{t}\triangleq\max\left\{\ell\in\left\{\widetilde{T}_{\epsilon},\widetilde{T}_{\epsilon}+1,\ldots,t-1\right\}:w_{\ell,i}\leq p_{i}^{*}+\epsilon\right\}.$
    Then,
    $w_{t,i}=\frac{\Psi_{t,i}}{t}=\frac{\Psi_{L_{t},i}}{t}\leq\frac{\Psi_{L_{t},i}}{L_{t}}=w_{L_{t},i}\leq p_{i}^{*}+\epsilon.$
  2. 2.
    $\forall\ell\in\left\{\widetilde{T}_{\epsilon},\widetilde{T}_{\epsilon}+1,\ldots,t-1\right\},w_{\ell,i}>p_{i}^{*}+\epsilon$.
    We have
    $w_{t,i}=\frac{\Psi_{t,i}}{t}=\frac{\Psi_{\widetilde{T}_{\epsilon},i}}{t}.$
    If $t\geq T_{\epsilon,i}\triangleq\max\left\{\widetilde{T}_{\epsilon},\frac{\Psi_{\widetilde{T}_{\epsilon},i}}{p_{i}^{*}+\epsilon}\right\}$, then $w_{t,i}\leq p_{i}^{*}+\epsilon$.
Taking $T_{\epsilon}=\max_{i\in[K]}T_{\epsilon,i}$ completes the proof. ∎
We now complete the proof of Theorem 6.
Proof of Theorem 6.
The preceding result establishes that no arm is over-sampled for sufficiently large $t$. Since the sampling proportions must sum to one, this also ensures that no arm is under-sampled for sufficiently large $t$. This completes the proof. ∎
Appendix G Miscellaneous Results and Proofs
In this appendix, we collect auxiliary proofs and useful tools used in our discussion.
G.1 Facts about the Exponential Family
We restrict our attention to reward distribution following a one-dimensional natural exponential family. In particular, a natural exponential family has identity natural statistic in canonical form, i.e.,
$p(y\mid\eta)=b(y)\exp\{\eta y-A(\eta)\},\quad\eta\in\mathcal{T}=\left\{\eta:\int b(y)e^{\eta y}\mathrm{d}y<\infty\right\}$
[…]
$\theta(\eta)=\mathbb{E}_{\eta}[Y_{t,i}]=A^{\prime}(\eta),\quad\text{and}\quad A^{\prime\prime}(\eta)=\text{Var}_{\eta}(Y_{t,i})>0.$
We immediately have the following lemma.
Lemma 20.
At any $\eta\in\mathcal{T}^{o}$, $\theta(\eta)$ is strictly increasing and $A(\eta)$ is strictly convex.
Lemma 20 implies that $\theta(\cdot)$ is a one-to-one mapping. Consequently, we can use $\theta$ to denote the distribution $p(\cdot\mid\eta)$ for convenience. Let $\theta_{1}$ and $\theta_{2}$ be the mean reward of the distribution $p(\cdot\mid\eta_{1})$ and $p(\cdot\mid\eta_{2})$. Denote by $d(\theta_{1},\theta_{2})$ the Kullback-Leibler (KL) divergence from $p(\cdot\mid\eta_{1})$ to $p(\cdot\mid\eta_{2})$. We have convenient closed-form expressions of $d(\cdot,\cdot)$ and its partial derivative.
Lemma 21.
The partial derivatives of the KL divergense of a one-dimensional natural exponential family satisfies
$d(\theta_{1},\theta_{2})=A(\eta_{2})-A(\eta_{1})-\theta_{1}(\eta_{2}-\eta_{1}),$
$\frac{\partial d(\theta_{1},\theta_{2})}{\partial\theta_{1}}=\eta_{1}-\eta_{2},\quad\frac{\partial d(\theta_{1},\theta_{2})}{\partial\theta_{2}}=(\theta_{2}-\theta_{1})\frac{\mathrm{d}\eta(\theta_{2})}{\mathrm{d}\theta_{2}}.$
G.2 Proof of Lemma 1
Proof.
When the reward distribution follows a single-parameter exponential family, the KL divergence $d(m,m^{\prime})$ is continuous in $(m,m^{\prime})$ and strictly convex in $m^{\prime}$. A direct application of Wang et al. [2021, Theorem 4 and Lemma 7] yields a unique $\bm{\vartheta}^{x}$ that achieves the infimum in (8), as well as the differentiability of $C_{x}$. For the concavity, note that $C_{x}$ is the infimum over a family of linear functions in $\bm{p}$. This directly implies the concavity of $C_{x}$ in $\bm{p}$; see Boyd and Vandenberghe [2004, Chapter 3.2].
We say a function $f$ is homogeneous of degree one if it satisfies $f(s\bm{p})=sf(\bm{p})$ for any $s\geq 0$. The fact that $C_{x}(\bm{p})$ is homogeneous of degree one follows directly from its definition in (8). Combined with differentiability, the PDE characterization follows from Euler’s theorem for homogeneous functions; see Apostol [1969, p. 287] for example. ∎
G.3 Properties of $C_{x}(\cdot)$ in BAI and Best-$k$-Arm Identification
We provide useful properties of the functions $C_{j}(\bm{p})$ for BAI and the functions $C_{i,j}(\bm{p})$ for best-$k$ identification; see Appendix A for explicit formula of these functions.
Lemma 22.
For BAI and any component-wise strictly positive $\bm{p}$, the minimizer $\bm{\vartheta}^{j}$ for (8) is given by
$\vartheta_{j}^{j}=\vartheta_{I^{*}}^{j}=\bar{\theta}_{I^{*},j}(\bm{p})\triangleq\frac{p_{I^{*}}\theta_{I^{*}}+p_{j}\theta_{j}}{p_{I^{*}}+p_{j}},\quad\forall j\neq I^{*},\quad\text{and}\quad\vartheta_{k}^{j}=0,\quad\forall k\neq I^{*},j.$
In particular, $C_{j}(\bm{p})$ is concave in $\bm{p}$, strictly increasing in $p_{I^{*}}$ and $p_{j}$, and
$\frac{\partial C_{j}(\bm{p})}{\partial p_{I^{*}}}=d(\theta_{I^{*}},\bar{\theta}_{I^{*},j})\quad\text{and}\quad\frac{\partial C_{j}(\bm{p})}{\partial p_{j}}=d(\theta_{j},\bar{\theta}_{I^{*},j}).$
For best-$k$ identification, the minimizer $\bm{\vartheta}^{(i,j)}$ for (8) is given by
$\vartheta_{i}^{(i,j)}=\vartheta_{j}^{(i,j)}=\bar{\theta}_{i,j}(\bm{p})\triangleq\frac{p_{i}\theta_{i}+p_{j}\theta_{j}}{p_{i}+p_{j}},\quad\forall(i,j)\in\mathcal{I}\times\mathcal{I}^{c},\quad\text{and}\quad\vartheta_{k}^{(i,j)}=0,\quad\forall k\neq i,j.$
In particular, $C_{i,j}(\bm{p})$ is concave in $\bm{p}$, strictly increasing in $p_{i}$ and $p_{j}$, and
$\frac{\partial C_{i,j}(\bm{p})}{\partial p_{i}}=d(\theta_{i},\bar{\theta}_{i,j})\quad\text{and}\quad\frac{\partial C_{i,j}(\bm{p})}{\partial p_{j}}=d(\theta_{j},\bar{\theta}_{i,j}).$
Proof.
Note that $C_{j}$ for BAI is equivalent to $C_{I^{*},j}$ for best-$k$ identification, where $I^{*}$ is the best arm. To avoid redundancy, we focus on the functions $C_{i,j}$. To find the minimizer, we note that
$C_{i,j}(\bm{p})\triangleq\inf_{\bm{\vartheta}\in\mathrm{Alt}_{i,j}}\sum_{i}p_{i}d(\theta_{i},\vartheta_{i})$ $=\inf_{\vartheta_{j}>\vartheta_{i}}\left\{p_{i}d(\theta_{i},\vartheta_{i})+p_{j}d(\theta_{j},\vartheta_{j})\right\}$
$=\inf_{\widetilde{\theta}}\left\{p_{i}d(\theta_{i},\widetilde{\theta})+p_{j}d(\theta_{j},\widetilde{\theta})\right\}$
$=p_{i}d(\theta_{i},\bar{\theta}_{i,j})+p_{j}d(\theta_{j},\bar{\theta}_{i,j}),\quad\hbox{where}\quad\bar{\theta}_{i,j}\triangleq\frac{p_{i}\theta_{i}+p_{j}\theta_{j}}{p_{i}+p_{j}}.$ (27)
In the second line, we set $\vartheta_{k}=\theta_{k}$ for all $k\neq i,j$; the third line follows from the monotonicity of the KL divergence, and the last equality results from analytically solving the one-dimensional convex optimization. The partial derivatives then follows directly from Lemma 1.
Alternatively, one may calculate the derivatives directly using Lemma 21 and the definition of $C_{i,j}$ in (27), as follows
$\frac{\partial C_{i,j}(\bm{p})}{\partial p_{i}}$ $=d(\theta_{i},\bar{\theta}_{i,j})+p_{i}\frac{\partial d(\theta_{i},\bar{\theta})}{\partial p_{i}}+p_{j}\frac{\partial d(\theta_{j},\bar{\theta})}{\partial p_{i}}$
$=d(\theta_{i},\bar{\theta}_{i,j})+\frac{\partial\bar{\theta}}{\partial p_{i}}\cdot\left[p_{i}\frac{\partial d(\theta_{i},\bar{\theta})}{\partial\bar{\theta}}+p_{j}\frac{\partial d(\theta_{j},\bar{\theta})}{\partial\bar{\theta}}\right]=d(\theta_{i},\bar{\theta}_{i,j}).$
The last equality follows from the first-order optimality condition for solving the convex optimization problem in the third line of (27). The calculation of $\frac{\partial C_{i,j}(\bm{p})}{\partial p_{j}}$ proceeds similarly. ∎
G.4 Proof of Corollary 1
Assuming strictly positiveness of $\bm{p}^{*}$, recall from (18b) that
$\lambda-\sum_{x\in\mathcal{X}}\mu_{x}\frac{\partial C_{x}(\bm{p})}{\partial p_{i}}=0,\quad\forall i\in[K],$
where we used the fact that $\iota_{i}=0$ due to complementary slackness. By the definition of
$\mathcal{X}_{i}\triangleq\left\{x\in\mathcal{X}(\bm{\theta}):\frac{\partial C_{x}(\bm{p})}{\partial p_{i}}>0\right\},$
we have
$\lambda-\sum_{x\in\mathcal{X}_{i}}\mu_{x}\frac{\partial C_{x}(\bm{p})}{\partial p_{i}}=0,\quad\forall i\in[K].$
Notice that $\lambda=\phi=\Gamma^{*}_{\bm{\theta}}>0$ from the proof of Theorem 4 and that $\frac{\partial C_{x}(\bm{p})}{\partial p_{i}}>0$, there must exist a $x_{0}\in\mathcal{X}_{i}$ such that $\mu_{x}>0$. The complementary-slackness condition then implies that $C_{x_{0}}(\bm{p}^{*})=\phi=\Gamma^{*}_{\bm{\theta}}$. Consequently,
$\Gamma_{\bm{\theta}}^{*}=\min_{x\in\mathcal{X}_{i}}C_{x}(\bm{p}^{*}),\quad\hbox{for all }i\in[K].$
Appendix H Existing Algorithms in the Literature
In this appendix, we briefly review existing algorithms designed for the best-$k$ identification problem in the fixed-confidence settings. These methods serves as benchmarks for our experiments.
UGapE [Gabillon et al., 2012].
At each round $t$, the algorithm computes the upper confidence bound $U_{i}(t)$ and lower confidence bound $L_{i}(t)$ of the mean reward for each arm $i$:
$U_{i}(t)$ $=\max\{q\in[\theta_{t,i},\infty):T_{t,i}d(\theta_{t,i},q)\leq\beta_{t,\delta}\},$ (28)
$L_{i}(t)$ $=\min\{q\in(-\infty,\theta_{t,i}]:T_{t,i}d(\theta_{t,i},q)\leq\beta_{t,\delta}\},$
where $\beta_{t,\delta}$ is the exploration rate we specify to be $\log\left((\log t+1)/\delta\right)$. Then it computes the gap index $B_{i}(t)=\max_{i^{\prime}\neq i}^{k}U_{i^{\prime}}(t)-L_{i}(t)$ for each arm $i$, where $\max^{k}$ denotes the operator returning the $k$-th largest value. Based on $B_{i}(t)$, it evaluates a set of $k$ arms: $J(t)=\{j:\operatorname*{argmin}^{j}B_{j}(t)\}$. From $J(t)$, it computes two critical arms $u_{t}=\operatorname*{argmax}_{i\notin J_{t}}U_{i}(t)$ and $l_{t}=\operatorname*{argmin}_{i\in J_{t}}L_{i}(t),$ and samples the one with less allocation. The algorithm stops when $\max_{i\in J(t)}B_{i}(t)>0$.
KL-LUCB [Kaufmann and Kalyanakrishnan, 2013].
At each round $t$, the algorithm first computes the upper and lower confidence bounds $U_{i}(t)$ and $L_{i}(t)$ as in (28). It then samples two arms, defined by
$u_{t}=\arg\max_{i\notin\widehat{\mathcal{I}}_{t}}U_{i}(t),\quad l_{t}=\arg\min_{i\in\widehat{\mathcal{I}}_{t}}L_{i}(t).$
The algorithm stops when $U_{u_{t}}(t)>L_{l_{t}}(t).$
Lazy-Mirror-Ascent [Ménard, 2019].
The algorithm treats the lower-bound problem as a generic maximization problem $\max_{\bm{w}}F(\bm{w})$. It uses an allocation rule based on online lazy mirror ascent. At each round $t$, it maintains allocation weights $\bm{w}(t)$ and updates them by
$\widetilde{\bm{w}}(t+1)=\arg\max_{\bm{w}}\Bigl{[}\eta_{t+1}\sum_{s=K}^{t}\bm{w}\cdot\nabla F\bigl{(}\widetilde{\bm{w}}(s);\bm{\mu}(s)\bigr{)}-\mathrm{KL}(\bm{w},\bm{\pi})\Bigr{]},$
and
$\bm{w}(t+1)=(1-\gamma_{t})\widetilde{\bm{w}}(t+1)+\gamma_{t}\bm{\pi},$
where $\bm{\mu}(s)$ is the vector of empirical mean estimates and $\bm{\pi}$ is the uniform distribution. There are two tuning parameters, $\eta_{t}$ and $\gamma_{t}$. Based on the code by Wouter M. Koolen (https://bitbucket.org/wmkoolen/tidnabbil) and the suggestion in Ménard [2019], we choose
$\eta_{t}=\frac{\sqrt{\log(K/t)}}{F^{*}},\quad\gamma_{t}=\frac{1}{4\sqrt{t}}.$
Wouter M. Koolen comments that, although setting $F^{*}$ requires knowledge of the true arm means, it is critical for performance. In the appendix of Wang et al. [2021], the authors also note that this method is very sensitive to the learning rate $\eta_{t}$; the current choice is only for experimental comparison and cannot be used in real-world scenarios.
$m$-LinGapE [Réda et al., 2021a].
At each round $t$, the algorithm first computes the same upper and lower confidence bounds $U_{i}(t)$ and $L_{i}(t)$ as in (28). It then computes the gap index
$B_{i,j}(t)=U_{i}(t)-L_{j}(t)$
for each pair of arms $(i,j)$. It chooses $J(t)$ as the set of the $k$ arms with the largest empirical means. From $J(t)$, it identifies two critical arms,
$u_{t}=\operatorname*{argmax}_{j\in J(t)}\max_{i\notin J(t)}B_{i,j}(t),\quad l_{t}=\operatorname*{argmax}_{i\notin J(t)}B_{i,u_{t}}(t),$
and samples the arm that would most reduce the variance of the empirical mean difference between them (the “greedy selection rule” in Réda et al. 2021a). The algorithm stops when $B_{l_{t},u_{t}}(t)>0$.
MisLid [Réda et al., 2021b].
The algorithm is derived from the lower-bound maximin problem and achieves asymptotic optimality. It maintains a learner over the probability simplex of arms. At each round $t$, it computes a gain function and updates the learner, then applies a tracking rule to sample an arm. The algorithm adopts AdaHedge as the learner and defines the empirical gain function
$g_{t}=\operatorname*{argmin}_{\bm{\lambda}\in\mathrm{Alt}(\bm{\mu}^{t})}\sum_{i=1}^{K}p^{t}_{i}\bigl{(}\mu^{t}_{i}-\lambda\bigr{)}^{2},$
where $\bm{p}^{t}$ is the probability distribution returned by AdaHedge, $\bm{\mu}^{t}$ is the vector of empirical mean estimates, and $\mathrm{Alt}(\bm{\mu}^{t})$ is the set of parameters under which the top-$k$ arm sets differ from that of $\bm{\mu}^{t}$. As suggested by Réda et al. [2021b], the empirical gain function reduces the sample complexity in practice compared to the optimistic gain function proposed by theory. For the tracking rule, the algorithm uses C-Tracking. For the stopping rule, it uses the GLRT statistic with threshold $\log\bigl{(}(\log t+1)/\delta\bigr{)}.$
FWS [Wang et al., 2021].
The algorithm utilizes the lower-bound maximin problem. At each round $t$, it performs either forced exploration or a Frank–Wolfe update. Due to the nonsmoothness of the objective function, it constructs an $r_{t}$-subdifferential subspace and then solves a zero-sum game. Here, $r_{t}$ is a tuning parameter; we choose $r_{t}=t^{-0.9}/K$ as suggested in Wang et al. [2021]. For the stopping rule, we adopt the GLRT statistic with threshold $\log\bigl{(}(\log t+1)/\delta\bigr{)}$.
Appendix I Additional Numerical Experiments
In this appendix, we provide additional numerical experiments, comparing our algorithms with the following algorithms from the literature: UGapE [Gabillon et al., 2012], SAR [Bubeck et al., 2013], OCBAss [Gao and Chen, 2015] and OCBASS [Gao and Chen, 2016] in the fixed-budget setting. We consider the best-$5$-arm identification problem for a Gaussian bandit with $\theta_{i}=1-0.05\times(i-1)$ and $\sigma_{i}^{2}=1/4$ for $i=[20]$.
Figure 3: Fix-budget performance estimated over $10^{7}$ independent replications.
In Figure 3, we present the fixed-budget performance of our proposed TS-KKT-IDS algorithm, along with benchmark methods. In the subplot of Figure 3, we provide a zoomed-out illustration on a logarithmic scale, which suggests a superior convergence rate of the PICS under our algorithm.
Figure 4: Compare the optimal allocation rate with the empirical allocation rate of the algorithms.
In Figure 4, we compare the optimal allocation rate with the empirical allocation rate of the algorithms, given a budget of $10^{6}$. We observe that both TS-KKT-IDS and OCBAss (omitted) allocate nearly exactly at the optimal rate. In this simpler case, OCBAss satisfies the correct sufficient condition and thus also converges to the optimal rate. In contrast, OCBASS, lacking the key step of balancing the sum-of-squared allocations, does not converge to the optimal allocation rate, though it remains sufficiently close. The UGapE and SAR algorithms deviate more significantly from the optimal allocation rate, although their empirical performance in terms of PICS is still respectable.
Figure 5: The probability of false selection for varying values of $k$ with a fixed budget of $1500$, estimated over $30000$ independent replications.
In Figure 5, we consider the case where the number of best arms to be identified varies from $k=1$ to $19$, displaying the PICS for each $k$ with a fixed budget of $1500$ samples. We observed the our TS-KKT-IDS algorithm consistently outperforms the alternatives across the range of $k$.